### v1

# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

Abstract
The Partition Principle $\mathsf{PP}$ asserts that whenever there is a surjection $A\twoheadrightarrow B$, there is an injection $B\hookrightarrow A$. Russell conjectured in 1906 that $\mathsf{PP}$ is equivalent to the Axiom of Choice $\mathsf{AC}$; while $\mathsf{AC}\Rightarrow\mathsf{PP}$ is immediate, the converse has remained open. We show that $\mathsf{PP}$ does not imply $\mathsf{AC}$ by constructing a transitive model of $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$.
Starting from a Cohen symmetric model $\mathcal{N}$ of $\operatorname{Add}(\omega,\omega_{1})$ built with a countable-support symmetry filter, we fix parameters $S:=A^{\omega}$ and $T:=\mathcal{P}(S)$ and perform a class-length countable-support symmetric iteration. At successor stages we use orbit-symmetrized packages that split targeted surjections, yielding $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$, while preserving $\mathsf{DC}$ and ensuring that $A$ remains non-well-orderable. A diagonal-cancellation/diagonal-lift infrastructure supplies a proper $\omega_{1}$-complete normal filter at limit stages. Finally, Ryan–Smith localization shows that under $\mathsf{SVC}^{+}(T)$, $\mathsf{PP}$ is equivalent to $\mathsf{PP}\!\restriction T\wedge\mathsf{AC}_{\mathsf{WO}}$, so the final model satisfies $\mathsf{PP}$ but not $\mathsf{AC}$.
1 Introduction
Whether the Partition Principle implies the Axiom of Choice is sometimes called “the oldest open problem in set theory.” The Partition Principle ($\mathsf{PP}$) asserts that whenever $A$ surjects onto $B$, one has $|B|\leq|A|$. Russell conjectured in 1906 that $\mathsf{PP}$ is equivalent to $\mathsf{AC}$, but no proof was given [1]. The converse is immediate: choice provides sections for all surjections. The question of whether $\mathsf{PP}$ and $\mathsf{AC}$ are equivalent over $\mathsf{ZF}$ has remained open.
Main Result. We resolve this question in the negative by constructing a model of
$\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC},$
thereby establishing that the Partition Principle does not imply the Axiom of Choice.
Theorem.
The theory $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$ is consistent relative to $\mathsf{ZF}$.
Strategy.
The construction proceeds in two stages.
Stage 1: Seed model. We build a Cohen symmetric model $\mathcal{N}$ over $\operatorname{Add}(\omega,\omega_{1})$ using countable supports, producing a model of $\mathsf{ZF}+\mathsf{DC}$ containing a non-well-orderable set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$ of Cohen reals. The fixed set $S$ witnesses $\mathsf{SVC}(S)$, which will drive the localization argument.
Stage 2: Symmetric iteration. We perform a class-length countable-support symmetric iteration over $\mathcal{N}$. At each successor stage, we force sections for surjections $f:Y\twoheadrightarrow X$ with $X,Y\subseteq T:=\mathcal{P}(S)$, as well as right inverses witnessing $\mathsf{AC}_{\mathsf{WO}}$. The iteration preserves $\mathsf{DC}$ and the non-well-orderability of $A$.
Localization. The global Partition Principle follows from Ryan–Smith’s localization theorem [13], which shows that under $\mathsf{SVC}^{+}(T)$, the principle $\mathsf{PP}$ is equivalent to the conjunction of $\mathsf{PP}\!\restriction\!T$ (the restriction to surjections between subsets of $T$) and $\mathsf{AC}_{\mathsf{WO}}$.
Technical innovations.
Two devices are essential to making the iteration work.
(1) Orbit-symmetrized packages. A naïve attempt to force a section for a single surjection $f$ fails because the action of $\operatorname{Sym}(\omega_{1})$ on conditions does not extend to an action on the forcing poset (automorphisms would need to map $\mathbb{Q}_{f}$ to $\mathbb{Q}_{\pi\cdot f}$, a different poset). We resolve this by forcing with the full orbit package
$\mathbb{Q}_{[f]}:=\prod^{\mathrm{cs}}_{g\in[f]}\mathbb{Q}_{g},$
the countable-support product over all $g$ in the $\operatorname{Sym}(\omega_{1})$-orbit of $f$. This ensures that every $\pi\in\operatorname{Sym}(\omega_{1})$ acts by permuting coordinates and transporting structure, yielding a genuine automorphism of $\mathbb{Q}_{[f]}$.
(2) Diagonal-cancellation symmetries. Orbit-symmetrization alone does not make the generic section $s_{f}$ hereditarily symmetric: the naïve action of $\pi\in\operatorname{Fix}(E)$ on conditions still moves the $f$-coordinate generic. We introduce diagonal-cancellation automorphisms $\widehat{\pi}^{\,D}$ that act as the identity on a protected set $D$ of coordinates (including $f$) while acting as $\pi$ elsewhere. The stage filter includes subgroups $\Delta^{\uparrow}(E,\{f\})$ generated by these diagonal lifts, ensuring that the generic section is stabilized by a filter subgroup and hence belongs to the symmetric model. Crucially, diagonal lifts still act as $\pi$ on the Cohen base (stage 0), so they swap Cohen reals $c_{\beta}\leftrightarrow c_{\gamma}$ for $\beta,\gamma\notin E$, preserving the $\neg\mathsf{AC}$ witness.
Conventions.
Throughout, $\mathsf{DC}$ denotes $\mathsf{DC}_{\omega}$, i.e., dependent choice for $\omega$-sequences. We write $\mathsf{PP}\!\restriction\!T$ for the restriction of $\mathsf{PP}$ to surjections $f:Y\twoheadrightarrow X$ with both $X,Y\subseteq T$.
External tools.
We rely on three external frameworks, imported as black boxes.
(1) Symmetric iteration machinery. For the countable-support symmetric iteration framework—including successor and limit constructions, filter generation, $\omega_{1}$-completeness at limits, and the hereditarily-symmetric calculus—we cite Gilson [3]. See Remark 2.1 for the specific API items we invoke.
(2) DC preservation. For preservation of Dependent Choice in symmetric extensions, we follow Karagila [6] and Karagila–Schilhan [8], building on classical work of Pincus [10]. The criterion we use is: $\omega_{1}$-closed forcing with $\omega_{1}$-complete filter preserves $\mathsf{DC}$.
(3) Localization of $P$. We use Ryan–Smith’s local reflection framework [13], which provides the equivalence $\mathsf{PP}\Longleftrightarrow\mathsf{PP}\!\restriction\!T+\mathsf{AC}_{\mathsf{WO}}$ under $\mathsf{SVC}^{+}(T)$. For background on $\mathsf{SVC}$ and Hartogs–Lindenbaum spectra, see [12, 2].
For general background on forcing and symmetric extensions we follow Jech [4] and Kunen [9].
Outline.
Section 3 constructs the seed model $\mathcal{N}$ and verifies $\mathsf{DC}$, $\neg\mathsf{AC}$, and $\mathsf{SVC}(S)$. Section 4 recalls the Ryan–Smith localization theorem and defines the target principles $\mathsf{PP}\!\restriction\!T$ and $\mathsf{AC}_{\mathsf{WO}}$. Subsection 4.1 and Subsection 4.2 define the package forcings $\mathbb{Q}_{f}$ and $\mathbb{R}_{f}$ for forcing sections and right inverses. Section 5 constructs the symmetric iteration, including the orbit-symmetrization and diagonal-cancellation machinery, and assembles the final model $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$.
2 Preliminaries and black-box iteration technology
We use the general symmetric-iteration machinery from [3] (and ultimately from Karagila’s iteration framework [5]). Here we only record the “API” we will use later. We treat the following as black boxes; the precise hypotheses and proofs are available in [3].
Remark 2.1 (Iteration API to be used later).
We will use, as needed, the following results from [3].
  1. (a)
    Set-length and class-length versions of symmetric iterations (cf. [3, §2.1]).
  2. (b)
    Definitions and basic properties of countable-support symmetric iterations (cf. [3, Def. 3.1]).
  3. (c)
    A $\mathsf{DC}$-preservation theorem for the relevant support (cf. [3, Thm. 3.20]).
  4. (d)
    Generation and completeness/normality of limit filters in the relevant cofinality cases (cf. [3, Thm. 5.5]).
  5. (e)
    A characterization of hereditarily symmetric names by “small support” (cf. [3, Lem. 3.14]).
We emphasize that at each stage $\lambda$ the iteration produces a single ambient group $\mathscr{G}_{\lambda}$ acting by automorphisms on the full initial-segment forcing $\mathbb{P}_{\lambda}$ (API item (b)), so the action continues to move the original Cohen coordinates; we are not iterating by repeatedly forcing over intermediate symmetric models in isolation.
Our main tool for reducing the global Partition Principle to a local instance is the "Small Violations of Choice" framework.
Definition 2.2 (Small Violations of Choice, [13]).
Let $S$ be a set.
  1. 1.
    We say that $\mathsf{SVC}(S)$ holds if for every set $X$, there exists an ordinal $\eta$ such that there is a surjection $f\colon S\times\eta\twoheadrightarrow X$ (equivalently, $|X|\leq^{*}|S\times\eta|$).
  2. 2.
    We say that $\mathsf{SVC}^{+}(S)$ holds if for every set $X$, there exists an ordinal $\eta$ such that there is an injection $f\colon X\hookrightarrow S\times\eta$ (equivalently, $|X|\leq|S\times\eta|$).
The following facts from [13] drive our localization strategy:
Fact 2.3.
For any set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$.
Theorem 2.4 (Ryan–Smith Localization, [13]).
Assume $\mathsf{SVC}^{+}(T)$. Then $\mathsf{PP}$ is equivalent to the conjunction of $\mathsf{PP}\!\restriction\!T$ and $\mathsf{AC}_{\mathsf{WO}}$.
Our strategy is to construct a model satisfying $\mathsf{SVC}(S)$ (via a Cohen seed) and then force $\mathsf{PP}\!\restriction\!\mathcal{P}(S)$ and $\mathsf{AC}_{\mathsf{WO}}$ while preserving $\mathsf{SVC}(S)$.
3 The Cohen symmetric seed model
We fix a ground model $V\models\mathsf{ZFC}$. Let
$\mathbb{P}\;=\;\operatorname{Add}(\omega,\omega_{1})\;:=\;\operatorname{Fn}(\omega_{1}\times\omega,2,{<}\omega),$
$\dot{c}_{\alpha}:=\{\langle\check{n},p\rangle:p(\alpha,n)=1\},c_{\alpha}=\dot{c}^{G}_{\alpha}$
the standard forcing to add $\omega_{1}$-many Cohen reals, presented as finite partial functions $p:\omega_{1}\times\omega\to 2$.
Lemma 3.1 (Cohen reals are pairwise distinct).
For $\alpha\neq\beta<\omega_{1}$, the Cohen reals $c_{\alpha}$ and $c_{\beta}$ are distinct.
Proof.
Fix $\alpha\neq\beta<\omega_{1}$. Let
$D_{\alpha,\beta}=\{p\in\operatorname{Add}(\omega,\omega_{1}):\exists n\,(p(\alpha,n)\neq p(\beta,n))\}.$
Given any $p$, choose $n$ such that neither $(\alpha,n)$ nor $(\beta,n)$ is in $\operatorname{dom}(p)$, and extend $p$ to $q\leq p$ with $q(\alpha,n)=0$ and $q(\beta,n)=1$. Then $q\in D_{\alpha,\beta}$. Hence $D_{\alpha,\beta}$ is dense, so any generic $G$ meets it and therefore $c_{\alpha}\neq c_{\beta}$. ∎
3.1 The forcing, group action, and the countable-support filter
Definition 3.2 (Automorphisms of $\operatorname{Add}(\omega,\omega_{1})$).
Let $\mathscr{G}:=\operatorname{Sym}(\omega_{1})$ be the full permutation group of $\omega_{1}$. Each $\pi\in\mathscr{G}$ acts on $\mathbb{P}$ by
$(\pi p)(\pi(\alpha),n)\;=\;p(\alpha,n),$
equivalently $(\pi p)=\{(\pi(\alpha),n,i):(\alpha,n,i)\in p\}$. This extends (recursively) to an action on $\mathbb{P}$-names in the usual way.
Definition 3.3 (Countable-support filter of subgroups).
For $E\subseteq\omega_{1}$, write
$\operatorname{Fix}(E)\;:=\;\{\pi\in\mathscr{G}:\pi\!\restriction E=\mathrm{id}_{E}\}.$
Let $\mathscr{F}$ be the normal filter of subgroups of $\mathscr{G}$ generated by the family $\{\operatorname{Fix}(E):E\in[\omega_{1}]^{\leq\omega}\}$.
A (normal) filter of subgroups is upward closed and closed under finite intersections (and conjugation).
Lemma 3.4 (Basis for the Cohen filter).
If $H\in\mathscr{F}$, then there is $E\in[\omega_{1}]^{\leq\omega}$ such that $\operatorname{Fix}(E)\leq H$.
Proof.
Unwinding $\mathscr{F}$ is the normal filter generated by $\{\operatorname{Fix}(E)\}$”, there are $E_{0},\dots,E_{k-1}\in[\omega_{1}]^{\leq\omega}$ and $\pi_{0},\dots,\pi_{k-1}\in\mathscr{G}$ such that
$\bigcap_{i<k}\pi_{i}\,\operatorname{Fix}(E_{i})\,\pi_{i}^{-1}\ \leq\ H.$
But $\pi_{i}\operatorname{Fix}(E_{i})\pi_{i}^{-1}=\operatorname{Fix}(\pi_{i}[E_{i}])$, and finite intersections of $\operatorname{Fix}(\cdot)$’s are $\operatorname{Fix}(\bigcup\cdot)$. Let $E:=\bigcup_{i<k}\pi_{i}[E_{i}]$ (countable). Then $\operatorname{Fix}(E)\leq H$. ∎
Corollary 3.5 ($\omega_{1}$-completeness of $\mathscr{F}$).
$\mathscr{F}$ is $\omega_{1}$-complete.
Proof.
Let $\langle H_{n}:n<\omega\rangle\subseteq\mathscr{F}$. By Lemma 3.4 pick countable $E_{n}$ with $\operatorname{Fix}(E_{n})\leq H_{n}$. Put $E=\bigcup_{n<\omega}E_{n}$ (countable), so $\operatorname{Fix}(E)=\bigcap_{n<\omega}\operatorname{Fix}(E_{n})\leq\bigcap_{n<\omega}H_{n}$, hence $\bigcap_{n<\omega}H_{n}\in\mathscr{F}$. ∎
Definition 3.6 (Symmetric and hereditarily symmetric names).
For a $\mathbb{P}$-name $\dot{x}$, define its stabilizer
$\operatorname{Stab}(\dot{x})\;:=\;\{\pi\in\mathscr{G}:\pi\dot{x}=\dot{x}\}.$
We call $\dot{x}$ symmetric if $\operatorname{Stab}(\dot{x})\in\mathscr{F}$. Let $\operatorname{HS}$ be the class of hereditarily symmetric names: $\dot{x}\in\operatorname{HS}$ iff $\dot{x}$ is symmetric and every name appearing in $\dot{x}$ is in $\operatorname{HS}$.
Fix $G\subseteq\mathbb{P}$ $V$-generic. The associated symmetric extension is
$\mathcal{N}\;:=\;\operatorname{HS}^{G}\;=\;\{\dot{x}^{G}:\dot{x}\in\operatorname{HS}\}.$
3.2 CCC and dependent choice
Lemma 3.7 (CCC).
$\operatorname{Add}(\omega,\omega_{1})$ is ccc. In particular, it preserves all cardinals and cofinalities.
Proof.
Let $\{p_{\xi}:\xi<\omega_{1}\}\subseteq\operatorname{Add}(\omega,\omega_{1})$ be uncountable. Write $D_{\xi}=\operatorname{dom}(p_{\xi})\subseteq\omega\times\omega_{1}$, so each $D_{\xi}$ is finite.
By thinning out, assume $|D_{\xi}|=n$ for all $\xi$. By the $\Delta$–system lemma (applied to the finite sets $D_{\xi}$), thin out again to an uncountable set $I\subseteq\omega_{1}$ such that $\{D_{\xi}:\xi\in I\}$ forms a $\Delta$–system with root $R$.
There are only finitely many possible values for $p_{\xi}\!\restriction R$. Thin out once more so that $p_{\xi}\!\restriction R=p_{\eta}\!\restriction R$ for all $\xi,\eta\in I$.
Now for $\xi\neq\eta$ in $I$, the conditions $p_{\xi}$ and $p_{\eta}$ agree on $D_{\xi}\cap D_{\eta}=R$, hence $p_{\xi}\cup p_{\eta}$ is a condition extending both. Thus $p_{\xi}$ and $p_{\eta}$ are compatible. Therefore there is no uncountable antichain, i.e. $\operatorname{Add}(\omega,\omega_{1})$ is ccc. ∎
Lemma 3.8 (ZF for symmetric extensions).
Let $\langle\mathbb{P},\mathscr{G},\mathscr{F}\rangle$ be a symmetric system, i.e. $\mathscr{G}\leq\operatorname{Aut}(\mathbb{P})$ and $\mathscr{F}$ is a normal filter of subgroups of $\mathscr{G}$. If $G\subseteq\mathbb{P}$ is $V$-generic, then $\operatorname{HS}^{G}=\{\dot{x}^{G}:\dot{x}\in\operatorname{HS}\}$ is a transitive model of $\mathsf{ZF}$ and $V\subseteq\operatorname{HS}^{G}\subseteq V[G]$.
Proof.
This is the standard symmetric extension theorem; see, e.g., Jech chapter 15, Lemma 15.51[4]. ∎
Lemma 3.9 (HS witnesses from HS-sets).
Let $\dot{A}\in\operatorname{HS}$ and let $\varphi(v,\vec{\dot{b}})$ be any formula with parameters $\vec{\dot{b}}\in\operatorname{HS}$. If $p\Vdash\exists v\in\dot{A}\,\varphi(v,\vec{\dot{b}})$, then there exist $q\leq p$ and $\dot{a}\in\operatorname{HS}$ such that $q\Vdash\dot{a}\in\dot{A}\wedge\varphi(\dot{a},\vec{\dot{b}})$.
Proof.
Fix $s\leq p$. Since $p\Vdash\exists v\in\dot{A}\,\varphi(v,\vec{\dot{b}})$, there are $t\leq s$ and a name $\dot{v}$ with $t\Vdash\dot{v}\in\dot{A}\wedge\varphi(\dot{v},\vec{\dot{b}})$. From $t\Vdash\dot{v}\in\dot{A}$, by the definition of forcing for membership there is $u\leq t$ and $\langle\dot{a},r\rangle\in\dot{A}$ such that $u\leq r$ and $u\Vdash\dot{v}=\dot{a}$. Then $u\Vdash\varphi(\dot{a},\vec{\dot{b}})$ as well, and since $\dot{A}\in\operatorname{HS}$ every name appearing in $\dot{A}$ (in particular $\dot{a}$) is hereditarily symmetric. Thus $u$ witnesses the conclusion. ∎
Theorem 3.10 ($\mathsf{DC}$ in the symmetric model).
$\mathcal{N}\models\mathsf{ZF}+\mathsf{DC}$.
Proof.
We verify $\mathsf{ZF}$ and $\mathsf{DC}$ separately.
$\mathsf{ZF}$. By Lemma 3.8, the symmetric extension $\mathcal{N}=\operatorname{HS}^{G}$ is a transitive model of $\mathsf{ZF}$ (containing all ordinals, as usual for symmetric extensions). We record the following symmetry lemma for the action of $\mathscr{G}$ on $\mathbb{P}$ for later use.
Claim 1. (Symmetry lemma) For every formula $\varphi$ in the forcing language, every $\mathbb{P}$-condition $p$, every $\pi\in\mathscr{G}$, and every tuple of $\mathbb{P}$-names $\vec{\dot{x}}$,
$p\Vdash\varphi(\vec{\dot{x}})\quad\Longleftrightarrow\quad\pi p\Vdash\varphi(\pi\vec{\dot{x}}).$
Proof.
By induction on the complexity of $\varphi$. The atomic cases reduce to the facts that $\pi$ is an automorphism of $\mathbb{P}$ and that the action on names is defined recursively so as to commute with membership and equality. Boolean connectives and quantifiers are then handled by the induction hypothesis in the usual way. ∎
$\mathsf{DC}$. Let $A,R\in\mathcal{N}$ with $A\neq\emptyset$ and $\forall x\in A\,\exists y\in A\,(xRy)$. Fix $\dot{A},\dot{R}\in\operatorname{HS}$ with $\dot{A}^{G}=A$ and $\dot{R}^{G}=R$. Choose $p_{\ast}\in\mathbb{P}$ such that
$p_{\ast}\Vdash\dot{A}\neq\emptyset\ \wedge\ \forall x\in\dot{A}\ \exists y\in\dot{A}\ ((x,y)\in\dot{R}).$
We recursively build, for each $n<\omega$,
  • a countable maximal antichain $D_{n}\subseteq\mathbb{P}$ below $p_{\ast}$, and
  • for each $p\in D_{n}$ a name $\dot{a}_{n,p}\in\operatorname{HS}$,
so that:
  1. (1)
    $p\Vdash\dot{a}_{n,p}\in\dot{A}$ for all $p\in D_{n}$;
  2. (2)
    for $n\geq 1$, if $p\in D_{n}$ and $p\leq q\in D_{n-1}$, then $p\Vdash(\dot{a}_{n-1,q},\dot{a}_{n,p})\in\dot{R}$.
For $n=0$: since $p_{\ast}\Vdash\dot{A}\neq\emptyset$, there is a maximal antichain $D_{0}$ below $p_{\ast}$ such that for each $p\in D_{0}$ there is a name $\dot{a}_{0,p}\in\operatorname{HS}$ with $p\Vdash\dot{a}_{0,p}\in\dot{A}$. (Existence of such conditions is ensured by Lemma 3.9, applied to $p_{\ast}\Vdash\dot{A}\neq\emptyset$.) Because $\mathbb{P}$ is ccc (Lemma 3.7), we may assume $D_{0}$ is countable.
[…]
$q\Vdash\dot{a}_{n+1,q}\in\dot{A}\ \wedge\ (\dot{a}_{n,p},\dot{a}_{n+1,q})\in\dot{R}.$
(This is justified by Lemma 3.9, applied below $p$ to the formula $v\in\dot{A}\wedge(\dot{a}_{n,p},v)\in\dot{R}$.)
Let $D_{n+1}:=\bigcup_{p\in D_{n}}D_{n+1,p}$. Then $D_{n+1}$ is countable (a countable union of countable sets) and is a maximal antichain below $p_{\ast}$: given any $r\leq p_{\ast}$, pick $p\in D_{n}$ compatible with $r$ (maximality of $D_{n}$), let $r^{\prime}\leq r,p$, and then pick $q\in D_{n+1,p}$ compatible with $r^{\prime}$ (maximality of $D_{n+1,p}$).
Now define, for each $n<\omega$, the mixed name
$\dot{a}_{n}\;:=\;\bigl\{\langle\dot{a}_{n,p},p\rangle:p\in D_{n}\bigr\}.$
Then $p_{\ast}\Vdash\dot{a}_{n}\in\dot{A}$ for all $n$, and $p_{\ast}\Vdash(\dot{a}_{n},\dot{a}_{n+1})\in\dot{R}$ for all $n$, by the coherence of the antichains.
It remains to see that each $\dot{a}_{n}$ is hereditarily symmetric and that the sequence name is hereditarily symmetric. Since each $\dot{a}_{n,p}\in\operatorname{HS}$, its stabilizer lies in $\mathscr{F}$; by definition of $\mathscr{F}$, for each such name choose a countable set $E(\dot{a}_{n,p})\subseteq\omega_{1}$ with $\operatorname{Fix}(E(\dot{a}_{n,p}))\leq\operatorname{Stab}(\dot{a}_{n,p})$. Also, for each $p\in D_{n}$, let $\operatorname{supp}(p):=\{\alpha:\exists m\,(\alpha,m)\in\operatorname{dom}(p)\}$, which is finite. Set
$E_{n}\;:=\;E(\dot{A})\ \cup\ E(\dot{R})\ \cup\ \bigcup_{p\in D_{n}}\Bigl(\operatorname{supp}(p)\ \cup\ E(\dot{a}_{n,p})\Bigr),$
a countable union of countable sets of ordinals, hence countable. Then every $\pi\in\operatorname{Fix}(E_{n})$ fixes each $p\in D_{n}$, fixes $\dot{A}$ and $\dot{R}$, and fixes each $\dot{a}_{n,p}$. Therefore $\pi\dot{a}_{n}=\dot{a}_{n}$, i.e. $\operatorname{Fix}(E_{n})\leq\operatorname{Stab}(\dot{a}_{n})$, so $\dot{a}_{n}$ is symmetric; hereditariness follows since the names appearing in $\dot{a}_{n}$ are among the $\dot{a}_{n,p}\in\operatorname{HS}$. Thus $\dot{a}_{n}\in\operatorname{HS}$ for all $n$.
Let $E:=\bigcup_{n<\omega}E_{n}$, again countable, and define the sequence name
$\dot{s}\;:=\;\bigl\{\langle\langle\check{n},\dot{a}_{n}\rangle,\mathbbm{1}_{\mathbb{P}}\rangle:n<\omega\bigr\}.$
[…]
Therefore $\mathcal{N}\models\mathsf{DC}$, and we have shown $\mathcal{N}\models\mathsf{ZF}+\mathsf{DC}$. ∎
3.3 Countable supports for $\operatorname{HS}$-names
Notation 3.11 (Restrictions to $E$).
For $E\subseteq\omega_{1}$ and $p\in\operatorname{Add}(\omega,\omega_{1})$, let
$p\!\restriction E\;:=\;p\cap(E\times\omega\times 2),$
be viewed as a condition in $\mathbb{P}_{E}:=\operatorname{Add}(\omega,E)$.
For a $\mathbb{P}$-name $\tau$, define recursively the restricted name $\tau\!\restriction E$ by
$\tau\!\restriction E\;:=\;\{\langle\sigma\!\restriction E,\;p\!\restriction E\rangle:\langle\sigma,p\rangle\in\tau\}.$
Then $\tau\!\restriction E$ is a $\mathbb{P}_{E}$-name (via the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$).
Definition 3.12 (Support).
If $\dot{x}\in\operatorname{HS}$, we say that $E\subseteq\omega_{1}$ is a support for $\dot{x}$ if $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$.
Lemma 3.13 (Every $\operatorname{HS}$-name has a countable support).
For every $\dot{x}\in\operatorname{HS}$ there exists $E\in[\omega_{1}]^{\leq\omega}$ such that $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$.
Proof.
If $\dot{x}\in\operatorname{HS}$, then in particular $\dot{x}$ is $\mathscr{F}$–symmetric, so $\operatorname{Stab}(\dot{x})\in\mathscr{F}$. Since $\operatorname{Stab}(\dot{x})\in\mathscr{F}$, Lemma 3.4 yields a countable $E$ with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$. This is exactly the assertion that $\dot{x}$ has countable support. ∎
Lemma 3.14 ($\mathbb{P}_{E}$-names are hereditarily symmetric).
Let $E\in[\omega_{1}]^{\leq\omega}$ and $\mathbb{P}_{E}=\operatorname{Add}(\omega,E)\leq\mathbb{P}=\operatorname{Add}(\omega,\omega_{1})$. If $\tau$ is a $\mathbb{P}_{E}$-name, view $\tau$ as a $\mathbb{P}$-name via the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$ (i.e. the same set-theoretic name, whose conditions all lie in $\mathbb{P}_{E}\subseteq\mathbb{P}$). Then $\tau\in\operatorname{HS}$, and $E$ is a support for $\tau$ (i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$).
Proof.
Fix $\pi\in\operatorname{Fix}(E)$. Since $\pi$ fixes $E$ pointwise, it acts trivially on $\mathbb{P}_{E}$: for every $p\in\mathbb{P}_{E}$ we have $\pi p=p$. We show by induction on the rank of $\tau$ that $\pi\tau=\tau$. Indeed, if $\langle\sigma,p\rangle\in\tau$, then $p\in\mathbb{P}_{E}$ so $\pi p=p$, and by induction $\pi\sigma=\sigma$, hence $\langle\pi\sigma,\pi p\rangle=\langle\sigma,p\rangle\in\tau$; similarly every member of $\pi\tau$ lies in $\tau$. Thus $\pi\tau=\tau$.
Therefore $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$. Since $E$ is countable, $\operatorname{Fix}(E)\in\mathscr{F}$, so $\tau$ is symmetric. The same argument applies to every subname of $\tau$ (all of whose conditions also lie in $\mathbb{P}_{E}$), so $\tau$ is hereditarily symmetric, i.e. $\tau\in\operatorname{HS}$. ∎
Lemma 3.15 (Support localization for $\operatorname{Add}(\omega,\omega_{1})$).
Let $E\subseteq\omega_{1}$ and let $\tau$ be a $\mathbb{P}$-name such that $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$. Let $\varphi$ be any formula whose parameters are $\mathbb{P}$-names also fixed by $\operatorname{Fix}(E)$. If $p\Vdash\varphi(\tau)$, then $p\!\restriction E\Vdash\varphi(\tau)$.
Proof.
Suppose $p\Vdash\varphi(\tau)$ but $p\!\restriction E\nVdash\varphi(\tau)$. Then there is $q\leq p\!\restriction E$ with $q\Vdash\neg\varphi(\tau)$.
Let $F_{p}\subseteq\omega_{1}\setminus E$ be the finite set of ordinals mentioned by $p$ outside $E$ (i.e. those $\alpha\notin E$ with $(\alpha,n)\in\operatorname{dom}(p)$ for some $n$), and define $F_{q}$ similarly. Choose a permutation $\pi\in\operatorname{Fix}(E)$ such that $\pi[F_{p}]$ is disjoint from $F_{q}$ and such that $\pi$ fixes every element of $F_{q}$. Then $\pi p$ is compatible with $q$ (they agree on $E$ and have disjoint requirements outside $E$), so let $r\leq\pi p,q$.
By the symmetry lemma (Claim 3.2) and $\pi\in\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$, we have $\pi p\Vdash\varphi(\pi\tau)=\varphi(\tau)$, hence $r\Vdash\varphi(\tau)$. But also $r\leq q$ and $q\Vdash\neg\varphi(\tau)$, so $r\Vdash\neg\varphi(\tau)$, a contradiction. ∎
Corollary 3.16 (Supported names live in the corresponding subextension).
Let $E\subseteq\omega_{1}$ and $\tau$ a $\mathbb{P}$-name with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$. Then $\mathbbm{1}_{\mathbb{P}}\Vdash\tau=\tau\!\restriction E$. In particular, for any $V$-generic $G\subseteq\mathbb{P}$ with $G_{E}:=G\cap\mathbb{P}_{E}$,
$\tau^{G}=(\tau\!\restriction E)^{G_{E}}\in V[G_{E}].$
Proof.
We prove by induction on rank that $\mathbbm{1}_{\mathbb{P}}\Vdash\tau=\tau\!\restriction E$. Fix a formula of the form $\sigma\in\tau$” or $\sigma=\tau$” and use Lemma 3.15 to show that whenever $p$ decides such a statement about $\tau$, its restriction $p\!\restriction E$ decides it the same way. This yields that membership in $\tau^{G}$ is determined by $G_{E}$ and the recursive definition of $\tau\!\restriction E$, and the induction closes. ∎
3.4 The canonical Cohen set $A$ and failure of choice
For each $\alpha<\omega_{1}$, define the canonical $\mathbb{P}$-name $\dot{c}_{\alpha}$ for the $\alpha$-th Cohen real by
$\dot{c}_{\alpha}\;:=\;\{(\check{n},p):p(\alpha,n)=1\}.$
Let $c_{\alpha}:=\dot{c}_{\alpha}^{G}\in 2^{\omega}$.
Definition 3.17 (The Cohen set).
$A\;:=\;\{c_{\alpha}:\alpha<\omega_{1}\}.$
Lemma 3.18.
$A\in\mathcal{N}$.
Proof.
For each $\alpha<\omega_{1}$, the canonical name $\dot{c}_{\alpha}$ for the $\alpha$-th Cohen real has support $\{\alpha\}$, hence is hereditarily symmetric. Let
$\dot{A}=\{(\dot{c}_{\alpha},\mathbbm{1}_{\mathbb{P}}):\alpha<\omega_{1}\}.$
Any $\pi\in\mathscr{G}$ permutes the $\dot{c}_{\alpha}$’s, so $\pi(\dot{A})=\dot{A}$. Thus $\operatorname{Stab}(\dot{A})=\mathscr{G}\in\mathscr{F}$, and $\dot{A}$ is symmetric. Since its members are hereditarily symmetric, $\dot{A}\in\operatorname{HS}$, hence $A=\operatorname{val}(\dot{A},G)\in\mathcal{N}$. Indeed $\dot{c}_{\alpha}$ is a $\mathbb{P}_{\{\alpha\}}$-name, hence $\dot{c}_{\alpha}\in\operatorname{HS}$ by Lemma 3.14. ∎
Proposition 3.19 (Non-well-orderability).
In $\mathcal{N}$, the set $A$ is not well-orderable. In particular, $\mathcal{N}\models\neg\mathsf{WO}$ and hence $\neg\mathsf{AC}$.
Proof.
We show that in $\mathcal{N}$ there is no injection from $A$ into any ordinal. Since $\mathsf{AC}$ implies that every set is well-orderable (equivalently, injects into an ordinal), this will imply that $A$ is not well-orderable in $\mathcal{N}$, hence $\mathcal{N}\models\neg\mathsf{WO}$ and therefore $\mathcal{N}\models\neg\mathsf{AC}$.
Suppose toward a contradiction that there are an ordinal $\theta$ and $\dot{f}\in\operatorname{HS}$ such that $\dot{f}^{G}$ is an injection $A\to\theta$ in $\mathcal{N}$. Choose $p_{0}\in G$ and $\theta$ with
$p_{0}\Vdash\text{``$\dot{f}:\dot{A}\to\check{\theta}$ is an injection.''}$
By Lemma 3.13 fix a countable support $E\in[\omega_{1}]^{\leq\omega}$ for $\dot{f}$, i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{f})$.
Let $\operatorname{supp}(p)\subseteq\omega_{1}$ denote the finite set of ordinals $\alpha$ such that $(\alpha,n)\in\operatorname{dom}(p)$ for some $n\in\omega$. Choose $\alpha\in\omega_{1}\setminus(E\cup\operatorname{supp}(p_{0}))$. Since $p_{0}$ forces that $\dot{f}(\dot{c}_{\alpha})$ is an ordinal $\theta$, there are $q\leq p_{0}$ and $\xi<\theta$ such that
$q\Vdash\dot{f}(\dot{c}_{\alpha})=\check{\xi}.$
Now choose $\beta\in\omega_{1}\setminus\bigl(E\cup\operatorname{supp}(q)\bigr)$ with $\beta\neq\alpha$, and let $\pi\in\operatorname{Sym}(\omega_{1})$ be the transposition swapping $\alpha$ and $\beta$ and fixing every other ordinal. Then $\pi\in\operatorname{Fix}(E)$, so $\pi\dot{f}=\dot{f}$, and also $\pi\dot{c}_{\alpha}=\dot{c}_{\beta}$.
By the standard automorphism invariance of the forcing relation,
$q\Vdash\dot{f}(\dot{c}_{\alpha})=\check{\xi}\quad\Longrightarrow\quad\pi q\Vdash(\pi\dot{f})(\pi\dot{c}_{\alpha})=\check{\xi},$
so using $\pi\dot{f}=\dot{f}$ and $\pi\dot{c}_{\alpha}=\dot{c}_{\beta}$ we obtain
$\pi q\Vdash\dot{f}(\dot{c}_{\beta})=\check{\xi}.$
Because $q$ mentions $\alpha$ but not $\beta$, and $\pi$ fixes every ordinal in $\operatorname{supp}(q)\setminus\{\alpha\}$, the conditions $q$ and $\pi q$ are compatible; indeed, they have disjoint requirements on the $\alpha$-row versus the $\beta$-row, and agree on their common domain. Let $r:=q\cup\pi q$, so $r\in\mathbb{P}$ and $r\leq q,\pi q$. Then
$r\Vdash\dot{f}(\dot{c}_{\alpha})=\check{\xi}\ \wedge\ \dot{f}(\dot{c}_{\beta})=\check{\xi}.$
Finally, strengthen $r$ to a condition $r^{\prime}\leq r$ forcing $\dot{c}_{\alpha}\neq\dot{c}_{\beta}$: choose $n\in\omega$ with $(\alpha,n),(\beta,n)\notin\operatorname{dom}(r)$ and put
$r^{\prime}:=r\cup\{((\alpha,n),0),((\beta,n),1)\}.$
Then $r^{\prime}\Vdash\dot{c}_{\alpha}(n)\neq\dot{c}_{\beta}(n)$, hence $r^{\prime}\Vdash\dot{c}_{\alpha}\neq\dot{c}_{\beta}$. But $r^{\prime}\leq r$ still forces $\dot{f}(\dot{c}_{\alpha})=\dot{f}(\dot{c}_{\beta})$, contradicting the injectivity of $\dot{f}$ on $\dot{A}$.
This contradiction shows that no injection $A\to\theta$ exists in $\mathcal{N}$ for any ordinal $\theta$. Therefore $A$ is not well-orderable in $\mathcal{N}$, and consequently $\mathcal{N}\models\neg\mathsf{WO}$ and $\mathcal{N}\models\neg\mathsf{AC}$. ∎
Remark 3.20 (Dedekind-infinitude).
Unlike the finite-support generalized Cohen model (which produces a Dedekind-finite $A$; cf. [11, §2.4]), the present countable-support filter allows many countable subsets of $A$: for each countable $E\subseteq\omega_{1}$ in $V$, the name $\dot{A}_{E}:=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha\in E\}$ is hereditarily symmetric, so $A_{E}:=\{c_{\alpha}:\alpha\in E\}\in\mathcal{N}$.
Lemma 3.21 (Well-orders are rigid).
If $(X,\prec)$ is a well-order and $h:X\to X$ is a bijection such that
$x\prec y\iff h(x)\prec h(y)\qquad(x,y\in X),$
then $h=\mathrm{id}_{X}$.
Proof.
Suppose $\sigma$ is an order automorphism of $(X,<)$ and $\sigma\neq\mathrm{id}_{X}$. Let $x_{0}$ be the <<-least element of $X$ with $\sigma(x_{0})\neq x_{0}$.
If $\sigma(x_{0})<x_{0}$, then $\sigma(x_{0})<x_{0}$ implies (by minimality of $x_{0}$) that $\sigma(\sigma(x_{0}))=\sigma(x_{0})$. But order preservation yields $\sigma(\sigma(x_{0}))<\sigma(x_{0})$, a contradiction.
If $\sigma(x_{0})>x_{0}$, then apply the previous case to $\sigma^{-1}$: since $\sigma^{-1}$ is also an order automorphism and $\sigma^{-1}(x_{0})<x_{0}$ (because $x_{0}<\sigma(x_{0})$), we get a contradiction.
Hence $\sigma=\mathrm{id}_{X}$. ∎
3.5 A safe $\mathsf{SVC}$ seed in $\mathcal{N}$
Proposition 3.22 ($\mathsf{SVC}(A^{\omega})$ in the Cohen symmetric model).
Work in the Cohen symmetric model $\mathcal{N}$, and let $S:=A^{\omega}$. Then $\mathcal{N}\models\mathsf{SVC}(S)$.
Proof.
Fix a nonempty $X\in\mathcal{N}$ and choose $\dot{X}\in\operatorname{HS}$ with $X=\dot{X}^{G}$.
By Lemma 3.13, fix a countable support $E\subseteq\omega_{1}$ for $\dot{X}$, i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{X})$. Let $\dot{X}\!\restriction E$ be the restricted $\mathbb{P}_{E}$-name from Notation above. By Corollary 3.16,
$X=\dot{X}^{G}=(\dot{X}\!\restriction E)^{G_{E}}\in V[G_{E}],$
where $G_{E}:=G\cap\mathbb{P}_{E}$.
Since $V[G_{E}]\models\mathsf{ZFC}$, there is an ordinal $\eta$ and a surjection $h:\eta\twoheadrightarrow X$ in $V[G_{E}]$. Let $\dot{h}$ be a $\mathbb{P}_{E}$-name for $h$. Every permutation $\pi\in\operatorname{Sym}(\omega_{1})$ which fixes $E$ pointwise acts trivially on $\mathbb{P}_{E}$ and hence fixes $\dot{h}$, so $\dot{h}\in\operatorname{HS}$ (indeed, it has support $E$). Therefore $h=\dot{h}^{G}\in\mathcal{N}$.
Finally, since $S=A^{\omega}$ is nonempty in $\mathcal{N}$, the map $f:S\times\eta\to X$ defined by $f(s,\xi)=h(\xi)$ is a surjection in $\mathcal{N}$. This proves $\mathcal{N}\models\mathsf{SVC}(S)$. ∎
Definition 3.23 (Fixed seed parameters: $S$ and $T$).
Work in $\mathcal{N}$ and let
$S:=A^{\omega},\qquad T:=\mathcal{P}(S).$
4 Reduction blueprint: global $\mathsf{PP}$ from localized $\mathsf{PP}$
We now pin down the exact hypotheses and the parameter set to which we will localize $\mathsf{PP}$.
Definition 4.1 (Localized $\mathsf{PP}$).
Let $T$ be a set.
  1. 1.
    $\mathsf{PP}\!\restriction T$ is the statement that for all $X,Y\subseteq T$, if there is a surjection $f:Y\twoheadrightarrow X$, then there is an injection $i:X\hookrightarrow Y$.
  2. 2.
    $\mathsf{PP}(T)$ is the statement that for every set $X$, if there is a surjection $f:T\twoheadrightarrow X$, then there is an injection $i:X\hookrightarrow T$.
Definition 4.2 ($\mathsf{SVC}^{+}$).
For a set $T$, $\mathsf{SVC}^{+}(T)$ denotes the statement that for every set $X$ there exists an ordinal $\eta$ and an injection $j:X\hookrightarrow T\times\eta$.
Fact 4.3 (Ryan–Smith).
For every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$. [13, §2.1, Fact]
Remark 4.4.
Let $S,T$ be as in Definition 3.23. By Fact 4.3 we have $\mathcal{N}\models\mathsf{SVC}^{+}(T)$, and we will keep this $T$ fixed throughout the construction.
Theorem 4.5 (Ryan–Smith localization of $\mathsf{PP}$).
Assume $\mathsf{SVC}^{+}(T)$. Then
$\mathsf{PP}\quad\Longleftrightarrow\quad\bigl(\mathsf{PP}\!\restriction T\ \wedge\ \mathsf{AC}_{\mathsf{WO}}\bigr).$
[13, Proposition 3.17] (stated there with parameter “SS”; we apply it with $S:=T$)
Remark 4.6.
Ryan–Smith formulate $\mathsf{PP}$ as the equivalence $|X|\leq|Y|\iff|X|\leq^{*}|Y|$. Since $|X|\leq|Y|\Rightarrow|X|\leq^{*}|Y|$ holds in $\mathsf{ZF}$, this is equivalent to our surjection-to-injection formulation of $\mathsf{PP}$.
Proposition 4.7 (The $\mathsf{AC}_{\mathsf{WO}}$ conjunct cannot be omitted).
There is a model $M\models\mathsf{ZF}$ and a set $S^{\ast}\in M$ such that
$M\models\mathsf{SVC}^{+}(S^{\ast})\ \wedge\ \mathsf{PP}(S^{\ast})\ \wedge\ \neg\mathsf{PP}.$
In particular, global $\mathsf{PP}$ cannot be recovered from a small-choice hypothesis together with a single-parameter instance $\mathsf{PP}(S^{\ast})$.
Reference.
This is witnessed by Cohen’s first model; see [13, Proposition 3.22]. ∎
Proposition 4.8 (The localized $\mathsf{PP}\!\restriction S$ component cannot be omitted).
There is a model $M\models\mathsf{ZF}$ and a set $S^{\ast}\in M$ such that
$M\models\mathsf{AC}_{\mathsf{WO}}\ \wedge\ \mathsf{SVC}^{+}(S^{\ast})\ \wedge\ \mathsf{PP}(S^{\ast})\ \wedge\ \neg\mathsf{PP}.$
Consequently, even assuming $\mathsf{AC}_{\mathsf{WO}}$ and small-choice hypotheses, one cannot in general replace the localized requirement $\mathsf{PP}\!\restriction S^{\ast}$ by the weaker single-parameter assertion $\mathsf{PP}(S^{\ast})$ and still obtain global $\mathsf{PP}$.
Reference.
This is witnessed by the Feferman-style model $N_{\aleph_{1}}$ of Truss; see [13, Proposition 3.23]. ∎
Remark 4.9 (Why we keep both conjuncts).
The preceding propositions are included only to justify (conceptually) why the localization theorem is stated with both components present: the “choice” side $\mathsf{AC}_{\mathsf{WO}}$ and the “localized comparability” side $\mathsf{PP}\!\restriction T$. We will force both conjuncts $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ in Theorem 4.5. In Ryan–Smith’s analysis the corresponding reduction is sharp: in general one cannot drop the displayed conjuncts (nor substantially weaken the hypotheses) without losing the equivalence; see [13, Props. 3.22 and 3.23].
4.1 The localized $\mathsf{PP}$-package forcing $\mathbb{Q}_{f}$
Fix a transitive model $M\models\mathsf{ZF}+\mathsf{DC}$ extending the Cohen seed $\mathcal{N}$ (in practice, $M$ will be an intermediate stage of our symmetric iteration). Work with the fixed parameter $T=\mathcal{P}(S)$ from Definition 3.23.
Definition 4.10 (Localized splitting principle).
Write $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ for the assertion that for all $X,Y\subseteq T$ and all surjections $f:Y\twoheadrightarrow X$, there exists a right inverse $s:X\to Y$ with $f\circ s=\operatorname{id}_{X}$ (necessarily injective: if $s(x_{1})=s(x_{2})$ then $x_{1}=f(s(x_{1}))=f(s(x_{2}))=x_{2}$.)
Remark 4.11 (Why we allow localized splittings).
The global statement “every surjection splits” (i.e. every surjection has a right inverse) is equivalent to $\mathsf{AC}$ (a right inverse to $\bigsqcup_{i\in I}A_{i}\twoheadrightarrow I$ is exactly a choice function). So we are not attempting to force global splittings.
Moreover, $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ only concerns surjections whose domain and codomain lie inside the fixed parameter $T=\mathcal{P}(S)$, and is therefore far weaker than the global splitting scheme equivalent to $\mathsf{AC}$.
However, for our Ryan–Smith reduction we only need to secure localized $\mathsf{PP}$ over the fixed parameter $T=\mathcal{P}(S)$. For that purpose it is convenient to force the stronger localized principle $\mathsf{PP}^{\mathrm{split}}\!\restriction T$, which implies $\mathsf{PP}\!\restriction T$ outright. This choice buys us a canonical, highly closed package poset at each stage.
For the choice principles on well-orderable sets, Ryan–Smith prove that assuming $\mathsf{SVC}(S)$,
$\mathsf{AC}_{\mathsf{WO}}\quad\Longleftrightarrow\quad\mathsf{AC}_{<\aleph^{*}(S)}(S)\quad\Longleftrightarrow\quad\text{every function }g:S\to\lambda\text{ splits (for each }\lambda\text{)}$
(see [13, Prop. 3.2]). This is the equivalence we invoke in Proposition 5.71.
(We also note that Ryan–Smith’s localization for $\mathsf{PP}$ uses both conjuncts: $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$; see [13, Prop. 3.17] and Question 3.18 therein, and see [13, Props. 3.22 and 3.23] for counterexamples showing the conjuncts cannot in general be omitted.)
Definition 4.12 ($\mathsf{PP}$-package for a fixed surjection).
Let $X,Y\subseteq T$ and let $f:Y\twoheadrightarrow X$ be a surjection in $M$. Define $\mathbb{Q}_{f}$ to be the poset whose conditions are countable partial injective sections of $f$:
[…]
The order is extension: $q\leq p$ iff $q\supseteq p$.
Remark 4.13 (Size and chain condition).
The forcing $\mathbb{Q}_{f}$ can be large and is not expected to be ccc in general (e.g. $|\mathbb{Q}_{f}|$ may be comparable to $|X|^{\leq\omega}\cdot|Y|^{\leq\omega}$). This is harmless for our purposes: what we use is countable closure (hence no new reals and $\mathsf{DC}$ preservation), not any chain condition.
Lemma 4.14 (Basic density).
For each $x\in X$, the set
$D_{x}:=\{p\in\mathbb{Q}_{f}:x\in\operatorname{dom}(p)\}$
is dense in $\mathbb{Q}_{f}$.
Proof.
Fix $p\in\mathbb{Q}_{f}$ and $x\in X$ with $x\notin\operatorname{dom}(p)$. Choose any $y\in Y$ with $f(y)=x$ (exists since $f$ is surjective). We claim $y\notin\operatorname{ran}(p)$: if $y=p(x^{\prime})$ for some $x^{\prime}\in\operatorname{dom}(p)$ then $x=f(y)=f(p(x^{\prime}))=x^{\prime}$, contradicting $x\notin\operatorname{dom}(p)$. Thus $q:=p\cup\{(x,y)\}$ is still a partial function, still injective, and still satisfies $f(q(x))=x$; hence $q\in\mathbb{Q}_{f}$ and $q\leq p$ with $x\in\operatorname{dom}(q)$. ∎
Lemma 4.15 (Countable closure).
$\mathbb{Q}_{f}$ is Countably-closed (sometimes called $\omega_{1}$-closed).
Proof.
Let $\langle p_{n}:n<\omega\rangle$ be a decreasing sequence in $\mathbb{Q}_{f}$. Put $p:=\bigcup_{n<\omega}p_{n}$. Since each $p_{n+1}\supseteq p_{n}$, this union is a function. It is injective, and for every $x\in\operatorname{dom}(p)$ we have $f(p(x))=x$ because this holds already in some $p_{n}$ containing $x$.
It remains to see that $\operatorname{dom}(p)=\bigcup_{n<\omega}\operatorname{dom}(p_{n})$ is countable in $M$. Here we use $M\models\mathsf{DC}$, hence $M\models\mathsf{AC}_{\omega}$, so a countable union of countable sets is countable in $M$. (This is the only place in the closure argument where $\mathsf{DC}$ is used.) Therefore $p\in\mathbb{Q}_{f}$, and clearly $p\leq p_{n}$ for all $n$. ∎
Remark 4.16 (Closure depends on $\mathsf{DC}$).
The $\omega_{1}$-closure of $\mathbb{Q}_{f}$ is proved inside $M$ using $\mathsf{DC}$ (equivalently $\mathsf{AC}_{\omega}$), via “countable union of countable sets is countable”. In a bare $\mathsf{ZF}$ context this need not hold, so the closure claim should be read as relative to our standing hypothesis $M\models\mathsf{ZF}+\mathsf{DC}$.
Proposition 4.17 (What $\mathbb{Q}_{f}$ forces).
If $G\subseteq\mathbb{Q}_{f}$ is $M$-generic, then
$s_{G}:=\bigcup G$
is a total injective function $s_{G}:X\to Y$ with $f\circ s_{G}=\operatorname{id}_{X}$. In particular, $M[G]\models``X\hookrightarrow Y$".
Proof.
By Lemma 4.14, $G$ meets $D_{x}$ for every $x\in X$, so $x\in\operatorname{dom}(s_{G})$ for all $x$ and $s_{G}$ is total on $X$. Injectivity and $f(s_{G}(x))=x$ are preserved under unions of compatible conditions, hence hold for $s_{G}$. ∎
Lemma 4.18 (Countably closed forcing adds no new $\omega$-sequences and preserves $\mathsf{DC}$).
Let $M\models\mathsf{ZF}+\mathsf{DC}$ be transitive and let $\mathbb{P}\in M$ be countably closed in $M$ (i.e. every decreasing $\omega$-sequence in $M$ has a lower bound in $M$). Then:
  1. (1)
    $\mathbb{P}$ adds no new $\omega$-sequences of ordinals.
  2. (2)
    Forcing with $\mathbb{P}$ over $M$ preserves $\mathsf{DC}$.
Proof.
(1) No new $\omega$-sequences of ordinals. Let $p_{0}\in\mathbb{P}$ and $\dot{a}\in M$ with $p_{0}\Vdash\dot{a}:\check{\omega}\to\mathrm{Ord}$. For each $n<\omega$, let
$D_{n}:=\Bigl\{p\leq p_{0}:\exists\alpha\in\mathrm{Ord}\ (p\Vdash\dot{a}(\check{n})=\check{\alpha})\Bigr\}.$
Each $D_{n}$ is dense below $p_{0}$.
Define the set of states
$\Sigma:=\Bigl\{(n,p,\vec{\alpha}):n<\omega,\ p\in\mathbb{P},\ p\leq p_{0},\ \vec{\alpha}\in\mathrm{Ord}^{n},\ p\Vdash\forall k<n\ \dot{a}(\check{k})=\check{\vec{\alpha}(k)}\Bigr\}.$
Let $R$ be the relation on $\Sigma$ given by
$(n,p,\vec{\alpha})\ R\ (n+1,q,\vec{\beta})$
iff $q\leq p$, $\vec{\beta}\in\mathrm{Ord}^{n+1}$ extends $\vec{\alpha}$, and $q\Vdash\dot{a}(\check{n})=\check{\vec{\beta}(n)}$. Then $R$ is total on $\Sigma$ by density of $D_{n}$.
By $\mathsf{DC}$ in $M$, there is an $R$-chain $\langle(n,p_{n},\vec{\alpha}_{n}):n<\omega\rangle\in M$ starting from some $(0,p_{0},\emptyset)\in\Sigma$ (note $\mathrm{Ord}^{0}=\{\emptyset\}$). By countable closure, there is $p_{\omega}\in\mathbb{P}$ with $p_{\omega}\leq p_{n}$ for all $n$.
Let $g:=\bigcup_{n<\omega}\vec{\alpha}_{n}$. Since the chain is an element of $M$, Replacement in $M$ yields $g\in M$ and $g:\omega\to\mathrm{Ord}$. By construction, $p_{\omega}\Vdash\forall n\ \dot{a}(\check{n})=\check{g(n)}$, hence $p_{\omega}\Vdash\dot{a}=\check{g}$. Thus $\mathbb{P}$ adds no new $\omega$-sequences of ordinals.
(2) Preservation of $\mathsf{DC}$. Let $p_{0}\in\mathbb{P}$ and $\dot{X},\dot{R},\dot{x}_{0}\in M$ with
$p_{0}\Vdash\ \dot{x}_{0}\in\dot{X}\ \wedge\ (\forall x\in\dot{X})(\exists y\in\dot{X})\ x\,\dot{R}\,y.$
Define the state space
$\Gamma:=\Bigl\{(n,p,\dot{x}):n<\omega,\ p\in\mathbb{P},\ p\leq p_{0},\ \dot{x}\in M,\ p\Vdash\dot{x}\in\dot{X}\Bigr\}.$
Define $S$ on $\Gamma$ by
$(n,p,\dot{x})\ S\ (n+1,q,\dot{y})$
iff $q\leq p$ and $q\Vdash\dot{y}\in\dot{X}\ \wedge\ \dot{x}\,\dot{R}\,\dot{y}$. Totality of $S$ follows from the premise forced by $p_{0}$.
By $\mathsf{DC}$ in $M$, pick an $S$-chain $\langle(n,p_{n},\dot{x}_{n}):n<\omega\rangle\in M$ starting with $(0,p_{0},\dot{x}_{0})\in\Gamma$. By countable closure, let $p_{\omega}\leq p_{n}$ for all $n$.
Let
$\dot{f}:=\bigl\{\langle\langle\check{n},\dot{x}_{n}\rangle,\mathbbm{1}_{\mathbb{P}}\rangle:n<\omega\bigr\}.$
Since the chain is in $M$, Replacement gives $\dot{f}\in M$. Then $p_{\omega}\Vdash\dot{f}:\check{\omega}\to\dot{X}$ and $p_{\omega}\Vdash\forall n\ \dot{f}(\check{n})\,\dot{R}\,\dot{f}(\check{n+1})$, which is exactly $\mathsf{DC}$ in the extension. ∎
Corollary 4.19 (No new $\omega$-sequences; $\mathsf{DC}$ preservation).
$\mathbb{Q}_{f}$ adds no new $\omega$-sequences of ordinals (in particular, no new reals). Moreover, over a $\mathsf{ZF}+\mathsf{DC}$ ground, forcing with $\mathbb{Q}_{f}$ preserves $\mathsf{DC}$.
Proof.
Apply Lemma 4.18 with $\mathbb{P}=\mathbb{Q}_{f}$. Countable closure holds by Lemma 4.15. ∎
4.2 The $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{f}$
Fix a transitive model $M\models\mathsf{ZF}+\mathsf{DC}$ extending the Cohen seed $\mathcal{N}$ (again, in practice $M$ will be an intermediate stage of our iteration). Retain the fixed seed parameter $S$ from Remark 4.4.
Definition 4.20 (The Lindenbaum bound $\aleph^{*}(S)$).
For a set $S$ in $M$, let $\aleph^{*}(S)^{M}$ denote the least ordinal $\kappa$ such that there is no surjection $S\twoheadrightarrow\kappa$ in $M$. Equivalently, $\kappa=\sup\{\alpha:\exists\text{ a surjection }S\twoheadrightarrow\alpha\}+1$.
Remark 4.21.
In $\mathsf{ZF}$ the ordinal $\aleph^{*}(S)$ exists for every set $S$ (it is the Lindenbaum number of $S$); see e.g. Jech [4, §(Hartogs/Lindenbaum)].
Remark 4.22 (Why the bound $\lambda<\aleph^{*}(S)$ is the correct target).
Assuming $\mathsf{SVC}(S)$, Ryan–Smith shows that $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the bounded scheme $\mathsf{AC}_{<\aleph^{*}(S)}$ (more precisely, $\mathsf{AC}_{<\aleph^{*}(S)}(S)$ in his notation); see [13, Cor. 3.3]. For fixed $\lambda$, the corresponding “splitting” reformulations of $\mathsf{AC}_{\lambda}$ appear in [13, Prop. 3.2]. Thus, in the presence of $\mathsf{SVC}(S)$ it suffices to schedule splittings only for surjections onto ordinals $\lambda<\aleph^{*}(S)$.
Definition 4.23 ($\mathsf{AC}_{\mathsf{WO}}$-package for a fixed surjection onto an ordinal).
Let $\lambda$ be an ordinal with $\lambda<\aleph^{*}(S)^{M}$, let $Y$ be a set in $M$, and let $f:Y\twoheadrightarrow\lambda$ be a surjection in $M$. Define $\mathbb{R}_{f}$ to be the poset of countable partial right inverses to $f$:
$r\in\mathbb{R}_{f}\iff\begin{cases}r\text{ is a function }r:\operatorname{dom}(r)\to Y,\\ \operatorname{dom}(r)\subseteq\lambda\text{ is countable},\\ \forall\xi\in\operatorname{dom}(r)\ \bigl(f(r(\xi))=\xi\bigr).\end{cases}$
The order is extension: $s\leq r$ iff $s\supseteq r$.
Remark 4.24 (Total sections are automatically injective).
If $s:\lambda\to Y$ satisfies $f\circ s=\operatorname{id}_{\lambda}$, then $s$ is injective: if $s(\xi_{1})=s(\xi_{2})$ then $\xi_{1}=f(s(\xi_{1}))=f(s(\xi_{2}))=\xi_{2}$.
Remark 4.25 (No localization to $T$ for $\mathsf{AC}_{\mathsf{WO}}$).
Unlike the $\mathsf{PP}$-packages $\mathbb{Q}_{f}$, which are explicitly localized to $X,Y\subseteq T=\mathcal{P}(S)$, the $\mathsf{AC}_{\mathsf{WO}}$-package must be global in nature: $\mathsf{AC}_{\mathsf{WO}}$ speaks about all well-ordered index sets. The point of Remark 4.22 is that, assuming $\mathsf{SVC}(S)$, it is enough to schedule splittings for surjections onto $\lambda<\aleph^{*}(S)$; this reduction is precisely what makes the bookkeeping manageable.
Lemma 4.26 (Basic density).
For each $\xi<\lambda$, the set
$D_{\xi}:=\{r\in\mathbb{R}_{f}:\xi\in\operatorname{dom}(r)\}$
is dense in $\mathbb{R}_{f}$.
Proof.
Fix $r\in\mathbb{R}_{f}$ and $\xi<\lambda$ with $\xi\notin\operatorname{dom}(r)$. Choose any $y\in Y$ with $f(y)=\xi$ (exists since $f$ is surjective). Then $s:=r\cup\{(\xi,y)\}$ is still a function, still satisfies $f(s(\xi))=\xi$, and has countable domain; hence $s\in\mathbb{R}_{f}$ and $s\leq r$ with $\xi\in\operatorname{dom}(s)$. ∎
Lemma 4.27 (Countable closure).
$\mathbb{R}_{f}$ is Countably closed.
Proof.
Let $\langle r_{n}:n<\omega\rangle$ be a decreasing sequence in $\mathbb{R}_{f}$. Put $r:=\bigcup_{n<\omega}r_{n}$. This is a function, and it satisfies $f(r(\xi))=\xi$ for every $\xi\in\operatorname{dom}(r)$ since this already holds in some $r_{n}$ containing $\xi$.
Finally, $\operatorname{dom}(r)=\bigcup_{n<\omega}\operatorname{dom}(r_{n})\subseteq\lambda$ is a countable union of countable sets of ordinals, hence countable in $M$ (this uses only $\mathsf{ZF}$ since $\lambda$ is well-ordered and each countable subset of an ordinal admits a canonical increasing enumeration). Therefore $r\in\mathbb{R}_{f}$, and clearly $r\leq r_{n}$ for all $n$. ∎
Proposition 4.28 (What $\mathbb{R}_{f}$ forces).
If $G\subseteq\mathbb{R}_{f}$ is $M$-generic, then
$s_{G}:=\bigcup G$
is a total function $s_{G}:\lambda\to Y$ with $f\circ s_{G}=\operatorname{id}_{\lambda}$ (hence $s_{G}$ is injective). In particular, $M[G]\models\text{``$f$ splits.''}$
Proof.
By Lemma 4.26, $G$ meets $D_{\xi}$ for every $\xi<\lambda$, so $\xi\in\operatorname{dom}(s_{G})$ for all $\xi<\lambda$ and $s_{G}$ is total on $\lambda$. The equation $f(s_{G}(\xi))=\xi$ is preserved under unions of compatible conditions, so $f\circ s_{G}=\operatorname{id}_{\lambda}$. ∎
Corollary 4.29 (No new reals; DC preservation (package-level)).
$\mathbb{R}_{f}$ adds no new reals (indeed, no new $\omega$-sequences of ordinals). Moreover, forcing with $\mathbb{R}_{f}$ over a $\mathsf{ZF}+\mathsf{DC}$ ground preserves $\mathsf{DC}$.
Proof.
Apply Lemma 4.18 with $\mathbb{P}=\mathbb{R}_{f}$. Countable closure holds by Lemma 4.27. ∎
Remark 4.30 (Package products).
When we say “finite product of package posets” we mean an arbitrary finite-support product of finitely many factors (not necessarily just two), typically consisting of one $\mathsf{AC}_{\mathsf{WO}}$-package $\mathbb{R}_{f}$ together with finitely many localized $\mathsf{PP}$-packages $\mathbb{Q}_{g_{0}},\dots,\mathbb{Q}_{g_{n-1}}$ scheduled at the same stage.
5 Iterating the packages over $\mathcal{N}$
Remark 5.1 (Roadmap of Section 5).
This section describes a class-length countable-support symmetric iteration built to force local splitting principles while preserving $\mathsf{DC}$ and killing $\mathsf{AC}$. Here is the dependency flow.
  1. (1)
    Coordinates and products at successor stages. At each successor stage $\alpha+1$ the forcing factor is a finite product of orbit packages, and package coordinates are tracked by the triple system of Definition 5.22. This is the bookkeeping substrate for all later “fix a coordinate” arguments.
  2. (2)
    Diagonal infrastructure. Subsection 5.2 defines iteration-level diagonal lifts $\widehat{\pi}^{\,D}$ and the associated subgroups $\Delta_{\lambda}^{\uparrow}(E,D)$, where $D\subseteq\mathrm{Coords}_{<\lambda}$ is countable and $(E,D)$ is required to be globally admissible (Definition 5.26). The successor “diagonal kernel” elements (Definition 5.23 and Lemma 5.25) supply the cancellation needed to define coherent lifts and to verify group membership.
  3. (3)
    The modified limit filter. Subsection 5.3 defines the limit-stage filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.48) generated by pushforwards together with globally admissible $\Delta^{\uparrow}$-subgroups. The key structural fact is the core lemma (Lemma 5.50): every $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$ contains some $\Delta_{\lambda}^{\uparrow}(E,D)$ with globally admissible parameters. Conjugation is handled by Lemma 5.38 together with the admissible-hull step (Lemma 5.40), so normal-filter generation never “escapes” admissibility.
  4. (4)
    Bookkeeping by names (not interpretations). Subsection 5.4 fixes the class bookkeeping $\mathcal{B}$ of codes for potential $\mathbb{P}_{\alpha}$-names (Definition 5.59). At stage $\alpha$ we test the current interpretation in $M_{\alpha}$ and add the corresponding orbit package factor(s) only when a code currently yields a relevant surjection instance not yet split.
  5. (5)
    Black-box symmetric iteration. Subsection 5.5 invokes the symmetric-iteration framework to assemble the class-length iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$, with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and successor stages $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$ where $\dot{\mathbb{S}}_{\alpha}$ is the finite product of the scheduled orbit package factors.
  6. (6)
    What the iteration yields. Subsection 5.6 extracts the main consequences in the final model $\mathcal{M}$: $\mathsf{ZF}$ and $\mathsf{DC}$ are preserved along the iteration; all surjections $Y\twoheadrightarrow X$ with $X,Y\subseteq T$ split (hence $\mathsf{PP}^{\mathrm{split}}\!\restriction T$); $\mathsf{SVC}(S)$ persists (hence $\mathsf{SVC}^{+}(T)$); $\mathsf{AC}_{\mathsf{WO}}$ holds via scheduled right inverses; the Ryan–Smith localization theorem upgrades $(\mathsf{PP}\!\restriction T\ \wedge\ \mathsf{AC}_{\mathsf{WO}})$ to full $\mathsf{PP}$; and finally $\neg\mathsf{AC}$ holds because the Cohen set $A$ is not well-orderable (Proposition 5.64) using the core lemma plus a nontrivial diagonal lift acting on $A$.
A reader who wants only the global theorem can skim Subsections 5.35.6 while treating the package/diagonal machinery as the engine ensuring the required stabilizers lie in $\tilde{\mathscr{F}}^{*}_{\lambda}$.
We now describe the symmetric iteration that forces the two conjuncts $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ over the Cohen symmetric seed model $\mathcal{N}$, while preserving $\neg\mathsf{AC}$ (witnessed by the non-well-orderability of $A$). All technical iteration infrastructure is imported via the API of Remark 2.1. The purpose of this section is to isolate the package systems, the bookkeeping targets, and the verification that the final symmetric model satisfies the desired principles.
Remark 5.2 (Modularity of the construction).
The argument is deliberately modular: the seed model technology (Section 3), the localized $\mathsf{PP}$-packages (Section 4.1), the $\mathsf{AC}_{\mathsf{WO}}$-packages (Section 4.2), and the symmetric-iteration “API” (Remark 2.1) are separated so that if a hypothesis later needs to be strengthened or a package definition adjusted, it can typically be done locally without rewriting unrelated parts of the paper.
5.1 Package symmetric systems at a stage
Fix a transitive $M\models\mathsf{ZF}+\mathsf{DC}$ extending $\mathcal{N}$ (in practice $M=M_{\alpha}$ at an intermediate stage of the iteration). Keep the parameters $S$ and $T=\mathcal{P}(S)$ as in Remark 4.4.
Remark 5.3 (An ambient $\operatorname{Sym}(\omega_{1})$-action on $T=\mathcal{P}(S)$).
We will often use the ambient action of $\operatorname{Sym}(\omega_{1})$ on the stage-0 seed parameter $S=A^{\omega}$ and on $T=\mathcal{P}(S)$ by permuting the Cohen reals $c_{\alpha}\in A$. Concretely, for $\pi\in\operatorname{Sym}(\omega_{1})$ we set $\pi\cdot c_{\alpha}:=c_{\pi(\alpha)}$, extend pointwise to sequences $s\in A^{\omega}$, and then define $\pi\cdot X:=\{\pi\cdot s:s\in X\}$ for $X\subseteq S$, i.e. $\pi\cdot X\in T$. We regard this as an action which fixes all ordinals pointwise; i.e. $\pi\cdot\xi=\xi$ for every ordinal $\xi$.
We furthermore extend this action to every set $x$ with $\operatorname{tc}(x)\subseteq V(T)$ by recursion on rank:
$\pi\cdot x\;:=\;\{\pi\cdot y:y\in x\}.$
In particular, if $f:X\to Y$ is a function with $\operatorname{tc}(f)\subseteq V(T)$ (e.g. $X,Y\subseteq T$), we identify $f$ with its graph and set
$\pi\cdot f\;:=\;\{(\pi\cdot a,\pi\cdot b):(a,b)\in f\}.$
Lemma 5.4 (Kernel acts trivially on $T$).
Let $\alpha$ be any stage and let $k\in\ker(\pi^{0}_{\alpha})$. Then $k$ fixes $A$, $S=A^{\omega}$, and $T=\mathcal{P}(S)$ pointwise for the ambient $\operatorname{Sym}(\omega_{1})$-action of Remark 5.3. Consequently, $k$ fixes every set/function whose transitive closure is contained in $V(T)\cup\mathrm{Ord}$; in particular, if $f:X\to Y$ with $X,Y\subseteq T$, or $f:S\times\eta\twoheadrightarrow\lambda$ with $\eta,\lambda\in\mathrm{Ord}$, then $k\cdot f=f$.
Proof.
Put $\sigma:=\pi^{0}_{\alpha}(k)\in\operatorname{Sym}(\omega_{1})$. Since $k\in\ker(\pi^{0}_{\alpha})$ we have $\sigma=\operatorname{id}$. By Remark 5.3, the ambient action on $A$, on $S=A^{\omega}$, and on $T=\mathcal{P}(S)$ is induced from the action $\alpha\mapsto\sigma(\alpha)$ on indices of the Cohen reals. Therefore $\sigma=\operatorname{id}$ implies that $k$ fixes $A$, and hence also $S$ and $T$, pointwise.
For the final assertion, use the recursive extension of the action to $V(T)$ from Remark 5.3 and argue by induction on rank that every $x$ with $\operatorname{tc}(x)\subseteq V(T)\cup\mathrm{Ord}$ is fixed pointwise (using that $k$ fixes all ordinals pointwise, and the recursive definition of the action on $V(T)$). In particular, if $f:X\to Y$ with $X,Y\subseteq T$, then $\operatorname{tc}(f)\subseteq V(T)\cup\mathrm{Ord}$ (viewing $f$ as its graph), so $k\cdot f=f$. ∎
Definition 5.5 (Orbit-symmetrized package systems).
Let $M\models\mathsf{ZF}+\mathsf{DC}$ be a transitive stage model and let $\mathscr{G}=\operatorname{Sym}(\omega_{1})$ act on $T$ as in Remark 5.3. Fix a surjection instance $f$ in $M$.
Write $f]:=\{\pi\cdot f:\pi\in\mathscr{G}\}$ for the $\mathscr{G}$-orbit of $f$ (a set in $M$ by Replacement).
  1. (a)
    Orbit $\mathsf{PP}^{\mathrm{split}}\!\restriction T$-package. If $f:Y\twoheadrightarrow X$ with $X,Y\subseteq T$, define
    $\mathbb{Q}_{[f]}:=\prod^{\mathrm{cs}}_{g\in[f]}\mathbb{Q}_{g},$
    where $\mathbb{Q}_{g}$ is the section forcing from Definition 4.12. A condition $\vec{p}\in\mathbb{Q}_{[f]}$ is a function with countable support $\operatorname{supp}(\vec{p}):=\operatorname{dom}(\vec{p})\subseteq[f]$ such that $\vec{p}(g)\in\mathbb{Q}_{g}$ for all $g\in\operatorname{dom}(\vec{p})$, ordered by coordinatewise extension.
  2. (b)
    Orbit $\mathsf{AC}_{\mathsf{WO}}$-package (in $S$-presentation). If $\eta$ is an ordinal, $\lambda<\aleph^{*}(S)^{M}$, and $f:S\times\eta\twoheadrightarrow\lambda$ is a surjection in $M$, define
    $\mathbb{R}_{[f]}:=\prod^{\mathrm{cs}}_{g\in[f]}\mathbb{R}_{g},$
    where $\mathbb{R}_{g}$ is the partial right-inverse forcing from Definition 4.23.
In both cases $\mathscr{G}$ acts on the orbit package by permuting the factors together with the canonical transport induced by $\pi$ on the defining parameters. More precisely, for $\pi\in\mathscr{G}$ and $g\in[f]$ let
$A_{\pi,g}:\mathbb{S}_{g}\to\mathbb{S}_{\pi\cdot g}$
be the transport-of-structure isomorphism induced by $\pi$ (i.e. apply $\pi$ to every ordinal and every element of $T$ appearing in a condition, using the ambient $\mathscr{G}$-action on $T$ from Remark 5.3 and the fact that $\mathbb{S}_{g}$ is defined functorially from $g$). Then the orbit action on conditions is
$(\pi\cdot\vec{p})(g)\ :=\ A_{\pi,\pi^{-1}\cdot g}\bigl(\vec{p}(\pi^{-1}\cdot g)\bigr).$
Lemma 5.6 (Orbit packages are $\omega_{1}$-closed and preserve $\mathsf{DC}$).
Assume $M\models\mathsf{ZF}+\mathsf{DC}$ and let $\mathbb{S}_{[f]}$ be either $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$. Then:
  1. (i)
    $\mathbb{S}_{[f]}$ is $\omega_{1}$-closed in $M$;
  2. (ii)
    $\mathbb{S}_{[f]}$ adds no new $\omega$-sequences of ordinals (in particular, no new reals);
  3. (iii)
    forcing with $\mathbb{S}_{[f]}$ over $M$ preserves $\mathsf{DC}$.
Proof.
(i) Let $\langle\vec{p}_{n}:n<\omega\rangle$ be decreasing in $\mathbb{S}_{[f]}$. For each $g\in[f]$ define $q_{g}:=\bigcup_{n<\omega}p_{n,g}$, where $p_{n,g}$ is the $g$-coordinate of $\vec{p}_{n}$ (taking $p_{n,g}=\emptyset$ if $g\notin\operatorname{supp}(\vec{p}_{n})$). Since each factor $\mathbb{Q}_{g}$ is $\omega_{1}$-closed by Lemma 4.15 (and each factor $\mathbb{R}_{g}$ by Lemma 4.27), $q_{g}$ is a condition in that factor. Let $\vec{q}$ have support $\operatorname{supp}(\vec{q}):=\bigcup_{n<\omega}\operatorname{supp}(\vec{p}_{n})$ and coordinates $q_{g}$ for $g\in\operatorname{supp}(\vec{q})$. Because $M\models\mathsf{DC}$ (hence $\mathsf{AC}_{\omega}$), a countable union of countable sets is countable, so $\operatorname{supp}(\vec{q})$ is countable. Thus $\vec{q}\in\mathbb{S}_{[f]}$ and $\vec{q}\leq\vec{p}_{n}$ for all $n$.
(ii) and (iii) now follow by applying Lemma 4.18 to $\mathbb{P}=\mathbb{S}_{[f]}$, using (i) for countable closure. ∎
Remark 5.7 (Limit-stage $\omega_{1}$-completeness).
At limit stages, $\omega_{1}$-completeness (and normality) of the limit filter is provided by the iteration machinery (API item (d), cf. [3, Thm. 5.5]).
Lemma 5.8 (Well-definedness of the orbit package actions).
In each case of Definition 5.5, if $\vec{p}$ is a condition of the orbit package $\mathbb{S}_{[f]}$ (i.e. $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$) and $\pi\in\operatorname{Sym}(\omega_{1})$, then $\pi\cdot\vec{p}$ is again a condition of the same forcing, and the action respects the order.
Proof.
By construction, each $A_{\pi,g}$ is a poset isomorphism $\mathbb{S}_{g}\cong\mathbb{S}_{\pi\cdot g}$, and the transport maps satisfy the cocycle law
$A_{\pi\circ\rho,g}\ =\ A_{\pi,\rho\cdot g}\circ A_{\rho,g}$
for all $\pi,\rho\in\operatorname{Sym}(\omega_{1})$ and $g\in[f]$. Therefore $(\pi\circ\rho)\cdot\vec{p}=\pi\cdot(\rho\cdot\vec{p})$ and $\operatorname{id}\cdot\vec{p}=\vec{p}$, so this is a well-defined action by poset automorphisms. Since the action only permutes coordinates and applies coordinatewise isomorphisms, it preserves the ideal of supports. ∎
Lemma 5.9 (Countable support of package coordinates).
For every package coordinate $d$ at any successor stage $\alpha+1$ of the iteration, there exists a countable set $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}(d)$, where $\operatorname{Stab}(d)=\{\pi\in\operatorname{Sym}(\omega_{1}):\pi\cdot d=d\}$ is the stabilizer of $d$ as an element of its orbit package.
Proof.
We proceed by induction on stages. At stage 0 there are no package coordinates, so the statement is vacuous.
At successor stage $\alpha+1$, package coordinates are elements of orbit packages $f]=\{\pi\cdot f:\pi\in\operatorname{Sym}(\omega_{1})\}$, where $f$ is a surjection in the symmetric model $M_{\alpha}$.
Step 1: The seed surjection $f$ has countable support.
Since $f\in M_{\alpha}$, there is a hereditarily symmetric $\mathbb{P}_{\alpha}$-name $\dot{f}$ with $f=\dot{f}^{G_{\alpha}}$. By the $\Delta$-support equivalence at stage $\alpha$ (which holds by the inductive hypothesis that the Task 1 filter infrastructure is established at stage $\alpha$), there exists a countable admissible $(E_{f},D_{f})$ such that
$\Delta_{\alpha}^{\uparrow}(E_{f},D_{f})\leq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{f}).$
Step 2: Bridge from $\mathscr{G}_{\alpha}$-stabilizer to $\operatorname{Sym}(\omega_{1})$-stabilizer.
We claim $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}(f)$. Let $\pi\in\operatorname{Fix}(E_{f})$.
By Lemma 5.43 (applied at stage $\alpha$, valid by the inductive hypothesis), there exists a diagonal lift $\widehat{\pi}^{\,D_{f}}\in\mathscr{G}_{\alpha}$ with $\pi^{0}_{\alpha}(\widehat{\pi}^{\,D_{f}})=\pi$ and which fixes every coordinate in $D_{f}$ pointwise. By Definition 5.27, this implies $\widehat{\pi}^{\,D_{f}}\in\Delta_{\alpha}^{\uparrow}(E_{f},D_{f})\leq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{f})$. Hence
$\widehat{\pi}^{\,D_{f}}\cdot\dot{f}=\dot{f}\quad\text{and therefore}\quad\widehat{\pi}^{\,D_{f}}\cdot f=f.$
Let $\widehat{\pi}^{\,\varnothing}\in\mathscr{G}_{\alpha}$ be the standard lift of $\pi$ (the diagonal lift with empty protection set), so $\pi^{0}_{\alpha}(\widehat{\pi}^{\,\varnothing})=\pi$. Set
$k:=(\widehat{\pi}^{\,\varnothing})^{-1}\cdot\widehat{\pi}^{\,D_{f}}\in\mathscr{G}_{\alpha}.$
Then $\pi^{0}_{\alpha}(k)=\operatorname{id}$, i.e. $k\in\ker(\pi^{0}_{\alpha})$. By Lemma 5.4, since $\operatorname{tc}(f)\subseteq V(T)\cup\mathrm{Ord}$, we have $k\cdot f=f$. Therefore,
$f=\widehat{\pi}^{\,D_{f}}\cdot f=\widehat{\pi}^{\,\varnothing}\cdot(k\cdot f)=\widehat{\pi}^{\,\varnothing}\cdot f.$
Finally, note that if $g,h\in\mathscr{G}_{\alpha}$ satisfy $\pi^{0}_{\alpha}(g)=\pi^{0}_{\alpha}(h)$, then $g^{-1}h\in\ker(\pi^{0}_{\alpha})$, so $g\cdot f=h\cdot f$ by Lemma 5.4. In particular, the action of the standard lift $\widehat{\pi}^{\,\varnothing}$ on $f$ depends only on $\pi=\pi^{0}_{\alpha}(\widehat{\pi}^{\,\varnothing})$, hence $\pi\cdot f=\widehat{\pi}^{\,\varnothing}\cdot f=f$. Hence
$\pi\cdot f=\widehat{\pi}^{\,\varnothing}\cdot f=f,$
and thus $\pi\in\operatorname{Stab}(f)$. This proves $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}(f)$.
Step 3: Orbit elements inherit countable support.
For any $g=\sigma\cdot f\in[f]$ (where $\sigma\in\operatorname{Sym}(\omega_{1})$), we have
$\operatorname{Stab}(g)=\operatorname{Stab}(\sigma\cdot f)=\sigma\cdot\operatorname{Stab}(f)\cdot\sigma^{-1}.$
Since $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}(f)$, conjugating gives
$\operatorname{Fix}(\sigma[E_{f}])=\sigma\cdot\operatorname{Fix}(E_{f})\cdot\sigma^{-1}\leq\sigma\cdot\operatorname{Stab}(f)\cdot\sigma^{-1}=\operatorname{Stab}(g).$
Thus $E_{g}:=\sigma[E_{f}]$ is a countable support for $g$. ∎
Remark 5.10 (No circularity).
The proof of Lemma 5.9 invokes Lemma 5.43 (existence of diagonal lifts) at stage $\alpha$. This does not create a circular dependency: the proof of Lemma 5.43 constructs diagonal lifts by direct recursion on stages using the package-level diagonal machinery of Section 5.1, and does not use Lemma 5.9, Lemma 5.40, or Lemma 5.50.
Definition 5.11 (Diagonal-cancellation automorphisms).
Fix an orbit package $\mathbb{S}_{[f]}$ (either $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$). Let $E\subseteq\omega_{1}$ be countable and let $D\subseteq[f]$ be countable such that
$\forall g\in D\ \bigl(\operatorname{Fix}(E)\leq\operatorname{Stab}(g)\bigr),$
i.e. every $\pi\in\operatorname{Fix}(E)$ fixes each $g\in D$ as an orbit element.
For $\pi\in\operatorname{Fix}(E)$ define an automorphism $\widehat{\pi}^{\,D}\in\operatorname{Aut}(\mathbb{S}_{[f]})$ by: for each $\vec{p}\in\mathbb{S}_{[f]}$ and each $g\in[f]$,
$(\widehat{\pi}^{\,D}\cdot\vec{p})(g):=\begin{cases}\vec{p}(g),&g\in D,\\ (\pi\cdot\vec{p})(g),&g\notin D,\end{cases}$
where $\pi\cdot\vec{p}$ is the orbit action from Definition 5.5 (cf. Lemma 5.8). Let
$\Delta(E,D):=\{\widehat{\pi}^{\,D}:\pi\in\operatorname{Fix}(E)\}\ \leq\ \operatorname{Aut}(\mathbb{S}_{[f]}).$
Remark 5.12 (Support convention for diagonal lifts).
Conditions $\vec{p}\in\mathbb{S}_{[f]}$ are partial functions with countable support $\operatorname{supp}(\vec{p})\subseteq[f]$. Fix an admissible pair $(E,D)$ and $\pi\in\operatorname{Fix}(E)$, and write $\vec{p}^{\prime}:=\widehat{\pi}^{\,D}\cdot\vec{p}$.
Then for each $g\in[f]$ we have the coordinate formula
$\vec{p}^{\prime}(g)=\begin{cases}\vec{p}(g),&g\in D,\\[2.0pt] A_{\pi,\pi^{-1}\cdot g}\bigl(\vec{p}(\pi^{-1}\cdot g)\bigr),&g\notin D,\end{cases}$
i.e. $\widehat{\pi}^{\,D}$ agrees with the standard orbit action off $D$ and is the identity on $D$.
Equivalently, for $h\in[f]\setminus D$ we may write
$\vec{p}^{\prime}(\pi\cdot h)=A_{\pi,h}\bigl(\vec{p}(h)\bigr).$
In particular,
$\operatorname{supp}(\vec{p}^{\prime})\;=\;(\operatorname{supp}(\vec{p})\cap D)\ \cup\ \pi\cdot(\operatorname{supp}(\vec{p})\setminus D),$
which is countable since $\pi$ is a bijection.
Definition 5.13 (Orbit-package stage filter (diagonal base)).
Let $\mathscr{G}_{[f]}:=\operatorname{Aut}(\mathbb{S}_{[f]})$.
For countable $E\subseteq\omega_{1}$ and countable $D\subseteq[f]$ such that $\operatorname{Fix}(E)\leq\operatorname{Stab}(g)$ for all $g\in D$, define
$\Delta^{\uparrow}(E,D)\ :=\ \big\langle\ \Delta(E^{\prime},D^{\prime})\ \bigm|\ \begin{array}[]{l}E\subseteq E^{\prime}\in[\omega_{1}]^{\leq\omega},\ D\subseteq D^{\prime}\subseteq[f]\ \text{countable, and}\\[2.0pt] \operatorname{Fix}(E^{\prime})\leq\operatorname{Stab}(g)\ \ \forall g\in D^{\prime}\end{array}\big\rangle,$
where $\Delta(E^{\prime},D^{\prime})$ is as in Definition 5.11.
Call a pair $(E,D)$ admissible if $E\subseteq\omega_{1}$ is countable, $D\subseteq[f]$ is countable, and $\operatorname{Fix}(E)\leq\operatorname{Stab}(g)$ for every $g\in D$.
Let $\mathscr{F}_{[f]}$ be the filter of subgroups of $\mathscr{G}_{[f]}$ generated by the family $\{\Delta^{\uparrow}(E,D)\mid(E,D)\text{ admissible}\}$. Since this family is downward directed (if $(E_{i},D_{i})$ are admissible then $(E_{0}\cup E_{1},\,D_{0}\cup D_{1})$ is admissible and $\Delta^{\uparrow}(E_{0}\cup E_{1},\,D_{0}\cup D_{1})\leq\Delta^{\uparrow}(E_{i},D_{i})$ for $i=0,1$), the generated filter is just the upward closure of this family; equivalently,
$H\in\mathscr{F}_{[f]}\quad\Longleftrightarrow\quad\exists\,\text{admissible}(E,D)\ \bigl(\Delta^{\uparrow}(E,D)\leq H\bigr).$
Remark 5.14.
Lemmas 5.18 and 5.19 show that these diagonal generators are closed under conjugation by the successor-stage groups (generated by standard and diagonal lifts), hence yield the required normal filters in the symmetric-iteration construction.
Lemma 5.15 (Monotonicity of $\Delta^{\uparrow}$).
If $(E_{0},D_{0})$ and $(E_{1},D_{1})$ are admissible and $E_{0}\subseteq E_{1}$ and $D_{0}\subseteq D_{1}$, then
$\Delta^{\uparrow}(E_{1},D_{1})\ \leq\ \Delta^{\uparrow}(E_{0},D_{0}).$
Proof.
Any pair $(E^{\prime},D^{\prime})$ eligible in the definition of $\Delta^{\uparrow}(E_{1},D_{1})$ satisfies $E_{1}\subseteq E^{\prime}$ and $D_{1}\subseteq D^{\prime}$, hence also $E_{0}\subseteq E^{\prime}$ and $D_{0}\subseteq D^{\prime}$. Thus the generating family for $\Delta^{\uparrow}(E_{1},D_{1})$ is a subfamily of the generating family for $\Delta^{\uparrow}(E_{0},D_{0})$, so the generated subgroup is smaller. ∎
Lemma 5.16 ($\omega_{1}$-completeness of the diagonal stage filter).
$\mathscr{F}_{[f]}$ is $\omega_{1}$-complete.
Proof.
Let $\langle H_{n}:n<\omega\rangle$ be members of $\mathscr{F}_{[f]}$. By Definition 5.13, for each $n$ choose an admissible pair $(E_{n},D_{n})$ such that
$\Delta^{\uparrow}(E_{n},D_{n})\ \leq\ H_{n}.$
Let $E:=\bigcup_{n<\omega}E_{n}$ and $D:=\bigcup_{n<\omega}D_{n}$; both are countable. Moreover $(E,D)$ is admissible: if $g\in D$, pick $n$ with $g\in D_{n}$, then $\operatorname{Fix}(E)\subseteq\operatorname{Fix}(E_{n})\leq\operatorname{Stab}(g)$.
By Lemma 5.15, for each $n$ we have
$\Delta^{\uparrow}(E,D)\ \leq\ \Delta^{\uparrow}(E_{n},D_{n})\ \leq\ H_{n}.$
Hence $\Delta^{\uparrow}(E,D)\leq\bigcap_{n<\omega}H_{n}$, so $\bigcap_{n<\omega}H_{n}\in\mathscr{F}_{[f]}$. ∎
Lemma 5.17 (Diagonal-cancellation gives a subgroup of automorphisms).
In the setup of Definition 5.11:
  1. (i)
    For each $\pi\in\operatorname{Fix}(E)$, $\widehat{\pi}^{\,D}$ is an automorphism of $\mathbb{S}_{[f]}$.
  2. (ii)
    The map $\pi\mapsto\widehat{\pi}^{\,D}$ is a group monomorphism $\operatorname{Fix}(E)\hookrightarrow\operatorname{Aut}(\mathbb{S}_{[f]})$. In particular, $\Delta(E,D)$ is a subgroup of $\operatorname{Aut}(\mathbb{S}_{[f]})$.
Proof.
Fix $\pi\in\operatorname{Fix}(E)$. Since $\pi\cdot(\cdot)$ is an automorphism of $\mathbb{S}_{[f]}$ and $\widehat{\pi}^{\,D}$ agrees with it on all coordinates outside $D$ while acting as the identity on coordinates in $D$, it preserves supports, preserves coordinatewise extension, and is bijective with inverse $\widehat{\pi^{-1}}^{\,D}$. This proves (i).
For (ii), $\widehat{\pi}^{\,D}\circ\widehat{\sigma}^{\,D}=\widehat{\pi\sigma}^{\,D}$ follows coordinatewise from the definition (on $D$ both sides are the identity; off $D$ both sides agree with the rigidified orbit action (Lemma 5.8)). Injectivity is immediate. ∎
Lemma 5.18 (Conjugation of diagonal-cancellation groups by standard lifts).
Let $(E,D)$ be admissible. For $\sigma\in\operatorname{Sym}(\omega_{1})$ write
$\sigma\cdot E:=\{\sigma(\xi):\xi\in E\}\subseteq\omega_{1},\qquad\sigma\cdot D:=\{\sigma\cdot g:g\in D\}\subseteq[f].$
Then $(\sigma\cdot E,\sigma\cdot D)$ is admissible. Moreover, for every $\pi\in\operatorname{Fix}(E)$,
$\widehat{\sigma}^{\,\emptyset}\circ\widehat{\pi}^{\,D}\circ(\widehat{\sigma}^{\,\emptyset})^{-1}\;=\;\widehat{\sigma\pi\sigma^{-1}}^{\,\sigma\cdot D}.$
Consequently,
$\widehat{\sigma}^{\,\emptyset}\,\Delta(E,D)\,(\widehat{\sigma}^{\,\emptyset})^{-1}=\Delta(\sigma\cdot E,\sigma\cdot D),\qquad\widehat{\sigma}^{\,\emptyset}\,\Delta^{\uparrow}(E,D)\,(\widehat{\sigma}^{\,\emptyset})^{-1}=\Delta^{\uparrow}(\sigma\cdot E,\sigma\cdot D).$
Proof.
Admissibility. Let $\tau\in\operatorname{Fix}(\sigma\cdot E)$ and let $g\in\sigma\cdot D$, say $g=\sigma\cdot h$ with $h\in D$. Then $\sigma^{-1}\tau\sigma\in\operatorname{Fix}(E)\leq\operatorname{Stab}(h)$, hence
$\tau\cdot g=\tau\cdot(\sigma\cdot h)=\sigma\cdot\bigl((\sigma^{-1}\tau\sigma)\cdot h\bigr)=\sigma\cdot h=g.$
So $\operatorname{Fix}(\sigma\cdot E)\leq\operatorname{Stab}(g)$ for all $g\in\sigma\cdot D$.
Conjugation formula. Fix $\pi\in\operatorname{Fix}(E)$ and $\vec{p}\in\mathbb{S}_{[f]}$, and let $g\in[f]$. Put $h:=\sigma^{-1}\cdot g$. If $g\in\sigma\cdot D$ (equivalently $h\in D$), then $\widehat{\pi}^{\,D}$ acts as the identity at $h$, so the conjugate acts as the identity at $g$, i.e. it fixes the $g$-coordinate of $\vec{p}$.
If $g\notin\sigma\cdot D$ (equivalently $h\notin D$), then at $h$ the map $\widehat{\pi}^{\,D}$ agrees with the standard orbit action by $\pi$. Using that $\widehat{\sigma}^{\,\emptyset}$ is the standard orbit action by $\sigma$ (i.e. $D=\emptyset$), a direct coordinate computation gives that the conjugate agrees with the standard orbit action by $\sigma\pi\sigma^{-1}$ at $g$. Thus the conjugate acts as the identity on $\sigma\cdot D$ and as $\sigma\pi\sigma^{-1}$ off $\sigma\cdot D$, exactly $\widehat{\sigma\pi\sigma^{-1}}^{\,\sigma\cdot D}$.
The subgroup identity for $\Delta(E,D)$ follows by taking images of $\operatorname{Fix}(E)$ under $\pi\mapsto\widehat{\pi}^{\,D}$ and using $\sigma\operatorname{Fix}(E)\sigma^{-1}=\operatorname{Fix}(\sigma\cdot E)$.
Finally, for $\Delta^{\uparrow}(E,D)$: conjugating any generator $\Delta(E^{\prime},D^{\prime})$ occurring in the definition of $\Delta^{\uparrow}(E,D)$ yields $\Delta(\sigma\cdot E^{\prime},\sigma\cdot D^{\prime})$, which is a generator for $\Delta^{\uparrow}(\sigma\cdot E,\sigma\cdot D)$. Applying the same argument with $\sigma^{-1}$ gives equality. ∎
Lemma 5.19 (Conjugation of diagonal generators by diagonal lifts).
Fix an admissible pair $(E_{0},D_{0})$ and let $\sigma\in\operatorname{Fix}(E_{0})$. Write $\widehat{\sigma}^{\,D_{0}}\in\Delta(E_{0},D_{0})$ for the corresponding diagonal lift.
Then for every admissible pair $(E,D)$, letting
$E^{\ast}:=E\cup E_{0},\qquad D^{\ast}:=D\cup D_{0},$
we have
$\widehat{\sigma}^{\,D_{0}}\cdot\Delta^{\uparrow}(E^{\ast},D^{\ast})\cdot(\widehat{\sigma}^{\,D_{0}})^{-1}\;=\;\Delta^{\uparrow}(\sigma\cdot E^{\ast},\sigma\cdot D^{\ast}).$
In particular, $\widehat{\sigma}^{\,D_{0}}\,H\,(\widehat{\sigma}^{\,D_{0}})^{-1}\in\mathscr{F}_{[f]}$ for every $H\in\mathscr{F}_{[f]}$.
Proof.
Let $(E,D)$ be admissible and set $E^{\ast}:=E\cup E_{0}$ and $D^{\ast}:=D\cup D_{0}$. Then $(E^{\ast},D^{\ast})$ is admissible since $\operatorname{Fix}(E^{\ast})\leq\operatorname{Fix}(E)$ and $\operatorname{Fix}(E^{\ast})\leq\operatorname{Fix}(E_{0})$.
We first claim that if $(E^{\prime},D^{\prime})$ is admissible with $E^{\ast}\subseteq E^{\prime}$ and $D^{\ast}\subseteq D^{\prime}$, then
$\widehat{\sigma}^{\,D_{0}}\cdot\Delta(E^{\prime},D^{\prime})\cdot(\widehat{\sigma}^{\,D_{0}})^{-1}\;=\;\Delta(\sigma\cdot E^{\prime},\sigma\cdot D^{\prime}).$
Indeed, $D_{0}\subseteq D^{\prime}$ by assumption, so every $\widehat{\pi}^{\,D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ is the identity on all coordinates in $D_{0}$. Since $\widehat{\sigma}^{\,D_{0}}$ agrees with the standard lift $\widehat{\sigma}^{\,\emptyset}$ off $D_{0}$, the same coordinate computation as in Lemma 5.18 gives
$\widehat{\sigma}^{\,D_{0}}\cdot\widehat{\pi}^{\,D^{\prime}}\cdot(\widehat{\sigma}^{\,D_{0}})^{-1}\;=\;\widehat{\sigma\pi\sigma^{-1}}^{\,\sigma\cdot D^{\prime}}.$
Taking images of $\operatorname{Fix}(E^{\prime})$ under $\pi\mapsto\widehat{\pi}^{\,D^{\prime}}$ and using $\sigma\operatorname{Fix}(E^{\prime})\sigma^{-1}=\operatorname{Fix}(\sigma\cdot E^{\prime})$ yields the subgroup identity above.
Now conjugate the generating family for $\Delta^{\uparrow}(E^{\ast},D^{\ast})$: each generator $\Delta(E^{\prime},D^{\prime})$ (with $E^{\ast}\subseteq E^{\prime}$ and $D^{\ast}\subseteq D^{\prime}$) is sent to $\Delta(\sigma\cdot E^{\prime},\sigma\cdot D^{\prime})$, which is a generator for $\Delta^{\uparrow}(\sigma\cdot E^{\ast},\sigma\cdot D^{\ast})$. Applying the same argument with $\sigma^{-1}$ gives equality.
Finally, if $H\in\mathscr{F}_{[f]}$, choose admissible $(E,D)$ with $\Delta^{\uparrow}(E,D)\leq H$. Then $\Delta^{\uparrow}(E^{\ast},D^{\ast})\leq\Delta^{\uparrow}(E,D)\leq H$ by monotonicity, so
$\Delta^{\uparrow}(\sigma\cdot E^{\ast},\sigma\cdot D^{\ast})\;=\;\widehat{\sigma}^{\,D_{0}}\cdot\Delta^{\uparrow}(E^{\ast},D^{\ast})\cdot(\widehat{\sigma}^{\,D_{0}})^{-1}\;\leq\;\widehat{\sigma}^{\,D_{0}}\,H\,(\widehat{\sigma}^{\,D_{0}})^{-1},$
which shows $\widehat{\sigma}^{\,D_{0}}\,H\,(\widehat{\sigma}^{\,D_{0}})^{-1}\in\mathscr{F}_{[f]}$. ∎
5.2 Iteration-level diagonal automorphisms
The diagonal-cancellation construction of Definition 5.11 is formulated for a single orbit package $\mathbb{S}_{[f]}$. For the $\neg\mathsf{AC}$ argument at limit stages, we need diagonal lifts acting on the full iteration forcing $\mathbb{P}_{\lambda}$, with protection sets $D$ that may span multiple packages across multiple stages.
Remark 5.20 (Simultaneous transfinite induction).
Several results in this section—particularly Lemma 5.9, Lemma 5.40, and Lemma 5.50—are established by simultaneous transfinite induction on stages $\alpha$. At each stage, we assume:
  1. (i)
    the filter infrastructure (generation, completeness, properness) holds at all earlier stages;
  2. (ii)
    diagonal lifts exist in $\mathscr{G}_{\beta}$ for all $\beta<\alpha$ (Lemma 5.43);
  3. (iii)
    the $\Delta$-support equivalence holds at all earlier stages.
These hypotheses are then discharged at stage $\alpha$ using only the inductive assumptions at earlier stages, avoiding any circularity.
Lemma 5.21 (Successor decomposition and splitting).
Let $\alpha$ be an ordinal. In the two-step presentation $\mathbb{P}_{\alpha+1}\cong\mathbb{P}_{\alpha}\ast\dot{\mathbb{Q}}_{\alpha}$, the projection $\pi^{\alpha}_{\alpha+1}:\mathscr{G}_{\alpha+1}\to\mathscr{G}_{\alpha}$ is a surjective group homomorphism. Let $K_{\alpha+1}:=\ker(\pi^{\alpha}_{\alpha+1})$.
Moreover, there is a group-theoretic splitting map $\mathrm{spl}:\mathscr{G}_{\alpha}\to\mathscr{G}_{\alpha+1}$ (the the splitting section) such that $\pi^{\alpha}_{\alpha+1}\circ\mathrm{spl}=\mathrm{id}_{\mathscr{G}_{\alpha}}$, and every $g\in\mathscr{G}_{\alpha+1}$ can be written as $g=\mathrm{spl}(h)\cdot k$ with $h=\pi^{\alpha}_{\alpha+1}(g)$ and $k\in K_{\alpha+1}$.
Proof.
This is exactly the successor-step group structure supplied by the countable-support symmetric iteration framework (Remark 2.1, cf. [3, Def. 3.1]). Concretely:
  • $\pi^{\alpha}_{\alpha+1}$ is restriction to the $\mathbb{P}_{\alpha}$-part;
  • $K_{\alpha+1}$ consists of automorphisms acting trivially on $\mathbb{P}_{\alpha}$;
  • $\mathrm{spl}(h)$ is the splitting section that acts as $h$ on $\mathbb{P}_{\alpha}$ and trivially on $\dot{\mathbb{Q}}_{\alpha}$.
The decomposition $g=\mathrm{spl}(\pi^{\alpha}_{\alpha+1}(g))\cdot k$ with $k=\mathrm{spl}(\pi^{\alpha}_{\alpha+1}(g))^{-1}\cdot g$ is immediate; one checks $k\in K_{\alpha+1}$ since $\pi^{\alpha}_{\alpha+1}(k)=\mathrm{id}$. ∎
Definition 5.22 (Package coordinate sets).
We define by transfinite recursion the sets $\mathrm{Coords}_{\alpha}$ of package coordinates occurring below stage $\alpha$.
  • $\mathrm{Coords}_{0}:=\varnothing$.
  • At a successor stage $\alpha+1$, write $\mathbb{P}_{\alpha+1}\cong\mathbb{P}_{\alpha}*\dot{\mathbb{Q}}_{\alpha}$. In this paper, $\dot{\mathbb{Q}}_{\alpha}$ is (forced to be) a finite product of orbit packages
    $\dot{\mathbb{Q}}_{\alpha}\cong\prod_{i<m_{\alpha}}\dot{\mathbb{S}}_{[f_{\alpha,i}]},$
    where each $\dot{\mathbb{S}}_{[f_{\alpha,i}]}$ is either an orbit $\mathsf{PP}$-package $\mathbb{Q}_{[f]}$ or an orbit $\mathsf{AC}_{\mathsf{WO}}$-package $\mathbb{R}_{[f]}$ (and $m_{\alpha}<\omega$). Define the stage-$\alpha$ index set as the disjoint union
    $I_{\alpha}:=\bigsqcup_{i<m_{\alpha}}[f_{\alpha,i}].$
    We set
    $\begin{split}\mathrm{NewCoords}_{\alpha+1}&:=\{\langle\alpha,z\rangle:z\in I_{\alpha}\},\\ \mathrm{Coords}_{\alpha+1}&:=\mathrm{Coords}_{\alpha}\ \dot{\cup}\ \mathrm{NewCoords}_{\alpha+1}.\end{split}$
    When $z\in I_{\alpha}$ corresponds to $(i,g)$, we also write $\langle\alpha,i,g\rangle$ for $\langle\alpha,z\rangle$.
  • At a limit stage $\lambda$, set $\mathrm{Coords}_{\lambda}:=\bigcup_{\alpha<\lambda}\mathrm{Coords}_{\alpha}$. We also write $\mathrm{Coords}_{<\lambda}:=\mathrm{Coords}_{\lambda}$.
When working at a fixed successor stage $\alpha+1$, we suppress the stage tag and write $\langle i,g\rangle$ for $\langle\alpha,i,g\rangle$ (so $\langle i,g\rangle\in\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$” means $i<n_{\alpha}$ and $g\in[f_{\alpha,i}]$).
We let $\operatorname{Sym}(\omega_{1})$ act on new coordinates by
$\pi\cdot\langle\alpha,i,g\rangle:=\langle\alpha,i,\pi\cdot g\rangle,$
equivalently $\pi\cdot\langle\alpha,z\rangle:=\langle\alpha,\pi\cdot z\rangle$ when using $I_{\alpha}$.
Definition 5.23 (Successor diagonal kernel element).
Fix a successor stage $\alpha+1$. Let $\pi\in\operatorname{Sym}(\omega_{1})$ and let $D^{\mathrm{new}}\subseteq\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$ be a set of package coordinates newly introduced at stage $\alpha+1$.
Assume $\pi$ fixes each $d\in D^{\mathrm{new}}$ under the induced action on tagged coordinates, i.e., if $d=\langle\alpha,i,g\rangle$ then $\pi\cdot d=d$ equivalently $\pi\cdot g=g$. (Here $\pi\cdot\langle\alpha,i,g\rangle:=\langle\alpha,i,\pi\cdot g\rangle$ using the orbit action on $f_{\alpha,i}]$.)
Under this assumption, define $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})\in\operatorname{Aut}(\mathbb{P}_{\alpha+1})$ by:
  • $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ acts as the identity on the $\mathbb{P}_{\alpha}$-part (all coordinates at stages $\leq\alpha$);
  • on each new package coordinate $d\in\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$:
    • if $d\notin D^{\mathrm{new}}$, act by the package action induced by $\pi$ (Definition 5.5);
    • if $d\in D^{\mathrm{new}}$, act as the identity (fix the $d$-coordinate pointwise).
Remark 5.24 (Why the hypothesis is necessary).
The hypothesis $\pi\cdot d=d$ for all $d\in D^{\mathrm{new}}$” ensures $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ is a bijection on package coordinates. Without it, collisions can occur: if $d\notin D^{\mathrm{new}}$ but $\pi\cdot d\in D^{\mathrm{new}}$, then both $d$ and $\pi\cdot d$ would map to $\pi\cdot d$.
In our application (Lemma 5.43), global admissibility guarantees this hypothesis: if $(E,D)$ is globally admissible and $\pi\in\operatorname{Fix}(E)$, then $\pi\cdot d=d$ for all $d\in D$, hence for all $d\in D^{\mathrm{new}}=D_{\beta+1}\setminus D_{\beta}\subseteq D$.
Lemma 5.25 (Successor diagonal kernel element lies in $\mathscr{G}_{\alpha+1}$ and its kernel).
Under the hypothesis of Definition 5.23:
  1. (i)
    $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})\in\mathscr{G}_{\alpha+1}$;
  2. (ii)
    $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})\in K_{\alpha+1}=\ker(\pi^{\alpha}_{\alpha+1})$.
Moreover, for fixed $D^{\mathrm{new}}$, the map $\pi\mapsto\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ is a group homomorphism on its domain $\{\pi\in\operatorname{Sym}(\omega_{1}):\pi\cdot d=d\ \forall d\in D^{\mathrm{new}}\}$ (Here the action on new coordinates is $\pi\cdot\langle\alpha,i,g\rangle:=\langle\alpha,i,\pi\cdot g\rangle$.)
Proof.
(i) By Definition 5.23, the map $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ acts trivially on the $\mathbb{P}_{\alpha}$-part, and on the successor iterand $\dot{\mathbb{Q}}_{\alpha}$ it acts coordinatewise on the new package coordinates $\mathrm{NewCoords}_{\alpha+1}$: it is the identity on $D^{\mathrm{new}}$ and it agrees with the $\pi$-induced orbit action on $\mathrm{NewCoords}_{\alpha+1}\setminus D^{\mathrm{new}}$. If (as allowed by Remark 4.30) $\dot{\mathbb{Q}}_{\alpha}$ is a finite product of orbit packages, then $\mathrm{NewCoords}_{\alpha+1}$ is tagged by the factor index, and the action is defined independently on each factor, hence yields an automorphism of the full stage-$({\alpha+1})$ forcing.
By Lemma 5.17, each factor is a legitimate automorphism of its package forcing $\mathbb{S}_{[f]}$. By the successor-stage group structure (Remark 2.1), $\mathscr{G}_{\alpha+1}$ is generated by lifts from $\mathscr{G}_{\alpha}$ together with automorphisms of the successor iterand $\dot{\mathbb{Q}}_{\alpha}$ (equivalently, of each new orbit-package factor). Since $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ acts as identity on $\mathbb{P}_{\alpha}$ and as a (product of) diagonal-cancellation automorphism(s) on new packages, it lies in $\mathscr{G}_{\alpha+1}$.
(ii) By Definition 5.23, $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ acts as the identity on $\mathbb{P}_{\alpha}$. Hence $\pi^{\alpha}_{\alpha+1}(\kappa_{\alpha+1}(\pi,D^{\mathrm{new}}))=\mathrm{id}$, so $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})\in\ker(\pi^{\alpha}_{\alpha+1})=K_{\alpha+1}$.
The homomorphism property: for $\pi,\sigma$ both fixing $D^{\mathrm{new}}$ pointwise,
$\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})\circ\kappa_{\alpha+1}(\sigma,D^{\mathrm{new}})=\kappa_{\alpha+1}(\pi\sigma,D^{\mathrm{new}}),$
since on coordinates outside $D^{\mathrm{new}}$, composing the $\sigma$-action with the $\pi$-action yields the $(\pi\sigma)$-action, and on $D^{\mathrm{new}}$ both sides act as identity. ∎
Definition 5.26 (Global admissibility).
Let $\lambda$ be a stage of the iteration. A pair $(E,D)$ is globally admissible at stage $\lambda$ if:
  1. (i)
    $E\in[\omega_{1}]^{\leq\omega}$ is countable;
  2. (ii)
    $D\subseteq\mathrm{Coords}_{<\lambda}$ is countable;
  3. (iii)
    $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for every $d\in D$.
When $D=\emptyset$, condition (iii) is vacuously satisfied, and $(E,\emptyset)$ is globally admissible for any countable $E$.
Definition 5.27 (Iteration-level diagonal subgroups and diagonal lifts).
Fix a stage $\lambda$.
(A) Characterized diagonal subgroups. For any $E\subseteq\omega_{1}$ and any $D\subseteq\mathrm{Coords}_{<\lambda}$ define
$\Delta_{\lambda}^{\uparrow}(E,D):=\bigl\{\,g\in\mathscr{G}_{\lambda}:\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)\ \text{and $g$ fixes every coordinate in $D$ pointwise}\,\bigr\}.$
(B) Diagonal lifts (require global admissibility). If $(E,D)$ is globally admissible at stage $\lambda$ and $\pi\in\operatorname{Fix}(E)$, define $\widehat{\pi}^{\,D}\in\operatorname{Aut}(\mathbb{P}_{\lambda})$ by:
  • Stage 0 (Cohen base): act as $\pi$, i.e. $(\widehat{\pi}^{\,D}\cdot p)(\pi(\alpha),n)=p(\alpha,n)$ for $(\alpha,n)\in\omega_{1}\times\omega$.
  • Stages >0>0 (package coordinates): for each package coordinate $g$, act as the identity if $g\in D$, and as the $\pi$-orbit action if $g\notin D$.
Finally set
$\Delta_{\lambda}(E,D):=\{\widehat{\pi}^{\,D}:\pi\in\operatorname{Fix}(E)\}\leq\operatorname{Aut}(\mathbb{P}_{\lambda}).$
Remark 5.28.
If $(E,D)$ is globally admissible and $\pi\in\operatorname{Fix}(E)$, then $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$. Indeed, $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D})=\pi\in\operatorname{Fix}(E)$ and $\widehat{\pi}^{\,D}$ fixes every coordinate in $D$ pointwise.
Remark 5.29 (Characterization vs. generation).
Definition 5.27 defines $\Delta_{\lambda}^{\uparrow}(E,D)$ as a characterized subgroup: $g\in\Delta_{\lambda}^{\uparrow}(E,D)$ iff $g$’s stage-0 projection fixes $E$ pointwise and $g$ fixes every package coordinate in $D$ pointwise.
This “characterization” definition makes monotonicity immediate (Lemma 5.34). In Lemma 5.44, we show that diagonal lifts $\widehat{\pi}^{\,D^{\prime}}$ (for $\pi\in\operatorname{Fix}(E)$ and $D^{\prime}\supseteq D$) lie in $\Delta_{\lambda}^{\uparrow}(E,D)$—this is immediate from the characterization. We do not need to prove that $\Delta_{\lambda}^{\uparrow}(E,D)$ is generated by such diagonal lifts; the characterization suffices for all filter-theoretic purposes.
Lemma 5.30 (Group action on coordinates).
Let $g\in\mathscr{G}_{\lambda}$. Then:
  1. (i)
    The stage-0 projection $\pi^{0}_{\lambda}:\mathscr{G}_{\lambda}\to\operatorname{Sym}(\omega_{1})$ is a group homomorphism.
  2. (ii)
    The element $g$ induces a bijection on the set of package coordinates at each stage $\beta<\lambda$. In particular, if $D$ is a countable set of package coordinates, then $g\cdot D$ is countable.
  3. (iii)
    For $E\subseteq\omega_{1}$, define $g\cdot E:=\pi^{0}_{\lambda}(g)[E]$. Then $g\cdot E$ is countable whenever $E$ is countable.
Proof.
(i) This is part of the inverse/direct system structure of the iteration groups (cf. [3, Def. 3.1]) and the fact that $\pi^{0}_{\lambda}$ is defined stagewise and respects composition.
(ii) We prove by induction on stages that each $g\in\mathscr{G}_{\beta}$ determines a bijection of $\mathrm{Coords}_{\beta}$ (in the sense of Definition 5.22), and hence sends countable sets of coordinates to countable sets.
At stage 0 there are no package coordinates. At a successor stage $\beta+1$, write $g=\mathrm{st}(h)\cdot k$ as in Lemma 5.21, where $h=\pi^{\beta}_{\beta+1}(g)\in\mathscr{G}_{\beta}$ and $k\in K_{\beta+1}$ acts trivially on the $\mathbb{P}_{\beta}$-part. By the inductive hypothesis, $h$ induces a bijection of $\mathrm{Coords}_{\beta}$. The standard lift $\mathrm{st}(h)$ extends this action and fixes every element of $\mathrm{NewCoords}_{\beta+1}$. On the other hand, $k$ fixes $\mathrm{Coords}_{\beta}$ pointwise and acts on the successor forcing $\dot{\mathbb{Q}}_{\beta}$ by compositions of package automorphisms (Lemma 5.8) and diagonal-cancellation automorphisms (Definition 5.11), each of which induces a bijection on the new coordinate index set (and in particular cannot collapse distinct coordinates, by the admissibility hypothesis built into their definition). Therefore $g$ induces a bijection on $\mathrm{Coords}_{\beta+1}=\mathrm{Coords}_{\beta}\dot{\cup}\mathrm{NewCoords}_{\beta+1}$.
At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are (by definition) coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf. [3, Def. 3.1]). The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection.
Consequently, if $D$ is countable then so is its image under the induced bijection.
(iii) Immediate: $\pi^{0}_{\lambda}(g)\in\operatorname{Sym}(\omega_{1})$ is a bijection and hence preserves countability. ∎
Lemma 5.31 (Iteration diagonal lifts are automorphisms).
For $(E,D)$ globally admissible at stage $\lambda$ and $\pi\in\operatorname{Fix}(E)$, the map $\widehat{\pi}^{\,D}$ is an automorphism of $\mathbb{P}_{\lambda}$ satisfying:
  1. (i)
    $\widehat{\pi}^{\,D}$ restricts to $\pi$ on the Cohen base $\operatorname{Add}(\omega,\omega_{1})$;
  2. (ii)
    $\widehat{\pi}^{\,D}$ restricts to the identity on the $g$-coordinate factor for each $g\in D$;
  3. (iii)
    $\widehat{\pi}^{\,D}$ restricts to the standard $\pi$-action on all coordinates outside $D$.
Moreover, $\pi\mapsto\widehat{\pi}^{\,D}$ is a group monomorphism $\operatorname{Fix}(E)\hookrightarrow\operatorname{Aut}(\mathbb{P}_{\lambda})$.
Proof.
By global admissibility, $\pi\cdot g=g$ for all $g\in D$, so the protected coordinates are pointwise fixed and the diagonal definition cannot create coordinate-collisions. Order preservation and bijectivity follow coordinatewise: on the Cohen base this is the standard automorphism; on each package coordinate outside $D$, this is the orbit action from Lemma 5.8; on coordinates in $D$, this is the identity.
The group homomorphism property $\widehat{\pi}^{\,D}\circ\widehat{\sigma}^{\,D}=\widehat{\pi\sigma}^{\,D}$ follows coordinatewise: on $D$ both sides are the identity; on the Cohen base and on coordinates outside $D$, both sides agree with the standard action of $\pi\sigma$. ∎
Corollary 5.32 (Diagonal lifts swap Cohen reals).
Let $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ be a diagonal lift with $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D})=\pi$. Then for each $\alpha<\omega_{1}$:
$\widehat{\pi}^{\,D}\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)},$
where $\dot{c}_{\alpha}$ is the canonical $\mathbb{P}_{0}$-name for the $\alpha$-th Cohen real.
In particular, if $\pi=(\beta\ \gamma)$ is a transposition with $\beta\neq\gamma$, then $\widehat{\pi}^{\,D}$ swaps $c_{\beta}\leftrightarrow c_{\gamma}$ in the generic extension.
Proof.
View $\dot{c}_{\alpha}$ as a $\mathbb{P}_{\lambda}$-name via the canonical inclusion of $\mathbb{P}_{0}$-names into $\mathbb{P}_{\lambda}$-names.
By Lemma 5.31(i), the diagonal lift $\widehat{\pi}^{\,D}$ restricts to $\pi$ on the Cohen base $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$. The canonical name $\dot{c}_{\alpha}$ is defined by $\dot{c}_{\alpha}=\{(\check{n},p):p(\alpha,n)=1\}$. The action of $\pi$ on $\mathbb{P}_{0}$ is $(\pi\cdot p)(\pi(\alpha),n)=p(\alpha,n)$, which induces $\pi\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)}$. Since $\widehat{\pi}^{\,D}$ restricts to $\pi$ on stage 0, we have $\widehat{\pi}^{\,D}\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)}$. ∎
Remark 5.33 (Standard lifts as empty diagonal lifts).
The standard lift $\widehat{\pi}^{\,\emptyset}$ (Definition 5.27 with $D=\emptyset$) acts as $\pi$ on all coordinates, recovering the usual lifted action of $\operatorname{Sym}(\omega_{1})$ on the iteration. Thus the pushforward subgroups $\operatorname{Fix}(E)\leq\mathscr{G}_{\lambda}$ (acting via standard lifts) coincide with $\Delta_{\lambda}(E,\emptyset)$, and consequently $\Delta_{\lambda}^{\uparrow}(E,\emptyset)=\operatorname{Fix}(E)^{\uparrow}$.
This identification is crucial: all filter generators at limit stages—both Cohen-type and package-type—are of the form $\Delta_{\lambda}^{\uparrow}(E_{i},D_{i})$, with $D_{i}=\emptyset$ for Cohen generators and $D_{i}=\{f_{i}\}$ for package generators. The monotonicity lemma (Lemma 5.34) therefore applies uniformly to combine them.
Lemma 5.34 (Monotonicity for iteration-level diagonal groups).
If $E_{0}\subseteq E_{1}$ and $D_{0}\subseteq D_{1}$, then
$\Delta_{\lambda}^{\uparrow}(E_{1},D_{1})\leq\Delta_{\lambda}^{\uparrow}(E_{0},D_{0}).$
Proof.
Let $g\in\Delta_{\lambda}^{\uparrow}(E_{1},D_{1})$. By the characterization definition:
  • $\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E_{1})\subseteq\operatorname{Fix}(E_{0})$ (since $E_{0}\subseteq E_{1}$ implies $\operatorname{Fix}(E_{1})\leq\operatorname{Fix}(E_{0})$).
  • $g$ fixes every coordinate in $D_{1}$ pointwise, hence fixes every coordinate in $D_{0}\subseteq D_{1}$ pointwise.
Thus $g\in\Delta_{\lambda}^{\uparrow}(E_{0},D_{0})$. ∎
Corollary 5.35 (Intersection of $\Delta^{\uparrow}$-groups).
For any family $\{(E_{i},D_{i}):i\in I\}$ (with $I$ finite or countable):
$\bigcap_{i\in I}\Delta_{\lambda}^{\uparrow}(E_{i},D_{i})\supseteq\Delta_{\lambda}^{\uparrow}\left(\bigcup_{i\in I}E_{i},\ \bigcup_{i\in I}D_{i}\right).$
Proof.
By Lemma 5.34, for each $i\in I$: $\Delta_{\lambda}^{\uparrow}(\bigcup_{j}E_{j},\bigcup_{j}D_{j})\leq\Delta_{\lambda}^{\uparrow}(E_{i},D_{i})$. Hence the union-indexed group is contained in every term of the intersection. ∎
Lemma 5.36 (Conjugation of iteration-level diagonal groups).
Let $(E,D)$ be globally admissible at stage $\lambda$. For $\sigma\in\operatorname{Sym}(\omega_{1})$:
$\widehat{\sigma}^{\,\emptyset}\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot(\widehat{\sigma}^{\,\emptyset})^{-1}=\Delta_{\lambda}^{\uparrow}(\sigma\cdot E,\sigma\cdot D).$
In particular, conjugates of diagonal generators remain diagonal generators (with shifted parameters).
Proof.
The proof is identical to Lemma 5.18, now applied at the iteration level. ∎
Remark 5.37 (Admissibility under conjugation up to enlarging $E$).
Let $\lambda$ be a stage and $(E,D)$ be globally admissible at stage $\lambda$. For any $g\in\mathcal{G}_{\lambda}$, there is a countable $E^{\ast}\supseteq g\cdot E$ such that $(E^{\ast},\,g\cdot D)$ is globally admissible. Consequently $\Delta_{\lambda}^{\uparrow}(E^{\ast},\,g\cdot D)\subseteq g\,\Delta_{\lambda}^{\uparrow}(E,D)\,g^{-1}$.
Lemma 5.38 (General conjugation of $\Delta^{\uparrow}$-groups).
For any $g\in\mathscr{G}_{\lambda}$, any $E\subseteq\omega_{1}$, and any set $D$ of package coordinates:
$g\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot g^{-1}=\Delta_{\lambda}^{\uparrow}(g\cdot E,g\cdot D)$
where $g\cdot E:=\pi^{0}_{\lambda}(g)[E]$ and $g\cdot D:=\{g\cdot d:d\in D\}$ is the image of $D$ under the action of $g$ on package coordinates.
Proof.
Fix $k\in\mathscr{G}_{\lambda}$ and write $\sigma:=\pi^{0}_{\lambda}(k)\in\operatorname{Sym}(\omega_{1})$. Let $g\in\Delta^{\uparrow}_{\lambda}(E,D)$, and put $\tau:=\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)$.
Stage-0 part. Since $\pi^{0}_{\lambda}:\mathscr{G}_{\lambda}\to\operatorname{Sym}(\omega_{1})$ is a homomorphism, we have
$\pi^{0}_{\lambda}(kgk^{-1})=\sigma\tau\sigma^{-1}\in\operatorname{Fix}(\sigma[E])=\operatorname{Fix}(\pi^{0}_{\lambda}(k)[E]).$
Coordinate part. Let $d\in D$. By definition of $\Delta^{\uparrow}_{\lambda}(E,D)$, $g$ fixes $d$ pointwise. Using the coordinate action $\cdot$ of $\mathscr{G}_{\lambda}$ on $\mathrm{Coords}_{<\lambda}$ (Lemma 5.30), we compute
$(kgk^{-1})\cdot(k\cdot d)=k\cdot\bigl(g\cdot(k^{-1}\cdot(k\cdot d))\bigr)=k\cdot(g\cdot d)=k\cdot d,$
so $kgk^{-1}$ fixes every coordinate in $k\cdot D$ pointwise.
Thus $kgk^{-1}\in\Delta^{\uparrow}_{\lambda}(\pi^{0}_{\lambda}(k)[E],k\cdot D)$. The reverse inclusion follows by conjugating with $k^{-1}$, so equality holds. ∎
Corollary 5.39 (Admissibility preserved under unions).
If $(E_{1},D_{1})$ and $(E_{2},D_{2})$ are globally admissible, then $(E_{1}\cup E_{2},D_{1}\cup D_{2})$ is globally admissible.
Proof.
For $d\in D_{1}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{1})\leq\operatorname{Stab}(d)$. For $d\in D_{2}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{2})\leq\operatorname{Stab}(d)$. ∎
Lemma 5.40 (Admissible hull for countable coordinate sets).
Let $D$ be a countable set of package coordinates (at stages $\lambda$) and let $E\subseteq\omega_{1}$ be countable. Then there exists a countable set $E^{\ast}\subseteq\omega_{1}$ with $E\subseteq E^{\ast}$ such that $(E^{\ast},D)$ is globally admissible.
Consequently, $\Delta_{\lambda}^{\uparrow}(E^{\ast},D)\leq\Delta_{\lambda}^{\uparrow}(E,D)$.
Proof.
For each $d\in D$, Lemma 5.9 provides a countable $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}(d)$.
Let $E^{\ast}:=E\cup\bigcup_{d\in D}E_{d}$, which is a countable union of countable sets, hence countable.
For each $d\in D$, we have $\operatorname{Fix}(E^{\ast})\leq\operatorname{Fix}(E_{d})\leq\operatorname{Stab}(d)$, so $(E^{\ast},D)$ is globally admissible.
The final statement follows from monotonicity: $E\subseteq E^{\ast}$ implies $\Delta_{\lambda}^{\uparrow}(E^{\ast},D)\leq\Delta_{\lambda}^{\uparrow}(E,D)$ by Lemma 5.34. ∎
Lemma 5.41 (Projection of $\Delta^{\uparrow}$-groups).
Let $\beta<\lambda$ and let $D\subseteq\mathrm{Coords}_{\beta}$ (i.e., $D$ only uses coordinates introduced below stage $\beta$). Then:
$\pi^{\beta}_{\lambda}\big[\Delta_{\lambda}^{\uparrow}(E,D)\big]\subseteq\Delta_{\beta}^{\uparrow}(E,D).$
Proof.
Let $g\in\Delta_{\lambda}^{\uparrow}(E,D)$. Then:
  • $\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)$, and since $\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}$, we have $\pi^{0}_{\beta}(\pi^{\beta}_{\lambda}(g))=\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)$.
  • $g$ fixes every coordinate in $D$ pointwise. Since $D\subseteq\mathrm{Coords}_{\beta}$, these coordinates are visible at stage $\beta$, and the projection $\pi^{\beta}_{\lambda}(g)$ also fixes them pointwise.
Hence $\pi^{\beta}_{\lambda}(g)\in\Delta_{\beta}^{\uparrow}(E,D)$. ∎
Lemma 5.42 (Pullback contains $\Delta^{\uparrow}$).
Let $\beta<\lambda$ and let $D\subseteq\mathrm{Coords}_{\beta}$. If $H\in\tilde{\mathscr{F}}^{*}_{\beta}$ contains $\Delta_{\beta}^{\uparrow}(E,D)$, then $(\pi^{\beta}_{\lambda})^{-1}[H]$ contains $\Delta_{\lambda}^{\uparrow}(E,D)$.
Proof.
Let $g\in\Delta_{\lambda}^{\uparrow}(E,D)$. By Lemma 5.41, $\pi^{\beta}_{\lambda}(g)\in\Delta_{\beta}^{\uparrow}(E,D)\subseteq H$. Hence $g\in(\pi^{\beta}_{\lambda})^{-1}[H]$. ∎
Lemma 5.43 (Iteration-level diagonal lifts lie in $\mathscr{G}_{\lambda}$).
Let $\lambda$ be a stage of the iteration. Let $(E,D)$ be globally admissible at stage $\lambda$ and let $\pi\in\operatorname{Fix}(E)$. Then the iteration-level diagonal lift $\widehat{\pi}^{\,D}\in\operatorname{Aut}(\mathbb{P}_{\lambda})$ (from Definition 5.27) belongs to $\mathscr{G}_{\lambda}$.
Proof.
For each $\beta\leq\lambda$, define $D_{\beta}:=D\cap\mathrm{Coords}_{\beta}$, the set of protected coordinates at stages $\leq\beta$.
Key observation from admissibility: Since $(E,D)$ is globally admissible and $\pi\in\operatorname{Fix}(E)$, we have $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for all $d\in D$, hence $\pi\cdot d=d$ for all $d\in D$. In particular, for each successor $\beta+1\leq\lambda$, the set $D_{\beta+1}\setminus D_{\beta}\subseteq D$ is fixed pointwise by $\pi$. This ensures the successor diagonal kernel element $\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})$ is well-defined (Definition 5.23).
We construct by transfinite recursion a coherent sequence $\langle g_{\beta}:\beta\leq\lambda\rangle$ with $g_{\beta}\in\mathscr{G}_{\beta}$ such that $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$ as automorphisms of $\mathbb{P}_{\beta}$. More precisely, we maintain the inductive invariant that for all $\gamma<\beta\leq\lambda$, $\pi^{\gamma}_{\beta}(g_{\beta})=g_{\gamma}$. In particular, at a limit stage $\eta$ the family $\langle g_{\beta}:\beta<\eta\rangle$ is coherent, so a direct-limit element exists.
Stage 0: Define $g_{0}:=\pi\in\operatorname{Sym}(\omega_{1})=\mathscr{G}_{0}$. Since $\mathrm{Coords}_{0}=\varnothing$, we have $D_{0}=\varnothing$, so $\widehat{\pi}^{\,D_{0}}=\widehat{\pi}^{\,\varnothing}$, which is the standard $\pi$-action on the Cohen forcing (Definition 5.27, stage-0 clause). Hence $g_{0}=\widehat{\pi}^{\,D_{0}}$.
Successor $\beta+1$: Assume $g_{\beta}\in\mathscr{G}_{\beta}$ with $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$. Define
$g_{\beta+1}:=\mathrm{spl}(g_{\beta})\cdot\kappa_{\beta+1}\bigl(\pi,\,D_{\beta+1}\setminus D_{\beta}\bigr),$
where $\mathrm{spl}$ is the splitting section from Lemma 5.21 and $\kappa_{\beta+1}$ is the successor diagonal kernel element from Definition 5.23 (applicable by the key observation above).
By Lemma 5.21, $\mathrm{spl}(g_{\beta})\in\mathscr{G}_{\beta+1}$. By Lemma 5.25(i), $\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})\in\mathscr{G}_{\beta+1}$. Hence $g_{\beta+1}\in\mathscr{G}_{\beta+1}$.
We verify $g_{\beta+1}=\widehat{\pi}^{\,D_{\beta+1}}$ by comparing actions (cf. Definition 5.27):
  • On the $\mathbb{P}_{\beta}$-part: $\mathrm{spl}(g_{\beta})$ acts as $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$, and $\kappa_{\beta+1}$ acts as identity. The combined action is $\widehat{\pi}^{\,D_{\beta}}$, which agrees with $\widehat{\pi}^{\,D_{\beta+1}}$ restricted to stages $\leq\beta$ (since $D_{\beta+1}\cap\mathrm{Coords}_{\beta}=D_{\beta}$).
  • On new coordinates at stage $\beta+1$: $\mathrm{spl}(g_{\beta})$ acts trivially. The kernel element $\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})$:
    • acts as $\pi$ on $d\notin D_{\beta+1}\setminus D_{\beta}$, i.e., on new coordinates not in $D$;
    • acts as identity on $d\in D_{\beta+1}\setminus D_{\beta}$, i.e., on new coordinates in $D$.
    This matches Definition 5.27: “act as $\pi$ if $d\notin D$, identity if $d\in D$.”
Limit $\eta$: Define $g_{\eta}$ to be the direct-limit element of $\mathscr{G}_{\eta}$ determined by the family $\langle g_{\beta}:\beta<\eta\rangle$, which is coherent by the inductive invariant above. By the limit-stage group definition ([3, Def. 3.1]), $g_{\eta}\in\mathscr{G}_{\eta}$.
By the uniqueness property of the inverse/direct limit presentation: elements of $\mathscr{G}_{\eta}$ are determined by their projections to all earlier stages ([3, Def. 3.1]). Since $g_{\eta}$ and $\widehat{\pi}^{\,D_{\eta}}$ have identical restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$ (namely, $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$), we conclude $g_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
Coherence: We verify $\pi^{\gamma}_{\beta}(g_{\beta})=g_{\gamma}$ for $\gamma<\beta\leq\lambda$.
For successor $\beta+1$:
$\pi^{\beta}_{\beta+1}(g_{\beta+1})$ $=\pi^{\beta}_{\beta+1}\bigl(\mathrm{spl}(g_{\beta})\cdot\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})\bigr)$
$=\pi^{\beta}_{\beta+1}(\mathrm{spl}(g_{\beta}))\cdot\pi^{\beta}_{\beta+1}(\kappa_{\beta+1}(\ldots))$
$=g_{\beta}\cdot\mathrm{id}=g_{\beta},$
using $\pi^{\beta}_{\beta+1}\circ\mathrm{spl}=\mathrm{id}$ (Lemma 5.21) and $\kappa_{\beta+1}(\ldots)\in K_{\beta+1}=\ker(\pi^{\beta}_{\beta+1})$ (Lemma 5.25(ii)).
For limits $\eta$, coherence is automatic: $g_{\eta}$ is defined as the direct-limit element, so $\pi^{\gamma}_{\eta}(g_{\eta})=g_{\gamma}$ for all $\gamma<\eta$ by construction.
Conclusion: Taking $\beta=\lambda$, we have $g_{\lambda}=\widehat{\pi}^{\,D_{\lambda}}$. Since $D_{\lambda}=D\cap\mathrm{Coords}_{\lambda}=D$ (since $D\subseteq\mathrm{Coords}_{<\lambda}$ by Definition 5.26(ii)), we have $g_{\lambda}=\widehat{\pi}^{\,D}$ and $g_{\lambda}\in\mathscr{G}_{\lambda}$. Hence $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$. ∎
Lemma 5.44 (Diagonal lifts lie in $\Delta^{\uparrow}$).
Let $\pi\in\operatorname{Fix}(E)$ and let $D^{\prime}\supseteq D$ be such that the diagonal lift $\widehat{\pi}^{\,D^{\prime}}\in\mathscr{G}_{\lambda}$ exists. Then $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
Proof.
By the characterization definition of $\Delta_{\lambda}^{\uparrow}(E,D)$ (Definition 5.27), we verify:
  1. (i)
    $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D^{\prime}})\in\operatorname{Fix}(E)$: By construction of diagonal lifts (Lemma 5.31(i)), the stage-0 projection of $\widehat{\pi}^{\,D^{\prime}}$ is $\pi$, and $\pi\in\operatorname{Fix}(E)$ by assumption.
  2. (ii)
    $\widehat{\pi}^{\,D^{\prime}}$ fixes every coordinate in $D$ pointwise: By construction (Lemma 5.31(ii)), $\widehat{\pi}^{\,D^{\prime}}$ fixes every coordinate in $D^{\prime}$ pointwise, and $D\subseteq D^{\prime}$.
Hence $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$. ∎
Lemma 5.45 (Package generic sections are hereditarily symmetric via diagonal cancellation).
Let $f$ have countable support $E$ (so $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$) and let $\dot{s}_{f}$ be the canonical $\mathbb{S}_{[f]}$-name for the $f$-coordinate generic section. Then $\Delta^{\uparrow}(E,\{f\})\leq\operatorname{Stab}(\dot{s}_{f})$, hence $\dot{s}_{f}\in\operatorname{HS}$ (with respect to $\mathscr{F}_{[f]}$).
Proof.
Let $E\subseteq\omega_{1}$ witness that $f$ has countable support, i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$.
Fix any admissible pair $(E^{\prime},D^{\prime})$ such that $E\subseteq E^{\prime}$ and $\{f\}\subseteq D^{\prime}$. Let $\pi\in\operatorname{Fix}(E^{\prime})$. Then the diagonal lift $\hat{\pi}^{D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ fixes the $f$-coordinate pointwise: for every condition $\vec{p}\in\mathbb{S}_{[f]}$,
$(\hat{\pi}^{D^{\prime}}\cdot\vec{p})(f)=\vec{p}(f).$
Since $\dot{s}_{f}$ is definable (as a $\mathbb{S}_{[f]}$-name) purely from the $f$-coordinate generic section, it follows that $\hat{\pi}^{D^{\prime}}\cdot\dot{s}_{f}=\dot{s}_{f}$. Hence $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}(\dot{s}_{f})$.
As $\Delta^{\uparrow}(E,\{f\})$ is generated by all such subgroups $\Delta(E^{\prime},D^{\prime})$, we have $\Delta^{\uparrow}(E,\{f\})\leq\operatorname{Stab}(\dot{s}_{f})$. Therefore $\operatorname{Stab}(\dot{s}_{f})\in\mathscr{F}_{[f]}$, so $\dot{s}_{f}\in\operatorname{HS}$. ∎
Lemma 5.46 (Diagonal cancellation makes right-inverse generics hereditarily symmetric).
Let $g$ be a package with countable support witnessed by $E\subseteq\omega_{1}$, i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(g)$. Let $\dot{r}_{g}$ be the canonical $\mathbb{R}_{[g]}$-name for the $g$-coordinate generic right inverse. Then $\Delta^{\uparrow}(E,\{g\})\leq\operatorname{Stab}(\dot{r}_{g})$, and hence $\dot{r}_{g}\in\operatorname{HS}$ (with respect to $(\mathbb{R}_{[g]},\mathscr{G}_{[g]},\mathscr{F}_{[g]})$).
Proof.
The proof is identical to Lemma 5.45. Fix an admissible pair $(E^{\prime},D^{\prime})$ with $E\subseteq E^{\prime}$ and $\{g\}\subseteq D^{\prime}$. For $\pi\in\operatorname{Fix}(E^{\prime})$, the diagonal lift $\hat{\pi}^{D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ fixes the $g$-coordinate pointwise, and $\dot{r}_{g}$ is definable purely from the $g$-coordinate generic object. Hence $\hat{\pi}^{D^{\prime}}\cdot\dot{r}_{g}=\dot{r}_{g}$, so $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}(\dot{r}_{g})$. Therefore $\Delta^{\uparrow}(E,\{g\})\leq\operatorname{Stab}(\dot{r}_{g})$, as required. ∎
Lemma 5.47 (Successor-step HS lift for package generics).
Fix a stage $\beta$ and a package coordinate $d\in\mathrm{NewCoords}_{\beta+1}=\mathrm{Coords}_{\beta+1}\setminus\mathrm{Coords}_{\beta}$ scheduled at stage $\beta$. (e.g. $d=f_{\alpha}$ for a splitting package $\mathbb{Q}_{[f_{\alpha}]}$, or $d=g_{\alpha}$ for a right-inverse package $\mathbb{R}_{[g_{\alpha}]}$). Let $\dot{\tau}$ be a canonical $\dot{\mathbb{Q}}_{\beta}$-name definable purely from the $d$-coordinate generic object (e.g. $\dot{s}_{f_{\alpha}}$ or $\dot{r}_{g_{\alpha}}$).
Assume $d$ has countable support witnessed by $E\subseteq\omega_{1}$ (so $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ in $M_{\beta}$). Then the canonical $\mathbb{P}_{\beta+1}$-name $\dot{\tau}^{*}$ is hereditarily symmetric at stage $\beta+1$; in particular, $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
Proof.
Since $d$ has countable support witnessed by $E$ in $M_{\beta}$, we have $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$. Thus $(E,\{d\})$ is globally admissible at stage $\beta+1$ (Definition 5.26), so $\Delta_{\beta+1}^{\uparrow}(E,\{d\})$ is one of the $\Delta^{\uparrow}$-generators of $\tilde{\mathscr{F}}^{*}_{\beta+1}$ (Definition 5.48).
Now let $g\in\Delta_{\beta+1}^{\uparrow}(E,\{d\})$. By the characterization in Definition 5.27, $g$ fixes the coordinate $d$ pointwise (Here “fixes the coordinate $d$ pointwise” is exactly the clause in Definition 5.27: the induced automorphism on the $d$-factor forcing ($\dot{\mathbb{Q}}_{d}$ or $\dot{\mathbb{R}}_{d}$, depending on the package) is literally the identity, i.e. $g\cdot\dot{p}=\dot{p}$ for every canonical name $\dot{p}$ for a condition in the $d$-factor.)
Since $\dot{\tau}^{*}$ is definable purely from the $d$-coordinate generic object, it follows that $g\cdot\dot{\tau}^{*}=\dot{\tau}^{*}$. Hence $\Delta_{\beta+1}^{\uparrow}(E,\{d\})\leq\operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*})$.
Therefore $\operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*})\in\tilde{\mathscr{F}}^{*}_{\beta+1}$ by upward closure, so $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$ and $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$. ∎
5.3 The modified limit filter
We now define the filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ that governs hereditary symmetry at limit stages of the iteration.
Definition 5.48 (Modified limit filter).
Define the filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ on $\mathscr{G}_{\lambda}$ inductively:
Stage 0: $\tilde{\mathscr{F}}^{*}_{0}$ is the filter on $\mathscr{G}_{0}=\operatorname{Sym}(\omega_{1})$ generated by $\{\operatorname{Fix}(E):E\subseteq\omega_{1}\text{ countable}\}$. Note that $\operatorname{Fix}(E)=\Delta_{0}^{\uparrow}(E,\varnothing)$, and $(E,\varnothing)$ is trivially globally admissible.
Successor stage $\alpha+1$: $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter on $\mathscr{G}_{\alpha+1}$ generated by:
  1. (a)
    Pushforwards: $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$;
  2. (b)
    $\Delta^{\uparrow}$-subgroups: $\Delta_{\alpha+1}^{\uparrow}(E,D)$ for countable $E\subseteq\omega_{1}$ and countable $D\subseteq\mathrm{Coords}_{<\alpha+1}$ such that $(E,D)$ is globally admissible at stage $\alpha+1$.
Limit stage $\lambda$: $\tilde{\mathscr{F}}^{*}_{\lambda}$ is the normal filter on $\mathscr{G}_{\lambda}$ generated by:
  1. (a)
    Pushforwards: $(\pi^{\beta}_{\lambda})^{-1}[H]$ for $\beta<\lambda$ and $H\in\tilde{\mathscr{F}}^{*}_{\beta}$;
  2. (b)
    $\Delta^{\uparrow}$-subgroups: $\Delta_{\lambda}^{\uparrow}(E,D)$ for countable $E\subseteq\omega_{1}$ and countable $D$ (package coordinates at stages $\lambda$) such that $(E,D)$ is globally admissible at stage $\lambda$.
Remark 5.49 (Admissibility in the filter definition).
We require global admissibility in the $\Delta^{\uparrow}$-generator clause to ensure that the core lemma (Lemma 5.50) produces admissible pairs, which is needed for the properness proof (Lemma 5.52(iii)).
By Lemma 5.38, conjugation sends $\Delta_{\lambda}^{\uparrow}(E,D)$ to $\Delta_{\lambda}^{\uparrow}(g\cdot E,\,g\cdot D)$, and $g\cdot E$ and $g\cdot D$ remain countable (Lemma 5.30). The pair $(g\cdot E,\,g\cdot D)$ need not be globally admissible, but by Lemma 5.40 there is a countable $E^{\ast}\supseteq g\cdot E$ such that $(E^{\ast},\,g\cdot D)$ is globally admissible and $\Delta_{\lambda}^{\uparrow}(E^{\ast},\,g\cdot D)\leq\Delta_{\lambda}^{\uparrow}(g\cdot E,\,g\cdot D)$. Together with Corollary 5.39, this ensures that the normal filter generation process always contains $\Delta^{\uparrow}$-groups with globally admissible parameters.
Lemma 5.50 (Every filter element contains a $\Delta^{\uparrow}$-group).
For every $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$, there exist countable $E\subseteq\omega_{1}$ and countable $D$ (package coordinates at stages $\lambda$) with $(E,D)$ globally admissible such that:
$\Delta_{\lambda}^{\uparrow}(E,D)\leq H.$
Proof.
By definition of the normal filter generated by a family, every element $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$ contains a finite intersection of conjugates of generators. We show that such intersections contain $\Delta^{\uparrow}$-groups with globally admissible parameters.
Conjugates of $\Delta^{\uparrow}$-generators: Let $\Delta_{\lambda}^{\uparrow}(E,D)$ be a globally admissible generator and let $g\in\mathscr{G}_{\lambda}$. By Lemma 5.38:
$g\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot g^{-1}=\Delta_{\lambda}^{\uparrow}(g\cdot E,g\cdot D).$
Note that $g\cdot E$ and $g\cdot D$ are countable: $g\cdot E=\pi^{0}_{\lambda}(g)[E]$ is countable because bijections preserve cardinality, and similarly $g\cdot D$ is countable because the action on package coordinates is a bijection (Lemma 5.30(ii)). The parameters $(g\cdot E,g\cdot D)$ need not be globally admissible. However, by Lemma 5.40, there exists $E^{\ast}\supseteq g\cdot E$ such that $(E^{\ast},g\cdot D)$ is globally admissible and $\Delta_{\lambda}^{\uparrow}(E^{\ast},g\cdot D)\leq\Delta_{\lambda}^{\uparrow}(g\cdot E,g\cdot D)$.
Thus every conjugate of an admissible $\Delta^{\uparrow}$-generator contains a globally admissible $\Delta^{\uparrow}$-group.
Finite intersections: If $H_{1}\supseteq\Delta_{\lambda}^{\uparrow}(E_{1},D_{1})$ and $H_{2}\supseteq\Delta_{\lambda}^{\uparrow}(E_{2},D_{2})$ with $(E_{i},D_{i})$ globally admissible, then by Corollary 5.35:
$H_{1}\cap H_{2}\supseteq\Delta_{\lambda}^{\uparrow}(E_{1}\cup E_{2},D_{1}\cup D_{2}).$
By Corollary 5.39, $(E_{1}\cup E_{2},D_{1}\cup D_{2})$ is globally admissible.
Pullback generators: Let $H=(\pi^{\beta}_{\lambda})^{-1}[K]$ with $K\in\tilde{\mathscr{F}}^{*}_{\beta}$. By induction on $\beta$, there exist countable admissible $(E,D)$ with $D\subseteq\mathrm{Coords}_{\beta}$ and $\Delta_{\beta}^{\uparrow}(E,D)\leq K$. By Lemma 5.42, $\Delta_{\lambda}^{\uparrow}(E,D)\leq H$. The pair $(E,D)$ remains globally admissible at stage $\lambda$ (the admissibility condition $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ depends only on how $d$ was introduced, not on which stage we evaluate it). ∎
Corollary 5.51 (Admissibility preserved under countable unions).
If $\{(E_{n},D_{n}):n<\omega\}$ is a countable family of globally admissible pairs, then $(\bigcup_{n}E_{n},\bigcup_{n}D_{n})$ is globally admissible.
Proof.
For $d\in\bigcup_{n}D_{n}$, choose $k$ with $d\in D_{k}$. Then $\operatorname{Fix}(\bigcup_{n}E_{n})\leq\operatorname{Fix}(E_{k})\leq\operatorname{Stab}(d)$, where the last inequality holds by global admissibility of $(E_{k},D_{k})$. ∎
Lemma 5.52 ($\tilde{\mathscr{F}}^{*}_{\lambda}$ is a proper normal filter).
For each stage $\lambda$:
  1. (i)
    $\tilde{\mathscr{F}}^{*}_{\lambda}$ is a filter (upward closed, closed under finite intersections).
  2. (ii)
    $\tilde{\mathscr{F}}^{*}_{\lambda}$ is normal (closed under conjugation by $\mathscr{G}_{\lambda}$).
  3. (iii)
    $\tilde{\mathscr{F}}^{*}_{\lambda}$ is proper ($\{\mathrm{id}\}\notin\tilde{\mathscr{F}}^{*}_{\lambda}$).
  4. (iv)
    $\tilde{\mathscr{F}}^{*}_{\lambda}$ is $\omega_{1}$-complete.
Proof.
(i) and (ii) follow from the definition as a normal filter generated by the given families.
(iii) By Lemma 5.50, if $\{\mathrm{id}\}\in\tilde{\mathscr{F}}^{*}_{\lambda}$, then $\{\mathrm{id}\}\supseteq\Delta_{\lambda}^{\uparrow}(E,D)$ for some countable globally admissible $E,D$.
Since $E$ is countable and $\omega_{1}$ is uncountable, there exist distinct $\beta,\gamma\in\omega_{1}\setminus E$. Let $\pi=(\beta\ \gamma)$ (transposition). Then $\pi\in\operatorname{Fix}(E)$.
By Lemma 5.43, the diagonal lift $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ exists. By Lemma 5.44, $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
Since $\widehat{\pi}^{\,D}$ acts as $\pi$ on the Cohen base (Lemma 5.31(i)), it swaps coordinates $\beta$ and $\gamma$, so $\widehat{\pi}^{\,D}\neq\mathrm{id}$.
Hence $\Delta_{\lambda}^{\uparrow}(E,D)$ contains a nontrivial element, contradicting $\Delta_{\lambda}^{\uparrow}(E,D)\leq\{\mathrm{id}\}$.
(iv) $\omega_{1}$-completeness. Let $\{H_{n}:n<\omega\}\subseteq\tilde{\mathscr{F}}^{*}_{\lambda}$. By Lemma 5.50, for each $n$ there exists a globally admissible $(E_{n},D_{n})$ with $\Delta_{\lambda}^{\uparrow}(E_{n},D_{n})\leq H_{n}$.
By Corollary 5.35:
$\bigcap_{n<\omega}H_{n}\supseteq\Delta_{\lambda}^{\uparrow}\left(\bigcup_{n}E_{n},\bigcup_{n}D_{n}\right).$
By Corollary 5.51, $(\bigcup_{n}E_{n},\bigcup_{n}D_{n})$ is globally admissible. Since countable unions of countable sets are countable (in the metatheory, which has $\mathsf{AC}$), the right-hand side is a $\Delta^{\uparrow}$-generator of the filter.
Hence $\bigcap_{n}H_{n}\in\tilde{\mathscr{F}}^{*}_{\lambda}$. ∎
Definition 5.53 (Hereditarily symmetric names).
A $\mathbb{P}_{\lambda}$-name $\tau$ is hereditarily symmetric ($\tau\in\operatorname{HS}_{\lambda}$) iff:
  1. (i)
    $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}$, and
  2. (ii)
    for every $(\sigma,p)\in\tau$, we have $\sigma\in\operatorname{HS}_{\lambda}$.
The symmetric extension at stage $\lambda$ is $M_{\lambda}=\operatorname{HS}_{\lambda}^{G_{\lambda}}=\{\tau^{G_{\lambda}}:\tau\in\operatorname{HS}_{\lambda}\}$.
Lemma 5.54 ($\Delta$-support equivalence).
For any $\mathbb{P}_{\lambda}$-name $\tau$:
$\operatorname{Stab}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}\iff\exists\text{ countable globally admissible }(E,D)\text{ such that }\Delta_{\lambda}^{\uparrow}(E,D)\subseteq\operatorname{Stab}(\tau).$
Proof.
$(\Leftarrow)$ By upward closure, since $\Delta_{\lambda}^{\uparrow}(E,D)\in\tilde{\mathscr{F}}^{*}_{\lambda}$ for globally admissible $(E,D)$. $(\Rightarrow)$ By Lemma 5.50. ∎
Theorem 5.55 ($M_{\lambda}\models\mathsf{ZF}$).
For each stage $\lambda$, the symmetric extension $M_{\lambda}=\operatorname{HS}_{\lambda}^{G_{\lambda}}$ satisfies $\mathsf{ZF}$.
Proof.
By Lemma 5.52, the triple $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ is a symmetric system in the sense of [4, Ch. 15]. By the symmetric extension theorem [4, Lemma. 15.51], $M_{\lambda}\models\mathsf{ZF}$. ∎
Theorem 5.56 ($\mathcal{M}\models\mathsf{ZF}$).
The final model $\mathcal{M}=\bigcup_{\alpha\in\mathrm{Ord}}M_{\alpha}$ satisfies $\mathsf{ZF}$.
Proof.
$\mathcal{M}$ is the directed union of the transitive models $\{M_{\alpha}:\alpha\in\mathrm{Ord}\}$, each satisfying $\mathsf{ZF}$ by Theorem 5.55.
For any $\mathsf{ZF}$ axiom instance with parameters $a_{1},\ldots,a_{n}\in\mathcal{M}$: choose $\lambda$ large enough that $a_{1},\ldots,a_{n}\in M_{\lambda}$. Apply the axiom in $M_{\lambda}$ to obtain the required witness in $M_{\lambda}\subseteq\mathcal{M}$.
The inclusion $M_{\beta}\subseteq M_{\lambda}$ for $\beta<\lambda$ is guaranteed by the pushforward generators: if $\tau\in\operatorname{HS}_{\beta}$ with $\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\beta}$, then $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\supseteq(\pi^{\beta}_{\lambda})^{-1}[\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)]\in\tilde{\mathscr{F}}^{*}_{\lambda}$, so $\tau\in\operatorname{HS}_{\lambda}$. ∎
5.4 Bookkeeping: names versus interpretations
Remark 5.57 (Background metatheory for class-length constructions).
To speak formally about an $\mathrm{Ord}$-length bookkeeping function and an $\mathrm{Ord}$-length symmetric iteration as class objects, we work in a background theory of classes, e.g. $\mathsf{GBC}+\mathsf{ETR}$ (optionally with Global Choice). Its set universe is denoted $V$ and satisfies $\mathsf{ZFC}$. All symmetric models $\mathcal{N}$, intermediate stages $M_{\alpha}$, and the final model $\mathcal{M}$ are constructed inside this set universe; only the bookkeeping map $\mathcal{B}$ and the iteration itself are treated as class objects in the ambient background.
We now fix a bookkeeping device which ensures that every relevant surjection instance appearing in the final model is handled at some successor stage.
Remark 5.58 (Syntactic bookkeeping).
In a class-length iteration, the interpretation of a fixed name can change as the iteration proceeds. Therefore, we bookkeep names (codes) rather than interpreted objects: at stage $\alpha$ we take the $\alpha$-th code $\dot{f}$ from our fixed enumeration, and in the current intermediate model $M_{\alpha}$ we check whether $\dot{f}$ (as currently interpreted) is a surjection instance of the required type. If it is not, we force trivially at that stage (or include only the other package factors scheduled for $\alpha$). This is the standard bookkeeping convention for class-length forcing/iterations.
Definition 5.59 (Bookkeeping of orbit instances).
Fix in the metatheory a definable class function $\mathcal{B}:\mathrm{Ord}\to V$ that enumerates pairs $(\beta,c)$ where $c$ is a set-theoretic code for a potential $\mathbb{P}$-name (in some standard Gödel coding). The enumeration is arranged so that for each code $c$, the set
$\{\alpha\in\mathrm{Ord}:\exists\,\beta\ (\mathcal{B}(\alpha)=(\beta,c))\}$
is unbounded in $\mathrm{Ord}$ (e.g. enumerate $\mathrm{Ord}\times V$ using a definable pairing of ordinals together with a definable class well-order of $V$ in the background metatheory).
At stage $\alpha$, let $\mathcal{B}(\alpha)=(\beta,c)$. We test whether the code $c$:
  1. (i)
    decodes to a valid $\mathbb{P}_{\alpha}$-name $\dot{f}$,
  2. (ii)
    is hereditarily symmetric for the stage-$\alpha$ symmetric system $(\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha})$, and
  3. (iii)
    interprets in $M_{\alpha}$ to a surjection of one of the following types:
    1. (1)
      $f:Y\twoheadrightarrow X$ with $X,Y\subseteq T$ (for $\mathsf{PP}$-packages), or
    2. (2)
      $f:S\times\eta\twoheadrightarrow\lambda$ with $\eta,\lambda\in\mathrm{Ord}$ (for $\mathsf{AC}_{\mathsf{WO}}$-packages).
If all conditions are met and the surjection $f$ does not already split in $M_{\alpha}$ (i.e., no section/right-inverse for $f$ exists yet), the successor forcing at stage $\alpha+1$ includes the corresponding orbit package $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ (Definition 5.5); otherwise we force trivially at that code.
This arrangement ensures that every code is tested at unboundedly many stages, so a name that becomes a valid HS surjection at some stage $\beta$ will eventually be scheduled at some stage $\alpha>\beta$.
Remark 5.60 (The $\aleph^{*}(S)$ horizon).
For the $\mathsf{AC}_{\mathsf{WO}}$-component we schedule splittings for surjections $f:Y\twoheadrightarrow\lambda$ with $\lambda<\aleph^{*}(S)$ (cf. Remark 4.22). This is the natural bookkeeping horizon: $\aleph^{*}(S)$ is the first ordinal $\theta$ (in the ambient stage model under discussion) such that there is no surjection $S\twoheadrightarrow\theta$, so $\lambda<\aleph^{*}(S)$ is exactly the range of well-ordered indices reachable from $S$-parameterized data.
For our purposes we only use the following two points.
  1. (i)
    The seed parameter $S=A^{\omega}$ is preserved at successor stages: our package forcings add no new reals (Corollaries 4.19 and 4.29), hence add no new $\omega$-sequences of Cohen reals and therefore no new elements of $S$.
  2. (ii)
    The value of $\aleph^{*}(S)$ need not be constant along the iteration (new functions from $S$ may be added even if $S$ itself is unchanged), but this causes no bookkeeping gap: we test the bound $\lambda<\aleph^{*}(S)$ in the current stage model when a code is examined. If at some stage a given name $\dot{f}$ interprets to a surjection instance of the required form with $\lambda<\aleph^{*}(S)$ (as computed in that stage), then from that point onward the instance is eligible to be scheduled. Since every code is retested at unboundedly many stages (Definition 5.59), no a priori stability theorem for $\aleph^{*}(S)$ is needed.
5.5 The symmetric iteration (black box)
We now invoke the iteration framework via Remark 2.1. For the global conclusion (a final model of $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$), we take a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$. Concretely, working in the metatheory, we form an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [3, §2.1], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$.
At a successor stage $\alpha+1$, we consult the finite list $\mathcal{B}(\alpha)$ from Definition 5.59. In the current model $M_{\alpha}$, each code on this list either interprets to a genuine package instance (yielding a corresponding orbit package forcing $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ as in Definition 5.5), or it does not (in which case we take the trivial forcing for that code). Let $\dot{\mathbb{S}}_{\alpha}$ be a $\mathbb{P}_{\alpha}$-name for the finite-support product of the resulting orbit package posets. Then $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$.
At a successor stage $\alpha+1$, the group $\mathscr{G}_{\alpha+1}$ and filter $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ are those determined by the symmetric-iteration framework together with the diagonal infrastructure from Subsection 5.2 and the filter definition in Definition 5.48. Concretely, $\mathscr{G}_{\alpha+1}$ is generated by standard lifts of $\mathscr{G}_{\alpha}$ together with the (stage-$\alpha$) package automorphisms and the successor diagonal kernel elements $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ (Definition 5.23) acting on the new package coordinates, and $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter generated by pushforwards from $\tilde{\mathscr{F}}^{*}_{\alpha}$ together with the admissible subgroups $\Delta^{\uparrow}_{\alpha+1}(E,D)$. In particular, for each active package coordinate $d$ at stage $\alpha+1$ with countable support witnessed by $E_{d}$, the generator $\Delta^{\uparrow}_{\alpha+1}(E_{d},\{d\})$ belongs to $\tilde{\mathscr{F}}^{*}_{\alpha+1}$, so the canonical package generics are hereditarily symmetric by Lemmas 5.45 and 5.46.
Thus each successor filter explicitly includes the diagonal subgroups protecting the representative coordinates of the active packages.
At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group/filter as provided by the symmetric-iteration framework (API items (b)–(e), especially filter generation/completeness at limits, [3, Thm. 5.5]). We write $\mathcal{M}$ for the final symmetric model of the $\mathrm{Ord}$-length iteration.
Remark 5.61 (No one-step collapse in general).
Unlike ordinary forcing iterations in $\mathsf{ZFC}$, there is no general theorem showing that an iteration of symmetric extensions can always be presented as a single symmetric extension of the ground model. Indeed, the usual “two-step collapse” phenomena from forcing do not directly apply in the symmetric context, and the literature discusses serious obstacles to such a reduction. It remains open in general to what extent arbitrary symmetric iterations can be compressed to a single symmetric extension; see the discussion in [5, 14].
5.6 What the iteration yields
Let $\mathcal{M}$ denote the final symmetric model obtained from the iteration.
Proposition 5.62 (Final localized splitting $\mathsf{PP}^{\mathrm{split}}\!\restriction T$).
$\mathcal{M}\models\mathsf{PP}^{\mathrm{split}}\!\restriction T$.
Proof.
Work in the final symmetric model $\mathcal{M}$. Let $f:Y\twoheadrightarrow X$ be a surjection in $\mathcal{M}$ with $X,Y\subseteq T$.
Choose a hereditarily symmetric name $\dot{f}$ such that $f=\dot{f}^{\,G_{\infty}}$. Let $I\subseteq\mathrm{Ord}$ be the countable stage support of $\dot{f}$ (API item (e)), and let $\alpha:=\sup(I)+1$. Then $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name and $f_{\alpha}:=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}$. Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name, its interpretation is unchanged by further forcing: $\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}$, so $f=f_{\alpha}$ in $\mathcal{M}$.
If $f_{\alpha}$ already has a section in $M_{\alpha}$, then $f$ splits in $\mathcal{M}$ and we are done. Otherwise, $f_{\alpha}$ is eligible for scheduling by the bookkeeping.
By the bookkeeping (Definition 5.59), at some successor stage we schedule the orbit package $\mathbb{Q}_{[f_{\alpha}]}$. Let $\dot{s}_{f_{\alpha}}$ be the canonical $\mathbb{Q}_{[f_{\alpha}]}$-name for the generic section in the $f_{\alpha}$-coordinate. By Lemma 5.45, $\dot{s}_{f_{\alpha}}$ is hereditarily symmetric for the package symmetric system $(\mathbb{Q}_{[f_{\alpha}]},\mathscr{G}_{[f_{\alpha}]},\mathscr{F}_{[f_{\alpha}]})$ in $M_{\beta}$. By Lemma 5.47, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{s}_{f_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence $s_{f_{\alpha}}:=\dot{s}_{f_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
Since $s_{f_{\alpha}}$ is a section of $f_{\alpha}$, it remains a section of $f$ in the final model $\mathcal{M}$. Therefore every $f:Y\twoheadrightarrow X\subseteq T$ in $\mathcal{M}$ splits, i.e. $\mathcal{M}\models\mathsf{PP}^{\mathrm{split}}\!\restriction T$. ∎
Remark 5.63 (Countable‐support names at limits of cofinality $\omega$).
For limits $\lambda$ of countable cofinality, a name may have countable support cofinal in $\lambda$. The iteration framework of [3] handles this by block‐factorization and limit‐filter completeness (see [3, Rem. 3.9, Thm. 5.5]); our argument implicitly relies on that machinery.
Proposition 5.64 ($A$ is not well-orderable).
In the final model $\mathcal{M}$, the set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$ of Cohen reals is not well-orderable. Hence $\mathcal{M}\models\neg\mathsf{AC}$.
Proof.
Suppose for contradiction that $A$ is well-orderable in $\mathcal{M}$. Let $w$ be a well-ordering of $A$ in $\mathcal{M}$.
Step 1: Locate $w$ at a bounded stage.
Since $\mathcal{M}=\bigcup_{\lambda\in\mathrm{Ord}}M_{\lambda}$ (Theorem 5.56), there exists an ordinal $\lambda$ such that $w\in M_{\lambda}$. Fix a hereditarily symmetric $\mathbb{P}_{\lambda}$-name $\dot{w}\in\operatorname{HS}_{\lambda}$ with $\dot{w}^{G_{\lambda}}=w$.
Step 2: Apply the core lemma.
Since $\dot{w}\in\operatorname{HS}_{\lambda}$, we have $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\dot{w})\in\tilde{\mathscr{F}}^{*}_{\lambda}$. By Lemma 5.50, there exist countable $E\subseteq\omega_{1}$ and countable $D\subseteq\mathrm{Coords}_{<\lambda}$ with $(E,D)$ globally admissible such that
$\Delta_{\lambda}^{\uparrow}(E,D)\leq\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\dot{w}).$
Step 3: Choose a nontrivial transposition.
Since $E$ is countable and $\omega_{1}$ is uncountable, there exist distinct $\beta,\gamma\in\omega_{1}\setminus E$. Let $\pi=(\beta\ \gamma)\in\operatorname{Sym}(\omega_{1})$ be the transposition swapping $\beta$ and $\gamma$. Then $\pi\in\operatorname{Fix}(E)$.
Step 4: Construct the diagonal lift.
By Lemma 5.43, since $(E,D)$ is globally admissible and $\pi\in\operatorname{Fix}(E)$, there exists a diagonal lift $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ with $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D})=\pi$.
By the characterization of $\Delta^{\uparrow}$ (Definition 5.27):
  • $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D})=\pi\in\operatorname{Fix}(E)$, and
  • $\widehat{\pi}^{\,D}$ fixes every coordinate in $D$ pointwise.
Hence $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)\leq\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\dot{w})$.
Step 5: Derive the contradiction.
Since $\widehat{\pi}^{\,D}\in\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\dot{w})$, we have $\widehat{\pi}^{\,D}\cdot\dot{w}=\dot{w}$.
By equivariance of name interpretation ([4, Lemma 14.37]):
$\widehat{\pi}^{\,D}\cdot w=(\widehat{\pi}^{\,D}\cdot\dot{w})^{G_{\lambda}}=\dot{w}^{G_{\lambda}}=w.$
Thus $w$ is invariant under the action of $\widehat{\pi}^{\,D}$. This means that for all $x,y\in A$:
$x<_{w}y\quad\Longleftrightarrow\quad\widehat{\pi}^{\,D}\cdot x<_{w}\widehat{\pi}^{\,D}\cdot y.$
Moreover $\widehat{\pi}^{\,D}[A]=A$ (the diagonal lift permutes Cohen reals by Corollary 5.32), so the map $x\mapsto\widehat{\pi}^{\,D}\cdot x$ is a bijection $A\to A$. Therefore the map $x\mapsto\widehat{\pi}^{\,D}\cdot x$ is an order automorphism of the well-ordered set $(A,w)$.
By Lemma 3.21, any order automorphism of a well-ordered set is the identity. Hence $\widehat{\pi}^{\,D}\cdot x=x$ for all $x\in A$.
However, by Corollary 5.32, $\widehat{\pi}^{\,D}$ swaps $c_{\beta}\leftrightarrow c_{\gamma}$:
$\widehat{\pi}^{\,D}\cdot c_{\beta}=c_{\gamma}\neq c_{\beta}.$
(The Cohen reals $c_{\beta}$ and $c_{\gamma}$ are distinct by standard properties of Cohen forcing; see Lemma 3.1.)
This contradicts $\widehat{\pi}^{\,D}\cdot c_{\beta}=c_{\beta}$. Therefore $A$ is not well-orderable in $\mathcal{M}$, and hence $\mathcal{M}\models\neg\mathsf{AC}$. ∎
Proposition 5.65 ($\mathsf{DC}$ preserved through the iteration).
$\mathcal{M}\models\mathsf{DC}$.
Proof.
We argue by induction along the symmetric iteration. The base model $M_{0}=\mathcal{N}$ satisfies $\mathsf{DC}$ by Theorem 3.10.
For the successor step, assume $M_{\alpha}\models\mathsf{ZF}+\mathsf{DC}$. The successor stage is obtained by forcing with the stage symmetric system at $\alpha$ and then taking the hereditarily symmetric submodel $M_{\alpha+1}$.
By Lemma 5.6, each orbit package $\mathbb{S}_{[f]}$ used at stage $\alpha$ is $\omega_{1}$-closed (and hence adds no new $\omega$-sequences of ordinals), and the finite-support product $\dot{\mathbb{S}}_{\alpha}$ is again $\omega_{1}$-closed for the same reason. By Lemma 5.16, the relevant stage filter is $\omega_{1}$-complete. Therefore the $\mathsf{DC}$-preservation theorem for symmetric extensions applies (API item (c), cf. [3, Thm. 3.20]), yielding $M_{\alpha+1}\models\mathsf{DC}$.
At limit stages, $\mathsf{DC}$ is preserved by the countable-support symmetric iteration machinery (API item (c), and the limit-filter generation/completeness from API item (d); cf. [3, Thm. 3.20] and [3, Thm. 5.5] for the limit construction).
Thus every stage model satisfies $\mathsf{DC}$, in particular the final model $\mathcal{M}\models\mathsf{DC}$. ∎
Definition 5.66 (Supported subforcing and supported names).
Fix a stage $\lambda$ and a globally admissible pair $(E,D)$ at stage $\lambda$. Define
$\mathbb{P}_{\lambda}\restriction(E,D)=\{p\in\mathbb{P}_{\lambda}:p\restriction 0\in\operatorname{Add}(\omega,E)\ \text{and}\ \operatorname{supp}_{>0}(p)\subseteq D\},$
where $\operatorname{supp}_{>0}(p):=\{c\in\mathrm{Coords}_{<\lambda}:p(c)\neq 1_{c}\}$ is the set of nontrivial higher-stage coordinates of $p$.
A $\mathbb{P}_{\lambda}$-name $\dot{x}$ is $(E,D)$-supported if every condition appearing in the transitive closure $\operatorname{tc}(\dot{x})$ belongs to $\mathbb{P}_{\lambda}\restriction(E,D)$.
Lemma 5.67 (Generic restriction to a supported subforcing).
Let $(E,D)$ be globally admissible at stage $\lambda$, and let $\mathbb{Q}=\mathbb{P}_{\lambda}\restriction(E,D)$. Then every maximal antichain $A\subseteq\mathbb{Q}$ is also a maximal antichain in $\mathbb{P}_{\lambda}$. Consequently, if $G_{\lambda}\subseteq\mathbb{P}_{\lambda}$ is $V$-generic, then $G_{E,D}:=G_{\lambda}\cap\mathbb{Q}$ is $V$-generic for $\mathbb{Q}$, and $V[G_{E,D}]\subseteq V[G_{\lambda}]$.
Proof.
Let $A\subseteq\mathbb{Q}$ be maximal, and let $p\in\mathbb{P}_{\lambda}$. Put $p^{\prime}:=p\restriction(E,D)\in\mathbb{Q}$. Choose $q\in A$ compatible with $p^{\prime}$ in $\mathbb{Q}$, and let $r\in\mathbb{Q}$ witness this, i.e. $r\leq_{\mathbb{Q}}q,p^{\prime}$. Since $p$ extends $p^{\prime}$ on the coordinates that $r$ constrains, and $r$ is trivial off $E\cup D$, we can amalgamate $r$ with $p$ to obtain a common extension in $\mathbb{P}_{\lambda}$. Thus $q$ is compatible with $p$ in $\mathbb{P}_{\lambda}$, so $A$ is predense in $\mathbb{P}_{\lambda}$ and hence maximal there. Now $G_{\lambda}$ meets every maximal antichain in $\mathbb{P}_{\lambda}$, so $G_{\lambda}\cap A\neq\emptyset$. Since $A\subseteq\mathbb{Q}$, this gives $(G_{\lambda}\cap\mathbb{Q})\cap A\neq\emptyset$, proving $\mathbb{Q}$-genericity. ∎
Lemma 5.68 (Supported names are $\mathbb{Q}$-names and evaluate correctly).
Let $(E,D)$ be globally admissible at stage $\lambda$, let $\mathbb{Q}=\mathbb{P}_{\lambda}\restriction(E,D)$, and let $G_{\lambda}\subseteq\mathbb{P}_{\lambda}$ be $V$-generic with $G_{E,D}=G_{\lambda}\cap\mathbb{Q}$. If $\dot{x}$ is $(E,D)$-supported, then $\dot{x}$ is (canonically) a $\mathbb{Q}$-name and
$\dot{x}^{G_{\lambda}}=\dot{x}^{G_{E,D}}.$
Proof.
Since every condition appearing in $\operatorname{tc}(\dot{x})$ lies in $\mathbb{Q}$, we may view $\dot{x}$ as a $\mathbb{Q}$-name (by the usual identification of names with the same transitive closure). The valuation equality is proved by induction on name-rank: for each pair $(\dot{y},p)\in\dot{x}$ we have $p\in\mathbb{Q}$, hence $p\in G_{\lambda}$ iff $p\in G_{E,D}$, and the induction hypothesis applies to $\dot{y}$. ∎
Lemma 5.69 (Triviality on supported names).
If $\dot{x}$ is $(E,D)$-supported and $g\in\Delta_{\lambda}^{\uparrow}(E,D)$, then $g\cdot\dot{x}=\dot{x}$.
Proof.
By induction on $\operatorname{rank}(\dot{x})$. If $p\in\mathbb{P}_{\lambda}\restriction(E,D)$, then $g(p)=p$ since the stage–0 projection of $g$ fixes $E$ pointwise and $g$ fixes every coordinate in $D$ pointwise. Now apply the recursive definition of the action on names. ∎
Lemma 5.70 (Persistence of $\mathsf{SVC}(S)$).
Let $S$ and $T$ be as in Definition 3.23. Since $M_{0}=\mathcal{N}$ and $\mathcal{N}\models\mathsf{SVC}(S)$ by Proposition 3.22, we have $M_{0}\models\mathsf{SVC}(S)$. Then for every stage $\lambda$, $M_{\lambda}\models\mathsf{SVC}(S)$, and consequently $M_{\lambda}\models\mathsf{SVC}^{+}(T)$.
Proof.
We argue by transfinite induction on $\lambda$.
Base. $M_{0}=\mathcal{N}\models\mathsf{SVC}(S)$ by Proposition 3.22.
Successor. Assume $M_{\alpha}\models\mathsf{SVC}(S)$ and let $X\in M_{\alpha+1}$ be nonempty. Choose a hereditarily symmetric $\mathbb{P}_{\alpha+1}$-name $\dot{X}\in\operatorname{HS}_{\alpha+1}$ (so $\dot{X}\in M_{\alpha}$) with $X=\dot{X}^{G_{\alpha+1}}$. By $\mathsf{SVC}(S)$ in $M_{\alpha}$, there exist an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow\dot{X}$ in $M_{\alpha}$.
Fix $p_{0}\in\mathbb{P}_{\alpha+1}$ and a name $\dot{x}_{0}$ such that $p_{0}\Vdash\dot{x}_{0}\in\dot{X}$. Define a $\mathbb{P}_{\alpha+1}$-name $\dot{h}$ for a function $S\times\eta\to\dot{X}$ by: for $s\in S$ and $\xi<\eta$, write $e(s,\xi)=\langle\dot{y},p\rangle$ and set
$\dot{h}(s,\xi)=\begin{cases}\dot{y}&\text{if }p\in\dot{G}_{\alpha+1},\\ \dot{x}_{0}&\text{otherwise.}\end{cases}$
Let $h:=\dot{h}^{G_{\alpha+1}}$. Then $h:S\times\eta\to X$, and $h$ is surjective by construction.
Moreover, $\dot{h}\in\operatorname{HS}_{\alpha+1}$: it is definable from parameters in $M_{\alpha}$, and any automorphism fixing $\dot{X}$ (hence lying in $\tilde{\mathscr{F}}^{*}_{\alpha+1}$) fixes $\dot{h}$ as well. Thus $h\in M_{\alpha+1}$, so $M_{\alpha+1}\models\mathsf{SVC}(S)$.
Limit. Let $\lambda$ be limit and assume $M_{\beta}\models\mathsf{SVC}(S)$ for all $\beta<\lambda$. Let $X\in M_{\lambda}$ be nonempty. By API item (e), $X$ has a hereditarily symmetric name with countable stage support bounded in $\lambda$, so $X\in M_{\beta}$ for some $\beta<\lambda$. By the inductive hypothesis, $M_{\beta}\models\mathsf{SVC}(S)$, hence $M_{\lambda}\models\mathsf{SVC}(S)$ as well.
Finally, $M_{\lambda}\models\mathsf{SVC}^{+}(T)$ follows from Fact 4.3. ∎
Proposition 5.71 (Final $\mathsf{AC}_{\mathsf{WO}}$).
$\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$.
Proof.
Work in the final symmetric model $\mathcal{M}$. By Lemma 5.70, we have $\mathcal{M}\models\mathsf{SVC}(S)$. We show that $\mathcal{M}\models\mathsf{AC}_{<\aleph^{*}(S)^{M}}$; by the Ryan–Smith equivalences this implies $\mathsf{AC}_{\mathsf{WO}}$.
Let $\lambda<\aleph^{*}(S)^{M}$ be an ordinal and let $f:Y\twoheadrightarrow\lambda$ be a surjection in $\mathcal{M}$. By $\mathsf{SVC}(S)$, fix an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow Y$ in $\mathcal{M}$. Let $g:=f\circ e:S\times\eta\twoheadrightarrow\lambda$.
Let $\dot{g}$ be a hereditarily symmetric name such that $g=\dot{g}^{\,G_{\infty}}$. Let $I\subseteq\mathrm{Ord}$ be the countable stage support of $\dot{g}$ (API item (e)), and let $\alpha:=\sup(I)+1$. Then $\dot{g}$ is a $\mathbb{P}_{\alpha}$-name, and $g_{\alpha}:=\dot{g}^{\,G_{\alpha}}\in M_{\alpha}$. By the bookkeeping (Definition 5.59), there is some successor stage $\beta>\alpha$ at which the orbit package $\mathbb{R}_{[g_{\alpha}]}$ is scheduled. Let $\dot{r}_{g_{\alpha}}$ be the canonical $\mathbb{R}_{[g_{\alpha}]}$-name for the generic right inverse in the $g_{\alpha}$-coordinate. By Lemma 5.46, the canonical $\mathbb{R}_{[g_{\alpha}]}$-name $\dot{r}_{g_{\alpha}}$ is hereditarily symmetric for the package symmetric system in $M_{\beta}$. By Lemma 5.47, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{r}_{g_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence $r_{g_{\alpha}}:=\dot{r}_{g_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$, and $r_{g_{\alpha}}$ is a right inverse of $g$ in $\mathcal{M}$.
Now define $t:=e\circ r_{g_{\alpha}}:\lambda\to Y$. Then for all $\xi<\lambda$,
$f\bigl(t(\xi)\bigr)=f\bigl(e(r_{g_{\alpha}}(\xi))\bigr)=g\bigl(r_{g_{\alpha}}(\xi)\bigr)=\xi,$
so $t$ is a right inverse to $f$. Thus every surjection $Y\twoheadrightarrow\lambda$ with $\lambda<\aleph^{*}(S)^{M}$ splits in $\mathcal{M}$, and therefore $\mathcal{M}\models\mathsf{AC}_{<\aleph^{*}(S)^{M}}$.
Finally, since $\mathcal{M}\models\mathsf{SVC}(S)$, the Ryan–Smith theorem (see Remark 4.11) yields $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$. ∎
Theorem 5.72 (Final model).
There is a transitive model $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$.
Proof.
By Theorem 5.56, $\mathcal{M}\models\mathsf{ZF}$, and by Proposition 5.65, $\mathcal{M}\models\mathsf{DC}$.
By Proposition 5.62, $\mathcal{M}\models\mathsf{PP}^{\mathrm{split}}\!\restriction T$, hence $\mathcal{M}\models\mathsf{PP}\!\restriction T$. By Proposition 5.71, $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$. By Lemma 5.70, $\mathcal{M}\models\mathsf{SVC}^{+}(T)$ for $T=\mathcal{P}(S)$, so the hypotheses of the Ryan–Smith localization theorem (Theorem 4.5, i.e., [13, Prop. 3.17] applied with Ryan–Smith’s parameter $S$ taken to be our $T$) hold in $\mathcal{M}$. Therefore $\mathcal{M}\models\mathsf{PP}$.
Finally, Proposition 5.64 gives $\mathcal{M}\models\neg\mathsf{AC}$. ∎

### v3

@@
 The global Partition Principle follows from Ryan–Smith’s localization theorem [17], which shows that under $\mathsf{SVC}^{+}(T)$, the principle $\mathsf{PP}$ is equivalent to the conjunction of $\mathsf{PP}\!\restriction\!T$ (the restriction to surjections between subsets of $T$) and $\mathsf{AC}_{\mathsf{WO}}$.
+Remark 1.1 (The axiom $\mathsf{AC}_{\mathsf{WO}}$).
+Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes the axiom commonly called choice for well–orderable sets, in the sense of Howard–Rubin [7, Form 40]. That is:
+Every well–orderable set of nonempty sets admits a choice function.
+Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds: if $\neq\varnothing, Y\varnothing\notin Y$, and $|Y|\leq|\lambda|$, then there exists a function $f$ with domain $Y$ such that $f(y)\in y$ for all $y\in Y$.
+This axiom should be distinguished from the trivial principle asserting choice for families already equipped with specified well–orderings, which is provable in $\mathsf{ZF}$.
+The present form asserts the existence of a uniform choice function for arbitrary well–orderable families, without any fixed well–orderings given in advance.
+As shown by Pincus (see [7]), $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the statement that $\aleph(X)=\aleph^{*}(X)$ for all sets $X$.
+This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used implicitly in Ryan–Smith’s localization theorem [17].
 Technical innovations.
@@
   1. (a)
-    Set-length and class-length versions of symmetric iterations (cf. [3, §2.1]).
+    Set-length and class-length versions of symmetric iterations (cf. [5, §2.1]).
   2. (b)
-    Definitions and basic properties of countable-support symmetric iterations (cf. [3, Def. 3.1]).
+    Definitions and basic properties of countable-support symmetric iterations (cf. [5, Def. 3.1]).
   3. (c)
-    A $\mathsf{DC}$-preservation theorem for the relevant support (cf. [3, Thm. 3.20]).
+    A $\mathsf{DC}$-preservation theorem for the relevant support (cf. [5, Thm. 3.20]).
   4. (d)
-    Generation and completeness/normality of limit filters in the relevant cofinality cases (cf. [3, Thm. 5.5]).
+    Generation and completeness/normality of limit filters in the relevant cofinality cases (cf. [5, Thm. 5.5]).
   5. (e)
-    A characterization of hereditarily symmetric names by “small support” (cf. [3, Lem. 3.14]).
+    A characterization of HS names by “small support” assuming the stage model satisfies $\mathsf{DC}$ (cf. [5, Lem. 3.14]).
+In our application, every stage model satisfies $\mathsf{DC}$ (Proposition 5.65), so the support characterization may be used freely throughout the iteration.
 We emphasize that at each stage $\lambda$ the iteration produces a single ambient group $\mathscr{G}_{\lambda}$ acting by automorphisms on the full initial-segment forcing $\mathbb{P}_{\lambda}$ (API item (b)), so the action continues to move the original Cohen coordinates; we are not iterating by repeatedly forcing over intermediate symmetric models in isolation.
@@
     We say that $\mathsf{SVC}^{+}(S)$ holds if for every set $X$, there exists an ordinal $\eta$ such that there is an injection $f\colon X\hookrightarrow S\times\eta$ (equivalently, $|X|\leq|S\times\eta|$).
-The following facts from [13] drive our localization strategy:
-Fact 2.3.
-For any set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$.
-Theorem 2.4 (Ryan–Smith Localization, [13]).
-Assume $\mathsf{SVC}^{+}(T)$. Then $\mathsf{PP}$ is equivalent to the conjunction of $\mathsf{PP}\!\restriction\!T$ and $\mathsf{AC}_{\mathsf{WO}}$.
-Our strategy is to construct a model satisfying $\mathsf{SVC}(S)$ (via a Cohen seed) and then force $\mathsf{PP}\!\restriction\!\mathcal{P}(S)$ and $\mathsf{AC}_{\mathsf{WO}}$ while preserving $\mathsf{SVC}(S)$.
+Ryan–Smith localization ingredients (used later). We will repeatedly appeal to two results from [17], which we record in Section 4 in the exact forms used in this paper:
+  • Injective upgrade: for every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$ (Fact 4.3).
+  • Localization of $\mathsf{PP}$: assuming $\mathsf{SVC}^{+}(T)$, one has
+    $\mathsf{PP}\ \Longleftrightarrow\ \bigl(\mathsf{PP}\!\restriction T\wedge\mathsf{AC}_{\mathsf{WO}}\bigr)$
+    (Theorem 4.5, i.e. [17, Proposition 3.17]).
+Accordingly, our strategy is to start from a Cohen symmetric seed $S$ with $\mathsf{SVC}(S)$, take $T:=\mathcal{P}(S)$, force $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ while preserving $\mathsf{SVC}(S)$, and then invoke Theorem 4.5 to conclude global $\mathsf{PP}$.
 3 The Cohen symmetric seed model
@@
 But $\pi_{i}\operatorname{Fix}(E_{i})\pi_{i}^{-1}=\operatorname{Fix}(\pi_{i}[E_{i}])$, and finite intersections of $\operatorname{Fix}(\cdot)$’s are $\operatorname{Fix}(\bigcup\cdot)$.
-Let $E:=\bigcup_{i<k}\pi_{i}[E_{i}]$ (countable).
+Let $E:=\bigcup_{i<k}\pi_{i}[E_{i}]$ (finite union of countable sets is countable).
 Then $\operatorname{Fix}(E)\leq H$.
@@
 Proof.
-Let $\langle H_{n}:n<\omega\rangle\subseteq\mathscr{F}$.
-By Lemma 3.4 pick countable $E_{n}$ with $\operatorname{Fix}(E_{n})\leq H_{n}$.
-Put $E=\bigcup_{n<\omega}E_{n}$ (countable), so $\operatorname{Fix}(E)=\bigcap_{n<\omega}\operatorname{Fix}(E_{n})\leq\bigcap_{n<\omega}H_{n}$, hence $\bigcap_{n<\omega}H_{n}\in\mathscr{F}$.
+By Lemma 3.4 (in $V$) choose $E_{n}\in[\omega_{1}]^{\leq\omega}$ with $\operatorname{Fix}(E_{n})\leq H_{n}$.
+Let $E:=\bigcup_{n<\omega}E_{n}$, which is countable in $V$ since $V\models\mathsf{ZFC}$.
+Then $\operatorname{Fix}(E)=\bigcap_{n<\omega}\operatorname{Fix}(E_{n})\leq\bigcap_{n<\omega}H_{n}$, hence $\bigcap_{n<\omega}H_{n}\in\mathscr{F}$.
 ∎
@@
 ∎
+Remark 3.10 (Metatheoretic countability vs. internal countability).
+Throughout the construction of HS names, “countable” means countable in the ground model $V$ (e.g.
+$E\in[\omega_{1}]^{\leq\omega}$ is evaluated in $V$).
+This notion is used only to ensure that $\operatorname{Fix}(E)\in\mathscr{F}$ (Definition 3.3), hence that the relevant names are symmetric.
+In particular, when we form supports such as $E=\bigcup_{n<\omega}E_{n}$ in the metatheory, we use that $V\models\mathsf{ZFC}$, so a countable union of countable sets (in $V$) is countable (in $V$).
+We do not claim that these supports are countable in the symmetric model $\mathcal{N}$.
 Theorem 3.11 ($\mathsf{DC}$ in the symmetric model).
@@
 Let $D_{n+1}:=\bigcup_{p\in D_{n}}D_{n+1,p}$.
-Then $D_{n+1}$ is countable (a countable union of countable sets) and is a maximal antichain below $p_{\ast}$: given any $r\leq p_{\ast}$, pick $p\in D_{n}$ compatible with $r$ (maximality of $D_{n}$), let $r^{\prime}\leq r,p$, and then pick $q\in D_{n+1,p}$ compatible with $r^{\prime}$ (maximality of $D_{n+1,p}$).
+Since $D_{n}$ is an antichain, if $q\in D_{n+1,p}$ and $q^{\prime}\in D_{n+1,p^{\prime}}$ with $p\neq p^{\prime}$, then $q\perp q^{\prime}$ (otherwise $p$ and $p^{\prime}$ would be compatible).
+Hence $D_{n+1}$ is an antichain.
+Moreover, working in the metatheory $V\models\mathsf{ZFC}$ (Remark 3.10), $D_{n+1}$ is countable because it is a countable union (indexed by the countable set $D_{n}$) of countable sets $D_{n+1,p}$.
+Finally, $D_{n+1}$ is a maximal antichain below $p_{\ast}$: given any $r\leq p_{\ast}$, pick $p\in D_{n}$ compatible with $r$ (maximality of $D_{n}$), let $r^{\prime}\leq r,p$, and then pick $q\in D_{n+1,p}$ compatible with $r^{\prime}$ (maximality of $D_{n+1,p}$).
 Now define, for each $n<\omega$, the mixed name
@@
 $E_{n}\;:=\;E(\dot{A})\ \cup\ E(\dot{R})\ \cup\ \bigcup_{p\in D_{n}}\Bigl(\operatorname{supp}(p)\ \cup\ E(\dot{a}_{n,p})\Bigr),$
-a countable union of countable sets of ordinals, hence countable.
+a countable union of $V$-countable subsets of $\omega_{1}$, hence $V$-countable (Remark 3.10).
+Therefore $\operatorname{Fix}(E_{n})\in\mathscr{F}$ and $\operatorname{Fix}(E_{n})\leq\operatorname{Stab}(\dot{a}_{n})$.
 Then every $\pi\in\operatorname{Fix}(E_{n})$ fixes each $p\in D_{n}$, fixes $\dot{A}$ and $\dot{R}$, and fixes each $\dot{a}_{n,p}$.
@@
 Thus $\dot{a}_{n}\in\operatorname{HS}$ for all $n$.
-Let $E:=\bigcup_{n<\omega}E_{n}$, again countable, and define the sequence name
+Let $E:=\bigcup_{n<\omega}E_{n}$. Then $E$ is $V$-countable (Remark 3.10), hence $\operatorname{Fix}(E)\in\mathscr{F}$, and define the sequence name
 $\dot{s}\;:=\;\bigl\{\langle\langle\check{n},\dot{a}_{n}\rangle,\mathbbm{1}_{\mathbb{P}}\rangle:n<\omega\bigr\}.$
@@
 3.3 Countable supports for $\operatorname{HS}$-names
-Notation 3.11 (Restrictions to $E$).
+Notation 3.12 (Restrictions to $E$).
 For $E\subseteq\omega_{1}$ and $p\in\operatorname{Add}(\omega,\omega_{1})$, let
@@
 ∎
-Corollary 3.16 (Supported names live in the corresponding subextension).
-Let $E\subseteq\omega_{1}$ and $\tau$ a $\mathbb{P}$-name with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$.
-Then $\mathbbm{1}_{\mathbb{P}}\Vdash\tau=\tau\!\restriction E$.
-In particular, for any $V$-generic $G\subseteq\mathbb{P}$ with $G_{E}:=G\cap\mathbb{P}_{E}$,
-$\tau^{G}=(\tau\!\restriction E)^{G_{E}}\in V[G_{E}].$
+Remark 3.17 (Symmetry support vs. forcing support).
+The condition $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$ ($\tau$ has symmetry support EE”) means that $\tau$ is invariant under permutations fixing $E$ pointwise.
+This does not imply that $\tau$ is a $\mathbb{P}_{E}$–name, nor that $\tau^{G}$ lies in $V[G_{E}]$.
+Indeed, letting $\dot{A}=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha<\omega_{1}\}$ be the canonical name for the Cohen set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$, we have $\operatorname{Stab}(\dot{A})=\operatorname{Sym}(\omega_{1})$, so $\dot{A}$ has symmetry support $\emptyset$.
+Nevertheless, $\dot{A}$ is not a $\mathbb{P}_{E}$–name for any countable $E$, and $A\notin V[G_{E}]$ for every countable $E\subseteq\omega_{1}$.
+Lemma 3.18 (Subextension evaluation for $\mathbb{P}_{E}$-names).
+Let $E\subseteq\omega_{1}$, let $\mathbb{P}_{E}=\operatorname{Add}(\omega,E)\leq\mathbb{P}=\operatorname{Add}(\omega,\omega_{1})$, and let $\tau$ be a $\mathbb{P}_{E}$–name (viewed as a $\mathbb{P}$–name via the inclusion).
+If $G\subseteq\mathbb{P}$ is VV–generic and $G_{E}:=G\cap\mathbb{P}_{E}$, then
+$\tau^{G}=\tau^{G_{E}}\in V[G_{E}].$
 Proof.
-We prove by induction on rank that $\mathbbm{1}_{\mathbb{P}}\Vdash\tau=\tau\!\restriction E$.
-Fix a formula of the form $\sigma\in\tau$” or $\sigma=\tau$” and use Lemma 3.15 to show that whenever $p$ decides such a statement about $\tau$, its restriction $p\!\restriction E$ decides it the same way.
-This yields that membership in $\tau^{G}$ is determined by $G_{E}$ and the recursive definition of $\tau\!\restriction E$, and the induction closes.
+By induction on the rank of $\tau$.
+Using the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$, the valuation recursion for $\tau^{G}$ only consults conditions from $G$ that appear in $\tau$, and these conditions all lie in $\mathbb{P}_{E}$, hence belong to $G_{E}$.
+Applying the inductive hypothesis to subnames of $\tau$ yields $\tau^{G}=\tau^{G_{E}}$.
 ∎
@@
 Unlike the finite-support generalized Cohen model (which produces a Dedekind-finite $A$; cf.
-[11, §2.4]), the present countable-support filter allows many countable subsets of $A$: for each countable $E\subseteq\omega_{1}$ in $V$, the name $\dot{A}_{E}:=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha\in E\}$ is hereditarily symmetric, so $A_{E}:=\{c_{\alpha}:\alpha\in E\}\in\mathcal{N}$.
+[15, §2.4]), the present countable-support filter allows many countable subsets of $A$: for each countable $E\subseteq\omega_{1}$ in $V$, the name $\dot{A}_{E}:=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha\in E\}$ is hereditarily symmetric, so $A_{E}:=\{c_{\alpha}:\alpha\in E\}\in\mathcal{N}$.
+Lemma 3.23 (Coding a countable set of Cohen reals by an element of $A^{\omega}$).
+Let $E\in[\omega_{1}]^{\leq\omega}$ be countable in $V$. Fix in $V$ a surjection $e:\omega\twoheadrightarrow E$. Then the sequence
+$s_{E}:=\langle c_{e(n)}:n\in\omega\rangle$
+is an element of $A^{\omega}$ in $\mathcal{N}$.
+Proof.
+Each $\dot{c}_{e(n)}$ is hereditarily symmetric with support $\{e(n)\}$, hence the sequence name $\langle\dot{c}_{e(n)}:n\in\omega\rangle$ is hereditarily symmetric with support $E$.
+Therefore $s_{E}\in\mathcal{N}$, and clearly $s_{E}\in A^{\omega}$.
+∎
+Lemma 3.24 (A forced nonempty name yields a ground-coded witness).
+Let $\dot{X}$ be a $\mathbb{P}$–name and suppose $p\Vdash``\dot{X}\neq\varnothing$". Then there exist $q\leq p$ and a $\mathbb{P}$–name $\dot{x}$ such that $q\Vdash``\dot{x}\in\dot{X}$".
+Proof.
+This is the standard forcing-theorem witness extraction for $\exists$. From $p\Vdash\exists x\,(x\in\dot{X})$ we obtain $q\leq p$ and a name $\dot{x}$ with $q\Vdash\dot{x}\in\dot{X}$. ∎
 Lemma 3.25 (Well-orders are rigid).
@@
 Proof.
-Fix a nonempty $X\in\mathcal{N}$ and choose $\dot{X}\in\operatorname{HS}$ with $X=\dot{X}^{G}$.
-By Lemma 3.13, fix a countable support $E\subseteq\omega_{1}$ for $\dot{X}$, i.e.
-$\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{X})$.
-Let $\dot{X}\!\restriction E$ be the restricted $\mathbb{P}_{E}$-name from Notation above.
-By Corollary 3.16,
-$X=\dot{X}^{G}=(\dot{X}\!\restriction E)^{G_{E}}\in V[G_{E}],$
-where $G_{E}:=G\cap\mathbb{P}_{E}$.
-Since $V[G_{E}]\models\mathsf{ZFC}$, there is an ordinal $\eta$ and a surjection $h:\eta\twoheadrightarrow X$ in $V[G_{E}]$.
-Let $\dot{h}$ be a $\mathbb{P}_{E}$-name for $h$.
-Every permutation $\pi\in\operatorname{Sym}(\omega_{1})$ which fixes $E$ pointwise acts trivially on $\mathbb{P}_{E}$ and hence fixes $\dot{h}$, so $\dot{h}\in\operatorname{HS}$ (indeed, it has support $E$).
-Therefore $h=\dot{h}^{G}\in\mathcal{N}$.
-Finally, since $S=A^{\omega}$ is nonempty in $\mathcal{N}$, the map $f:S\times\eta\to X$ defined by $f(s,\xi)=h(\xi)$ is a surjection in $\mathcal{N}$.
-This proves $\mathcal{N}\models\mathsf{SVC}(S)$.
+We use the standard “continuity” analysis of Cohen-type symmetric models.
+In the basic Cohen model $V(A)$ generated by an unordered set $A$ of Cohen reals, a key lemma (sometimes called a continuity or indiscernibility lemma) shows that statements over $V(A)$ with parameters from $V$ are already decided in a subextension generated by finitely many members of $A$; see, e.g., Felgner [4, p.
+133] or Blass [3, Prop.
+1.2] (and the discussion in [6, Lem.
+2.8]).
+More precisely, for each $x\in X$ there is a $\mathbb{P}$–name $\tau_{x}\in V$ and a sequence $s_{x}\in A^{\omega}$ (coding the countable symmetry support relevant to $\tau_{x}$) such that $x$ is recovered uniformly from the pair $(\tau_{x},s_{x})$.
+Enumerating the set of $\mathbb{P}$–names in $V$ by ordinals yields the required surjection from $A^{\omega}\times\eta$ onto $X$ for some ordinal $\eta$.
+The same argument applies to our Cohen symmetric model $\mathcal{N}$ built from $\operatorname{Add}(\omega,\omega_{1})$ with the countable-support symmetry filter: every set $X\in\mathcal{N}$ is obtained (uniformly in $X$) from $V$ and a countable sequence of members of $A$, and therefore is a surjective image of $A^{\omega}\times\eta$ for some ordinal $\eta$.
+Equivalently, $\mathcal{N}\models\mathsf{SVC}(A^{\omega})$.
 ∎
@@
 Fact 4.3 (Ryan–Smith).
-For every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$. [13, §2.1, Fact]
+For every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$. [17, §2.1, Fact]
 Remark 4.4.
@@
 $\mathsf{PP}\quad\Longleftrightarrow\quad\bigl(\mathsf{PP}\!\restriction T\ \wedge\ \mathsf{AC}_{\mathsf{WO}}\bigr).$
-[13, Proposition 3.17] (stated there with parameter “SS”; we apply it with $S:=T$)
+[17, Proposition 3.17] (stated there with parameter “SS”; we apply it with $S:=T$). As this argument is critical to our goals, we demonstrate a proof here.
+Proof.
+($\Rightarrow$) If $\mathsf{PP}$ holds, then $\mathsf{PP}\!\restriction T$ is immediate.
+Moreover, $\mathsf{PP}$ implies $\mathsf{AC}_{\mathsf{WO}}$ (e.g.
+via the consequence $(\forall X)\,\aleph(X)=\aleph^{\ast}(X)$; see [17, Proposition 3.17 and footnote 4] for references).
+($\Leftarrow$) Assume $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$, and let $\pi:Y\twoheadrightarrow X$ be a surjection. We must produce an injection $X\hookrightarrow Y$.
+Apply $\mathsf{SVC}^{+}(T)$ to the set
+$Z\;:=\;(X\times\{0\})\ \cup\ (Y\times\{1\}),$
+obtaining an ordinal $\eta$ and an injection $j:Z\hookrightarrow T\times\eta$. Define injections
+$\pi_{X}:X\to T\times\eta,\quad\pi_{X}(x)=j(\langle x,0\rangle),\qquad\pi_{Y}:Y\to T\times\eta,\quad\pi_{Y}(y)=j(\langle y,1\rangle),$
+and let
+$A:=\operatorname{ran}(\pi_{X})\subseteq T\times\eta,\qquad B:=\operatorname{ran}(\pi_{Y})\subseteq T\times\eta.$
+Then $\pi_{X}:X\to A$ and $\pi_{Y}:Y\to B$ are bijections, hence have inverses $\pi_{X}^{-1}:A\to X$ and $\pi_{Y}^{-1}:B\to Y$. Define
+$f\;:=\;\pi_{X}\circ\pi\circ\pi_{Y}^{-1}\;:\;B\twoheadrightarrow A.$
+It therefore suffices to show that whenever $A,B\subseteq T\times\eta$ and $f:B\twoheadrightarrow A$, there is an injection $g:A\hookrightarrow B$.
+Claim 2.
+(Slicing lemma for $T\times\eta$) Let $\eta$ be an ordinal.
+Assume $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$.
+If $A,B\subseteq T\times\eta$ and $f:B\twoheadrightarrow A$, then there is an injection $g:A\hookrightarrow B$.
+Proof of Claim 4.
+Treat $f$ as a partial map on $T\times\eta$ with domain $B$. For each $\langle t,\alpha\rangle\in A$, define
+$\varepsilon_{t,\alpha}\;:=\;\min\bigl\{\varepsilon<\eta:(\exists s\in T)\ \bigl(\langle s,\varepsilon\rangle\in B\ \wedge\ f(\langle s,\varepsilon\rangle)=\langle t,\alpha\rangle\bigr)\bigr\}.$
+(Nonempty since $f$ is surjective and $\min$ exists since $\eta$ is an ordinal.)
+For $\varepsilon,\alpha<\eta$, define the slices
+$A^{\varepsilon}_{\alpha}\;:=\;\{\,t\in T:\langle t,\alpha\rangle\in A\ \wedge\ \varepsilon_{t,\alpha}=\varepsilon\,\}\ \subseteq\ T,$
+and
+$B^{\varepsilon}_{\alpha}\;:=\;\{\,s\in T:\langle s,\varepsilon\rangle\in B\ \wedge\ f(\langle s,\varepsilon\rangle)=\langle t,\alpha\rangle\text{ for some }t\in A^{\varepsilon}_{\alpha}\,\}\ \subseteq\ T.$
+Define $h^{\varepsilon}_{\alpha}:B^{\varepsilon}_{\alpha}\to A^{\varepsilon}_{\alpha}$ by
+$h^{\varepsilon}_{\alpha}(s)=t\quad\text{iff}\quad f(\langle s,\varepsilon\rangle)=\langle t,\alpha\rangle.$
+This is well-defined (since $f$ is a function), and it is a surjection: if $t\in A^{\varepsilon}_{\alpha}$, then by definition of $\varepsilon_{t,\alpha}=\varepsilon$ there exists $s$ with $\langle s,\varepsilon\rangle\in B$ and $f(\langle s,\varepsilon\rangle)=\langle t,\alpha\rangle$, hence $s\in B^{\varepsilon}_{\alpha}$ and $h^{\varepsilon}_{\alpha}(s)=t$.
+By $\mathsf{PP}\!\restriction T$ applied to the surjection $h^{\varepsilon}_{\alpha}:B^{\varepsilon}_{\alpha}\twoheadrightarrow A^{\varepsilon}_{\alpha}$, there exists an injection $i^{\varepsilon}_{\alpha}:A^{\varepsilon}_{\alpha}\hookrightarrow B^{\varepsilon}_{\alpha}$ whenever $A^{\varepsilon}_{\alpha}\neq\varnothing$.
+Let
+$E\;:=\;\{\,\langle\varepsilon,\alpha\rangle\in\eta\times\eta:A^{\varepsilon}_{\alpha}\neq\varnothing\,\}.$
+Then $E$ is well-orderable (e.g. by the lexicographic order on $\eta\times\eta$), so by $\mathsf{AC}_{\mathsf{WO}}$ we may choose a function $c$ with
+$c(\varepsilon,\alpha)\in\{\,i:i\text{ is an injection }A^{\varepsilon}_{\alpha}\hookrightarrow B^{\varepsilon}_{\alpha}\,\}\quad\text{for each }\langle\varepsilon,\alpha\rangle\in E.$
+Define $g:A\to T\times\eta$ by
+$g(\langle t,\alpha\rangle)\;:=\;\bigl\langle\,c(\varepsilon_{t,\alpha},\alpha)(t),\ \varepsilon_{t,\alpha}\,\bigr\rangle.$
+Then $g(\langle t,\alpha\rangle)\in B$: indeed, $c(\varepsilon_{t,\alpha},\alpha)(t)\in B^{\varepsilon_{t,\alpha}}_{\alpha}$, so by definition of $B^{\varepsilon_{t,\alpha}}_{\alpha}$ we have $\bigl\langle c(\varepsilon_{t,\alpha},\alpha)(t),\varepsilon_{t,\alpha}\bigr\rangle\in B$.
+Finally, $g$ is injective.
+If $g(\langle t,\alpha\rangle)=g(\langle t^{\prime},\alpha^{\prime}\rangle)$, then $\varepsilon_{t,\alpha}=\varepsilon_{t^{\prime},\alpha^{\prime}}=:\varepsilon$ from equality of second coordinates.
+Let $s:=c(\varepsilon,\alpha)(t)=c(\varepsilon,\alpha^{\prime})(t^{\prime})$ be the common first coordinate.
+Then $s\in B^{\varepsilon}_{\alpha}\cap B^{\varepsilon}_{\alpha^{\prime}}$.
+But if $\alpha\neq\alpha^{\prime}$ this is impossible: $s\in B^{\varepsilon}_{\alpha}$ implies $f(\langle s,\varepsilon\rangle)$ has second coordinate $\alpha$, while $s\in B^{\varepsilon}_{\alpha^{\prime}}$ implies the second coordinate is $\alpha^{\prime}$, contradicting that $f$ is a function.
+Hence $\alpha=\alpha^{\prime}$, and then $c(\varepsilon,\alpha)$ is injective, so $t=t^{\prime}$.
+Thus $\langle t,\alpha\rangle=\langle t^{\prime},\alpha^{\prime}\rangle$.
+∎
+By Claim 4 there is an injection $g:A\hookrightarrow B$. Then
+$i\;:=\;\pi_{Y}^{-1}\circ g\circ\pi_{X}\;:\;X\to Y$
+is an injection, as required. Therefore $\mathsf{PP}$ holds. ∎
 Remark 4.6.
@@
 Reference.
-This is witnessed by Cohen’s first model; see [13, Proposition 3.22]. ∎
+This is witnessed by Cohen’s first model; see [17, Proposition 3.22]. ∎
 Proposition 4.8 (The localized $\mathsf{PP}\!\restriction S$ component cannot be omitted).
@@
 Reference.
-This is witnessed by the Feferman-style model $N_{\aleph_{1}}$ of Truss; see [13, Proposition 3.23]. ∎
+This is witnessed by the Feferman-style model $N_{\aleph_{1}}$ of Truss; see [17, Proposition 3.23]. ∎
 Remark 4.9 (Why we keep both conjuncts).
@@
 We will force both conjuncts $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ in Theorem 4.5.
-In Ryan–Smith’s analysis the corresponding reduction is sharp: in general one cannot drop the displayed conjuncts (nor substantially weaken the hypotheses) without losing the equivalence; see [13, Props.
+In Ryan–Smith’s analysis the corresponding reduction is sharp: in general one cannot drop the displayed conjuncts (nor substantially weaken the hypotheses) without losing the equivalence; see [17, Props.
 3.22 and 3.23].
@@
 $\mathsf{AC}_{\mathsf{WO}}\quad\Longleftrightarrow\quad\mathsf{AC}_{<\aleph^{*}(S)}(S)\quad\Longleftrightarrow\quad\text{every function }g:S\to\lambda\text{ splits (for each }\lambda\text{)}$
-(see [13, Prop. 3.2]). This is the equivalence we invoke in Proposition 5.71.
-(We also note that Ryan–Smith’s localization for $\mathsf{PP}$ uses both conjuncts: $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$; see [13, Prop.
-3.17] and Question 3.18 therein, and see [13, Props.
+(see [17, Prop. 3.2]). This is the equivalence we invoke in Proposition 5.71.
+(We also note that Ryan–Smith’s localization for $\mathsf{PP}$ uses both conjuncts: $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$; see [17, Prop.
+3.17] and Question 3.18 therein, and see [17, Props.
 3.22 and 3.23] for counterexamples showing the conjuncts cannot in general be omitted.)
@@
 Fix $p\in\mathbb{Q}_{f}$ and $x\in X$ with $x\notin\operatorname{dom}(p)$.
-Choose any $y\in Y$ with $f(y)=x$ (exists since $f$ is surjective).
+Let $y\in Y$ be such that $f(y)=x$ (exists since $f$ is surjective).
 We claim $y\notin\operatorname{ran}(p)$: if $y=p(x^{\prime})$ for some $x^{\prime}\in\operatorname{dom}(p)$ then $x=f(y)=f(p(x^{\prime}))=x^{\prime}$, contradicting $x\notin\operatorname{dom}(p)$.
@@
 Remark 4.21.
-In $\mathsf{ZF}$ the ordinal $\aleph^{*}(S)$ exists for every set $S$ (it is the Lindenbaum number of $S$); see e.g. Jech [4, §(Hartogs/Lindenbaum)].
+In $\mathsf{ZF}$ the ordinal $\aleph^{*}(S)$ exists for every set $S$ (it is the Lindenbaum number of $S$); see e.g. Jech [8, §(Hartogs/Lindenbaum)].
 Remark 4.22 (Why the bound $\lambda<\aleph^{*}(S)$ is the correct target).
-Assuming $\mathsf{SVC}(S)$, Ryan–Smith shows that $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the bounded scheme $\mathsf{AC}_{<\aleph^{*}(S)}$ (more precisely, $\mathsf{AC}_{<\aleph^{*}(S)}(S)$ in his notation); see [13, Cor.
+Assuming $\mathsf{SVC}(S)$, Ryan–Smith shows that $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the bounded scheme $\mathsf{AC}_{<\aleph^{*}(S)}$ (more precisely, $\mathsf{AC}_{<\aleph^{*}(S)}(S)$ in his notation); see [17, Cor.
 3.3].
-For fixed $\lambda$, the corresponding “splitting” reformulations of $\mathsf{AC}_{\lambda}$ appear in [13, Prop.
+For fixed $\lambda$, the corresponding “splitting” reformulations of $\mathsf{AC}_{\lambda}$ appear in [17, Prop.
 3.2].
@@
 Lemma 4.27 (Countable closure).
-$\mathbb{R}_{f}$ is Countably closed.
+Assume $M\models\mathsf{ZF}+\mathsf{DC}$. Then $\mathbb{R}_{f}$ is countably closed.
 Proof.
@@
 This is a function, and it satisfies $f(r(\xi))=\xi$ for every $\xi\in\operatorname{dom}(r)$ since this already holds in some $r_{n}$ containing $\xi$.
-Finally, $\operatorname{dom}(r)=\bigcup_{n<\omega}\operatorname{dom}(r_{n})\subseteq\lambda$ is a countable union of countable sets of ordinals, hence countable in $M$ (this uses only $\mathsf{ZF}$ since $\lambda$ is well-ordered and each countable subset of an ordinal admits a canonical increasing enumeration).
+Finally, $\operatorname{dom}(r)=\bigcup_{n<\omega}\operatorname{dom}(r_{n})$ is a countable union of countable sets.
+Since $M\models\mathsf{DC}$ (hence $M\models AC_{\omega}$), $\operatorname{dom}(r)$ is countable in $M$.
 Therefore $r\in\mathbb{R}_{f}$, and clearly $r\leq r_{n}$ for all $n$.
 ∎
+Remark 4.28 (Closure depends on $\mathsf{DC}$).
+The $\omega_{1}$-closure of $\mathbb{R}_{f}$ is proved inside $M$ using $\mathsf{DC}$ (equivalently $AC_{\omega}$), via “countable union of countable sets is countable”.
+In bare $\mathsf{ZF}$ this need not hold.
 Proposition 4.29 (What $\mathbb{R}_{f}$ forces).
@@
 Remark 5.7 (Limit-stage $\omega_{1}$-completeness).
-At limit stages, $\omega_{1}$-completeness (and normality) of the limit filter is provided by the iteration machinery (API item (d), cf. [3, Thm. 5.5]).
+At limit stages, $\omega_{1}$-completeness (and normality) of the limit filter is provided by the iteration machinery (API item (d), cf. [5, Thm. 5.5]).
 Lemma 5.8 (Well-definedness of the orbit package actions).
@@
 Proof.
-This is exactly the successor-step group structure supplied by the countable-support symmetric iteration framework (Remark 2.1, cf. [3, Def. 3.1]). Concretely:
+This is exactly the successor-step group structure supplied by the countable-support symmetric iteration framework (Remark 2.1, cf. [5, Def. 3.1]). Concretely:
   • $\pi^{\alpha}_{\alpha+1}$ is restriction to the $\mathbb{P}_{\alpha}$-part;
@@
 Proof.
-(i) This is part of the inverse/direct system structure of the iteration groups (cf. [3, Def. 3.1]) and the fact that $\pi^{0}_{\lambda}$ is defined stagewise and respects composition.
+(i) This is part of the inverse/direct system structure of the iteration groups (cf. [5, Def. 3.1]) and the fact that $\pi^{0}_{\lambda}$ is defined stagewise and respects composition.
 (ii) We prove by induction on stages that each $g\in\mathscr{G}_{\beta}$ determines a bijection of $\mathrm{Coords}_{\beta}$ (in the sense of Definition 5.22), and hence sends countable sets of coordinates to countable sets.
@@
 At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are (by definition) coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf.
-[3, Def.
+[5, Def.
 3.1]).
@@
 Limit $\eta$: Define $g_{\eta}$ to be the direct-limit element of $\mathscr{G}_{\eta}$ determined by the family $\langle g_{\beta}:\beta<\eta\rangle$, which is coherent by the inductive invariant above.
-By the limit-stage group definition ([3, Def.
+By the limit-stage group definition ([5, Def.
 3.1]), $g_{\eta}\in\mathscr{G}_{\eta}$.
-By the uniqueness property of the inverse/direct limit presentation: elements of $\mathscr{G}_{\eta}$ are determined by their projections to all earlier stages ([3, Def.
+By the uniqueness property of the inverse/direct limit presentation: elements of $\mathscr{G}_{\eta}$ are determined by their projections to all earlier stages ([5, Def.
 3.1]).
@@
 Proof.
-By Lemma 5.52, the triple $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ is a symmetric system in the sense of [4, Ch.
+By Lemma 5.52, the triple $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ is a symmetric system in the sense of [8, Ch.
 15].
-By the symmetric extension theorem [4, Lemma.
+By the symmetric extension theorem [8, Lemma.
 15.51], $M_{\lambda}\models\mathsf{ZF}$.
@@
   1. (i)
-    The seed parameter $S=A^{\omega}$ is preserved at successor stages: our package forcings add no new reals (Corollaries 4.19 and 4.29), hence add no new $\omega$-sequences of Cohen reals and therefore no new elements of $S$.
+    The seed parameter $S=A^{\omega}$ is preserved at successor stages: our package forcings add no new reals (Corollaries 4.19 and 4.30), hence add no new $\omega$-sequences of Cohen reals and therefore no new elements of $S$.
   2. (ii)
@@
 For the global conclusion (a final model of $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$), we take a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$.
-Concretely, working in the metatheory, we form an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [3, §2.1], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$.
+Concretely, working in the metatheory, we form an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [5, §2.1], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$.
 At a successor stage $\alpha+1$, we consult the finite list $\mathcal{B}(\alpha)$ from Definition 5.59.
@@
 Thus each successor filter explicitly includes the diagonal subgroups protecting the representative coordinates of the active packages.
-At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group/filter as provided by the symmetric-iteration framework (API items (b)–(e), especially filter generation/completeness at limits, [3, Thm.
+At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group/filter as provided by the symmetric-iteration framework (API items (b)–(e), especially filter generation/completeness at limits, [5, Thm.
 5.5]).
@@
 Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name, its interpretation is unchanged by further forcing: $\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}$, so $f=f_{\alpha}$ in $\mathcal{M}$.
+Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$–name of bounded rank, and no forcing past stage $\alpha$ adds sets of that rank, Theorem 9.2 of Karagila [9] implies that $\dot{f}^{G_{\lambda}}=\dot{f}^{G_{\alpha}}$.
+Hence $f=f_{\alpha}$ in $\mathcal{M}$.
 If $f_{\alpha}$ already has a section in $M_{\alpha}$, then $f$ splits in $\mathcal{M}$ and we are done. Otherwise, $f_{\alpha}$ is eligible for scheduling by the bookkeeping.
@@
 ∎
-Remark 5.63 (Countable‐support names at limits of cofinality $\omega$).
-For limits $\lambda$ of countable cofinality, a name may have countable support cofinal in $\lambda$.
-The iteration framework of [3] handles this by block‐factorization and limit‐filter completeness (see [3, Rem.
+Remark 5.63 (Countable-support names at limits of cofinality $\omega$).
+At limits $\lambda$ with $\operatorname{cf}(\lambda)=\omega$, a hereditarily symmetric name may have countable support that is cofinal in $\lambda$, so stage-bounding arguments do not apply.
+The iteration machinery of [5] handles such limits using limit-filter $\omega_{1}$-completeness together with the appropriate block/factorization analysis (cf.
+[5, Rem.
 3.9, Thm.
-5.5]); our argument implicitly relies on that machinery.
+5.5]), and we rely on those limit-stage conclusions.
 Proposition 5.64 ($A$ is not well-orderable).
@@
 Since $\widehat{\pi}^{\,D}\in\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\dot{w})$, we have $\widehat{\pi}^{\,D}\cdot\dot{w}=\dot{w}$.
-By equivariance of name interpretation ([4, Lemma 14.37]):
+By equivariance of name interpretation ([8, Lemma 14.37]):
 $\widehat{\pi}^{\,D}\cdot w=(\widehat{\pi}^{\,D}\cdot\dot{w})^{G_{\lambda}}=\dot{w}^{G_{\lambda}}=w.$
@@
 Therefore the $\mathsf{DC}$-preservation theorem for symmetric extensions applies (API item (c), cf.
-[3, Thm.
+[5, Thm.
 3.20]), yielding $M_{\alpha+1}\models\mathsf{DC}$.
 At limit stages, $\mathsf{DC}$ is preserved by the countable-support symmetric iteration machinery (API item (c), and the limit-filter generation/completeness from API item (d); cf.
-[3, Thm.
-3.20] and [3, Thm.
+[5, Thm.
+3.20] and [5, Thm.
 5.5] for the limit construction).
@@
 By Proposition 5.71, $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$.
-By Lemma 5.70, $\mathcal{M}\models\mathsf{SVC}^{+}(T)$ for $T=\mathcal{P}(S)$, so the hypotheses of the Ryan–Smith localization theorem (Theorem 4.5, i.e., [13, Prop.
+By Lemma 5.70, $\mathcal{M}\models\mathsf{SVC}^{+}(T)$ for $T=\mathcal{P}(S)$, so the hypotheses of the Ryan–Smith localization theorem (Theorem 4.5, i.e., [17, Prop.
 3.17] applied with Ryan–Smith’s parameter $S$ taken to be our $T$) hold in $\mathcal{M}$.

### v4

@@
-Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework
+A countable-support symmetric iteration separating $\mathsf{PP}$ from $\mathsf{AC}$
 Abstract
+We construct, from a ground model of $\mathsf{ZFC}$, a transitive symmetric model $\mathcal{M}$ satisfying $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\mathsf{AC}_{\mathsf{WO}}+\neg\mathsf{AC}$.
+The construction starts with a Cohen symmetric seed model $\mathcal{N}$ over $\operatorname{Add}(\omega,\omega_{1})$ and performs an $\mathrm{Ord}$-length countable-support symmetric iteration.
+For fixed parameters $S:=A^{\omega}$ and $T:=\mathcal{P}(S)$ (as computed in $\mathcal{N}$), successor stages add orbit-symmetrized packages which force the localized splitting principle $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ (hence $\mathsf{PP}\restriction T$) and the choice principle $\mathsf{AC}_{\mathsf{WO}}$, while preserving $\mathsf{DC}$ and keeping $A$ non-well-orderable.
+A diagonal-lift/diagonal-cancellation scheme produces $\omega_{1}$-complete normal limit filters.
+A persistence argument yields $\mathsf{SVC}^{+}(T)$ in $\mathcal{M}$, and Ryan–Smith localization then upgrades $\mathsf{PP}\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ to $\mathsf{PP}$.
+1 Introduction
 The Partition Principle $\mathsf{PP}$ asserts that whenever there is a surjection $A\twoheadrightarrow B$, there is an injection $B\hookrightarrow A$.
-Russell conjectured in 1906 that $\mathsf{PP}$ is equivalent to the Axiom of Choice $\mathsf{AC}$; while $\mathsf{AC}\Rightarrow\mathsf{PP}$ is immediate, the converse has remained open.
-We show that $\mathsf{PP}$ does not imply $\mathsf{AC}$ by constructing a transitive model of $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$.
-Starting from a Cohen symmetric model $\mathcal{N}$ of $\operatorname{Add}(\omega,\omega_{1})$ built with a countable-support symmetry filter, we fix parameters $S:=A^{\omega}$ and $T:=\mathcal{P}(S)$ and perform a class-length countable-support symmetric iteration.
-At successor stages we use orbit-symmetrized packages that split targeted surjections, yielding $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$, while preserving $\mathsf{DC}$ and ensuring that $A$ remains non-well-orderable.
-A diagonal-cancellation/diagonal-lift infrastructure supplies a proper $\omega_{1}$-complete normal filter at limit stages.
-Finally, Ryan–Smith localization shows that under $\mathsf{SVC}^{+}(T)$, $\mathsf{PP}$ is equivalent to $\mathsf{PP}\!\restriction T\wedge\mathsf{AC}_{\mathsf{WO}}$, so the final model satisfies $\mathsf{PP}$ but not $\mathsf{AC}$.
-1 Introduction
-Whether the Partition Principle implies the Axiom of Choice is sometimes called “the oldest open problem in set theory.” The Partition Principle ($\mathsf{PP}$) asserts that whenever $A$ surjects onto $B$, one has $|B|\leq|A|$.
-Russell conjectured in 1906 that $\mathsf{PP}$ is equivalent to $\mathsf{AC}$, but no proof was given [1].
-The converse is immediate: choice provides sections for all surjections.
-The question of whether $\mathsf{PP}$ and $\mathsf{AC}$ are equivalent over $\mathsf{ZF}$ has remained open.
-Main Result. We resolve this question in the negative by constructing a model of
-$\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC},$
-thereby establishing that the Partition Principle does not imply the Axiom of Choice.
+Over $\mathsf{ZFC}$ this is immediate, since every surjection admits a right inverse.
+The aim of this manuscript is to separate $\mathsf{PP}$ from the full Axiom of Choice by a symmetric-iteration construction.
+Main theorem.
+The following is proved as Theorem 5.73.
 Theorem.
-The theory $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$ is consistent relative to $\mathsf{ZF}$.
-Strategy.
-The construction proceeds in two stages.
-Stage 1: Seed model.
-We build a Cohen symmetric model $\mathcal{N}$ over $\operatorname{Add}(\omega,\omega_{1})$ using countable supports, producing a model of $\mathsf{ZF}+\mathsf{DC}$ containing a non-well-orderable set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$ of Cohen reals.
-The fixed set $S$ witnesses $\mathsf{SVC}(S)$, which will drive the localization argument.
-Stage 2: Symmetric iteration.
-We perform a class-length countable-support symmetric iteration over $\mathcal{N}$.
-At each successor stage, we force sections for surjections $f:Y\twoheadrightarrow X$ with $X,Y\subseteq T:=\mathcal{P}(S)$, as well as right inverses witnessing $\mathsf{AC}_{\mathsf{WO}}$.
-The iteration preserves $\mathsf{DC}$ and the non-well-orderability of $A$.
-Localization.
-The global Partition Principle follows from Ryan–Smith’s localization theorem [17], which shows that under $\mathsf{SVC}^{+}(T)$, the principle $\mathsf{PP}$ is equivalent to the conjunction of $\mathsf{PP}\!\restriction\!T$ (the restriction to surjections between subsets of $T$) and $\mathsf{AC}_{\mathsf{WO}}$.
+Assume $V\models\mathsf{ZFC}$.
+In a class forcing extension $V[G]$ obtained by an $\mathrm{Ord}$-length countable-support symmetric iteration over a Cohen symmetric seed model $\mathcal{N}$, the final symmetric model $\mathcal{M}$ satisfies
+$\mathsf{ZF}\ +\ \mathsf{DC}\ +\ \mathsf{PP}\ +\ \mathsf{AC}_{\mathsf{WO}}\ +\ \neg\mathsf{AC}.$
+Metatheory.
+All forcing and symmetry constructions are carried out over a fixed ground universe $V$.
+Set-sized stages are treated in the usual set-theoretic background.
+To define and analyze the $\mathrm{Ord}$-length iteration as a definable class forcing (and to use the forcing theorem for $\mathbb{P}_{\mathrm{Ord}}$), we work in a background class theory supporting the relevant recursions, namely $\mathsf{GBC}+\mathsf{ETR}$, using a class well-order of $V$ only where it is explicitly invoked for bookkeeping; see Section 2 and Remark 5.59.
 Remark 1.1 (The axiom $\mathsf{AC}_{\mathsf{WO}}$).
-Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes the axiom commonly called choice for well–orderable sets, in the sense of Howard–Rubin [7, Form 40]. That is:
-Every well–orderable set of nonempty sets admits a choice function.
-Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds: if $\neq\varnothing, Y\varnothing\notin Y$, and $|Y|\leq|\lambda|$, then there exists a function $f$ with domain $Y$ such that $f(y)\in y$ for all $y\in Y$.
-This axiom should be distinguished from the trivial principle asserting choice for families already equipped with specified well–orderings, which is provable in $\mathsf{ZF}$.
-The present form asserts the existence of a uniform choice function for arbitrary well–orderable families, without any fixed well–orderings given in advance.
-As shown by Pincus (see [7]), $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the statement that $\aleph(X)=\aleph^{*}(X)$ for all sets $X$.
-This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used implicitly in Ryan–Smith’s localization theorem [17].
-Technical innovations.
-Two devices are essential to making the iteration work.
-(1) Orbit-symmetrized packages.
-A naïve attempt to force a section for a single surjection $f$ fails because the action of $\operatorname{Sym}(\omega_{1})$ on conditions does not extend to an action on the forcing poset (automorphisms would need to map $\mathbb{Q}_{f}$ to $\mathbb{Q}_{\pi\cdot f}$, a different poset).
-We resolve this by forcing with the full orbit package
-$\mathbb{Q}_{[f]}:=\prod^{\mathrm{cs}}_{g\in[f]}\mathbb{Q}_{g},$
-the countable-support product over all $g$ in the $\operatorname{Sym}(\omega_{1})$-orbit of $f$.
-This ensures that every $\pi\in\operatorname{Sym}(\omega_{1})$ acts by permuting coordinates and transporting structure, yielding a genuine automorphism of $\mathbb{Q}_{[f]}$.
-(2) Diagonal-cancellation symmetries.
-Orbit-symmetrization alone does not make the generic section $s_{f}$ hereditarily symmetric: the naïve action of $\pi\in\operatorname{Fix}(E)$ on conditions still moves the $f$-coordinate generic.
-We introduce diagonal-cancellation automorphisms $\widehat{\pi}^{\,D}$ that act as the identity on a protected set $D$ of coordinates (including $f$) while acting as $\pi$ elsewhere.
-The stage filter includes subgroups $\Delta^{\uparrow}(E,\{f\})$ generated by these diagonal lifts, ensuring that the generic section is stabilized by a filter subgroup and hence belongs to the symmetric model.
-Crucially, diagonal lifts still act as $\pi$ on the Cohen base (stage 0), so they swap Cohen reals $c_{\beta}\leftrightarrow c_{\gamma}$ for $\beta,\gamma\notin E$, preserving the $\neg\mathsf{AC}$ witness.
-Conventions.
-Throughout, $\mathsf{DC}$ denotes $\mathsf{DC}_{\omega}$, i.e., dependent choice for $\omega$-sequences.
-We write $\mathsf{PP}\!\restriction\!T$ for the restriction of $\mathsf{PP}$ to surjections $f:Y\twoheadrightarrow X$ with both $X,Y\subseteq T$.
-External tools.
-We rely on three external frameworks, imported as black boxes.
-(1) Symmetric iteration machinery.
-For the countable-support symmetric iteration framework—including successor and limit constructions, filter generation, $\omega_{1}$-completeness at limits, and the hereditarily-symmetric calculus—we cite Gilson [5].
-See Remark 2.1 for the specific API items we invoke.
-(2) DC preservation.
-For preservation of Dependent Choice in symmetric extensions, we follow Karagila [10] and Karagila–Schilhan [12], building on classical work of Pincus [14].
-The criterion we use is: $\omega_{1}$-closed forcing with $\omega_{1}$-complete filter preserves $\mathsf{DC}$.
-(3) Localization of $P$.
-We use Ryan–Smith’s local reflection framework [17], which provides the equivalence $\mathsf{PP}\Longleftrightarrow\mathsf{PP}\!\restriction\!T+\mathsf{AC}_{\mathsf{WO}}$ under $\mathsf{SVC}^{+}(T)$.
-For background on $\mathsf{SVC}$ and Hartogs–Lindenbaum spectra, see [16, 2].
-For general background on forcing and symmetric extensions we follow Jech [8] and Kunen [13].
-Outline.
-Section 3 constructs the seed model $\mathcal{N}$ and verifies $\mathsf{DC}$, $\neg\mathsf{AC}$, and $\mathsf{SVC}(S)$.
-Section 4 recalls the Ryan–Smith localization theorem and defines the target principles $\mathsf{PP}\!\restriction\!T$ and $\mathsf{AC}_{\mathsf{WO}}$.
-Subsection 4.1 and Subsection 4.2 define the package forcings $\mathbb{Q}_{f}$ and $\mathbb{R}_{f}$ for forcing sections and right inverses.
-Section 5 constructs the symmetric iteration, including the orbit-symmetrization and diagonal-cancellation machinery, and assembles the final model $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$.
-2 Preliminaries and black-box iteration technology
-We use the general symmetric-iteration machinery from [5] (and ultimately from Karagila’s iteration framework [9]).
-Here we only record the “API” we will use later.
-We treat the following as black boxes; the precise hypotheses and proofs are available in [5].
-Remark 2.1 (Iteration API to be used later).
-We will use, as needed, the following results from [5].
+Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes choice for well-orderable sets: every well-orderable set of nonempty sets admits a choice function.
+Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds.
+This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used when applying Ryan–Smith localization (Theorem 4.5).
+How $\mathsf{PP}$ is obtained.
+Let $\mathcal{N}$ be the Cohen symmetric seed model over $\operatorname{Add}(\omega,\omega_{1})$ constructed in Section 3, and let $A=\{c_{\alpha}:\alpha<\omega_{1}\}$ be the canonical set of Cohen reals in $\mathcal{N}$.
+Fix $S:=A^{\omega}$ and $T:=\mathcal{P}(S)$ as computed in $\mathcal{N}$ (Definition 3.27); these parameters are held fixed throughout the iteration.
+Successor stages of the iteration schedule orbit-symmetrized package forcings which (in the resulting stage models) add:
+  • witnesses for the localized splitting principle $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ (Definition 4.8); and
+  • right inverses witnessing $\mathsf{AC}_{\mathsf{WO}}$ for well-orderable targets.
+We force splitting because it is the local property that admits a uniform, highly closed "package forcing" $\mathbb{Q}_{f}$ per surjection, and it immediately implies the $\mathsf{PP}\restriction T$ hypothesis needed.
+At limit stages, diagonal-lift/diagonal-cancellation symmetries are used to define $\omega_{1}$-complete normal limit filters compatible with countable support.
+In the final model $\mathcal{M}$ we prove the local splitting statement (Proposition 5.65) and $\mathsf{AC}_{\mathsf{WO}}$ (Proposition 5.72), and we show that $\mathsf{SVC}(S)$ persists through the iteration (Lemma 5.69 and Corollary 5.70), hence $\mathsf{SVC}^{+}(T)$ holds in $\mathcal{M}$.
+Therefore, Ryan–Smith localization (Theorem 4.5) yields $\mathsf{PP}$ in $\mathcal{M}$ from $\mathsf{PP}\restriction T$ together with $\mathsf{AC}_{\mathsf{WO}}$; we obtain $\mathsf{PP}\restriction T$ because we force $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ (Definition 4.8).
+The remaining clauses of the main theorem are established by the corresponding final-stage propositions: $\mathsf{ZF}$ (Theorem 5.58), $\mathsf{DC}$ (Proposition 5.68), and $\neg\mathsf{AC}$ via non-well-orderability of $A$ (Proposition 5.67).
+Consequences recorded here.
+From $\mathsf{SVC}(S)$ and $\mathsf{PP}$ in $\mathcal{M}$, we derive the Ordering Principle (Corollary 5.74).
+We also record, conservatively, the Kinna–Wagner consequences proved in this manuscript (Remark 5.75).
+Relation to countable-support symmetric iterations.
+The iteration is an instance of the general countable-support symmetric-iteration framework developed in [5]; we use its successor and limit constructions, the $\omega_{1}$-completeness and normality facts for the generated filters, and the $\mathsf{ZF}/\mathsf{DC}$ preservation theorems.
+What is specific to the present manuscript is the package bookkeeping for $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ and the orbit-symmetrization required to obtain genuine automorphisms of the successor iterands (Remark 2.4).
+What is not claimed.
+Beyond the theory displayed above and the explicit corollaries proved in the final section (such as Corollary 5.74 and Remark 5.75), we do not claim additional choiceless consequences.
+Organization.
+Section 2 fixes the metatheoretic conventions and records the countable-support symmetric-iteration API items imported from [5].
+Section 3 constructs the seed model $\mathcal{N}$ and establishes $\mathsf{DC}$, $\neg\mathsf{AC}$, and $\mathsf{SVC}(S)$ at stage 0.
+Section 4 recalls Ryan–Smith localization and isolates the fixed parameter $T$ together with the localized principles.
+Subsections 4.1 and 4.2 define the package forcings used at successor stages.
+Section 5 defines the class-length symmetric iteration and verifies the final-stage propositions, culminating in Theorem 5.73.
+2 Preliminaries and metatheoretic conventions
+We follow the forcing, automorphism, symmetric-extension, and countable-support symmetric-iteration framework developed in [5], based on Karagila’s iteration scheme [9].
+This section fixes the metatheory and notation used throughout, and records precisely which results are imported from [5].
+2.1 Background theory and conventions
+Remark 2.1 (Metatheory vs. object theory).
+All forcing notions, names, automorphism groups, filters, and supports are constructed in a fixed background universe $V$.
+  • For the set-forcing and symmetric-extension statements used at bounded stages, we work over $V\models\mathsf{ZF}$.
+  • Whenever we invoke preservation of dependent choice for a bounded stage symmetric model, we assume the background satisfies $V\models\mathsf{ZFC}$.
+Throughout this paper, $\mathsf{DC}$ means $\mathsf{DC}_{\omega}$ (dependent choice for $\omega$-sequences, in the standard set formulation used in [5, §3.5]).
+  • For the class-length recursion through $\mathrm{Ord}$, we assume a background theory adequate for class recursion (e.g.
+$V\models\mathsf{GBC}+\mathsf{ETR}$), and we assume Global Choice only when explicitly invoked for bookkeeping.
+When we say “countable”, $\omega_{1}$-complete”, etc., this is always with respect to the background universe $V$.
+2.2 Restrictions, supports, and stage generics
+Remark 2.2 (Forcing vs. group restrictions).
+We follow the convention of [5, Remark 2.2] and do not overload restriction/projection notation.
+  • For forcing, $p\upharpoonright\beta$ denotes the usual restriction of a condition $p\in\mathbb{P}_{\lambda}$ to an initial segment $\beta\leq\lambda$ (and when needed, $\pi_{\beta,\lambda}:\mathbb{P}_{\lambda}\to\mathbb{P}_{\beta}$ denotes the canonical projection map).
+  • For groups, $\rho_{\beta,\lambda}:\mathscr{G}_{\lambda}\to\mathscr{G}_{\beta}$ denotes the restriction homomorphism.
+Notation 2.3 (Stage generics).
+Let $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic for the class-length iteration forcing.
+For each $\alpha\leq\mathrm{Ord}$, write $G_{\alpha}:=G\cap\mathbb{P}_{\alpha}$ for the induced $V$-generic filter on the initial segment forcing.
+2.3 Imported iteration results from [5]
+Remark 2.4 (Imported iteration facts).
+We will use the following results from [5]. Each time we invoke an item below, we will explicitly point to the corresponding hypothesis in the current construction.
   1. (a)
-    Set-length and class-length versions of symmetric iterations (cf. [5, §2.1]).
+    Existence of set- and class-length iterations.
+The countable-support symmetric iteration recursion can be carried out through any set ordinal, and (under the background class-recursion assumptions used there) through $\mathrm{Ord}$; see [5, §Countable-support symmetric iterations and §Class-length iterations].
   2. (b)
-    Definitions and basic properties of countable-support symmetric iterations (cf. [5, Def. 3.1]).
+    Framework and actions.
+At each stage $\lambda$ the iteration provides a single ambient group $\mathscr{G}_{\lambda}\leq\operatorname{Aut}(\mathbb{P}_{\lambda})$ acting by automorphisms on the full initial-segment forcing $\mathbb{P}_{\lambda}$ (so automorphisms may continue to move earlier “seed” coordinates); see [5, §Framework and Setup].
   3. (c)
-    A $\mathsf{DC}$-preservation theorem for the relevant support (cf. [5, Thm. 3.20]).
+    Limit filters. In the relevant cofinality cases, the canonical limit filter generated from earlier-stage filters is normal and $\omega_{1}$-complete; see [5, §Limit filters].
   4. (d)
-    Generation and completeness/normality of limit filters in the relevant cofinality cases (cf. [5, Thm. 5.5]).
-  5. (e)
-    A characterization of HS names by “small support” assuming the stage model satisfies $\mathsf{DC}$ (cf. [5, Lem. 3.14]).
-In our application, every stage model satisfies $\mathsf{DC}$ (Proposition 5.65), so the support characterization may be used freely throughout the iteration.
-We emphasize that at each stage $\lambda$ the iteration produces a single ambient group $\mathscr{G}_{\lambda}$ acting by automorphisms on the full initial-segment forcing $\mathbb{P}_{\lambda}$ (API item (b)), so the action continues to move the original Cohen coordinates; we are not iterating by repeatedly forcing over intermediate symmetric models in isolation.
-Our main tool for reducing the global Partition Principle to a local instance is the "Small Violations of Choice" framework.
-Definition 2.2 (Small Violations of Choice, [17]).
-Let $S$ be a set.
-  1. 1.
-    We say that $\mathsf{SVC}(S)$ holds if for every set $X$, there exists an ordinal $\eta$ such that there is a surjection $f\colon S\times\eta\twoheadrightarrow X$ (equivalently, $|X|\leq^{*}|S\times\eta|$).
-  2. 2.
-    We say that $\mathsf{SVC}^{+}(S)$ holds if for every set $X$, there exists an ordinal $\eta$ such that there is an injection $f\colon X\hookrightarrow S\times\eta$ (equivalently, $|X|\leq|S\times\eta|$).
-Ryan–Smith localization ingredients (used later). We will repeatedly appeal to two results from [17], which we record in Section 4 in the exact forms used in this paper:
-  • Injective upgrade: for every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$ (Fact 4.3).
-  • Localization of $\mathsf{PP}$: assuming $\mathsf{SVC}^{+}(T)$, one has
-    $\mathsf{PP}\ \Longleftrightarrow\ \bigl(\mathsf{PP}\!\restriction T\wedge\mathsf{AC}_{\mathsf{WO}}\bigr)$
-    (Theorem 4.5, i.e. [17, Proposition 3.17]).
-Accordingly, our strategy is to start from a Cohen symmetric seed $S$ with $\mathsf{SVC}(S)$, take $T:=\mathcal{P}(S)$, force $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ while preserving $\mathsf{SVC}(S)$, and then invoke Theorem 4.5 to conclude global $\mathsf{PP}$.
+    Bounded-stage $\mathsf{DC}$.
+For set ordinals $\lambda$, assuming $V\models\mathsf{ZFC}$, if the stage filter is $\omega_{1}$-complete then the stage symmetric model $M_{\lambda}=V(\mathbb{P}_{\lambda})_{\operatorname{HS}}$ satisfies $\mathsf{DC}$; see [5, §Dependent Choice and Choice Fragments].
+We stress that [5] does not assert $\mathsf{DC}$ for the final class-length symmetric model in general.
+In this paper, $\mathsf{DC}$ for the final model is obtained by combining item (d) at bounded stages with the stage-localization lemma for set-sized $\mathbb{P}_{\mathrm{Ord}}$-names (Lemma 5.64).
+Choice principles and the Ryan–Smith localization theorem are stated (and used) in Section 4.
 3 The Cohen symmetric seed model
@@
 $\mathbb{P}\;=\;\operatorname{Add}(\omega,\omega_{1})\;:=\;\operatorname{Fn}(\omega_{1}\times\omega,2,{<}\omega),$
-$\dot{c}_{\alpha}:=\{\langle\check{n},p\rangle:p(\alpha,n)=1\},c_{\alpha}=\dot{c}^{G}_{\alpha}$
-the standard forcing to add $\omega_{1}$-many Cohen reals, presented as finite partial functions $p:\omega_{1}\times\omega\to 2$.
+ordered by reverse inclusion.
+For each $\alpha<\omega_{1}$, let $\dot{c}_{\alpha}$ be the canonical $\mathbb{P}$-name for the $\alpha$-th Cohen real,
+$\dot{c}_{\alpha}\;:=\;\{\langle\check{n},p\rangle:p\in\mathbb{P}\ \wedge\ (\alpha,n)\in\operatorname{dom}(p)\ \wedge\ p(\alpha,n)=1\}.$
+Fix $G_{0}\subseteq\mathbb{P}$ $V$-generic and write $c_{\alpha}:=\dot{c}_{\alpha}^{G_{0}}\in 2^{\omega}$.
 Lemma 3.1 (Cohen reals are pairwise distinct).
@@
 Fix $\alpha\neq\beta<\omega_{1}$. Let
-$D_{\alpha,\beta}=\{p\in\operatorname{Add}(\omega,\omega_{1}):\exists n\,(p(\alpha,n)\neq p(\beta,n))\}.$
+$D_{\alpha,\beta}=\Bigl\{p\in\mathbb{P}:\exists n\in\omega\ \bigl[(\alpha,n),(\beta,n)\in\operatorname{dom}(p)\ \wedge\ p(\alpha,n)\neq p(\beta,n)\bigr]\Bigr\}.$
 Given any $p$, choose $n$ such that neither $(\alpha,n)$ nor $(\beta,n)$ is in $\operatorname{dom}(p)$, and extend $p$ to $q\leq p$ with $q(\alpha,n)=0$ and $q(\beta,n)=1$.
 Then $q\in D_{\alpha,\beta}$.
-Hence $D_{\alpha,\beta}$ is dense, so any generic $G$ meets it and therefore $c_{\alpha}\neq c_{\beta}$.
+Hence $D_{\alpha,\beta}$ is dense, so $G_{0}$ meets it and therefore $c_{\alpha}\neq c_{\beta}$.
 ∎
@@
 Definition 3.2 (Automorphisms of $\operatorname{Add}(\omega,\omega_{1})$).
-Let $\mathscr{G}:=\operatorname{Sym}(\omega_{1})$ be the full permutation group of $\omega_{1}$. Each $\pi\in\mathscr{G}$ acts on $\mathbb{P}$ by
-$(\pi p)(\pi(\alpha),n)\;=\;p(\alpha,n),$
-equivalently $(\pi p)=\{(\pi(\alpha),n,i):(\alpha,n,i)\in p\}$. This extends (recursively) to an action on $\mathbb{P}$-names in the usual way.
+Let $\mathscr{G}:=\operatorname{Sym}(\omega_{1})$ be the full permutation group of $\omega_{1}$.
+Each $\pi\in\mathscr{G}$ induces an automorphism of $\mathbb{P}$ by permuting the $\omega_{1}$-coordinate:
+$\operatorname{dom}(\pi p)=\{(\pi(\alpha),n):(\alpha,n)\in\operatorname{dom}(p)\}\quad\text{and}\quad(\pi p)(\pi(\alpha),n)=p(\alpha,n).$
+This extends to $\mathbb{P}$-names by the standard recursion:
+$\pi\dot{x}\;:=\;\{\langle\pi\dot{y},\pi p\rangle:\langle\dot{y},p\rangle\in\dot{x}\}.$
 Definition 3.3 (Countable-support filter of subgroups).
@@
 $\operatorname{Stab}(\dot{x})\;:=\;\{\pi\in\mathscr{G}:\pi\dot{x}=\dot{x}\}.$
-We call $\dot{x}$ symmetric if $\operatorname{Stab}(\dot{x})\in\mathscr{F}$.
-Let $\operatorname{HS}$ be the class of hereditarily symmetric names: $\dot{x}\in\operatorname{HS}$ iff $\dot{x}$ is symmetric and every name appearing in $\dot{x}$ is in $\operatorname{HS}$.
-Fix $G\subseteq\mathbb{P}$ $V$-generic. The associated symmetric extension is
-$\mathcal{N}\;:=\;\operatorname{HS}^{G}\;=\;\{\dot{x}^{G}:\dot{x}\in\operatorname{HS}\}.$
+For $E\subseteq\omega_{1}$, we say that $E$ supports $\dot{x}$ if $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$.
+We call $\dot{x}$ symmetric if $\operatorname{Stab}(\dot{x})\in\mathscr{F}$, equivalently if $\dot{x}$ has some $V$-countable support $E\in[\omega_{1}]^{\leq\omega}$ (by Lemma 3.4).
+Let $\operatorname{HS}$ be the class of hereditarily symmetric names: $\dot{x}\in\operatorname{HS}\iff\dot{x}$ is symmetric and every name appearing in $\dot{x}$ lies in $\operatorname{HS}$.
+Fix $G_{0}\subseteq\mathbb{P}$ $V$-generic. The associated symmetric extension is
+$\mathcal{N}\;:=\;\operatorname{HS}^{G_{0}}\;=\;\{\dot{x}^{G_{0}}:\dot{x}\in\operatorname{HS}\}.$
 3.2 CCC and dependent choice
@@
 Proof.
-Let $\{p_{\xi}:\xi<\omega_{1}\}\subseteq\operatorname{Add}(\omega,\omega_{1})$ be uncountable. Write $D_{\xi}=\operatorname{dom}(p_{\xi})\subseteq\omega\times\omega_{1}$, so each $D_{\xi}$ is finite.
+Let $\{p_{\xi}:\xi<\omega_{1}\}\subseteq\operatorname{Add}(\omega,\omega_{1})$ be uncountable. Write $D_{\xi}=\operatorname{dom}(p_{\xi})\subseteq\omega_{1}\times\omega$, so each $D_{\xi}$ is finite.
 By thinning out, assume $|D_{\xi}|=n$ for all $\xi$.
@@
 $\mathscr{G}\leq\operatorname{Aut}(\mathbb{P})$ and $\mathscr{F}$ is a normal filter of subgroups of $\mathscr{G}$.
-If $G\subseteq\mathbb{P}$ is $V$-generic, then $\operatorname{HS}^{G}=\{\dot{x}^{G}:\dot{x}\in\operatorname{HS}\}$ is a transitive model of $\mathsf{ZF}$ and $V\subseteq\operatorname{HS}^{G}\subseteq V[G]$.
+If $H\subseteq\mathbb{P}$ is $V$-generic, then $\operatorname{HS}^{H}=\{\dot{x}^{H}:\dot{x}\in\operatorname{HS}\}$ is a transitive model of $\mathsf{ZF}$ and $V\subseteq\operatorname{HS}^{H}\subseteq V[H]$.
 Proof.
-This is the standard symmetric extension theorem; see, e.g., Jech chapter 15, Lemma 15.51[8]. ∎
+This is the standard symmetric extension theorem; see [8, Lemma 15.51]. ∎
 Lemma 3.9 (HS witnesses from HS-sets).
@@
 $\mathsf{ZF}$.
-By Lemma 3.8, the symmetric extension $\mathcal{N}=\operatorname{HS}^{G}$ is a transitive model of $\mathsf{ZF}$ (containing all ordinals, as usual for symmetric extensions).
+By Lemma 3.8, the symmetric extension $\mathcal{N}=\operatorname{HS}^{G_{0}}$ is a transitive model of $\mathsf{ZF}$ (containing all ordinals, as usual for symmetric extensions).
 We record the following symmetry lemma for the action of $\mathscr{G}$ on $\mathbb{P}$ for later use.
@@
 Let $A,R\in\mathcal{N}$ with $A\neq\emptyset$ and $\forall x\in A\,\exists y\in A\,(xRy)$.
-Fix $\dot{A},\dot{R}\in\operatorname{HS}$ with $\dot{A}^{G}=A$ and $\dot{R}^{G}=R$.
+Fix $\dot{A},\dot{R}\in\operatorname{HS}$ with $\dot{A}^{G_{0}}=A$ and $\dot{R}^{G_{0}}=R$.
 Choose $p_{\ast}\in\mathbb{P}$ such that
@@
 3.3 Countable supports for $\operatorname{HS}$-names
-Notation 3.12 (Restrictions to $E$).
-For $E\subseteq\omega_{1}$ and $p\in\operatorname{Add}(\omega,\omega_{1})$, let
-$p\!\restriction E\;:=\;p\cap(E\times\omega\times 2),$
-be viewed as a condition in $\mathbb{P}_{E}:=\operatorname{Add}(\omega,E)$.
-For a $\mathbb{P}$-name $\tau$, define recursively the restricted name $\tau\!\restriction E$ by
-$\tau\!\restriction E\;:=\;\{\langle\sigma\!\restriction E,\;p\!\restriction E\rangle:\langle\sigma,p\rangle\in\tau\}.$
-Then $\tau\!\restriction E$ is a $\mathbb{P}_{E}$-name (via the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$).
-Definition 3.13 (Support).
-If $\dot{x}\in\operatorname{HS}$, we say that $E\subseteq\omega_{1}$ is a support for $\dot{x}$ if $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$.
+Notation 3.12 (Coordinate restriction to $E$).
+Let $E\subseteq\omega_{1}$. Set
+$\mathbb{P}_{E}\;:=\;\operatorname{Add}(\omega,E)\;=\;\operatorname{Fn}(E\times\omega,2,{<}\omega),$
+viewed as a (complete) subposet of $\mathbb{P}=\operatorname{Add}(\omega,\omega_{1})$ via the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$.
+For $p\in\mathbb{P}$, define the restriction of $p$ to $E$ by
+$p\!\restriction E\;:=\;p\cap\bigl((E\times\omega)\times 2\bigr).$
+Equivalently, $\operatorname{dom}(p\!\restriction E)=\operatorname{dom}(p)\cap(E\times\omega)$ and $(p\!\restriction E)(\alpha,n)=p(\alpha,n)$ for $(\alpha,n)\in\operatorname{dom}(p)\cap(E\times\omega)$.
+Then $p\!\restriction E\in\mathbb{P}_{E}$ and $p\leq p\!\restriction E$ (reverse inclusion order).
+Remark 3.13 (Symmetry supports).
+Recall that $E\subseteq\omega_{1}$ supports a $\mathbb{P}$-name $\tau$ if $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$.
+We refer to such $E$ as a symmetry support to distinguish it from forcing supports.
 Lemma 3.14 (Every $\operatorname{HS}$-name has a countable support).
@@
 If $\dot{x}\in\operatorname{HS}$, then in particular $\dot{x}$ is $\mathscr{F}$–symmetric, so $\operatorname{Stab}(\dot{x})\in\mathscr{F}$.
-Since $\operatorname{Stab}(\dot{x})\in\mathscr{F}$, Lemma 3.4 yields a countable $E$ with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$.
+By Lemma 3.4, there exists a countable $E\in[\omega_{1}]^{\leq\omega}$ with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$.
 This is exactly the assertion that $\dot{x}$ has countable support.
@@
 ∎
+Lemma 3.16, Remark 3.17, and Lemma 3.18 record standard support and subextension facts for the Cohen seed, and will not be invoked explicitly later; they are included to fix conventions and prevent common confusions about symmetry supports versus $\mathbb{P}_{E}$-names.
 Lemma 3.16 (Support localization for $\operatorname{Add}(\omega,\omega_{1})$).
-Let $E\subseteq\omega_{1}$ and let $\tau$ be a $\mathbb{P}$-name such that $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$.
-Let $\varphi$ be any formula whose parameters are $\mathbb{P}$-names also fixed by $\operatorname{Fix}(E)$.
-If $p\Vdash\varphi(\tau)$, then $p\!\restriction E\Vdash\varphi(\tau)$.
+Let $E\subseteq\omega_{1}$ and let $\tau$ be a $\mathbb{P}$-name with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$.
+Let $\varphi$ be any formula and let $\vec{\sigma}$ be a tuple of $\mathbb{P}$-names such that $\operatorname{Fix}(E)\leq\operatorname{Stab}(\sigma_{i})$ for each $i$.
+Then for every $p\in\mathbb{P}$,
+$p\Vdash\varphi(\tau,\vec{\sigma})\quad\Longleftrightarrow\quad p\!\restriction E\Vdash\varphi(\tau,\vec{\sigma}).$
 Proof.
-Suppose $p\Vdash\varphi(\tau)$ but $p\!\restriction E\nVdash\varphi(\tau)$. Then there is $q\leq p\!\restriction E$ with $q\Vdash\neg\varphi(\tau)$.
-Let $F_{p}\subseteq\omega_{1}\setminus E$ be the finite set of ordinals mentioned by $p$ outside $E$ (i.e.
-those $\alpha\notin E$ with $(\alpha,n)\in\operatorname{dom}(p)$ for some $n$), and define $F_{q}$ similarly.
-Choose a permutation $\pi\in\operatorname{Fix}(E)$ such that $\pi[F_{p}]$ is disjoint from $F_{q}$ and such that $\pi$ fixes every element of $F_{q}$.
-Then $\pi p$ is compatible with $q$ (they agree on $E$ and have disjoint requirements outside $E$), so let $r\leq\pi p,q$.
-By the symmetry lemma (Claim 3.2) and $\pi\in\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$, we have $\pi p\Vdash\varphi(\pi\tau)=\varphi(\tau)$, hence $r\Vdash\varphi(\tau)$.
-But also $r\leq q$ and $q\Vdash\neg\varphi(\tau)$, so $r\Vdash\neg\varphi(\tau)$, a contradiction.
+The implication $\Leftarrow$” holds since $p\leq p\!\restriction E$.
+For $\Rightarrow$”, suppose $p\Vdash\varphi(\tau,\vec{\sigma})$ but $p\!\restriction E\nVdash\varphi(\tau,\vec{\sigma})$.
+Then there is $q\leq p\!\restriction E$ with $q\Vdash\neg\varphi(\tau,\vec{\sigma})$.
+Let
+$\begin{split}&F_{p}:=\{\alpha\in\omega_{1}\setminus E:\exists n\in\omega\ ((\alpha,n)\in\operatorname{dom}(p))\},\\ &F_{q}:=\{\alpha\in\omega_{1}\setminus E:\exists n\in\omega\ ((\alpha,n)\in\operatorname{dom}(q))\}.\end{split}$
+Choose $\pi\in\operatorname{Fix}(E)$ such that $\pi[F_{p}]\cap F_{q}=\varnothing$ and $\pi$ fixes every element of $F_{q}$. Then $\pi p$ is compatible with $q$, so pick $r\leq\pi p,q$.
+Since $\pi\in\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$ and $\operatorname{Fix}(E)\leq\operatorname{Stab}(\sigma_{i})$ for each $i$, we have $\pi\tau=\tau$ and $\pi\sigma_{i}=\sigma_{i}$ for all $i$.
+By automorphism invariance of the forcing relation, $\pi p\Vdash\varphi(\pi\tau,\pi\vec{\sigma})$, i.e.
+$\pi p\Vdash\varphi(\tau,\vec{\sigma})$, hence $r\Vdash\varphi(\tau,\vec{\sigma})$.
+But $r\leq q$ and $q\Vdash\neg\varphi(\tau,\vec{\sigma})$, contradiction.
 ∎
@@
 The condition $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$ ($\tau$ has symmetry support EE”) means that $\tau$ is invariant under permutations fixing $E$ pointwise.
-This does not imply that $\tau$ is a $\mathbb{P}_{E}$–name, nor that $\tau^{G}$ lies in $V[G_{E}]$.
-Indeed, letting $\dot{A}=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha<\omega_{1}\}$ be the canonical name for the Cohen set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$, we have $\operatorname{Stab}(\dot{A})=\operatorname{Sym}(\omega_{1})$, so $\dot{A}$ has symmetry support $\emptyset$.
+This does not imply that $\tau$ is a $\mathbb{P}_{E}$–name, nor that $\tau^{G_{0}}$ lies in $V[G_{E}]$.
+Indeed, letting $\dot{A}:=\{\langle\dot{c}_{\alpha},\mathbf{1}\rangle:\alpha<\omega_{1}\}$ be the canonical name for the Cohen set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$, we have $\operatorname{Stab}(\dot{A})=\operatorname{Sym}(\omega_{1})$, so $\dot{A}$ has symmetry support $\emptyset$.
 Nevertheless, $\dot{A}$ is not a $\mathbb{P}_{E}$–name for any countable $E$, and $A\notin V[G_{E}]$ for every countable $E\subseteq\omega_{1}$.
@@
 Let $E\subseteq\omega_{1}$, let $\mathbb{P}_{E}=\operatorname{Add}(\omega,E)\leq\mathbb{P}=\operatorname{Add}(\omega,\omega_{1})$, and let $\tau$ be a $\mathbb{P}_{E}$–name (viewed as a $\mathbb{P}$–name via the inclusion).
-If $G\subseteq\mathbb{P}$ is VV–generic and $G_{E}:=G\cap\mathbb{P}_{E}$, then
-$\tau^{G}=\tau^{G_{E}}\in V[G_{E}].$
+If $G_{0}\subseteq\mathbb{P}$ is VV–generic and $G_{E}:=G_{0}\cap\mathbb{P}_{E}$, then
+$\tau^{G_{0}}=\tau^{G_{E}}\in V[G_{E}].$
 Proof.
 By induction on the rank of $\tau$.
-Using the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$, the valuation recursion for $\tau^{G}$ only consults conditions from $G$ that appear in $\tau$, and these conditions all lie in $\mathbb{P}_{E}$, hence belong to $G_{E}$.
-Applying the inductive hypothesis to subnames of $\tau$ yields $\tau^{G}=\tau^{G_{E}}$.
+Using the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$, the valuation recursion for $\tau^{G_{0}}$ only consults conditions from $G_{0}$ that appear in $\tau$, and these conditions all lie in $\mathbb{P}_{E}$, hence belong to $G_{E}$.
+Applying the inductive hypothesis to subnames of $\tau$ yields $\tau^{G_{0}}=\tau^{G_{E}}$.
 ∎
 3.4 The canonical Cohen set $A$ and failure of choice
-For each $\alpha<\omega_{1}$, define the canonical $\mathbb{P}$-name $\dot{c}_{\alpha}$ for the $\alpha$-th Cohen real by
-$\dot{c}_{\alpha}\;:=\;\{(\check{n},p):p(\alpha,n)=1\}.$
-Let $c_{\alpha}:=\dot{c}_{\alpha}^{G}\in 2^{\omega}$.
+Recall that for each $\alpha<\omega_{1}$, $\dot{c}_{\alpha}$ denotes the canonical $\mathbb{P}$-name for the $\alpha$-th Cohen real and $c_{\alpha}:=\dot{c}_{\alpha}^{G_{0}}\in 2^{\omega}$.
 Definition 3.19 (The Cohen set).
-$A\;:=\;\{c_{\alpha}:\alpha<\omega_{1}\}.$
+Define the canonical $\mathbb{P}$-name
+$\dot{A}\;:=\;\{\langle\dot{c}_{\alpha},\mathbf{1}_{\mathbb{P}}\rangle:\alpha<\omega_{1}\},$
+and set
+$A\;:=\;\dot{A}^{G_{0}}\;=\;\{c_{\alpha}:\alpha<\omega_{1}\}\subseteq 2^{\omega}.$
 Lemma 3.20.
@@
 Proof.
-For each $\alpha<\omega_{1}$, the canonical name $\dot{c}_{\alpha}$ for the $\alpha$-th Cohen real has support $\{\alpha\}$, hence is hereditarily symmetric. Let
-$\dot{A}=\{(\dot{c}_{\alpha},\mathbbm{1}_{\mathbb{P}}):\alpha<\omega_{1}\}.$
-Any $\pi\in\mathscr{G}$ permutes the $\dot{c}_{\alpha}$’s, so $\pi(\dot{A})=\dot{A}$.
-Thus $\operatorname{Stab}(\dot{A})=\mathscr{G}\in\mathscr{F}$, and $\dot{A}$ is symmetric.
-Since its members are hereditarily symmetric, $\dot{A}\in\operatorname{HS}$, hence $A=\operatorname{val}(\dot{A},G)\in\mathcal{N}$.
-Indeed $\dot{c}_{\alpha}$ is a $\mathbb{P}_{\{\alpha\}}$-name, hence $\dot{c}_{\alpha}\in\operatorname{HS}$ by Lemma 3.15.
+By Lemma 3.15, each $\dot{c}_{\alpha}$ lies in $\operatorname{HS}$ (indeed it has symmetry support $\{\alpha\}$).
+For any $\pi\in\mathscr{G}$, $\pi\dot{A}=\dot{A}$, hence $\operatorname{Stab}(\dot{A})=\mathscr{G}\in\mathscr{F}$ and $\dot{A}$ is symmetric.
+Since all members of $\dot{A}$ are hereditarily symmetric, $\dot{A}\in\operatorname{HS}$.
+Therefore $A=\dot{A}^{G_{0}}\in\operatorname{HS}^{G_{0}}=\mathcal{N}$.
 ∎
@@
 Since $\mathsf{AC}$ implies that every set is well-orderable (equivalently, injects into an ordinal), this will imply that $A$ is not well-orderable in $\mathcal{N}$, hence $\mathcal{N}\models\neg\mathsf{WO}$ and therefore $\mathcal{N}\models\neg\mathsf{AC}$.
-Suppose toward a contradiction that there are an ordinal $\theta$ and $\dot{f}\in\operatorname{HS}$ such that $\dot{f}^{G}$ is an injection $A\to\theta$ in $\mathcal{N}$.
-Choose $p_{0}\in G$ and $\theta$ with
+Suppose toward a contradiction that there exist an ordinal $\theta$, a name $\dot{f}\in\operatorname{HS}$, and a condition $p_{0}\in\mathbb{P}$ such that
 $p_{0}\Vdash\text{``$\dot{f}:\dot{A}\to\check{\theta}$ is an injection.''}$
@@
 ∎
-Remark 3.22 (Dedekind-infinitude).
-Unlike the finite-support generalized Cohen model (which produces a Dedekind-finite $A$; cf.
-[15, §2.4]), the present countable-support filter allows many countable subsets of $A$: for each countable $E\subseteq\omega_{1}$ in $V$, the name $\dot{A}_{E}:=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha\in E\}$ is hereditarily symmetric, so $A_{E}:=\{c_{\alpha}:\alpha\in E\}\in\mathcal{N}$.
+Remark 3.22 (Countable subsets of $A$).
+In contrast with the finite-support generalized Cohen model (where the analogous $A$ can be Dedekind-finite; cf.
+[16, §2.4]), the present countable-support filter admits many canonical countable subsets of $A$.
+Fix $E\in[\omega_{1}]^{\leq\omega}\cap V$ and set
+$\dot{A}_{E}\;:=\;\{\langle\dot{c}_{\alpha},\mathbf{1}_{\mathbb{P}}\rangle:\alpha\in E\}.$
+Then $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{A}_{E})$, hence $\dot{A}_{E}\in\operatorname{HS}$ and
+$A_{E}:=\dot{A}_{E}^{G_{0}}=\{c_{\alpha}:\alpha\in E\}\in\mathcal{N}.$
+If $E$ is infinite and $b:\omega\to E$ is a bijection in $V$, then the sequence $\langle c_{b(n)}:n\in\omega\rangle\in\mathcal{N}$ (Lemma 3.23) witnesses that $A$ is Dedekind-infinite in $\mathcal{N}$.
 Lemma 3.23 (Coding a countable set of Cohen reals by an element of $A^{\omega}$).
@@
 ∎
-Lemma 3.24 (A forced nonempty name yields a ground-coded witness).
-Let $\dot{X}$ be a $\mathbb{P}$–name and suppose $p\Vdash``\dot{X}\neq\varnothing$". Then there exist $q\leq p$ and a $\mathbb{P}$–name $\dot{x}$ such that $q\Vdash``\dot{x}\in\dot{X}$".
-Proof.
-This is the standard forcing-theorem witness extraction for $\exists$. From $p\Vdash\exists x\,(x\in\dot{X})$ we obtain $q\leq p$ and a name $\dot{x}$ with $q\Vdash\dot{x}\in\dot{X}$. ∎
 Lemma 3.24 (Well-orders are rigid).
@@
 Proof.
-Suppose $\sigma$ is an order automorphism of $(X,<)$ and $\sigma\neq\mathrm{id}_{X}$. Let $x_{0}$ be the <<-least element of $X$ with $\sigma(x_{0})\neq x_{0}$.
-If $\sigma(x_{0})<x_{0}$, then $\sigma(x_{0})<x_{0}$ implies (by minimality of $x_{0}$) that $\sigma(\sigma(x_{0}))=\sigma(x_{0})$.
-But order preservation yields $\sigma(\sigma(x_{0}))<\sigma(x_{0})$, a contradiction.
-If $\sigma(x_{0})>x_{0}$, then apply the previous case to $\sigma^{-1}$: since $\sigma^{-1}$ is also an order automorphism and $\sigma^{-1}(x_{0})<x_{0}$ (because $x_{0}<\sigma(x_{0})$), we get a contradiction.
-Hence $\sigma=\mathrm{id}_{X}$. ∎
+Suppose $h$ is an order automorphism of $(X,\prec)$ and $h\neq\mathrm{id}_{X}$. Let $x_{0}$ be the $\prec$-least element of $X$ with $h(x_{0})\neq x_{0}$.
+If $h(x_{0})\prec x_{0}$, then by minimality of $x_{0}$ we have $h(h(x_{0}))=h(x_{0})$. But order preservation gives $h(h(x_{0}))\prec h(x_{0})$, contradiction.
+If $x_{0}\prec h(x_{0})$, apply the previous case to $h^{-1}$, which is also an order automorphism of $(X,\prec)$, and satisfies $h^{-1}(x_{0})\prec x_{0}$, again a contradiction.
+Therefore $h=\mathrm{id}_{X}$. ∎
 3.5 A safe $\mathsf{SVC}$ seed in $\mathcal{N}$
+Definition 3.25 ($\mathsf{SVC}(S)$).
+Let $S$ be a set. We write $\mathsf{SVC}(S)$ for the statement:
+$\forall X\,\exists\eta\in\mathrm{Ord}\,\exists e\ (e:S\times\eta\twoheadrightarrow X).$
 Proposition 3.26 ($\mathsf{SVC}(A^{\omega})$ in the Cohen symmetric model).
@@
 Proof.
-We use the standard “continuity” analysis of Cohen-type symmetric models.
-In the basic Cohen model $V(A)$ generated by an unordered set $A$ of Cohen reals, a key lemma (sometimes called a continuity or indiscernibility lemma) shows that statements over $V(A)$ with parameters from $V$ are already decided in a subextension generated by finitely many members of $A$; see, e.g., Felgner [4, p.
-133] or Blass [3, Prop.
-1.2] (and the discussion in [6, Lem.
-2.8]).
-More precisely, for each $x\in X$ there is a $\mathbb{P}$–name $\tau_{x}\in V$ and a sequence $s_{x}\in A^{\omega}$ (coding the countable symmetry support relevant to $\tau_{x}$) such that $x$ is recovered uniformly from the pair $(\tau_{x},s_{x})$.
-Enumerating the set of $\mathbb{P}$–names in $V$ by ordinals yields the required surjection from $A^{\omega}\times\eta$ onto $X$ for some ordinal $\eta$.
-The same argument applies to our Cohen symmetric model $\mathcal{N}$ built from $\operatorname{Add}(\omega,\omega_{1})$ with the countable-support symmetry filter: every set $X\in\mathcal{N}$ is obtained (uniformly in $X$) from $V$ and a countable sequence of members of $A$, and therefore is a surjective image of $A^{\omega}\times\eta$ for some ordinal $\eta$.
-Equivalently, $\mathcal{N}\models\mathsf{SVC}(A^{\omega})$.
-∎
+We work in $\mathcal{N}$.
+The conclusion $\mathsf{SVC}(A^{\omega})$ is a standard consequence of the “continuous reading/continuity” analysis for generalized Cohen symmetric models under a symmetry filter generated by $\operatorname{Fix}(E)$ for countable $E\subseteq\omega_{1}$.
+In the finite-support Cohen model $V(A)$, the corresponding analysis yields $\mathsf{SVC}(A^{<\omega})$ (see Felgner [4, p.
+133] and Blass [3, Prop.
+1.2]).
+In our present seed model $\mathcal{N}$, the symmetry filter is generated by $\operatorname{Fix}(E)$ for countable $E$, and thus the same argument runs with $\omega$-sequences of elements of $A$ in place of finite tuples.
+Concretely, the relevant input is exactly:
+  1. 1.
+    every hereditarily symmetric name admits a countable symmetry support (Lemma 3.14); and
+  2. 2.
+    every countable $E\subseteq\omega_{1}$ from $V$ can be coded by an element of $A^{\omega}$ in $\mathcal{N}$ (Lemma 3.23).
+A detailed presentation of this countable-support version is given in [6, Lem. 2.8]. Therefore $\mathcal{N}\models\mathsf{SVC}(A^{\omega})$. ∎
 Definition 3.27 (Fixed seed parameters: $S$ and $T$).
@@
     $\mathsf{PP}(T)$ is the statement that for every set $X$, if there is a surjection $f:T\twoheadrightarrow X$, then there is an injection $i:X\hookrightarrow T$.
+The notation $\mathsf{PP}\!\restriction T$ follows [18, §3.8] and is unrelated to forcing-condition restrictions $p\restriction E$.
 Definition 4.2 ($\mathsf{SVC}^{+}$).
@@
 Fact 4.3 (Ryan–Smith).
-For every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$. [17, §2.1, Fact]
+For every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$. [18, §2.1, Fact]
 Remark 4.4.
@@
 Theorem 4.5 (Ryan–Smith localization of $\mathsf{PP}$).
-Assume $\mathsf{SVC}^{+}(T)$. Then
+Assume $\mathsf{SVC}^{+}(T)$ (Definition 4.2). Then
 $\mathsf{PP}\quad\Longleftrightarrow\quad\bigl(\mathsf{PP}\!\restriction T\ \wedge\ \mathsf{AC}_{\mathsf{WO}}\bigr).$
-[17, Proposition 3.17] (stated there with parameter “SS”; we apply it with $S:=T$). As this argument is critical to our goals, we demonstrate a proof here.
+[18, Proposition 3.17] (stated there with parameter “SS”; we apply it with $S:=T$).
+Here $\mathsf{PP}\!\restriction T$ is as in Definition 4.1, and $\mathsf{AC}_{\mathsf{WO}}$ is as in Remark $1.$ As this argument is critical to our goals, we demonstrate a proof here.
 Proof.
@@
 Moreover, $\mathsf{PP}$ implies $\mathsf{AC}_{\mathsf{WO}}$ (e.g.
-via the consequence $(\forall X)\,\aleph(X)=\aleph^{\ast}(X)$; see [17, Proposition 3.17 and footnote 4] for references).
-($\Leftarrow$) Assume $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$, and let $\pi:Y\twoheadrightarrow X$ be a surjection. We must produce an injection $X\hookrightarrow Y$.
+via the consequence $(\forall X)\,\aleph(X)=\aleph^{\ast}(X)$; see [18, Proposition 3.17 and footnote 4] for references).
+($\Leftarrow$) Assume $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$, and let $\varphi:Y\twoheadrightarrow X$ be a surjection. We must produce an injection $X\hookrightarrow Y$.
 Apply $\mathsf{SVC}^{+}(T)$ to the set
@@
 obtaining an ordinal $\eta$ and an injection $j:Z\hookrightarrow T\times\eta$. Define injections
-$\pi_{X}:X\to T\times\eta,\quad\pi_{X}(x)=j(\langle x,0\rangle),\qquad\pi_{Y}:Y\to T\times\eta,\quad\pi_{Y}(y)=j(\langle y,1\rangle),$
+$\iota_{X}:X\to T\times\eta,\quad\iota_{X}(x)=j(\langle x,0\rangle),\qquad\iota_{Y}:Y\to T\times\eta,\quad\iota_{Y}(y)=j(\langle y,1\rangle),$
 and let
-$A:=\operatorname{ran}(\pi_{X})\subseteq T\times\eta,\qquad B:=\operatorname{ran}(\pi_{Y})\subseteq T\times\eta.$
-Then $\pi_{X}:X\to A$ and $\pi_{Y}:Y\to B$ are bijections, hence have inverses $\pi_{X}^{-1}:A\to X$ and $\pi_{Y}^{-1}:B\to Y$. Define
-$f\;:=\;\pi_{X}\circ\pi\circ\pi_{Y}^{-1}\;:\;B\twoheadrightarrow A.$
+$A:=\operatorname{ran}(\iota_{X})\subseteq T\times\eta,\qquad B:=\operatorname{ran}(\iota_{Y})\subseteq T\times\eta.$
+Then $\iota_{X}:X\to A$ and $\iota_{Y}:Y\to B$ are bijections, hence have inverses $\iota_{X}^{-1}:A\to X$ and $\iota_{Y}^{-1}:B\to Y$. Define
+$f\;:=\;\iota_{X}\circ\varphi\circ\iota_{Y}^{-1}\;:\;B\twoheadrightarrow A.$
 It therefore suffices to show that whenever $A,B\subseteq T\times\eta$ and $f:B\twoheadrightarrow A$, there is an injection $g:A\hookrightarrow B$.
@@
 By Claim 4 there is an injection $g:A\hookrightarrow B$. Then
-$i\;:=\;\pi_{Y}^{-1}\circ g\circ\pi_{X}\;:\;X\to Y$
+$i\;:=\;\iota_{Y}^{-1}\circ g\circ\iota_{X}\;:\;X\to Y$
 is an injection, as required. Therefore $\mathsf{PP}$ holds. ∎
@@
 Since $|X|\leq|Y|\Rightarrow|X|\leq^{*}|Y|$ holds in $\mathsf{ZF}$, this is equivalent to our surjection-to-injection formulation of $\mathsf{PP}$.
-Proposition 4.7 (The $\mathsf{AC}_{\mathsf{WO}}$ conjunct cannot be omitted).
-There is a model $M\models\mathsf{ZF}$ and a set $S^{\ast}\in M$ such that
-$M\models\mathsf{SVC}^{+}(S^{\ast})\ \wedge\ \mathsf{PP}(S^{\ast})\ \wedge\ \neg\mathsf{PP}.$
-In particular, global $\mathsf{PP}$ cannot be recovered from a small-choice hypothesis together with a single-parameter instance $\mathsf{PP}(S^{\ast})$.
-Reference.
-This is witnessed by Cohen’s first model; see [17, Proposition 3.22]. ∎
-Proposition 4.8 (The localized $\mathsf{PP}\!\restriction S$ component cannot be omitted).
-There is a model $M\models\mathsf{ZF}$ and a set $S^{\ast}\in M$ such that
-$M\models\mathsf{AC}_{\mathsf{WO}}\ \wedge\ \mathsf{SVC}^{+}(S^{\ast})\ \wedge\ \mathsf{PP}(S^{\ast})\ \wedge\ \neg\mathsf{PP}.$
-Consequently, even assuming $\mathsf{AC}_{\mathsf{WO}}$ and small-choice hypotheses, one cannot in general replace the localized requirement $\mathsf{PP}\!\restriction S^{\ast}$ by the weaker single-parameter assertion $\mathsf{PP}(S^{\ast})$ and still obtain global $\mathsf{PP}$.
-Reference.
-This is witnessed by the Feferman-style model $N_{\aleph_{1}}$ of Truss; see [17, Proposition 3.23]. ∎
-Remark 4.9 (Why we keep both conjuncts).
-The preceding propositions are included only to justify (conceptually) why the localization theorem is stated with both components present: the “choice” side $\mathsf{AC}_{\mathsf{WO}}$ and the “localized comparability” side $\mathsf{PP}\!\restriction T$.
-We will force both conjuncts $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ in Theorem 4.5.
-In Ryan–Smith’s analysis the corresponding reduction is sharp: in general one cannot drop the displayed conjuncts (nor substantially weaken the hypotheses) without losing the equivalence; see [17, Props.
-3.22 and 3.23].
+Remark 4.7 (On the hypotheses in Theorem 4.5).
+Theorem 4.5 is exactly [18, Proposition 3.17] (with parameter $T$). We therefore keep both target principles $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ explicit in the iteration.
+Ryan–Smith also considers the single-parameter instance $\mathsf{PP}(T)$ (Definition 4.1).
+In general, small-choice hypotheses together with $\mathsf{PP}(T)$ do not suffice to recover global $\mathsf{PP}$: Cohen’s first model satisfies $\mathsf{SVC}^{+}(\mathbb{R})\wedge\mathsf{PP}(\mathbb{R})\wedge\neg\mathsf{PP}$ [18, Proposition 3.22].
+Moreover, even adding $\mathsf{AC}_{\mathsf{WO}}$ does not allow one to replace a localized requirement of the form $\mathsf{PP}\!\restriction T$ by a single-parameter assertion $\mathsf{PP}(T)$: the Feferman-style model $N_{\aleph_{1}}$ of Truss satisfies $\mathsf{AC}_{\mathsf{WO}}\wedge\mathsf{SVC}^{+}(\mathcal{P}(\mathbb{R}))\wedge\mathsf{PP}(\mathcal{P}(\mathbb{R}))\wedge\neg\mathsf{PP}$ [18, Proposition 3.23].
+Finally, whether $\mathsf{AC}_{\mathsf{WO}}$ follows from $\mathsf{SVC}^{+}(T)\wedge\mathsf{PP}\!\restriction T$ is open in general; see [18, Question 3.18].
+This is why $\mathsf{AC}_{\mathsf{WO}}$ remains an explicit forcing target below.
 4.1 The localized $\mathsf{PP}$-package forcing $\mathbb{Q}_{f}$
 Fix a transitive model $M\models\mathsf{ZF}+\mathsf{DC}$ extending the Cohen seed $\mathcal{N}$ (in practice, $M$ will be an intermediate stage of our symmetric iteration).
-Work with the fixed parameter $T=\mathcal{P}(S)$ from Definition 3.27.
+Let $T$ be the fixed set defined in $\mathcal{N}$ by Definition 3.27.
+(Thus $T\in M$, but $T$ need not equal $\mathcal{P}(S)$ as computed in $M$.)
 Definition 4.8 (Localized splitting principle).
@@
 $\mathsf{AC}_{\mathsf{WO}}\quad\Longleftrightarrow\quad\mathsf{AC}_{<\aleph^{*}(S)}(S)\quad\Longleftrightarrow\quad\text{every function }g:S\to\lambda\text{ splits (for each }\lambda\text{)}$
-(see [17, Prop. 3.2]). This is the equivalence we invoke in Proposition 5.71.
-(We also note that Ryan–Smith’s localization for $\mathsf{PP}$ uses both conjuncts: $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$; see [17, Prop.
-3.17] and Question 3.18 therein, and see [17, Props.
-3.22 and 3.23] for counterexamples showing the conjuncts cannot in general be omitted.)
+(see [18, Prop. 3.2]). This is the equivalence we invoke in Proposition 5.72.
+(For discussion of the hypotheses in Ryan–Smith’s localization (including limitations on weakening them and the redundancy question for $\mathsf{AC}_{\mathsf{WO}}$), see [18, Prop.
+3.17, Question 3.18, and Props.
+3.22–3.23].)
 Definition 4.10 ($\mathsf{PP}$-package for a fixed surjection).
@@
 Lemma 4.13 (Countable closure).
-$\mathbb{Q}_{f}$ is Countably-closed (sometimes called $\omega_{1}$-closed).
+$\mathbb{Q}_{f}$ is countably closed (i.e. $\omega$-closed) in $M$.
 Proof.
@@
 Remark 4.14 (Closure depends on $\mathsf{DC}$).
-The $\omega_{1}$-closure of $\mathbb{Q}_{f}$ is proved inside $M$ using $\mathsf{DC}$ (equivalently $\mathsf{AC}_{\omega}$), via “countable union of countable sets is countable”.
-In a bare $\mathsf{ZF}$ context this need not hold, so the closure claim should be read as relative to our standing hypothesis $M\models\mathsf{ZF}+\mathsf{DC}$.
+The countable closure of $\mathbb{Q}_{f}$ is proved inside $M$ using $\mathsf{DC}$ (equivalently $\mathsf{AC}_{\omega}$), via “countable union of countable sets is countable”.
+In a bare $\mathsf{ZF}$ context this need not hold, so Lemma 4.13 should be read as relative to our standing hypothesis $M\models\mathsf{ZF}+\mathsf{DC}$.
 Proposition 4.15 (What $\mathbb{Q}_{f}$ forces).
@@
 Lemma 4.16 (Countably closed forcing adds no new $\omega$-sequences and preserves $\mathsf{DC}$).
-Let $M\models\mathsf{ZF}+\mathsf{DC}$ be transitive and let $\mathbb{P}\in M$ be countably closed in $M$ (i.e. every decreasing $\omega$-sequence in $M$ has a lower bound in $M$). Then:
+Let $M\models\mathsf{ZF}+\mathsf{DC}$ be transitive and let $\mathbb{U}\in M$ be countably closed in $M$ (i.e. every decreasing $\omega$-sequence in $M$ has a lower bound in $M$). Then:
   1. (1)
-    $\mathbb{P}$ adds no new $\omega$-sequences of ordinals.
+    $\mathbb{U}$ adds no new $\omega$-sequences of ordinals.
   2. (2)
-    Forcing with $\mathbb{P}$ over $M$ preserves $\mathsf{DC}$.
+    Forcing with $\mathbb{U}$ over $M$ preserves $\mathsf{DC}$.
 Proof.
-(1) No new $\omega$-sequences of ordinals. Let $p_{0}\in\mathbb{P}$ and $\dot{a}\in M$ with $p_{0}\Vdash\dot{a}:\check{\omega}\to\mathrm{Ord}$. For each $n<\omega$, let
-$D_{n}:=\Bigl\{p\leq p_{0}:\exists\alpha\in\mathrm{Ord}\ (p\Vdash\dot{a}(\check{n})=\check{\alpha})\Bigr\}.$
-Each $D_{n}$ is dense below $p_{0}$.
+(1) No new $\omega$-sequences of ordinals. Let $u_{0}\in\mathbb{U}$ and $\dot{a}\in M$ with $u_{0}\Vdash\dot{a}:\check{\omega}\to\mathrm{Ord}$. For each $n<\omega$, let
+$D_{n}:=\Bigl\{u\leq u_{0}:\exists\alpha\in\mathrm{Ord}\ (u\Vdash\dot{a}(\check{n})=\check{\alpha})\Bigr\}.$
+Each $D_{n}$ is dense below $u_{0}$.
 Define the set of states
-$\Sigma:=\Bigl\{(n,p,\vec{\alpha}):n<\omega,\ p\in\mathbb{P},\ p\leq p_{0},\ \vec{\alpha}\in\mathrm{Ord}^{n},\ p\Vdash\forall k<n\ \dot{a}(\check{k})=\check{\vec{\alpha}(k)}\Bigr\}.$
+$\Sigma:=\Bigl\{(n,u,\vec{\alpha}):n<\omega,\ u\in\mathbb{U},\ u\leq u_{0},\ \vec{\alpha}\in\mathrm{Ord}^{n},\ u\Vdash\forall k<n\ \dot{a}(\check{k})=\check{\vec{\alpha}(k)}\Bigr\}.$
 Let $R$ be the relation on $\Sigma$ given by
-$(n,p,\vec{\alpha})\ R\ (n+1,q,\vec{\beta})$
-iff $q\leq p$, $\vec{\beta}\in\mathrm{Ord}^{n+1}$ extends $\vec{\alpha}$, and $q\Vdash\dot{a}(\check{n})=\check{\vec{\beta}(n)}$. Then $R$ is total on $\Sigma$ by density of $D_{n}$.
-By $\mathsf{DC}$ in $M$, there is an $R$-chain $\langle(n,p_{n},\vec{\alpha}_{n}):n<\omega\rangle\in M$ starting from some $(0,p_{0},\emptyset)\in\Sigma$ (note $\mathrm{Ord}^{0}=\{\emptyset\}$).
-By countable closure, there is $p_{\omega}\in\mathbb{P}$ with $p_{\omega}\leq p_{n}$ for all $n$.
+$(n,u,\vec{\alpha})\ R\ (n+1,v,\vec{\beta})$
+iff $v\leq u$, $\vec{\beta}\in\mathrm{Ord}^{n+1}$ extends $\vec{\alpha}$, and $v\Vdash\dot{a}(\check{n})=\check{\vec{\beta}(n)}$. Then $R$ is total on $\Sigma$ by density of $D_{n}$.
+By $\mathsf{DC}$ in $M$, there is an $R$-chain $\langle(n,u_{n},\vec{\alpha}_{n}):n<\omega\rangle\in M$ starting from some $(0,u_{0},\emptyset)\in\Sigma$ (note $\mathrm{Ord}^{0}=\{\emptyset\}$).
+By countable closure, there is $u_{\omega}\in\mathbb{U}$ with $u_{\omega}\leq u_{n}$ for all $n$.
 Let $g:=\bigcup_{n<\omega}\vec{\alpha}_{n}$.
 Since the chain is an element of $M$, Replacement in $M$ yields $g\in M$ and $g:\omega\to\mathrm{Ord}$.
-By construction, $p_{\omega}\Vdash\forall n\ \dot{a}(\check{n})=\check{g(n)}$, hence $p_{\omega}\Vdash\dot{a}=\check{g}$.
-Thus $\mathbb{P}$ adds no new $\omega$-sequences of ordinals.
-(2) Preservation of $\mathsf{DC}$. Let $p_{0}\in\mathbb{P}$ and $\dot{X},\dot{R},\dot{x}_{0}\in M$ with
-$p_{0}\Vdash\ \dot{x}_{0}\in\dot{X}\ \wedge\ (\forall x\in\dot{X})(\exists y\in\dot{X})\ x\,\dot{R}\,y.$
+By construction, $u_{\omega}\Vdash\forall n\ \dot{a}(\check{n})=\check{g(n)}$, hence $u_{\omega}\Vdash\dot{a}=\check{g}$.
+Thus $\mathbb{U}$ adds no new $\omega$-sequences of ordinals.
+(2) Preservation of $\mathsf{DC}$. Let $u_{0}\in\mathbb{U}$ and $\dot{X},\dot{R},\dot{x}_{0}\in M$ with
+$u_{0}\Vdash\ \dot{x}_{0}\in\dot{X}\ \wedge\ (\forall x\in\dot{X})(\exists y\in\dot{X})\ x\,\dot{R}\,y.$
 Define the state space
-$\Gamma:=\Bigl\{(n,p,\dot{x}):n<\omega,\ p\in\mathbb{P},\ p\leq p_{0},\ \dot{x}\in M,\ p\Vdash\dot{x}\in\dot{X}\Bigr\}.$
+$\Gamma:=\Bigl\{(n,u,\dot{x}):n<\omega,\ u\in\mathbb{U},\ u\leq u_{0},\ \dot{x}\in M,\ u\Vdash\dot{x}\in\dot{X}\Bigr\}.$
 Define $S$ on $\Gamma$ by
-$(n,p,\dot{x})\ S\ (n+1,q,\dot{y})$
-iff $q\leq p$ and $q\Vdash\dot{y}\in\dot{X}\ \wedge\ \dot{x}\,\dot{R}\,\dot{y}$. Totality of $S$ follows from the premise forced by $p_{0}$.
-By $\mathsf{DC}$ in $M$, pick an $S$-chain $\langle(n,p_{n},\dot{x}_{n}):n<\omega\rangle\in M$ starting with $(0,p_{0},\dot{x}_{0})\in\Gamma$.
-By countable closure, let $p_{\omega}\leq p_{n}$ for all $n$.
+$(n,u,\dot{x})\ S\ (n+1,v,\dot{y})$
+iff $v\leq u$ and $v\Vdash\dot{y}\in\dot{X}\ \wedge\ \dot{x}\,\dot{R}\,\dot{y}$. Totality of $S$ follows from the premise forced by $u_{0}$.
+By $\mathsf{DC}$ in $M$, pick an $S$-chain $\langle(n,u_{n},\dot{x}_{n}):n<\omega\rangle\in M$ starting with $(0,u_{0},\dot{x}_{0})\in\Gamma$.
+By countable closure, let $u_{\omega}\leq u_{n}$ for all $n$.
 Let
-$\dot{f}:=\bigl\{\langle\langle\check{n},\dot{x}_{n}\rangle,\mathbbm{1}_{\mathbb{P}}\rangle:n<\omega\bigr\}.$
+$\dot{f}:=\bigl\{\langle\langle\check{n},\dot{x}_{n}\rangle,\mathbbm{1}_{\mathbb{U}}\rangle:n<\omega\bigr\}.$
 Since the chain is in $M$, Replacement gives $\dot{f}\in M$.
-Then $p_{\omega}\Vdash\dot{f}:\check{\omega}\to\dot{X}$ and $p_{\omega}\Vdash\forall n\ \dot{f}(\check{n})\,\dot{R}\,\dot{f}(\check{n+1})$, which is exactly $\mathsf{DC}$ in the extension.
+Then $u_{\omega}\Vdash\dot{f}:\check{\omega}\to\dot{X}$ and $u_{\omega}\Vdash\forall n\ \dot{f}(\check{n})\,\dot{R}\,\dot{f}(\check{n+1})$, which is exactly $\mathsf{DC}$ in the extension.
 ∎
@@
 Proof.
-Apply Lemma 4.18 with $\mathbb{P}=\mathbb{Q}_{f}$. Countable closure holds by Lemma 4.15. ∎
+Apply Lemma 4.16 with $\mathbb{U}=\mathbb{Q}_{f}$. Countable closure holds by Lemma 4.13. ∎
 4.2 The $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{f}$
 Fix a transitive model $M\models\mathsf{ZF}+\mathsf{DC}$ extending the Cohen seed $\mathcal{N}$ (again, in practice $M$ will be an intermediate stage of our iteration).
-Retain the fixed seed parameter $S$ from Remark 4.4.
+Retain the fixed seed parameters $S,T$ from Definition 3.27 (and recall from Remark 4.4 that $T$ is kept fixed throughout the iteration).
 Definition 4.18 (The Lindenbaum bound $\aleph^{*}(S)$).
@@
 Remark 4.20 (Why the bound $\lambda<\aleph^{*}(S)$ is the correct target).
-Assuming $\mathsf{SVC}(S)$, Ryan–Smith shows that $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the bounded scheme $\mathsf{AC}_{<\aleph^{*}(S)}$ (more precisely, $\mathsf{AC}_{<\aleph^{*}(S)}(S)$ in his notation); see [17, Cor.
+Assuming $\mathsf{SVC}(S)$, Ryan–Smith shows that $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the bounded scheme $\mathsf{AC}_{<\aleph^{*}(S)}$ (more precisely, $\mathsf{AC}_{<\aleph^{*}(S)}(S)$ in his notation); see [18, Cor.
 3.3].
-For fixed $\lambda$, the corresponding “splitting” reformulations of $\mathsf{AC}_{\lambda}$ appear in [17, Prop.
+For fixed $\lambda$, the corresponding “splitting” reformulations of $\mathsf{AC}_{\lambda}$ appear in [18, Prop.
 3.2].
-Thus, in the presence of $\mathsf{SVC}(S)$ it suffices to schedule splittings only for surjections onto ordinals $\lambda<\aleph^{*}(S)$.
+Thus, in the presence of $\mathsf{SVC}(S)$ it suffices to schedule splittings only for surjections onto ordinals $\lambda<\aleph^{*}(S)^{M}$.
+In our application $\mathsf{SVC}(S)$ holds in each stage model; see Lemma 5.69 below.
 Definition 4.21 ($\mathsf{AC}_{\mathsf{WO}}$-package for a fixed surjection onto an ordinal).
@@
 Remark 4.23 (No localization to $T$ for $\mathsf{AC}_{\mathsf{WO}}$).
-Unlike the $\mathsf{PP}$-packages $\mathbb{Q}_{f}$, which are explicitly localized to $X,Y\subseteq T=\mathcal{P}(S)$, the $\mathsf{AC}_{\mathsf{WO}}$-package must be global in nature: $\mathsf{AC}_{\mathsf{WO}}$ speaks about all well-ordered index sets.
-The point of Remark 4.22 is that, assuming $\mathsf{SVC}(S)$, it is enough to schedule splittings for surjections onto $\lambda<\aleph^{*}(S)$; this reduction is precisely what makes the bookkeeping manageable.
+Unlike the localized $\mathsf{PP}$-packages $\mathbb{Q}_{f}$ (whose definitions quantify only over $X,Y\subseteq T$, with $T$ the fixed seed parameter from $\mathcal{N}$), the $\mathsf{AC}_{\mathsf{WO}}$-package is necessarily global: $\mathsf{AC}_{\mathsf{WO}}$ concerns all well-ordered index sets.
+The reduction in Remark 4.20 shows that, assuming $\mathsf{SVC}(S)$, it suffices to schedule splittings only for surjections onto ordinals $\lambda<\aleph^{*}(S)^{M}$, which is what makes the bookkeeping manageable.
 Lemma 4.24 (Basic density).
@@
 Remark 4.26 (Closure depends on $\mathsf{DC}$).
-The $\omega_{1}$-closure of $\mathbb{R}_{f}$ is proved inside $M$ using $\mathsf{DC}$ (equivalently $AC_{\omega}$), via “countable union of countable sets is countable”.
+The countable closure of $\mathbb{R}_{f}$ (Lemma 4.25) is proved inside $M$ using $\mathsf{DC}$ (equivalently $\mathsf{AC}_{\omega}$), via “countable union of countable sets is countable”.
 In bare $\mathsf{ZF}$ this need not hold.
@@
 Proof.
-Apply Lemma 4.18 with $\mathbb{P}=\mathbb{R}_{f}$. Countable closure holds by Lemma 4.27. ∎
+Apply Lemma 4.16 with $\mathbb{U}=\mathbb{R}_{f}$. Countable closure holds by Lemma 4.25. ∎
 Remark 4.29 (Package products).
@@
 Remark 5.1 (Roadmap of Section 5).
-This section describes a class-length countable-support symmetric iteration built to force local splitting principles while preserving $\mathsf{DC}$ and killing $\mathsf{AC}$.
+This section describes a class-length countable-support symmetric iteration built to force local splitting principles while preserving $\mathsf{DC}$ and maintaining $\neg\mathsf{AC}$.
 Here is the dependency flow.
@@
 A reader who wants only the global theorem can skim Subsections 5.35.6 while treating the package/diagonal machinery as the engine ensuring the required stabilizers lie in $\tilde{\mathscr{F}}^{*}_{\lambda}$.
-We now describe the symmetric iteration that forces the two conjuncts $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ over the Cohen symmetric seed model $\mathcal{N}$, while preserving $\neg\mathsf{AC}$ (witnessed by the non-well-orderability of $A$).
-All technical iteration infrastructure is imported via the API of Remark 2.1.
+We now describe the symmetric iteration that forces the two conjuncts $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ over the Cohen symmetric seed model $\mathcal{N}$, while maintaining $\neg\mathsf{AC}$ (witnessed by the non-well-orderability of $A$).
+All technical iteration infrastructure is imported via the API of Remark 2.4 (in particular, existence of the class-length iteration is used exactly under the hypotheses of [5, Theorem 3.25]).
 The purpose of this section is to isolate the package systems, the bookkeeping targets, and the verification that the final symmetric model satisfies the desired principles.
 Remark 5.2 (Modularity of the construction).
-The argument is deliberately modular: the seed model technology (Section 3), the localized $\mathsf{PP}$-packages (Section 4.1), the $\mathsf{AC}_{\mathsf{WO}}$-packages (Section 4.2), and the symmetric-iteration “API” (Remark 2.1) are separated so that if a hypothesis later needs to be strengthened or a package definition adjusted, it can typically be done locally without rewriting unrelated parts of the paper.
+The argument is deliberately modular: the seed model technology (Section 3), the localized $\mathsf{PP}$-packages (Subsection 4.1), the $\mathsf{AC}_{\mathsf{WO}}$-packages (Subsection 4.2), and the symmetric-iteration “API” (Remark 2.4) are separated so that if a hypothesis later needs to be strengthened or a package definition adjusted, it can typically be done locally without rewriting unrelated parts of the paper.
 5.1 Package symmetric systems at a stage
 Fix a transitive $M\models\mathsf{ZF}+\mathsf{DC}$ extending $\mathcal{N}$ (in practice $M=M_{\alpha}$ at an intermediate stage of the iteration).
-Keep the parameters $S$ and $T=\mathcal{P}(S)$ as in Remark 4.4.
-Remark 5.3 (An ambient $\operatorname{Sym}(\omega_{1})$-action on $T=\mathcal{P}(S)$).
-We will often use the ambient action of $\operatorname{Sym}(\omega_{1})$ on the stage-0 seed parameter $S=A^{\omega}$ and on $T=\mathcal{P}(S)$ by permuting the Cohen reals $c_{\alpha}\in A$.
-Concretely, for $\pi\in\operatorname{Sym}(\omega_{1})$ we set $\pi\cdot c_{\alpha}:=c_{\pi(\alpha)}$, extend pointwise to sequences $s\in A^{\omega}$, and then define $\pi\cdot X:=\{\pi\cdot s:s\in X\}$ for $X\subseteq S$, i.e.
-$\pi\cdot X\in T$.
-We regard this as an action which fixes all ordinals pointwise; i.e.
-$\pi\cdot\xi=\xi$ for every ordinal $\xi$.
-We furthermore extend this action to every set $x$ with $\operatorname{tc}(x)\subseteq V(T)$ by recursion on rank:
-$\pi\cdot x\;:=\;\{\pi\cdot y:y\in x\}.$
-In particular, if $f:X\to Y$ is a function with $\operatorname{tc}(f)\subseteq V(T)$ (e.g. $X,Y\subseteq T$), we identify $f$ with its graph and set
+Keep the seed parameters $S$ and $T$ from Definition 3.27.
+Thus $S=A^{\omega}$, and $T=\mathcal{P}(S)^{\mathcal{N}}$ is computed in the seed model and then treated as a fixed parameter in all later stages (we do not re-interpret $T$ as $\mathcal{P}(S)^{M}$).
+Remark 5.3 (An ambient $\mathscr{G}$-action on $S$ and $T$).
+Recall $\mathscr{G}=\operatorname{Sym}(\omega_{1})$ (Definition 3.3).
+Using the Cohen reals $c_{\alpha}\in A$, we obtain an induced action of $\mathscr{G}$ on $A$ by $\pi\cdot c_{\alpha}:=c_{\pi(\alpha)}$, hence on $S=A^{\omega}$ by coordinatewise action, and on $\mathcal{P}(S)$ by $\pi\cdot X:=\{\pi\cdot s:s\in X\}$.
+Since $T=\mathcal{P}(S)^{\mathcal{N}}\subseteq\mathcal{P}(S)$, we use the restricted action on $T$.
+We regard this as an action fixing all ordinals pointwise.
+More generally, whenever $x$ is a set with $\operatorname{tc}(x)\subseteq V(T)\cup\mathrm{Ord}$, we extend the action recursively by $\pi\cdot x:=\{\pi\cdot y:y\in x\}$.
+In particular, for a function $f:X\to Y$ with $\operatorname{tc}(f)\subseteq V(T)\cup\mathrm{Ord}$ (e.g.
+$X,Y\subseteq T\times\eta$ for an ordinal $\eta$), we identify $f$ with its graph and set
 $\pi\cdot f:=\{(\pi\cdot a,\pi\cdot b):(a,b)\in f\}.$
 Lemma 5.4 (Kernel acts trivially on $T$).
-Let $\alpha$ be any stage and let $k\in\ker(\pi^{0}_{\alpha})$.
-Then $k$ fixes $A$, $S=A^{\omega}$, and $T=\mathcal{P}(S)$ pointwise for the ambient $\operatorname{Sym}(\omega_{1})$-action of Remark 5.3.
+Let $\alpha$ be any stage and let $k\in\ker(\pi^{0}_{\alpha})$, where $\pi^{0}_{\alpha}$ is the stage-0 projection of Definition 5.23.
+Then $k$ fixes $A$, $S=A^{\omega}$, and the fixed seed parameter $T=\mathcal{P}(S)^{\mathcal{N}}$ pointwise for the ambient $\mathscr{G}$-action of Remark 5.3.
 Consequently, $k$ fixes every set/function whose transitive closure is contained in $V(T)\cup\mathrm{Ord}$; in particular, if $f:X\to Y$ with $X,Y\subseteq T$, or $f:S\times\eta\twoheadrightarrow\lambda$ with $\eta,\lambda\in\mathrm{Ord}$, then $k\cdot f=f$.
 Proof.
-Put $\sigma:=\pi^{0}_{\alpha}(k)\in\operatorname{Sym}(\omega_{1})$.
+Put $\sigma:=\pi^{0}_{\alpha}(k)\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$.
 Since $k\in\ker(\pi^{0}_{\alpha})$ we have $\sigma=\operatorname{id}$.
-By Remark 5.3, the ambient action on $A$, on $S=A^{\omega}$, and on $T=\mathcal{P}(S)$ is induced from the action $\alpha\mapsto\sigma(\alpha)$ on indices of the Cohen reals.
+By Remark 5.3, the ambient action on $A$ and on $S=A^{\omega}$ is induced from the action $\alpha\mapsto\sigma(\alpha)$ on indices of the Cohen reals, and the induced action on $\mathcal{P}(S)$ restricts to an action on the fixed subset $T=\mathcal{P}(S)^{\mathcal{N}}$.
 Therefore $\sigma=\operatorname{id}$ implies that $k$ fixes $A$, and hence also $S$ and $T$, pointwise.
@@
 $(\pi\cdot\vec{p})(g)\ :=\ A_{\pi,\pi^{-1}\cdot g}\bigl(\vec{p}(\pi^{-1}\cdot g)\bigr).$
-Lemma 5.6 (Orbit packages are $\omega_{1}$-closed and preserve $\mathsf{DC}$).
-Assume $M\models\mathsf{ZF}+\mathsf{DC}$ and let $\mathbb{S}_{[f]}$ be either $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$. Then:
-  1. (i)
-    $\mathbb{S}_{[f]}$ is $\omega_{1}$-closed in $M$;
-  2. (ii)
-    $\mathbb{S}_{[f]}$ adds no new $\omega$-sequences of ordinals (in particular, no new reals);
-  3. (iii)
-    forcing with $\mathbb{S}_{[f]}$ over $M$ preserves $\mathsf{DC}$.
+Lemma 5.6 (Orbit packages are countably closed and preserve $\mathsf{DC}$).
+Work in $M\models\mathsf{ZF}+\mathsf{DC}$. Let $f\in M$ be a surjection instance and let $\mathbb{S}_{[f]}$ be its orbit package. Then:
+  1. 1.
+    $\mathbb{S}_{[f]}$ is countably closed;
+  2. 2.
+    forcing with $\mathbb{S}_{[f]}$ adds no new $\omega$-sequences of ordinals;
+  3. 3.
+    $M^{\mathbb{S}_{[f]}}\models\mathsf{DC}$.
 Proof.
-(i) Let $\langle\vec{p}_{n}:n<\omega\rangle$ be decreasing in $\mathbb{S}_{[f]}$.
-For each $g\in[f]$ define $q_{g}:=\bigcup_{n<\omega}p_{n,g}$, where $p_{n,g}$ is the $g$-coordinate of $\vec{p}_{n}$ (taking $p_{n,g}=\emptyset$ if $g\notin\operatorname{supp}(\vec{p}_{n})$).
-Since each factor $\mathbb{Q}_{g}$ is $\omega_{1}$-closed by Lemma 4.15 (and each factor $\mathbb{R}_{g}$ by Lemma 4.27), $q_{g}$ is a condition in that factor.
-Let $\vec{q}$ have support $\operatorname{supp}(\vec{q}):=\bigcup_{n<\omega}\operatorname{supp}(\vec{p}_{n})$ and coordinates $q_{g}$ for $g\in\operatorname{supp}(\vec{q})$.
-Because $M\models\mathsf{DC}$ (hence $\mathsf{AC}_{\omega}$), a countable union of countable sets is countable, so $\operatorname{supp}(\vec{q})$ is countable.
-Thus $\vec{q}\in\mathbb{S}_{[f]}$ and $\vec{q}\leq\vec{p}_{n}$ for all $n$.
-(ii) and (iii) now follow by applying Lemma 4.18 to $\mathbb{P}=\mathbb{S}_{[f]}$, using (i) for countable closure. ∎
+Let $\langle\vec{p}_{n}:n<\omega\rangle$ be a descending sequence in $\mathbb{S}_{[f]}$.
+Let $C:=\bigcup_{n<\omega}\operatorname{supp}(\vec{p}_{n})\subseteq[f]$.
+Each $\operatorname{supp}(\vec{p}_{n})$ is countable and $M\models\mathsf{DC}$, so $C$ is countable.
+For each $g\in C$, consider the descending sequence $\langle\vec{p}_{n}(g):n<\omega\rangle$ in the factor $\mathbb{S}_{g}$ (interpreting $\vec{p}_{n}(g)$ as the top condition whenever $g\notin\operatorname{supp}(\vec{p}_{n})$).
+By Lemma 4.13 and Lemma 4.25, each factor $\mathbb{S}_{g}$ is countably closed, so choose $q_{g}\in\mathbb{S}_{g}$ with $q_{g}\leq\vec{p}_{n}(g)$ for all $n$.
+Define $\vec{q}$ by $\operatorname{supp}(\vec{q})=C$ and $\vec{q}(g)=q_{g}$.
+Then $\vec{q}\in\mathbb{S}_{[f]}$ and $\vec{q}\leq\vec{p}_{n}$ for all $n$, proving countable closure.
+Items (2) and (3) follow from Lemma 4.16. ∎
 Remark 5.7 (Limit-stage $\omega_{1}$-completeness).
-At limit stages, $\omega_{1}$-completeness (and normality) of the limit filter is provided by the iteration machinery (API item (d), cf. [5, Thm. 5.5]).
+The countable completeness of the diagonal stage filters is what allows the CSI iteration theorem to handle limit stages.
+Concretely, the limit-step existence theorem for iterations of symmetric systems requires $\omega_{1}$-complete filters (CSI, [5, Theorem 3.13]), and we have imported this requirement as item (c) in Remark 2.4.
 Lemma 5.8 (Well-definedness of the orbit package actions).
@@
 ∎
-Lemma 5.9 (Countable support of package coordinates).
-For every package coordinate $d$ at any successor stage $\alpha+1$ of the iteration, there exists a countable set $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}(d)$, where $\operatorname{Stab}(d)=\{\pi\in\operatorname{Sym}(\omega_{1}):\pi\cdot d=d\}$ is the stabilizer of $d$ as an element of its orbit package.
+Lemma 5.9 (Countable $\mathscr{G}$-support of successor coordinates).
+Let $\alpha$ be a stage and let $d\in\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$ be a new package coordinate at the successor stage $\alpha+1$ (Definition 5.24).
+Then there exists a countable $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}_{\mathscr{G}}(d)$.
 Proof.
-We proceed by induction on stages. At stage 0 there are no package coordinates, so the statement is vacuous.
-At successor stage $\alpha+1$, package coordinates are elements of orbit packages $f]=\{\pi\cdot f:\pi\in\operatorname{Sym}(\omega_{1})\}$, where $f$ is a surjection in the symmetric model $M_{\alpha}$.
-Step 1: The seed surjection $f$ has countable support.
-Since $f\in M_{\alpha}$, there is a hereditarily symmetric $\mathbb{P}_{\alpha}$-name $\dot{f}$ with $f=\dot{f}^{G_{\alpha}}$.
-By the $\Delta$-support equivalence at stage $\alpha$ (which holds by the inductive hypothesis that the Task 1 filter infrastructure is established at stage $\alpha$), there exists a countable admissible $(E_{f},D_{f})$ such that
-$\Delta_{\alpha}^{\uparrow}(E_{f},D_{f})\leq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{f}).$
-Step 2: Bridge from $\mathscr{G}_{\alpha}$-stabilizer to $\operatorname{Sym}(\omega_{1})$-stabilizer.
-We claim $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}(f)$. Let $\pi\in\operatorname{Fix}(E_{f})$.
-By Lemma 5.43 (applied at stage $\alpha$, valid by the inductive hypothesis), there exists a diagonal lift $\widehat{\pi}^{\,D_{f}}\in\mathscr{G}_{\alpha}$ with $\pi^{0}_{\alpha}(\widehat{\pi}^{\,D_{f}})=\pi$ and which fixes every coordinate in $D_{f}$ pointwise.
-By Definition 5.27, this implies $\widehat{\pi}^{\,D_{f}}\in\Delta_{\alpha}^{\uparrow}(E_{f},D_{f})\leq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{f})$.
-Hence
-$\widehat{\pi}^{\,D_{f}}\cdot\dot{f}=\dot{f}\quad\text{and therefore}\quad\widehat{\pi}^{\,D_{f}}\cdot f=f.$
-Let $\widehat{\pi}^{\,\varnothing}\in\mathscr{G}_{\alpha}$ be the standard lift of $\pi$ (the diagonal lift with empty protection set), so $\pi^{0}_{\alpha}(\widehat{\pi}^{\,\varnothing})=\pi$. Set
-$k:=(\widehat{\pi}^{\,\varnothing})^{-1}\cdot\widehat{\pi}^{\,D_{f}}\in\mathscr{G}_{\alpha}.$
-Then $\pi^{0}_{\alpha}(k)=\operatorname{id}$, i.e. $k\in\ker(\pi^{0}_{\alpha})$. By Lemma 5.4, since $\operatorname{tc}(f)\subseteq V(T)\cup\mathrm{Ord}$, we have $k\cdot f=f$. Therefore,
-$f=\widehat{\pi}^{\,D_{f}}\cdot f=\widehat{\pi}^{\,\varnothing}\cdot(k\cdot f)=\widehat{\pi}^{\,\varnothing}\cdot f.$
-Finally, note that if $g,h\in\mathscr{G}_{\alpha}$ satisfy $\pi^{0}_{\alpha}(g)=\pi^{0}_{\alpha}(h)$, then $g^{-1}h\in\ker(\pi^{0}_{\alpha})$, so $g\cdot f=h\cdot f$ by Lemma 5.4.
-In particular, the action of the standard lift $\widehat{\pi}^{\,\varnothing}$ on $f$ depends only on $\pi=\pi^{0}_{\alpha}(\widehat{\pi}^{\,\varnothing})$, hence $\pi\cdot f=\widehat{\pi}^{\,\varnothing}\cdot f=f$.
-Hence
-$\pi\cdot f=\widehat{\pi}^{\,\varnothing}\cdot f=f,$
-and thus $\pi\in\operatorname{Stab}(f)$. This proves $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}(f)$.
-Step 3: Orbit elements inherit countable support.
-For any $g=\sigma\cdot f\in[f]$ (where $\sigma\in\operatorname{Sym}(\omega_{1})$), we have
-$\operatorname{Stab}(g)=\operatorname{Stab}(\sigma\cdot f)=\sigma\cdot\operatorname{Stab}(f)\cdot\sigma^{-1}.$
-Since $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}(f)$, conjugating gives
-$\operatorname{Fix}(\sigma[E_{f}])=\sigma\cdot\operatorname{Fix}(E_{f})\cdot\sigma^{-1}\leq\sigma\cdot\operatorname{Stab}(f)\cdot\sigma^{-1}=\operatorname{Stab}(g).$
-Thus $E_{g}:=\sigma[E_{f}]$ is a countable support for $g$. ∎
+Write $d=\langle\alpha,i,g\rangle$ with $i<m_{\alpha}$ and $g\in[f_{\alpha,i}]$.
+By the inductive hypothesis at stage $\alpha$ (see Remark 5.20), there is a countable admissible pair $(E_{f},D_{f})$ such that $\Delta^{\uparrow}_{\alpha}(E_{f},D_{f})\leq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{f})$, where $\dot{f}$ is the canonical $\operatorname{HS}_{\alpha}$-name for $f$.
+Fix $\pi\in\operatorname{Fix}(E_{f})$.
+By Lemma 5.45, the diagonal lift $\widehat{\pi}^{\,D_{f}}\in\mathscr{G}_{\alpha}$ is defined, and by the choice of $(E_{f},D_{f})$ it fixes $\dot{f}$.
+Let $\widehat{\pi}^{\,\varnothing}\in\mathscr{G}_{\alpha}$ denote the standard lift.
+Then
+$k:=\bigl(\widehat{\pi}^{\,\varnothing}\bigr)^{-1}\cdot\widehat{\pi}^{\,D_{f}}\in\ker(\pi^{0}_{\alpha}),$
+so by Lemma 5.4 the element $k$ fixes $f$ (since $\operatorname{tc}(f)\subseteq V(T)\cup\mathrm{Ord}$).
+As $\widehat{\pi}^{\,D_{f}}$ fixes $f$, it follows that $\widehat{\pi}^{\,\varnothing}$ fixes $f$ as well, i.e.
+$\pi\cdot f=f$.
+Hence $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}_{\mathscr{G}}(f)$.
+Now if $g=\sigma\cdot f$ for some $\sigma\in\mathscr{G}$, then $\operatorname{Fix}(\sigma``E_{f})\leq\operatorname{Stab}_{\mathscr{G}}(g)$.
+Taking $E_{d}:=\sigma``E_{f}$ yields the desired countable support for $d$.
+∎
 Remark 5.10 (No circularity).
-The proof of Lemma 5.9 invokes Lemma 5.43 (existence of diagonal lifts) at stage $\alpha$.
-This does not create a circular dependency: the proof of Lemma 5.43 constructs diagonal lifts by direct recursion on stages using the package-level diagonal machinery of Section 5.1, and does not use Lemma 5.9, Lemma 5.40, or Lemma 5.50.
+The proof uses only stage-$\alpha$ information: the inductive hypothesis that every relevant $\operatorname{HS}_{\alpha}$-name has a countable $\Delta$-support (Remark 5.20), together with the existence of diagonal lifts in $\mathscr{G}_{\alpha}$.
+No stage-$\alpha+1$ properties are used.
 Definition 5.11 (Diagonal-cancellation automorphisms).
@@
 Definition 5.13 (Orbit-package stage filter (diagonal base)).
-Let $\mathscr{G}_{[f]}:=\operatorname{Aut}(\mathbb{S}_{[f]})$.
+Let $\mathscr{G}_{[f]}\leq\operatorname{Aut}(\mathbb{S}_{[f]})$ be the subgroup generated by all diagonal-cancellation maps $\widehat{\pi}^{\,D}$ (equivalently, $\mathscr{G}_{[f]}=\langle\Delta(E,D):(E,D)\text{ admissible}\rangle$).
 For countable $E\subseteq\omega_{1}$ and countable $D\subseteq[f]$ such that $\operatorname{Fix}(E)\leq\operatorname{Stab}(g)$ for all $g\in D$, define
@@
 Let $K_{\alpha+1}:=\ker(\pi^{\alpha}_{\alpha+1})$.
-Moreover, there is a group-theoretic splitting map $\mathrm{spl}:\mathscr{G}_{\alpha}\to\mathscr{G}_{\alpha+1}$ (the the splitting section) such that $\pi^{\alpha}_{\alpha+1}\circ\mathrm{spl}=\mathrm{id}_{\mathscr{G}_{\alpha}}$, and every $g\in\mathscr{G}_{\alpha+1}$ can be written as $g=\mathrm{spl}(h)\cdot k$ with $h=\pi^{\alpha}_{\alpha+1}(g)$ and $k\in K_{\alpha+1}$.
+Moreover, there is a group-theoretic splitting map $\mathrm{spl}:\mathscr{G}_{\alpha}\to\mathscr{G}_{\alpha+1}$ (the splitting section) such that $\pi^{\alpha}_{\alpha+1}\circ\mathrm{spl}=\mathrm{id}_{\mathscr{G}_{\alpha}}$, and every $g\in\mathscr{G}_{\alpha+1}$ can be written as $g=\mathrm{spl}(h)\cdot k$ with $h=\pi^{\alpha}_{\alpha+1}(g)$ and $k\in K_{\alpha+1}$.
 Proof.
@@
 ∎
+Remark 5.22 (Projection maps are coherent).
+For $\gamma<\beta$ we write $\pi^{\gamma}_{\beta}:\mathscr{G}_{\beta}\to\mathscr{G}_{\gamma}$ for the canonical projection homomorphisms of the iteration groups (cf.
+[5, Def.
+3.1]).
+These satisfy coherence: for all $\gamma<\beta<\lambda$,
+$\pi^{\gamma}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{\gamma}_{\lambda}.$
+Definition 5.23 (Stage-0 projection).
+Set $\mathscr{G}_{0}:=\mathscr{G}=\operatorname{Sym}(\omega_{1})$. For each stage $\alpha$ define $\pi^{0}_{\alpha}:\mathscr{G}_{\alpha}\to\mathscr{G}$ by transfinite recursion:
+$\pi^{0}_{0}=\operatorname{id},\qquad\pi^{0}_{\alpha+1}=\pi^{0}_{\alpha}\circ\pi^{\alpha}_{\alpha+1},$
+and for limit $\lambda$ set
+$\pi^{0}_{\lambda}(g):=\pi^{0}_{\alpha}\bigl(\pi^{\alpha}_{\lambda}(g)\bigr)\quad\text{for any }\alpha<\lambda.$
+This is well-defined by Remark 5.22. In particular, for $\beta<\lambda$, $\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}$.
 Definition 5.24 (Package coordinate sets).
@@
   • $\mathrm{Coords}_{0}:=\varnothing$.
-  • At a successor stage $\alpha+1$, write $\mathbb{P}_{\alpha+1}\cong\mathbb{P}_{\alpha}*\dot{\mathbb{Q}}_{\alpha}$.
+  • At a successor stage $\alpha+1$, write $\mathbb{P}_{\alpha+1}\cong\mathbb{P}_{\alpha}*\dot{\mathbb{Q}}_{\alpha}$, where $\dot{\mathbb{Q}}_{\alpha}$ denotes the stage iterand (a finite product of orbit packages) and should not be confused with the localized $\mathsf{PP}$-package forcing $\mathbb{Q}_{f}$.
 In this paper, $\dot{\mathbb{Q}}_{\alpha}$ is (forced to be) a finite product of orbit packages
@@
   • At a limit stage $\lambda$, set $\mathrm{Coords}_{\lambda}:=\bigcup_{\alpha<\lambda}\mathrm{Coords}_{\alpha}$. We also write $\mathrm{Coords}_{<\lambda}:=\mathrm{Coords}_{\lambda}$.
-When working at a fixed successor stage $\alpha+1$, we suppress the stage tag and write $\langle i,g\rangle$ for $\langle\alpha,i,g\rangle$ (so $\langle i,g\rangle\in\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$” means $i<n_{\alpha}$ and $g\in[f_{\alpha,i}]$).
-We let $\operatorname{Sym}(\omega_{1})$ act on new coordinates by
+When working at a fixed successor stage $\alpha+1$, we suppress the stage tag and write $\langle i,g\rangle$ for $\langle\alpha,i,g\rangle$ (so $\langle i,g\rangle\in\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$” means $i<m_{\alpha}$ and $g\in[f_{\alpha,i}]$).
+We let $\mathscr{G}=\operatorname{Sym}(\omega_{1})$ act on new coordinates by
 $\pi\cdot\langle\alpha,i,g\rangle:=\langle\alpha,i,\pi\cdot g\rangle,$
@@
 Fix a successor stage $\alpha+1$.
-Let $\pi\in\operatorname{Sym}(\omega_{1})$ and let $D^{\mathrm{new}}\subseteq\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$ be a set of package coordinates newly introduced at stage $\alpha+1$.
-Assume $\pi$ fixes each $d\in D^{\mathrm{new}}$ under the induced action on tagged coordinates, i.e., if $d=\langle\alpha,i,g\rangle$ then $\pi\cdot d=d$ equivalently $\pi\cdot g=g$.
+Let $\pi\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$ and let $D^{\mathrm{new}}\subseteq\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$ be a set of package coordinates newly introduced at stage $\alpha+1$.
+Assume $\pi$ fixes each $d\in D^{\mathrm{new}}$ under the induced action on tagged coordinates, i.e., if $d=\langle\alpha,i,g\rangle$ then $\pi\cdot d=d$, equivalently $\pi\cdot g=g$.
 (Here $\pi\cdot\langle\alpha,i,g\rangle:=\langle\alpha,i,\pi\cdot g\rangle$ using the orbit action on $f_{\alpha,i}]$.)
@@
     $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})\in K_{\alpha+1}=\ker(\pi^{\alpha}_{\alpha+1})$.
-Moreover, for fixed $D^{\mathrm{new}}$, the map $\pi\mapsto\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ is a group homomorphism on its domain $\{\pi\in\operatorname{Sym}(\omega_{1}):\pi\cdot d=d\ \forall d\in D^{\mathrm{new}}\}$ (Here the action on new coordinates is $\pi\cdot\langle\alpha,i,g\rangle:=\langle\alpha,i,\pi\cdot g\rangle$.)
+Moreover, for fixed $D^{\mathrm{new}}$, the map
+$\pi\longmapsto\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$
+is a group homomorphism on its domain
+$\{\pi\in\mathscr{G}=\operatorname{Sym}(\omega_{1}):\forall d\in D^{\mathrm{new}}\ (\pi\cdot d=d)\},$
+where the action on new coordinates is the one fixed in Definition 5.24.
 Proof.
@@
 Remark 5.30.
-If $(E,D)$ is globally admissible and $\pi\in\operatorname{Fix}(E)$, then $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
-Indeed, $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D})=\pi\in\operatorname{Fix}(E)$ and $\widehat{\pi}^{\,D}$ fixes every coordinate in $D$ pointwise.
+Let $(E,D)$ be globally admissible at stage $\lambda$ and let $\pi\in\operatorname{Fix}(E)$.
+The diagonal lift $\widehat{\pi}^{\,D}$ is defined as an element of $\operatorname{Aut}(\mathbb{P}_{\lambda})$.
+Once we verify that $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ (Lemma 5.45), it follows immediately from the defining clause (A) that $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
 Remark 5.31 (Characterization vs. generation).
@@
 This “characterization” definition makes monotonicity immediate (Lemma 5.36).
-In Lemma 5.44, we show that diagonal lifts $\widehat{\pi}^{\,D^{\prime}}$ (for $\pi\in\operatorname{Fix}(E)$ and $D^{\prime}\supseteq D$) lie in $\Delta_{\lambda}^{\uparrow}(E,D)$—this is immediate from the characterization.
+In Lemma 5.46, we show that once diagonal lifts $\widehat{\pi}^{\,D^{\prime}}$ are known to belong to $\mathscr{G}_{\lambda}$ (Lemma 5.45), then for $\pi\in\operatorname{Fix}(E)$ and $D^{\prime}\supseteq D$ we have $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$; this is then immediate from the characterization.
 We do not need to prove that $\Delta_{\lambda}^{\uparrow}(E,D)$ is generated by such diagonal lifts; the characterization suffices for all filter-theoretic purposes.
 Lemma 5.32 (Group action on coordinates).
-Let $g\in\mathscr{G}_{\lambda}$. Then:
+Let $k\in\mathscr{G}_{\lambda}$. Then:
   1. (i)
-    The stage-0 projection $\pi^{0}_{\lambda}:\mathscr{G}_{\lambda}\to\operatorname{Sym}(\omega_{1})$ is a group homomorphism.
+    The stage-0 projection $\pi^{0}_{\lambda}:\mathscr{G}_{\lambda}\to\mathscr{G}$ is a group homomorphism.
   2. (ii)
-    The element $g$ induces a bijection on the set of package coordinates at each stage $\beta<\lambda$. In particular, if $D$ is a countable set of package coordinates, then $g\cdot D$ is countable.
+    The element $k$ induces a bijection on the set of package coordinates at each stage $\beta<\lambda$ (in the sense of Definition 5.24).
+In particular, if $D$ is a countable set of package coordinates, then $k\cdot D$ is countable.
   3. (iii)
-    For $E\subseteq\omega_{1}$, define $g\cdot E:=\pi^{0}_{\lambda}(g)[E]$. Then $g\cdot E$ is countable whenever $E$ is countable.
+    For $E\subseteq\omega_{1}$, define $k\cdot E:=\pi^{0}_{\lambda}(k)[E]$. Then $k\cdot E$ is countable whenever $E$ is countable.
 Proof.
-(i) This is part of the inverse/direct system structure of the iteration groups (cf. [5, Def. 3.1]) and the fact that $\pi^{0}_{\lambda}$ is defined stagewise and respects composition.
-(ii) We prove by induction on stages that each $g\in\mathscr{G}_{\beta}$ determines a bijection of $\mathrm{Coords}_{\beta}$ (in the sense of Definition 5.22), and hence sends countable sets of coordinates to countable sets.
+(i) This is immediate from Definition 5.23 together with the coherence of the projections (Remark 5.22); cf. also [5, Def. 3.1].
+(ii) We prove by induction on stages that each $k\in\mathscr{G}_{\beta}$ determines a bijection of $\mathrm{Coords}_{\beta}$, and hence sends countable sets of coordinates to countable sets.
 At stage 0 there are no package coordinates.
-At a successor stage $\beta+1$, write $g=\mathrm{st}(h)\cdot k$ as in Lemma 5.21, where $h=\pi^{\beta}_{\beta+1}(g)\in\mathscr{G}_{\beta}$ and $k\in K_{\beta+1}$ acts trivially on the $\mathbb{P}_{\beta}$-part.
+At a successor stage $\beta+1$, write $k=\mathrm{st}(h)\cdot u$ as in Lemma 5.21, where $h=\pi^{\beta}_{\beta+1}(k)\in\mathscr{G}_{\beta}$ and $u\in K_{\beta+1}$ acts trivially on the $\mathbb{P}_{\beta}$-part.
 By the inductive hypothesis, $h$ induces a bijection of $\mathrm{Coords}_{\beta}$.
 The standard lift $\mathrm{st}(h)$ extends this action and fixes every element of $\mathrm{NewCoords}_{\beta+1}$.
-On the other hand, $k$ fixes $\mathrm{Coords}_{\beta}$ pointwise and acts on the successor forcing $\dot{\mathbb{Q}}_{\beta}$ by compositions of package automorphisms (Lemma 5.8) and diagonal-cancellation automorphisms (Definition 5.11), each of which induces a bijection on the new coordinate index set (and in particular cannot collapse distinct coordinates, by the admissibility hypothesis built into their definition).
-Therefore $g$ induces a bijection on $\mathrm{Coords}_{\beta+1}=\mathrm{Coords}_{\beta}\dot{\cup}\mathrm{NewCoords}_{\beta+1}$.
-At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are (by definition) coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf.
+On the other hand, $u$ fixes $\mathrm{Coords}_{\beta}$ pointwise and acts on the successor iterand $\dot{\mathbb{Q}}_{\beta}$ by compositions of package automorphisms (Lemma 5.8) and diagonal-cancellation automorphisms (Definition 5.11), each of which induces a bijection on the new coordinate index set.
+Therefore $k$ induces a bijection on $\mathrm{Coords}_{\beta+1}=\mathrm{Coords}_{\beta}\dot{\cup}\mathrm{NewCoords}_{\beta+1}$.
+At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf.
 [5, Def.
@@
 The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection.
-Consequently, if $D$ is countable then so is its image under the induced bijection.
-(iii) Immediate: $\pi^{0}_{\lambda}(g)\in\operatorname{Sym}(\omega_{1})$ is a bijection and hence preserves countability. ∎
+Consequently, if $D$ is countable then so is $k\cdot D$.
+(iii) Immediate, since $\pi^{0}_{\lambda}(k)\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$ is a bijection and hence preserves countability. ∎
 Lemma 5.33 (Iteration diagonal lifts are automorphisms).
@@
 The group homomorphism property $\widehat{\pi}^{\,D}\circ\widehat{\sigma}^{\,D}=\widehat{\pi\sigma}^{\,D}$ follows coordinatewise: on $D$ both sides are the identity; on the Cohen base and on coordinates outside $D$, both sides agree with the standard action of $\pi\sigma$.
+Injectivity is immediate from (i): if $\widehat{\pi}^{\,D}=\widehat{\sigma}^{\,D}$ then their restrictions to the Cohen base agree, hence $\pi=\sigma$. ∎
+Corollary 5.34 (Stage-0 projection controls the Cohen reals).
+Let $g\in\mathscr{G}_{\lambda}$ and put $\pi:=\pi^{0}_{\lambda}(g)\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$. Then for each $\alpha<\omega_{1}$,
+$g\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)},$
+where $\dot{c}_{\alpha}$ is the canonical $\mathbb{P}_{0}$-name for the $\alpha$-th Cohen real.
+In particular, if $\pi=(\beta\ \gamma)$ is a transposition with $\beta\neq\gamma$, then $g$ swaps $c_{\beta}\leftrightarrow c_{\gamma}$ in the generic extension.
+Proof.
+View $\dot{c}_{\alpha}$ as a $\mathbb{P}_{\lambda}$-name via the canonical inclusion of $\mathbb{P}_{0}$-names into $\mathbb{P}_{\lambda}$-names. Recall
+$\dot{c}_{\alpha}:=\{\langle\check{n},p\rangle:p(\alpha,n)=1\}.$
+Since $\pi^{0}_{\lambda}(g)=\pi$, the restriction of $g$ to the Cohen base $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ is the usual $\pi$-action.
+Therefore $g\cdot\dot{c}_{\alpha}=\pi\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)}$.
 ∎
-Corollary 5.32 (Diagonal lifts swap Cohen reals).
-Let $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ be a diagonal lift with $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D})=\pi$. Then for each $\alpha<\omega_{1}$:
-$\widehat{\pi}^{\,D}\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)},$
-where $\dot{c}_{\alpha}$ is the canonical $\mathbb{P}_{0}$-name for the $\alpha$-th Cohen real.
-In particular, if $\pi=(\beta\ \gamma)$ is a transposition with $\beta\neq\gamma$, then $\widehat{\pi}^{\,D}$ swaps $c_{\beta}\leftrightarrow c_{\gamma}$ in the generic extension.
-Proof.
-View $\dot{c}_{\alpha}$ as a $\mathbb{P}_{\lambda}$-name via the canonical inclusion of $\mathbb{P}_{0}$-names into $\mathbb{P}_{\lambda}$-names.
-By Lemma 5.31(i), the diagonal lift $\widehat{\pi}^{\,D}$ restricts to $\pi$ on the Cohen base $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$.
-The canonical name $\dot{c}_{\alpha}$ is defined by $\dot{c}_{\alpha}=\{(\check{n},p):p(\alpha,n)=1\}$.
-The action of $\pi$ on $\mathbb{P}_{0}$ is $(\pi\cdot p)(\pi(\alpha),n)=p(\alpha,n)$, which induces $\pi\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)}$.
-Since $\widehat{\pi}^{\,D}$ restricts to $\pi$ on stage 0, we have $\widehat{\pi}^{\,D}\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)}$.
-∎
-Remark 5.33 (Standard lifts as empty diagonal lifts).
-The standard lift $\widehat{\pi}^{\,\emptyset}$ (Definition 5.27 with $D=\emptyset$) acts as $\pi$ on all coordinates, recovering the usual lifted action of $\operatorname{Sym}(\omega_{1})$ on the iteration.
-Thus the pushforward subgroups $\operatorname{Fix}(E)\leq\mathscr{G}_{\lambda}$ (acting via standard lifts) coincide with $\Delta_{\lambda}(E,\emptyset)$, and consequently $\Delta_{\lambda}^{\uparrow}(E,\emptyset)=\operatorname{Fix}(E)^{\uparrow}$.
-This identification is crucial: all filter generators at limit stages—both Cohen-type and package-type—are of the form $\Delta_{\lambda}^{\uparrow}(E_{i},D_{i})$, with $D_{i}=\emptyset$ for Cohen generators and $D_{i}=\{f_{i}\}$ for package generators.
-The monotonicity lemma (Lemma 5.34) therefore applies uniformly to combine them.
+Remark 5.35 (Standard lifts and Cohen-type diagonal groups).
+The standard lift $\widehat{\pi}^{\,\emptyset}$ from Definition 5.29(B) is the diagonal lift with no protected coordinates, hence it acts as $\pi$ on the Cohen base and as the standard $\pi$-action on every package coordinate.
+For a countable $E\subseteq\omega_{1}$, the subgroup of standard lifts
+$\operatorname{Fix}(E)^{\uparrow}:=\{\widehat{\pi}^{\,\emptyset}:\pi\in\operatorname{Fix}(E)\}\leq\mathscr{G}_{\lambda}$
+is contained in the characterized diagonal group
+$\Delta_{\lambda}^{\uparrow}(E,\emptyset)=\{g\in\mathscr{G}_{\lambda}:\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)\},$
+and in general this inclusion may be proper.
+In the limit-stage diagonal filters, both Cohen-type constraints and package-coordinate constraints are handled uniformly using groups of the form $\Delta_{\lambda}^{\uparrow}(E,D)$ for countable parameters $E$ and countable sets of coordinates $D\subseteq\mathrm{Coords}_{<\lambda}$.
 Lemma 5.36 (Monotonicity for iteration-level diagonal groups).
@@
 Proof.
-Let $g\in\Delta_{\lambda}^{\uparrow}(E_{1},D_{1})$. By the characterization definition:
-  • $\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E_{1})\subseteq\operatorname{Fix}(E_{0})$ (since $E_{0}\subseteq E_{1}$ implies $\operatorname{Fix}(E_{1})\leq\operatorname{Fix}(E_{0})$).
-  • $g$ fixes every coordinate in $D_{1}$ pointwise, hence fixes every coordinate in $D_{0}\subseteq D_{1}$ pointwise.
-Thus $g\in\Delta_{\lambda}^{\uparrow}(E_{0},D_{0})$. ∎
-Corollary 5.35 (Intersection of $\Delta^{\uparrow}$-groups).
-For any family $\{(E_{i},D_{i}):i\in I\}$ (with $I$ finite or countable):
+Let $k\in\Delta_{\lambda}^{\uparrow}(E_{1},D_{1})$. By Definition 5.29(A):
+  • $\pi^{0}_{\lambda}(k)\in\operatorname{Fix}(E_{1})\subseteq\operatorname{Fix}(E_{0})$ (since $E_{0}\subseteq E_{1}$ implies $\operatorname{Fix}(E_{1})\leq\operatorname{Fix}(E_{0})$);
+  • $k$ fixes every coordinate in $D_{1}$ pointwise, hence fixes every coordinate in $D_{0}\subseteq D_{1}$ pointwise.
+Thus $k\in\Delta_{\lambda}^{\uparrow}(E_{0},D_{0})$. ∎
+Corollary 5.37 (Intersections contain a uniform diagonal subgroup).
+Let $\{(E_{i},D_{i}):i\in I\}$ be a family with $I$ finite or countable, and assume each $E_{i}$ and each $D_{i}$ is countable. Then
 $\bigcap_{i\in I}\Delta_{\lambda}^{\uparrow}(E_{i},D_{i})\supseteq\Delta_{\lambda}^{\uparrow}\left(\bigcup_{i\in I}E_{i},\ \bigcup_{i\in I}D_{i}\right).$
+In particular, under $M\models\mathsf{DC}$ the unions on the right are countable, so the right-hand group is again of the form used to generate the diagonal stage filters.
 Proof.
@@
 Lemma 5.38 (Conjugation of iteration-level diagonal groups).
-Let $(E,D)$ be globally admissible at stage $\lambda$. For $\sigma\in\operatorname{Sym}(\omega_{1})$:
-$\widehat{\sigma}^{\,\emptyset}\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot(\widehat{\sigma}^{\,\emptyset})^{-1}=\Delta_{\lambda}^{\uparrow}(\sigma\cdot E,\sigma\cdot D).$
-In particular, conjugates of diagonal generators remain diagonal generators (with shifted parameters).
+Let $E\subseteq\omega_{1}$ and let $D\subseteq\mathrm{Coords}_{<\lambda}$. For $\sigma\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$,
+$\widehat{\sigma}^{\,\emptyset}\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot(\widehat{\sigma}^{\,\emptyset})^{-1}=\Delta_{\lambda}^{\uparrow}\bigl(\sigma[E],\ \widehat{\sigma}^{\,\emptyset}\cdot D\bigr),$
+where $\widehat{\sigma}^{\,\emptyset}\cdot D:=\{\widehat{\sigma}^{\,\emptyset}\cdot d:d\in D\}$.
 Proof.
-The proof is identical to Lemma 5.18, now applied at the iteration level. ∎
+Let $k\in\Delta_{\lambda}^{\uparrow}(E,D)$ and set $h:=\widehat{\sigma}^{\,\emptyset}\,k\,(\widehat{\sigma}^{\,\emptyset})^{-1}$. By Lemma 5.32(i), $\pi^{0}_{\lambda}$ is a homomorphism, hence
+$\pi^{0}_{\lambda}(h)=\sigma\,\pi^{0}_{\lambda}(k)\,\sigma^{-1}.$
+Since $\pi^{0}_{\lambda}(k)\in\operatorname{Fix}(E)$, the conjugate fixes $\sigma[E]$ pointwise, so $\pi^{0}_{\lambda}(h)\in\operatorname{Fix}(\sigma[E])$.
+Next, for $d\in D$ we have $k\cdot d=d$, hence $h\cdot(\widehat{\sigma}^{\,\emptyset}\cdot d)=\widehat{\sigma}^{\,\emptyset}\cdot d$.
+Therefore $h$ fixes every coordinate in $\widehat{\sigma}^{\,\emptyset}\cdot D$ pointwise.
+By Definition 5.29(A), $h\in\Delta_{\lambda}^{\uparrow}(\sigma[E],\widehat{\sigma}^{\,\emptyset}\cdot D)$.
+The reverse inclusion follows by conjugating with $(\widehat{\sigma}^{\,\emptyset})^{-1}$.
+∎
 Remark 5.39 (Admissibility under conjugation up to enlarging $E$).
-Let $\lambda$ be a stage and $(E,D)$ be globally admissible at stage $\lambda$.
-For any $g\in\mathcal{G}_{\lambda}$, there is a countable $E^{\ast}\supseteq g\cdot E$ such that $(E^{\ast},\,g\cdot D)$ is globally admissible.
-Consequently $\Delta_{\lambda}^{\uparrow}(E^{\ast},\,g\cdot D)\subseteq g\,\Delta_{\lambda}^{\uparrow}(E,D)\,g^{-1}$.
+Let $\lambda$ be a stage and let $(E,D)$ be globally admissible at stage $\lambda$.
+For any $k\in\mathscr{G}_{\lambda}$, there is a countable $E^{\ast}\supseteq k\cdot E$ such that $(E^{\ast},\,k\cdot D)$ is globally admissible.
+Consequently,
+$\Delta_{\lambda}^{\uparrow}(E^{\ast},\,k\cdot D)\subseteq k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}.$
+Indeed, by Lemma 5.40 we have $k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}=\Delta_{\lambda}^{\uparrow}(k\cdot E,k\cdot D)$, and since $k\cdot E\subseteq E^{\ast}$ monotonicity (Lemma 5.36) yields $\Delta_{\lambda}^{\uparrow}(E^{\ast},k\cdot D)\leq\Delta_{\lambda}^{\uparrow}(k\cdot E,k\cdot D)$.
 Lemma 5.40 (General conjugation of $\Delta^{\uparrow}$-groups).
-For any $g\in\mathscr{G}_{\lambda}$, any $E\subseteq\omega_{1}$, and any set $D$ of package coordinates:
-$g\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot g^{-1}=\Delta_{\lambda}^{\uparrow}(g\cdot E,g\cdot D)$
-where $g\cdot E:=\pi^{0}_{\lambda}(g)[E]$ and $g\cdot D:=\{g\cdot d:d\in D\}$ is the image of $D$ under the action of $g$ on package coordinates.
+Let $k\in\mathscr{G}_{\lambda}$, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{<\lambda}$. Then
+$k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}=\Delta_{\lambda}^{\uparrow}(k\cdot E,k\cdot D),$
+where $k\cdot E:=\pi^{0}_{\lambda}(k)[E]$ and $k\cdot D:=\{k\cdot d:d\in D\}$ (using the $\mathscr{G}_{\lambda}$-action on coordinates from Lemma 5.32).
 Proof.
-Fix $k\in\mathscr{G}_{\lambda}$ and write $\sigma:=\pi^{0}_{\lambda}(k)\in\operatorname{Sym}(\omega_{1})$.
-Let $g\in\Delta^{\uparrow}_{\lambda}(E,D)$, and put $\tau:=\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)$.
-Stage-0 part. Since $\pi^{0}_{\lambda}:\mathscr{G}_{\lambda}\to\operatorname{Sym}(\omega_{1})$ is a homomorphism, we have
-$\pi^{0}_{\lambda}(kgk^{-1})=\sigma\tau\sigma^{-1}\in\operatorname{Fix}(\sigma[E])=\operatorname{Fix}(\pi^{0}_{\lambda}(k)[E]).$
+Fix $k\in\mathscr{G}_{\lambda}$ and put $\sigma:=\pi^{0}_{\lambda}(k)\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$.
+Let $h\in\Delta^{\uparrow}_{\lambda}(E,D)$, and put $\tau:=\pi^{0}_{\lambda}(h)\in\operatorname{Fix}(E)$.
+Stage-0 part. By Lemma 5.32(i), $\pi^{0}_{\lambda}$ is a homomorphism, hence
+$\pi^{0}_{\lambda}(khk^{-1})=\sigma\tau\sigma^{-1}\in\operatorname{Fix}(\sigma[E])=\operatorname{Fix}(\pi^{0}_{\lambda}(k)[E]).$
 Coordinate part.
 Let $d\in D$.
-By definition of $\Delta^{\uparrow}_{\lambda}(E,D)$, $g$ fixes $d$ pointwise.
-Using the coordinate action $\cdot$ of $\mathscr{G}_{\lambda}$ on $\mathrm{Coords}_{<\lambda}$ (Lemma 5.30), we compute
-$(kgk^{-1})\cdot(k\cdot d)=k\cdot\bigl(g\cdot(k^{-1}\cdot(k\cdot d))\bigr)=k\cdot(g\cdot d)=k\cdot d,$
-so $kgk^{-1}$ fixes every coordinate in $k\cdot D$ pointwise.
-Thus $kgk^{-1}\in\Delta^{\uparrow}_{\lambda}(\pi^{0}_{\lambda}(k)[E],k\cdot D)$. The reverse inclusion follows by conjugating with $k^{-1}$, so equality holds. ∎
+By definition of $\Delta^{\uparrow}_{\lambda}(E,D)$, $h$ fixes $d$ pointwise.
+Using the coordinate action of $\mathscr{G}_{\lambda}$ on $\mathrm{Coords}_{<\lambda}$ (Lemma 5.32),
+$(khk^{-1})\cdot(k\cdot d)=k\cdot\bigl(h\cdot(k^{-1}\cdot(k\cdot d))\bigr)=k\cdot(h\cdot d)=k\cdot d,$
+so $khk^{-1}$ fixes every coordinate in $k\cdot D$ pointwise.
+Thus $khk^{-1}\in\Delta^{\uparrow}_{\lambda}(k\cdot E,k\cdot D)$, proving $k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}\subseteq\Delta^{\uparrow}_{\lambda}(k\cdot E,k\cdot D)$.
+The reverse inclusion follows by conjugating with $k^{-1}$.
+∎
 Corollary 5.41 (Admissibility preserved under unions).
@@
 Lemma 5.43 (Projection of $\Delta^{\uparrow}$-groups).
-Let $\beta<\lambda$ and let $D\subseteq\mathrm{Coords}_{\beta}$ (i.e., $D$ only uses coordinates introduced below stage $\beta$). Then:
+Let $\beta<\lambda$, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{\beta}$ (i.e., $D$ only uses coordinates introduced below stage $\beta$). Then
 $\pi^{\beta}_{\lambda}\big[\Delta_{\lambda}^{\uparrow}(E,D)\big]\subseteq\Delta_{\beta}^{\uparrow}(E,D).$
 Proof.
-Let $g\in\Delta_{\lambda}^{\uparrow}(E,D)$. Then:
-  • $\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)$, and since $\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}$, we have $\pi^{0}_{\beta}(\pi^{\beta}_{\lambda}(g))=\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)$.
-  • $g$ fixes every coordinate in $D$ pointwise.
-Since $D\subseteq\mathrm{Coords}_{\beta}$, these coordinates are visible at stage $\beta$, and the projection $\pi^{\beta}_{\lambda}(g)$ also fixes them pointwise.
-Hence $\pi^{\beta}_{\lambda}(g)\in\Delta_{\beta}^{\uparrow}(E,D)$. ∎
+Let $k\in\Delta_{\lambda}^{\uparrow}(E,D)$. Then:
+  • $\pi^{0}_{\lambda}(k)\in\operatorname{Fix}(E)$, and by Definition 5.23 (using coherence from Remark 5.22) we have $\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}$. Hence
+    $\pi^{0}_{\beta}(\pi^{\beta}_{\lambda}(k))=\pi^{0}_{\lambda}(k)\in\operatorname{Fix}(E).$
+  • $k$ fixes every coordinate in $D$ pointwise.
+Since $D\subseteq\mathrm{Coords}_{\beta}$, these coordinates are already present at stage $\beta$, and the projection $\pi^{\beta}_{\lambda}(k)$ acts on $\mathrm{Coords}_{\beta}$ as the restriction of $k$ (cf.
+Lemma 5.32(ii)), so $\pi^{\beta}_{\lambda}(k)$ fixes every coordinate in $D$ pointwise.
+Hence $\pi^{\beta}_{\lambda}(k)\in\Delta_{\beta}^{\uparrow}(E,D)$ by Definition 5.29(A). ∎
 Lemma 5.44 (Pullback contains $\Delta^{\uparrow}$).
-Let $\beta<\lambda$ and let $D\subseteq\mathrm{Coords}_{\beta}$.
+Let $\beta<\lambda$, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{\beta}$.
 If $H\in\tilde{\mathscr{F}}^{*}_{\beta}$ contains $\Delta_{\beta}^{\uparrow}(E,D)$, then $(\pi^{\beta}_{\lambda})^{-1}[H]$ contains $\Delta_{\lambda}^{\uparrow}(E,D)$.
 Proof.
-Let $g\in\Delta_{\lambda}^{\uparrow}(E,D)$. By Lemma 5.41, $\pi^{\beta}_{\lambda}(g)\in\Delta_{\beta}^{\uparrow}(E,D)\subseteq H$. Hence $g\in(\pi^{\beta}_{\lambda})^{-1}[H]$. ∎
+Let $k\in\Delta_{\lambda}^{\uparrow}(E,D)$. By Lemma 5.43, $\pi^{\beta}_{\lambda}(k)\in\Delta_{\beta}^{\uparrow}(E,D)\subseteq H$. Hence $k\in(\pi^{\beta}_{\lambda})^{-1}[H]$. ∎
 Lemma 5.45 (Iteration-level diagonal lifts lie in $\mathscr{G}_{\lambda}$).
@@
 Proof.
-For each $\beta\leq\lambda$, define $D_{\beta}:=D\cap\mathrm{Coords}_{\beta}$, the set of protected coordinates at stages $\leq\beta$.
-Key observation from admissibility: Since $(E,D)$ is globally admissible and $\pi\in\operatorname{Fix}(E)$, we have $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for all $d\in D$, hence $\pi\cdot d=d$ for all $d\in D$.
-In particular, for each successor $\beta+1\leq\lambda$, the set $D_{\beta+1}\setminus D_{\beta}\subseteq D$ is fixed pointwise by $\pi$.
-This ensures the successor diagonal kernel element $\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})$ is well-defined (Definition 5.23).
-We construct by transfinite recursion a coherent sequence $\langle g_{\beta}:\beta\leq\lambda\rangle$ with $g_{\beta}\in\mathscr{G}_{\beta}$ such that $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$ as automorphisms of $\mathbb{P}_{\beta}$.
-More precisely, we maintain the inductive invariant that for all $\gamma<\beta\leq\lambda$, $\pi^{\gamma}_{\beta}(g_{\beta})=g_{\gamma}$.
-In particular, at a limit stage $\eta$ the family $\langle g_{\beta}:\beta<\eta\rangle$ is coherent, so a direct-limit element exists.
-Stage 0: Define $g_{0}:=\pi\in\operatorname{Sym}(\omega_{1})=\mathscr{G}_{0}$.
-Since $\mathrm{Coords}_{0}=\varnothing$, we have $D_{0}=\varnothing$, so $\widehat{\pi}^{\,D_{0}}=\widehat{\pi}^{\,\varnothing}$, which is the standard $\pi$-action on the Cohen forcing (Definition 5.27, stage-0 clause).
-Hence $g_{0}=\widehat{\pi}^{\,D_{0}}$.
-Successor $\beta+1$: Assume $g_{\beta}\in\mathscr{G}_{\beta}$ with $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$. Define
-$g_{\beta+1}:=\mathrm{spl}(g_{\beta})\cdot\kappa_{\beta+1}\bigl(\pi,\,D_{\beta+1}\setminus D_{\beta}\bigr),$
-where $\mathrm{spl}$ is the splitting section from Lemma 5.21 and $\kappa_{\beta+1}$ is the successor diagonal kernel element from Definition 5.23 (applicable by the key observation above).
-By Lemma 5.21, $\mathrm{spl}(g_{\beta})\in\mathscr{G}_{\beta+1}$.
-By Lemma 5.25(i), $\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})\in\mathscr{G}_{\beta+1}$.
-Hence $g_{\beta+1}\in\mathscr{G}_{\beta+1}$.
-We verify $g_{\beta+1}=\widehat{\pi}^{\,D_{\beta+1}}$ by comparing actions (cf. Definition 5.27):
-  • On the $\mathbb{P}_{\beta}$-part: $\mathrm{spl}(g_{\beta})$ acts as $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$, and $\kappa_{\beta+1}$ acts as identity.
-The combined action is $\widehat{\pi}^{\,D_{\beta}}$, which agrees with $\widehat{\pi}^{\,D_{\beta+1}}$ restricted to stages $\leq\beta$ (since $D_{\beta+1}\cap\mathrm{Coords}_{\beta}=D_{\beta}$).
-  • On new coordinates at stage $\beta+1$: $\mathrm{spl}(g_{\beta})$ acts trivially. The kernel element $\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})$:
-    • acts as $\pi$ on $d\notin D_{\beta+1}\setminus D_{\beta}$, i.e., on new coordinates not in $D$;
-    • acts as identity on $d\in D_{\beta+1}\setminus D_{\beta}$, i.e., on new coordinates in $D$.
-    This matches Definition 5.27: “act as $\pi$ if $d\notin D$, identity if $d\in D$.”
-Limit $\eta$: Define $g_{\eta}$ to be the direct-limit element of $\mathscr{G}_{\eta}$ determined by the family $\langle g_{\beta}:\beta<\eta\rangle$, which is coherent by the inductive invariant above.
+For each $\beta\leq\lambda$ let $D_{\beta}:=D\cap\mathrm{Coords}_{\beta}$.
+By Definition 5.28, $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for every $d\in D$, hence $\pi\cdot d=d$ for all $d\in D$.
+In particular, for each successor $\beta+1\leq\lambda$ the set
+$D^{\mathrm{new}}_{\beta+1}:=D_{\beta+1}\setminus D_{\beta}$
+is pointwise fixed by $\pi$, so $\kappa_{\beta+1}(\pi,D^{\mathrm{new}}_{\beta+1})$ is well-defined (Definition 5.25).
+We build by transfinite recursion a coherent sequence $\langle k_{\beta}:\beta\leq\lambda\rangle$ with $k_{\beta}\in\mathscr{G}_{\beta}$ such that $k_{\beta}=\widehat{\pi}^{\,D_{\beta}}$ as automorphisms of $\mathbb{P}_{\beta}$, and $\pi^{\gamma}_{\beta}(k_{\beta})=k_{\gamma}$ for all $\gamma<\beta\leq\lambda$.
+Stage 0. Set $k_{0}:=\pi\in\mathscr{G}_{0}=\mathscr{G}$. Since $D_{0}=\varnothing$, this agrees with $\widehat{\pi}^{\,D_{0}}=\widehat{\pi}^{\,\emptyset}$ on $\mathbb{P}_{0}$.
+Successor $\beta+1$. Assume $k_{\beta}\in\mathscr{G}_{\beta}$ with $k_{\beta}=\widehat{\pi}^{\,D_{\beta}}$. Define
+$k_{\beta+1}:=\mathrm{spl}(k_{\beta})\cdot\kappa_{\beta+1}\!\left(\pi,\,D^{\mathrm{new}}_{\beta+1}\right),$
+where $\mathrm{spl}$ is the splitting section from Lemma 5.21.
+By Lemma 5.21, $\mathrm{spl}(k_{\beta})\in\mathscr{G}_{\beta+1}$, and by Lemma 5.27(i) we have $\kappa_{\beta+1}(\pi,D^{\mathrm{new}}_{\beta+1})\in\mathscr{G}_{\beta+1}$, hence $k_{\beta+1}\in\mathscr{G}_{\beta+1}$.
+Moreover, $k_{\beta+1}$ agrees with $\widehat{\pi}^{\,D_{\beta+1}}$ on $\mathbb{P}_{\beta+1}$: on the $\mathbb{P}_{\beta}$-part, $\mathrm{spl}(k_{\beta})$ acts as $k_{\beta}=\widehat{\pi}^{\,D_{\beta}}$ and the kernel element acts trivially; on the new coordinates at stage $\beta+1$, $\mathrm{spl}(k_{\beta})$ acts trivially and $\kappa_{\beta+1}(\pi,D^{\mathrm{new}}_{\beta+1})$ acts as $\pi$ outside $D^{\mathrm{new}}_{\beta+1}$ and as the identity on $D^{\mathrm{new}}_{\beta+1}$, which is exactly the defining clause of $\widehat{\pi}^{\,D_{\beta+1}}$ (Definition 5.29).
+Finally,
+$\pi^{\beta}_{\beta+1}(k_{\beta+1})=\pi^{\beta}_{\beta+1}(\mathrm{spl}(k_{\beta}))\cdot\pi^{\beta}_{\beta+1}(\kappa_{\beta+1}(\ldots))=k_{\beta}\cdot\mathrm{id}=k_{\beta},$
+using $\pi^{\beta}_{\beta+1}\circ\mathrm{spl}=\mathrm{id}$ (Lemma 5.21) and $\kappa_{\beta+1}(\ldots)\in K_{\beta+1}=\ker(\pi^{\beta}_{\beta+1})$ (Lemma 5.27(ii)).
+Limit $\eta\leq\lambda$.
+Let $\eta$ be limit and assume $\langle k_{\beta}:\beta<\eta\rangle$ has been constructed coherently.
 By the limit-stage group definition ([5, Def.
-3.1]), $g_{\eta}\in\mathscr{G}_{\eta}$.
-By the uniqueness property of the inverse/direct limit presentation: elements of $\mathscr{G}_{\eta}$ are determined by their projections to all earlier stages ([5, Def.
-3.1]).
-Since $g_{\eta}$ and $\widehat{\pi}^{\,D_{\eta}}$ have identical restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$ (namely, $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$), we conclude $g_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
-Coherence: We verify $\pi^{\gamma}_{\beta}(g_{\beta})=g_{\gamma}$ for $\gamma<\beta\leq\lambda$.
-For successor $\beta+1$:
-$\pi^{\beta}_{\beta+1}(g_{\beta+1})$ $=\pi^{\beta}_{\beta+1}\bigl(\mathrm{spl}(g_{\beta})\cdot\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})\bigr)$
-$=\pi^{\beta}_{\beta+1}(\mathrm{spl}(g_{\beta}))\cdot\pi^{\beta}_{\beta+1}(\kappa_{\beta+1}(\ldots))$
-$=g_{\beta}\cdot\mathrm{id}=g_{\beta},$
-using $\pi^{\beta}_{\beta+1}\circ\mathrm{spl}=\mathrm{id}$ (Lemma 5.21) and $\kappa_{\beta+1}(\ldots)\in K_{\beta+1}=\ker(\pi^{\beta}_{\beta+1})$ (Lemma 5.25(ii)).
-For limits $\eta$, coherence is automatic: $g_{\eta}$ is defined as the direct-limit element, so $\pi^{\gamma}_{\eta}(g_{\eta})=g_{\gamma}$ for all $\gamma<\eta$ by construction.
-Conclusion: Taking $\beta=\lambda$, we have $g_{\lambda}=\widehat{\pi}^{\,D_{\lambda}}$.
-Since $D_{\lambda}=D\cap\mathrm{Coords}_{\lambda}=D$ (since $D\subseteq\mathrm{Coords}_{<\lambda}$ by Definition 5.26(ii)), we have $g_{\lambda}=\widehat{\pi}^{\,D}$ and $g_{\lambda}\in\mathscr{G}_{\lambda}$.
-Hence $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$.
+3.1]), there is a unique direct-limit element $k_{\eta}\in\mathscr{G}_{\eta}$ with $\pi^{\beta}_{\eta}(k_{\eta})=k_{\beta}$ for all $\beta<\eta$.
+Since $\widehat{\pi}^{\,D_{\eta}}$ has the same restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$, we have $k_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
+Conclusion.
+At $\beta=\lambda$ we obtain $k_{\lambda}=\widehat{\pi}^{\,D_{\lambda}}\in\mathscr{G}_{\lambda}$.
+Because $D\subseteq\mathrm{Coords}_{<\lambda}$ (Definition 5.28(ii)), we have $D_{\lambda}=D$, hence $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$.
 ∎
 Lemma 5.46 (Diagonal lifts lie in $\Delta^{\uparrow}$).
+Let $\lambda$ be a stage, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{<\lambda}$.
 Let $\pi\in\operatorname{Fix}(E)$ and let $D^{\prime}\supseteq D$ be such that the diagonal lift $\widehat{\pi}^{\,D^{\prime}}\in\mathscr{G}_{\lambda}$ exists.
-Then $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
+Then
+$\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D).$
 Proof.
-By the characterization definition of $\Delta_{\lambda}^{\uparrow}(E,D)$ (Definition 5.27), we verify:
-  1. (i)
-    $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D^{\prime}})\in\operatorname{Fix}(E)$: By construction of diagonal lifts (Lemma 5.31(i)), the stage-0 projection of $\widehat{\pi}^{\,D^{\prime}}$ is $\pi$, and $\pi\in\operatorname{Fix}(E)$ by assumption.
-  2. (ii)
-    $\widehat{\pi}^{\,D^{\prime}}$ fixes every coordinate in $D$ pointwise: By construction (Lemma 5.31(ii)), $\widehat{\pi}^{\,D^{\prime}}$ fixes every coordinate in $D^{\prime}$ pointwise, and $D\subseteq D^{\prime}$.
-Hence $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$. ∎
+Since $\widehat{\pi}^{\,D^{\prime}}\in\mathscr{G}_{\lambda}$ and $\widehat{\pi}^{\,D^{\prime}}$ restricts to $\pi$ on the Cohen base (Lemma 5.33(i)), we have $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D^{\prime}})=\pi\in\operatorname{Fix}(E)$ (e.g.
+by Corollary 5.34).
+Moreover, by Lemma 5.33(ii) the automorphism $\widehat{\pi}^{\,D^{\prime}}$ fixes every coordinate in $D^{\prime}$ pointwise, hence fixes every coordinate in $D\subseteq D^{\prime}$ pointwise.
+Therefore $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$ by Definition 5.29(A).
+∎
 Lemma 5.47 (Package generic sections are hereditarily symmetric via diagonal cancellation).
-Let $f$ have countable support $E$ (so $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$) and let $\dot{s}_{f}$ be the canonical $\mathbb{S}_{[f]}$-name for the $f$-coordinate generic section.
-Then $\Delta^{\uparrow}(E,\{f\})\leq\operatorname{Stab}(\dot{s}_{f})$, hence $\dot{s}_{f}\in\operatorname{HS}$ (with respect to $\mathscr{F}_{[f]}$).
+Let $f$ have countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$ (so $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$), and let $\dot{s}_{f}$ be the canonical $\mathbb{S}_{[f]}$-name for the $f$-coordinate generic section.
+Then
+$\Delta^{\uparrow}(E,\{f\})\ \leq\ \operatorname{Stab}_{\mathscr{G}_{[f]}}(\dot{s}_{f}),$
+and hence $\dot{s}_{f}$ is hereditarily symmetric for the orbit package system $(\mathbb{S}_{[f]},\mathscr{G}_{[f]},\mathscr{F}_{[f]})$.
 Proof.
-Let $E\subseteq\omega_{1}$ witness that $f$ has countable support, i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$.
-Fix any admissible pair $(E^{\prime},D^{\prime})$ such that $E\subseteq E^{\prime}$ and $\{f\}\subseteq D^{\prime}$.
-Let $\pi\in\operatorname{Fix}(E^{\prime})$.
-Then the diagonal lift $\hat{\pi}^{D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ fixes the $f$-coordinate pointwise: for every condition $\vec{p}\in\mathbb{S}_{[f]}$,
-$(\hat{\pi}^{D^{\prime}}\cdot\vec{p})(f)=\vec{p}(f).$
-Since $\dot{s}_{f}$ is definable (as a $\mathbb{S}_{[f]}$-name) purely from the $f$-coordinate generic section, it follows that $\hat{\pi}^{D^{\prime}}\cdot\dot{s}_{f}=\dot{s}_{f}$.
-Hence $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}(\dot{s}_{f})$.
-As $\Delta^{\uparrow}(E,\{f\})$ is generated by all such subgroups $\Delta(E^{\prime},D^{\prime})$, we have $\Delta^{\uparrow}(E,\{f\})\leq\operatorname{Stab}(\dot{s}_{f})$.
-Therefore $\operatorname{Stab}(\dot{s}_{f})\in\mathscr{F}_{[f]}$, so $\dot{s}_{f}\in\operatorname{HS}$.
+Fix $E\in[\omega_{1}]^{\leq\omega}$ with $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$.
+Let $(E^{\prime},D^{\prime})$ be any admissible pair in the sense of Definition 5.11 such that $E\subseteq E^{\prime}$ and $f\in D^{\prime}$.
+Let $\pi\in\operatorname{Fix}(E^{\prime})$ and consider the diagonal-cancellation automorphism $\widehat{\pi}^{\,D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ (Definition 5.11).
+Because $f\in D^{\prime}$, this automorphism fixes the $f$-coordinate pointwise, i.e.
+for every $\vec{p}\in\mathbb{S}_{[f]}$,
+$(\widehat{\pi}^{\,D^{\prime}}\cdot\vec{p})(f)=\vec{p}(f).$
+Since $\dot{s}_{f}$ is the canonical name definable purely from the $f$-coordinate generic section, we have $\widehat{\pi}^{\,D^{\prime}}\cdot\dot{s}_{f}=\dot{s}_{f}$.
+Thus $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}_{\mathscr{G}_{[f]}}(\dot{s}_{f})$.
+By Definition 5.13, $\Delta^{\uparrow}(E,\{f\})$ is generated by all such $\Delta(E^{\prime},D^{\prime})$ with $E\subseteq E^{\prime}$ and $f\in D^{\prime}$, so $\Delta^{\uparrow}(E,\{f\})\leq\operatorname{Stab}_{\mathscr{G}_{[f]}}(\dot{s}_{f})$.
+Therefore $\operatorname{Stab}_{\mathscr{G}_{[f]}}(\dot{s}_{f})\in\mathscr{F}_{[f]}$ by Definition 5.13, and $\dot{s}_{f}$ is hereditarily symmetric.
 ∎
 Lemma 5.48 (Diagonal cancellation makes right-inverse generics hereditarily symmetric).
-Let $g$ be a package with countable support witnessed by $E\subseteq\omega_{1}$, i.e.
+Let $g$ be a package with countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$, i.e.
 $\operatorname{Fix}(E)\leq\operatorname{Stab}(g)$.
 Let $\dot{r}_{g}$ be the canonical $\mathbb{R}_{[g]}$-name for the $g$-coordinate generic right inverse.
-Then $\Delta^{\uparrow}(E,\{g\})\leq\operatorname{Stab}(\dot{r}_{g})$, and hence $\dot{r}_{g}\in\operatorname{HS}$ (with respect to $(\mathbb{R}_{[g]},\mathscr{G}_{[g]},\mathscr{F}_{[g]})$).
+Then
+$\Delta^{\uparrow}(E,\{g\})\ \leq\ \operatorname{Stab}_{\mathscr{G}_{[g]}}(\dot{r}_{g}),$
+and hence $\dot{r}_{g}$ is hereditarily symmetric for $(\mathbb{R}_{[g]},\mathscr{G}_{[g]},\mathscr{F}_{[g]})$.
 Proof.
-The proof is identical to Lemma 5.45.
-Fix an admissible pair $(E^{\prime},D^{\prime})$ with $E\subseteq E^{\prime}$ and $\{g\}\subseteq D^{\prime}$.
-For $\pi\in\operatorname{Fix}(E^{\prime})$, the diagonal lift $\hat{\pi}^{D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ fixes the $g$-coordinate pointwise, and $\dot{r}_{g}$ is definable purely from the $g$-coordinate generic object.
-Hence $\hat{\pi}^{D^{\prime}}\cdot\dot{r}_{g}=\dot{r}_{g}$, so $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}(\dot{r}_{g})$.
-Therefore $\Delta^{\uparrow}(E,\{g\})\leq\operatorname{Stab}(\dot{r}_{g})$, as required.
+Fix $E\in[\omega_{1}]^{\leq\omega}$ with $\operatorname{Fix}(E)\leq\operatorname{Stab}(g)$.
+Let $(E^{\prime},D^{\prime})$ be any admissible pair (Definition 5.11) such that $E\subseteq E^{\prime}$ and $g\in D^{\prime}$.
+For $\pi\in\operatorname{Fix}(E^{\prime})$, the diagonal-cancellation map $\widehat{\pi}^{\,D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ fixes the $g$-coordinate pointwise (since $g\in D^{\prime}$).
+Because $\dot{r}_{g}$ is the canonical name definable purely from the $g$-coordinate generic object, we have $\widehat{\pi}^{\,D^{\prime}}\cdot\dot{r}_{g}=\dot{r}_{g}$.
+Hence $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}_{\mathscr{G}_{[g]}}(\dot{r}_{g})$.
+By Definition 5.13, $\Delta^{\uparrow}(E,\{g\})$ is generated by all such $\Delta(E^{\prime},D^{\prime})$ with $E\subseteq E^{\prime}$ and $g\in D^{\prime}$, so $\Delta^{\uparrow}(E,\{g\})\leq\operatorname{Stab}_{\mathscr{G}_{[g]}}(\dot{r}_{g})$.
+Therefore $\operatorname{Stab}_{\mathscr{G}_{[g]}}(\dot{r}_{g})\in\mathscr{F}_{[g]}$ (Definition 5.13), and $\dot{r}_{g}$ is hereditarily symmetric.
 ∎
-Lemma 5.47 (Successor-step HS lift for package generics).
-Fix a stage $\beta$ and a package coordinate $d\in\mathrm{NewCoords}_{\beta+1}=\mathrm{Coords}_{\beta+1}\setminus\mathrm{Coords}_{\beta}$ scheduled at stage $\beta$.
-(e.g.
-$d=f_{\alpha}$ for a splitting package $\mathbb{Q}_{[f_{\alpha}]}$, or $d=g_{\alpha}$ for a right-inverse package $\mathbb{R}_{[g_{\alpha}]}$).
-Let $\dot{\tau}$ be a canonical $\dot{\mathbb{Q}}_{\beta}$-name definable purely from the $d$-coordinate generic object (e.g.
-$\dot{s}_{f_{\alpha}}$ or $\dot{r}_{g_{\alpha}}$).
-Assume $d$ has countable support witnessed by $E\subseteq\omega_{1}$ (so $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ in $M_{\beta}$).
-Then the canonical $\mathbb{P}_{\beta+1}$-name $\dot{\tau}^{*}$ is hereditarily symmetric at stage $\beta+1$; in particular, $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
+Lemma 5.49 (Successor-stage diagonal cancellation yields $\operatorname{HS}$-lifts).
+Let $\beta$ be an ordinal and consider the successor stage $\beta+1$ with $\mathbb{P}_{\beta+1}\cong\mathbb{P}_{\beta}*\dot{\mathbb{Q}}_{\beta}$.
+Let $d$ be a package coordinate of $\dot{\mathbb{Q}}_{\beta}$ (e.g.
+a $\mathbb{Q}$-package coordinate or a $\mathbb{R}$-package coordinate), and let $\dot{\tau}$ be a canonical $\dot{\mathbb{Q}}_{\beta}$-name definable purely from the $d$-coordinate generic object (e.g.
+$\dot{s}_{f_{\beta}}$ or $\dot{r}_{g_{\beta}}$).
+Let $\dot{\tau}^{*}$ denote the corresponding canonical $\mathbb{P}_{\beta+1}$-name under this two-step presentation.
+Assume $d$ has countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$ in $M_{\beta}$, so $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ in $M_{\beta}$.
+Then $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$; in particular, $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
 Proof.
-Since $d$ has countable support witnessed by $E$ in $M_{\beta}$, we have $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$.
-Thus $(E,\{d\})$ is globally admissible at stage $\beta+1$ (Definition 5.26), so $\Delta_{\beta+1}^{\uparrow}(E,\{d\})$ is one of the $\Delta^{\uparrow}$-generators of $\tilde{\mathscr{F}}^{*}_{\beta+1}$ (Definition 5.48).
-Now let $g\in\Delta_{\beta+1}^{\uparrow}(E,\{d\})$.
-By the characterization in Definition 5.27, $g$ fixes the coordinate $d$ pointwise (Here “fixes the coordinate $d$ pointwise” is exactly the clause in Definition 5.27: the induced automorphism on the $d$-factor forcing ($\dot{\mathbb{Q}}_{d}$ or $\dot{\mathbb{R}}_{d}$, depending on the package) is literally the identity, i.e.
-$g\cdot\dot{p}=\dot{p}$ for every canonical name $\dot{p}$ for a condition in the $d$-factor.)
-Since $\dot{\tau}^{*}$ is definable purely from the $d$-coordinate generic object, it follows that $g\cdot\dot{\tau}^{*}=\dot{\tau}^{*}$.
-Hence $\Delta_{\beta+1}^{\uparrow}(E,\{d\})\leq\operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*})$.
-Therefore $\operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*})\in\tilde{\mathscr{F}}^{*}_{\beta+1}$ by upward closure, so $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$ and $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
+From $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ we get that $(E,\{d\})$ is globally admissible at stage $\beta+1$ (Definition 5.28).
+Hence $\Delta_{\beta+1}^{\uparrow}(E,\{d\})$ is one of the designated $\Delta^{\uparrow}$-generators of $\tilde{\mathscr{F}}^{*}_{\beta+1}$ (Definition 5.50).
+Let $k\in\Delta_{\beta+1}^{\uparrow}(E,\{d\})$.
+By Definition 5.29, $k$ fixes the coordinate $d$ pointwise (i.e.
+its induced action on the $d$-factor forcing is the identity).
+Since $\dot{\tau}^{*}$ is a canonical name definable purely from the $d$-coordinate generic object, it follows that $k\cdot\dot{\tau}^{*}=\dot{\tau}^{*}$.
+Therefore
+$\Delta_{\beta+1}^{\uparrow}(E,\{d\})\ \leq\ \operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*}).$
+By upward closure of $\tilde{\mathscr{F}}^{*}_{\beta+1}$, we get $\operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*})\in\tilde{\mathscr{F}}^{*}_{\beta+1}$, hence $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$.
+The interpretation $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$ follows.
 ∎
@@
 We now define the filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ that governs hereditary symmetry at limit stages of the iteration.
-Definition 5.48 (Modified limit filter).
-Define the filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ on $\mathscr{G}_{\lambda}$ inductively:
-Stage 0: $\tilde{\mathscr{F}}^{*}_{0}$ is the filter on $\mathscr{G}_{0}=\operatorname{Sym}(\omega_{1})$ generated by $\{\operatorname{Fix}(E):E\subseteq\omega_{1}\text{ countable}\}$.
-Note that $\operatorname{Fix}(E)=\Delta_{0}^{\uparrow}(E,\varnothing)$, and $(E,\varnothing)$ is trivially globally admissible.
+Definition 5.50 (Modified stage filters).
+Define by recursion on stages $\alpha$ a normal filter $\tilde{\mathscr{F}}^{*}_{\alpha}$ on $\mathscr{G}_{\alpha}$ as follows.
+Stage 0: $\tilde{\mathscr{F}}^{*}_{0}$ is the normal filter on $\mathscr{G}_{0}=\mathscr{G}=\operatorname{Sym}(\omega_{1})$ generated by $\{\operatorname{Fix}(E):E\in[\omega_{1}]^{\leq\omega}\}$.
+(Note that $\operatorname{Fix}(E)=\Delta^{\uparrow}_{0}(E,\varnothing)$, and $(E,\varnothing)$ is trivially globally admissible.)
 Successor stage $\alpha+1$: $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter on $\mathscr{G}_{\alpha+1}$ generated by:
   1. (a)
-    Pushforwards: $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$;
+    Pullbacks: $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$;
   2. (b)
-    $\Delta^{\uparrow}$-subgroups: $\Delta_{\alpha+1}^{\uparrow}(E,D)$ for countable $E\subseteq\omega_{1}$ and countable $D\subseteq\mathrm{Coords}_{<\alpha+1}$ such that $(E,D)$ is globally admissible at stage $\alpha+1$.
+    $\Delta^{\uparrow}$-subgroups: $\Delta_{\alpha+1}^{\uparrow}(E,D)$ for countable $(E,D)$ globally admissible at stage $\alpha+1$.
 Limit stage $\lambda$: $\tilde{\mathscr{F}}^{*}_{\lambda}$ is the normal filter on $\mathscr{G}_{\lambda}$ generated by:
   1. (a)
-    Pushforwards: $(\pi^{\beta}_{\lambda})^{-1}[H]$ for $\beta<\lambda$ and $H\in\tilde{\mathscr{F}}^{*}_{\beta}$;
+    Pullbacks: $(\pi^{\beta}_{\lambda})^{-1}[H]$ for $\beta<\lambda$ and $H\in\tilde{\mathscr{F}}^{*}_{\beta}$;
   2. (b)
-    $\Delta^{\uparrow}$-subgroups: $\Delta_{\lambda}^{\uparrow}(E,D)$ for countable $E\subseteq\omega_{1}$ and countable $D$ (package coordinates at stages $\lambda$) such that $(E,D)$ is globally admissible at stage $\lambda$.
+    $\Delta^{\uparrow}$-subgroups: $\Delta_{\lambda}^{\uparrow}(E,D)$ for countable $(E,D)$ globally admissible at stage $\lambda$.
 Remark 5.51 (Admissibility in the filter definition).
-We require global admissibility in the $\Delta^{\uparrow}$-generator clause to ensure that the core lemma (Lemma 5.50) produces admissible pairs, which is needed for the properness proof (Lemma 5.52(iii)).
-By Lemma 5.38, conjugation sends $\Delta_{\lambda}^{\uparrow}(E,D)$ to $\Delta_{\lambda}^{\uparrow}(g\cdot E,\,g\cdot D)$, and $g\cdot E$ and $g\cdot D$ remain countable (Lemma 5.30).
-The pair $(g\cdot E,\,g\cdot D)$ need not be globally admissible, but by Lemma 5.40 there is a countable $E^{\ast}\supseteq g\cdot E$ such that $(E^{\ast},\,g\cdot D)$ is globally admissible and $\Delta_{\lambda}^{\uparrow}(E^{\ast},\,g\cdot D)\leq\Delta_{\lambda}^{\uparrow}(g\cdot E,\,g\cdot D)$.
-Together with Corollary 5.39, this ensures that the normal filter generation process always contains $\Delta^{\uparrow}$-groups with globally admissible parameters.
+We restrict the $\Delta^{\uparrow}$-generator clause in Definition 5.50 to globally admissible pairs so that later arguments can uniformly recover $\Delta^{\uparrow}$-subgroups with admissible parameters (notably Lemma 5.52, used in Lemma 5.54(iii),(iv)).
+Normality interacts with admissibility as follows. If $(E,D)$ is globally admissible at stage $\lambda$ and $k\in\mathscr{G}_{\lambda}$, then Lemma 5.40 gives
+$k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}=\Delta_{\lambda}^{\uparrow}(k\cdot E,\ k\cdot D),$
+where $k\cdot E=\pi^{0}_{\lambda}(k)[E]$ and $k\cdot D$ is the induced action on coordinates (Lemma 5.32).
+The pair $(k\cdot E,\ k\cdot D)$ need not be globally admissible, but by Lemma 5.42 there is a countable $E^{\ast}\supseteq k\cdot E$ such that $(E^{\ast},\ k\cdot D)$ is globally admissible and
+$\Delta_{\lambda}^{\uparrow}(E^{\ast},\ k\cdot D)\ \leq\ \Delta_{\lambda}^{\uparrow}(k\cdot E,\ k\cdot D).$
+Together with Corollary 5.41 (and its countable variant Corollary 5.53), this ensures that the normal-filter generation process always contains $\Delta^{\uparrow}$-groups with globally admissible parameters.
 Lemma 5.52 (Every filter element contains a $\Delta^{\uparrow}$-group).
-For every $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$, there exist countable $E\subseteq\omega_{1}$ and countable $D$ (package coordinates at stages $\lambda$) with $(E,D)$ globally admissible such that:
+For each stage $\lambda$ and every $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$, there exist countable $E\subseteq\omega_{1}$ and a countable $D\subseteq\mathrm{Coords}_{<\lambda}$ such that $(E,D)$ is globally admissible at stage $\lambda$ and
 $\Delta_{\lambda}^{\uparrow}(E,D)\ \leq\ H.$
 Proof.
-By definition of the normal filter generated by a family, every element $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$ contains a finite intersection of conjugates of generators.
-We show that such intersections contain $\Delta^{\uparrow}$-groups with globally admissible parameters.
-Conjugates of $\Delta^{\uparrow}$-generators: Let $\Delta_{\lambda}^{\uparrow}(E,D)$ be a globally admissible generator and let $g\in\mathscr{G}_{\lambda}$. By Lemma 5.38:
-$g\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot g^{-1}=\Delta_{\lambda}^{\uparrow}(g\cdot E,g\cdot D).$
-Note that $g\cdot E$ and $g\cdot D$ are countable: $g\cdot E=\pi^{0}_{\lambda}(g)[E]$ is countable because bijections preserve cardinality, and similarly $g\cdot D$ is countable because the action on package coordinates is a bijection (Lemma 5.30(ii)).
-The parameters $(g\cdot E,g\cdot D)$ need not be globally admissible.
-However, by Lemma 5.40, there exists $E^{\ast}\supseteq g\cdot E$ such that $(E^{\ast},g\cdot D)$ is globally admissible and $\Delta_{\lambda}^{\uparrow}(E^{\ast},g\cdot D)\leq\Delta_{\lambda}^{\uparrow}(g\cdot E,g\cdot D)$.
-Thus every conjugate of an admissible $\Delta^{\uparrow}$-generator contains a globally admissible $\Delta^{\uparrow}$-group.
-Finite intersections: If $H_{1}\supseteq\Delta_{\lambda}^{\uparrow}(E_{1},D_{1})$ and $H_{2}\supseteq\Delta_{\lambda}^{\uparrow}(E_{2},D_{2})$ with $(E_{i},D_{i})$ globally admissible, then by Corollary 5.35:
-$H_{1}\cap H_{2}\supseteq\Delta_{\lambda}^{\uparrow}(E_{1}\cup E_{2},D_{1}\cup D_{2}).$
-By Corollary 5.39, $(E_{1}\cup E_{2},D_{1}\cup D_{2})$ is globally admissible.
-Pullback generators: Let $H=(\pi^{\beta}_{\lambda})^{-1}[K]$ with $K\in\tilde{\mathscr{F}}^{*}_{\beta}$.
-By induction on $\beta$, there exist countable admissible $(E,D)$ with $D\subseteq\mathrm{Coords}_{\beta}$ and $\Delta_{\beta}^{\uparrow}(E,D)\leq K$.
-By Lemma 5.42, $\Delta_{\lambda}^{\uparrow}(E,D)\leq H$.
-The pair $(E,D)$ remains globally admissible at stage $\lambda$ (the admissibility condition $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ depends only on how $d$ was introduced, not on which stage we evaluate it).
-∎
+We argue by induction on the stage $\lambda$.
+Stage 0.
+By Definition 5.50, $\tilde{\mathscr{F}}^{*}_{0}$ is the normal filter on $\mathscr{G}_{0}=\operatorname{Sym}(\omega_{1})$ generated by $\{\operatorname{Fix}(E):E\in[\omega_{1}]^{\leq\omega}\}$.
+Every conjugate of $\operatorname{Fix}(E)$ is again of the form $\operatorname{Fix}(E^{\prime})$ for some countable $E^{\prime}$, and finite intersections satisfy $\operatorname{Fix}(E_{1})\cap\operatorname{Fix}(E_{2})\supseteq\operatorname{Fix}(E_{1}\cup E_{2})$.
+Hence every $H\in\tilde{\mathscr{F}}^{*}_{0}$ contains $\operatorname{Fix}(E)=\Delta^{\uparrow}_{0}(E,\varnothing)$ for some countable $E$, and $(E,\varnothing)$ is globally admissible.
+This proves the claim at stage 0.
+Inductive step.
+Fix $\lambda>0$ and assume the statement holds for every $\beta<\lambda$.
+Let $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$.
+By definition of “normal filter generated by a family”, there exist generators $G_{0},\dots,G_{n-1}$ from Definition 5.50 and elements $k_{0},\dots,k_{n-1}\in\mathscr{G}_{\lambda}$ such that
+$\bigcap_{i<n}k_{i}G_{i}k_{i}^{-1}\ \subseteq\ H.$
+We show that each conjugate $k_{i}G_{i}k_{i}^{-1}$ contains a globally admissible $\Delta^{\uparrow}_{\lambda}$-group.
+Case 1: $G_{i}=\Delta_{\lambda}^{\uparrow}(E,D)$ with $(E,D)$ globally admissible at stage $\lambda$. By Lemma 5.40,
+$k_{i}\,\Delta_{\lambda}^{\uparrow}(E,D)\,k_{i}^{-1}=\Delta_{\lambda}^{\uparrow}(k_{i}\cdot E,\ k_{i}\cdot D).$
+Apply Lemma 5.42 to the countable set $k_{i}\cdot D$ and the countable set $k_{i}\cdot E$ to obtain a countable $E_{i}^{\ast}\supseteq k_{i}\cdot E$ such that $(E_{i}^{\ast},\ k_{i}\cdot D)$ is globally admissible.
+Then
+$\Delta_{\lambda}^{\uparrow}(E_{i}^{\ast},\ k_{i}\cdot D)\ \leq\ \Delta_{\lambda}^{\uparrow}(k_{i}\cdot E,\ k_{i}\cdot D)=k_{i}G_{i}k_{i}^{-1}.$
+Set $(E_{i},D_{i}):=(E_{i}^{\ast},\ k_{i}\cdot D)$.
+Case 2: $G_{i}$ is a pullback generator.
+Then $G_{i}=(\pi^{\beta}_{\lambda})^{-1}[K]$ for some $\beta<\lambda$ and some $K\in\tilde{\mathscr{F}}^{*}_{\beta}$ (as in Definition 5.50).
+Put $\sigma:=\pi^{\beta}_{\lambda}(k_{i})\in\mathscr{G}_{\beta}$.
+Since $\pi^{\beta}_{\lambda}$ is a group homomorphism (Remark 5.22), one has
+$k_{i}\,(\pi^{\beta}_{\lambda})^{-1}[K]\,k_{i}^{-1}=(\pi^{\beta}_{\lambda})^{-1}\!\big[\sigma K\sigma^{-1}\big].$
+Because $\tilde{\mathscr{F}}^{*}_{\beta}$ is normal, $\sigma K\sigma^{-1}\in\tilde{\mathscr{F}}^{*}_{\beta}$.
+By the inductive hypothesis at stage $\beta$, choose a globally admissible pair $(E,D)$ at stage $\beta$ with $\Delta_{\beta}^{\uparrow}(E,D)\leq\sigma K\sigma^{-1}$.
+Then Lemma 5.44 yields
+$\Delta_{\lambda}^{\uparrow}(E,D)\ \leq\ (\pi^{\beta}_{\lambda})^{-1}\!\big[\Delta_{\beta}^{\uparrow}(E,D)\big]\ \leq\ (\pi^{\beta}_{\lambda})^{-1}\!\big[\sigma K\sigma^{-1}\big]=k_{i}G_{i}k_{i}^{-1}.$
+Since $\beta<\lambda$, the same pair $(E,D)$ is globally admissible at stage $\lambda$ as well. Set $(E_{i},D_{i}):=(E,D)$.
+Assembling the finite intersection.
+Let $E^{\ast}:=\bigcup_{i<n}E_{i}$ and $D^{\ast}:=\bigcup_{i<n}D_{i}$.
+Since $n$ is finite, both unions are countable, and by iterating Corollary 5.41 the pair $(E^{\ast},D^{\ast})$ is globally admissible at stage $\lambda$.
+Moreover, by Corollary 5.37 (finite case),
+$\Delta_{\lambda}^{\uparrow}(E^{\ast},D^{\ast})\ \leq\ \bigcap_{i<n}\Delta_{\lambda}^{\uparrow}(E_{i},D_{i})\ \leq\ \bigcap_{i<n}k_{i}G_{i}k_{i}^{-1}\ \subseteq\ H.$
+This completes the induction and proves the lemma. ∎
 Corollary 5.53 (Admissibility preserved under countable unions).
@@
 Proof.
-For $d\in\bigcup_{n}D_{n}$, choose $k$ with $d\in D_{k}$.
-Then $\operatorname{Fix}(\bigcup_{n}E_{n})\leq\operatorname{Fix}(E_{k})\leq\operatorname{Stab}(d)$, where the last inequality holds by global admissibility of $(E_{k},D_{k})$.
+Put $E:=\bigcup_{n<\omega}E_{n}$ and $D:=\bigcup_{n<\omega}D_{n}$.
+Since we are working in the ambient universe in which the iteration is constructed, countable unions of countable sets are countable; hence $E\in[\omega_{1}]^{\leq\omega}$ and $D$ is countable.
+(No claim is made that such unions are countable inside the eventual symmetric model.)
+To verify clause (iii) of Definition 5.28, fix $d\in D$ and choose $k$ with $d\in D_{k}$.
+Then $\operatorname{Fix}(E)\leq\operatorname{Fix}(E_{k})\leq\operatorname{Stab}(d)$, where the last inequality holds because $(E_{k},D_{k})$ is globally admissible.
+Therefore $(E,D)$ is globally admissible.
 ∎
@@
   4. (iv)
-    $\tilde{\mathscr{F}}^{*}_{\lambda}$ is $\omega_{1}$-complete.
+    $\tilde{\mathscr{F}}^{*}_{\lambda}$ is $\omega_{1}$-complete (i.e. closed under countable intersections).
 Proof.
@@
 (iii) By Lemma 5.52, if $\{\mathrm{id}\}\in\tilde{\mathscr{F}}^{*}_{\lambda}$, then $\{\mathrm{id}\}\supseteq\Delta_{\lambda}^{\uparrow}(E,D)$ for some countable globally admissible $E,D$.
-Since $E$ is countable and $\omega_{1}$ is uncountable, there exist distinct $\beta,\gamma\in\omega_{1}\setminus E$. Let $\pi=(\beta\ \gamma)$ (transposition). Then $\pi\in\operatorname{Fix}(E)$.
-By Lemma 5.43, the diagonal lift $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ exists. By Lemma 5.44, $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
-Since $\widehat{\pi}^{\,D}$ acts as $\pi$ on the Cohen base (Lemma 5.31(i)), it swaps coordinates $\beta$ and $\gamma$, so $\widehat{\pi}^{\,D}\neq\mathrm{id}$.
-Hence $\Delta_{\lambda}^{\uparrow}(E,D)$ contains a nontrivial element, contradicting $\Delta_{\lambda}^{\uparrow}(E,D)\leq\{\mathrm{id}\}$.
-(iv) $\omega_{1}$-completeness.
-Let $\{H_{n}:n<\omega\}\subseteq\tilde{\mathscr{F}}^{*}_{\lambda}$.
-By Lemma 5.50, for each $n$ there exists a globally admissible $(E_{n},D_{n})$ with $\Delta_{\lambda}^{\uparrow}(E_{n},D_{n})\leq H_{n}$.
-By Corollary 5.35:
-$\bigcap_{n<\omega}H_{n}\supseteq\Delta_{\lambda}^{\uparrow}\left(\bigcup_{n}E_{n},\bigcup_{n}D_{n}\right).$
-By Corollary 5.51, $(\bigcup_{n}E_{n},\bigcup_{n}D_{n})$ is globally admissible.
-Since countable unions of countable sets are countable (in the metatheory, which has $\mathsf{AC}$), the right-hand side is a $\Delta^{\uparrow}$-generator of the filter.
-Hence $\bigcap_{n}H_{n}\in\tilde{\mathscr{F}}^{*}_{\lambda}$. ∎
+Since $E$ is countable and $\omega_{1}$ is uncountable, choose distinct $\beta,\gamma\in\omega_{1}\setminus E$ and let $\pi=(\beta\ \gamma)$. Then $\pi\in\operatorname{Fix}(E)$.
+By Lemma 5.45, the diagonal lift $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ exists, and by Lemma 5.46 we have $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
+By Corollary 5.34, $\widehat{\pi}^{\,D}$ swaps $c_{\beta}\leftrightarrow c_{\gamma}$, so $\widehat{\pi}^{\,D}\neq\mathrm{id}$, contradicting $\Delta_{\lambda}^{\uparrow}(E,D)\leq\{\mathrm{id}\}$.
+(iv) Let $\{H_{n}:n<\omega\}\subseteq\tilde{\mathscr{F}}^{*}_{\lambda}$.
+By Lemma 5.52, for each $n$ choose a globally admissible $(E_{n},D_{n})$ with $\Delta_{\lambda}^{\uparrow}(E_{n},D_{n})\leq H_{n}$.
+By Corollary 5.37,
+$\bigcap_{n<\omega}H_{n}\ \supseteq\ \Delta_{\lambda}^{\uparrow}\Bigl(\bigcup_{n}E_{n},\ \bigcup_{n}D_{n}\Bigr).$
+By Corollary 5.53, the pair $\bigl(\bigcup_{n}E_{n},\bigcup_{n}D_{n}\bigr)$ is globally admissible, and in the ambient universe where the iteration and filters are defined these unions are countable.
+Hence the right-hand side is one of the $\Delta^{\uparrow}$-generators in Definition 5.50, so by upward closure $\bigcap_{n<\omega}H_{n}\in\tilde{\mathscr{F}}^{*}_{\lambda}$.
+∎
 Definition 5.55 (Hereditarily symmetric names).
@@
   2. (ii)
-    for every $(\sigma,p)\in\tau$, we have $\sigma\in\operatorname{HS}_{\lambda}$.
+    for every $\langle\sigma,p\rangle\in\tau$, we have $\sigma\in\operatorname{HS}_{\lambda}$.
 The symmetric extension at stage $\lambda$ is $M_{\lambda}=\operatorname{HS}_{\lambda}^{G_{\lambda}}=\{\tau^{G_{\lambda}}:\tau\in\operatorname{HS}_{\lambda}\}$.
 Lemma 5.56 ($\Delta$-support equivalence).
-For any $\mathbb{P}_{\lambda}$-name $\tau$:
-$\operatorname{Stab}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}\iff\exists\text{ countable globally admissible }(E,D)\text{ such that }\Delta_{\lambda}^{\uparrow}(E,D)\subseteq\operatorname{Stab}(\tau).$
+For any $\mathbb{P}_{\lambda}$-name $\tau$,
+$\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}\iff\exists(E,D)\ \bigl[(E,D)\text{ is globally admissible at stage }\lambda\ \wedge\ \Delta_{\lambda}^{\uparrow}(E,D)\subseteq\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\bigr].$
 Proof.
-$(\Leftarrow)$ By upward closure, since $\Delta_{\lambda}^{\uparrow}(E,D)\in\tilde{\mathscr{F}}^{*}_{\lambda}$ for globally admissible $(E,D)$. $(\Rightarrow)$ By Lemma 5.50. ∎
+$(\Leftarrow)$ If $\Delta_{\lambda}^{\uparrow}(E,D)\subseteq\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)$ and $(E,D)$ is globally admissible, then $\Delta_{\lambda}^{\uparrow}(E,D)\in\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.50), so by upward closure $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}$.
+$(\Rightarrow)$ Apply Lemma 5.52 to $H=\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}$. ∎
 Theorem 5.57 ($M_{\lambda}\models\mathsf{ZF}$).
@@
 Proof.
-By Lemma 5.52, the triple $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ is a symmetric system in the sense of [8, Ch.
+By Lemma 5.54(i),(ii), the triple $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ forms a symmetric system in the sense of [8, Ch.
 15].
-By the symmetric extension theorem [8, Lemma.
-15.51], $M_{\lambda}\models\mathsf{ZF}$.
+Therefore the associated symmetric extension $M_{\lambda}=\operatorname{HS}_{\lambda}^{G_{\lambda}}$ satisfies $\mathsf{ZF}$ by the symmetric extension theorem [8, Lemma 15.51].
 ∎
 Theorem 5.58 ($\mathcal{M}\models\mathsf{ZF}$).
-The final model $\mathcal{M}=\bigcup_{\alpha\in\mathrm{Ord}}M_{\alpha}$ satisfies $\mathsf{ZF}$.
+Let $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic and $G_{\lambda}:=G_{\infty}\cap\mathbb{P}_{\lambda}$.
+Let $\operatorname{HS}_{\mathrm{Ord}}$ be the class of hereditarily symmetric (cf.
+Definition 5.55) $\mathbb{P}_{\mathrm{Ord}}$-names (class-names) with respect to $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$, and define
+$\mathcal{M}\ :=\ \{\dot{x}^{G_{\infty}}:\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}\}.$
+Then $\mathcal{M}$ is a transitive class and $\mathcal{M}\models\mathsf{ZF}$.
 Proof.
-$\mathcal{M}$ is the directed union of the transitive models $\{M_{\alpha}:\alpha\in\mathrm{Ord}\}$, each satisfying $\mathsf{ZF}$ by Theorem 5.55.
-For any $\mathsf{ZF}$ axiom instance with parameters $a_{1},\ldots,a_{n}\in\mathcal{M}$: choose $\lambda$ large enough that $a_{1},\ldots,a_{n}\in M_{\lambda}$.
-Apply the axiom in $M_{\lambda}$ to obtain the required witness in $M_{\lambda}\subseteq\mathcal{M}$.
-The inclusion $M_{\beta}\subseteq M_{\lambda}$ for $\beta<\lambda$ is guaranteed by the pushforward generators: if $\tau\in\operatorname{HS}_{\beta}$ with $\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\beta}$, then $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\supseteq(\pi^{\beta}_{\lambda})^{-1}[\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)]\in\tilde{\mathscr{F}}^{*}_{\lambda}$, so $\tau\in\operatorname{HS}_{\lambda}$.
+Work in the background metatheory used to formalize class-length iterations (cf.
+[5, §(Class-length iterations)] and Remark 5.59).
+In this setting, the forcing theorem holds for the definable class forcing $\mathbb{P}_{\mathrm{Ord}}$, and the class system $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$ supports the usual recursion defining hereditarily symmetric class-names.
+For each set ordinal $\lambda$, the restriction to stage $\lambda$ yields the set-length symmetric system $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ and hence $M_{\lambda}\models\mathsf{ZF}$ by Theorem 5.57.
+Moreover, every set-sized $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{x}$ is supported in some stage $\lambda$, so $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\lambda}}\in M_{\lambda}$.
+Consequently,
+$\mathcal{M}=\{\dot{x}^{G_{\infty}}:\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}\}=\bigcup_{\lambda\in\mathrm{Ord}}M_{\lambda},$
+so $\mathcal{M}$ is a transitive class.
+The verification of the $\mathsf{ZF}$ axioms in $\mathcal{M}$ follows by the standard class-length symmetric iteration argument (ibid.): each axiom instance with set parameters from $\mathcal{M}$ is witnessed at some stage $\lambda$ and hence holds in $M_{\lambda}$, and therefore holds in $\mathcal{M}$.
 ∎
@@
 Remark 5.59 (Background metatheory for class-length constructions).
-To speak formally about an $\mathrm{Ord}$-length bookkeeping function and an $\mathrm{Ord}$-length symmetric iteration as class objects, we work in a background theory of classes, e.g.
-$\mathsf{GBC}+\mathsf{ETR}$ (optionally with Global Choice).
+To speak formally about an $\mathrm{Ord}$-length bookkeeping function and an $\mathrm{Ord}$-length symmetric iteration as class objects, we work in a background theory of classes supporting definable class recursion for the iteration, e.g.
+$\mathsf{GBC}+\mathsf{ETR}$, and we fix a class well-order of $V$ (equivalently, assume Global Choice in the background).
 Its set universe is denoted $V$ and satisfies $\mathsf{ZFC}$.
@@
 If it is not, we force trivially at that stage (or include only the other package factors scheduled for $\alpha$).
+We also use the canonical inclusion of $\mathbb{P}_{\beta}$-names into $\mathbb{P}_{\alpha}$-names for $\beta<\alpha$, so a code intended for an earlier stage can be retested later.
 This is the standard bookkeeping convention for class-length forcing/iterations.
@@
 is unbounded in $\mathrm{Ord}$ (e.g. enumerate $\mathrm{Ord}\times V$ using a definable pairing of ordinals together with a definable class well-order of $V$ in the background metatheory).
-At stage $\alpha$, let $\mathcal{B}(\alpha)=(\beta,c)$. We test whether the code $c$:
+At stage $\alpha$, let $\mathcal{B}(\alpha)=(\beta,c)$.
+We attempt to decode $c$ as a $\mathbb{P}_{\beta}$-name $\dot{f}$ (in some fixed coding of names), and when $\beta\leq\alpha$ we view $\dot{f}$ as a $\mathbb{P}_{\alpha}$-name via the canonical inclusion of $\mathbb{P}_{\beta}$-names into $\mathbb{P}_{\alpha}$-names.
+We then test whether (in the current stage model $M_{\alpha}$):
   1. (i)
-    decodes to a valid $\mathbb{P}_{\alpha}$-name $\dot{f}$,
+    $\dot{f}$ is defined (the decoding succeeds) and $\beta\leq\alpha$,
   2. (ii)
-    is hereditarily symmetric for the stage-$\alpha$ symmetric system $(\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha})$, and
+    $\dot{f}\in\operatorname{HS}_{\alpha}$ for the stage-$\alpha$ symmetric system $(\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha})$, and
   3. (iii)
-    interprets in $M_{\alpha}$ to a surjection of one of the following types:
+    $f:=\dot{f}^{G_{\alpha}}$ is a surjection instance of one of the following types:
     1. (1)
@@
     2. (2)
-      $f:S\times\eta\twoheadrightarrow\lambda$ with $\eta,\lambda\in\mathrm{Ord}$ (for $\mathsf{AC}_{\mathsf{WO}}$-packages).
-If all conditions are met and the surjection $f$ does not already split in $M_{\alpha}$ (i.e., no section/right-inverse for $f$ exists yet), the successor forcing at stage $\alpha+1$ includes the corresponding orbit package $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ (Definition 5.5); otherwise we force trivially at that code.
-This arrangement ensures that every code is tested at unboundedly many stages, so a name that becomes a valid HS surjection at some stage $\beta$ will eventually be scheduled at some stage $\alpha>\beta$.
+      $f:S\times\eta\twoheadrightarrow\lambda$ with $\eta,\lambda\in\mathrm{Ord}$ (for $\mathsf{AC}_{\mathsf{WO}}$-packages),
+and moreover $f$ does not already split in $M_{\alpha}$ (i.e., no section/right-inverse for $f$ exists in $M_{\alpha}$).
+If the test succeeds, we define the stage-$\alpha$ iterand $\dot{\mathbb{Q}}_{\alpha}$ so that in $M_{\alpha}$ it is the corresponding orbit package $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ (Definition 5.5); otherwise we set $\dot{\mathbb{Q}}_{\alpha}$ to be the trivial forcing.
+Thus $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{Q}}_{\alpha}$.
+This arrangement ensures that every code is tested at unboundedly many stages, so a name that becomes a valid HS surjection instance at some stage $\beta$ will eventually be scheduled at some stage $\alpha>\beta$.
 Remark 5.62 (The $\aleph^{*}(S)$ horizon).
-For the $\mathsf{AC}_{\mathsf{WO}}$-component we schedule splittings for surjections $f:Y\twoheadrightarrow\lambda$ with $\lambda<\aleph^{*}(S)$ (cf.
+For the $\mathsf{AC}_{\mathsf{WO}}$-component we schedule splittings for surjections $f:S\times\eta\twoheadrightarrow\lambda$ with $\lambda<\aleph^{*}(S)$ (cf.
 Remark 4.20).
@@
   1. (i)
-    The seed parameter $S=A^{\omega}$ is preserved at successor stages: our package forcings add no new reals (Corollaries 4.19 and 4.30), hence add no new $\omega$-sequences of Cohen reals and therefore no new elements of $S$.
+    The seed parameter $S=A^{\omega}$ is preserved at successor stages: our package forcings add no new reals (Corollaries 4.17 and 4.28), hence add no new $\omega$-sequences of Cohen reals and therefore no new elements of $S$.
   2. (ii)
@@
 We now invoke the iteration framework via Remark 2.4.
-For the global conclusion (a final model of $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$), we take a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$.
-Concretely, working in the metatheory, we form an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [5, §2.1], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$.
-At a successor stage $\alpha+1$, we consult the finite list $\mathcal{B}(\alpha)$ from Definition 5.59.
-In the current model $M_{\alpha}$, each code on this list either interprets to a genuine package instance (yielding a corresponding orbit package forcing $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ as in Definition 5.5), or it does not (in which case we take the trivial forcing for that code).
-Let $\dot{\mathbb{S}}_{\alpha}$ be a $\mathbb{P}_{\alpha}$-name for the finite-support product of the resulting orbit package posets.
+Working in the background metatheory, we form a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$.
+Concretely, we build an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [5, §3.6], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$.
+The next subsection states the properties of the final symmetric model that we prove.
+At a successor stage $\alpha+1$, we consult the bookkeeping output $\mathcal{B}(\alpha)=(\beta,c)$ from Definition 5.61.
+In the current model $M_{\alpha}$, the code $c$ either decodes (via the canonical inclusion $\mathbb{P}_{\beta}$-names $\subseteq$ $\mathbb{P}_{\alpha}$-names) to a genuine package instance, yielding a corresponding orbit package forcing $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ as in Definition 5.5, or it does not, in which case we take trivial forcing for that stage.
+Let $\dot{\mathbb{S}}_{\alpha}$ be the resulting $\mathbb{P}_{\alpha}$-name for the stage iterand (orbit package or trivial).
 Then $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$.
-At a successor stage $\alpha+1$, the group $\mathscr{G}_{\alpha+1}$ and filter $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ are those determined by the symmetric-iteration framework together with the diagonal infrastructure from Subsection 5.2 and the filter definition in Definition 5.48.
-Concretely, $\mathscr{G}_{\alpha+1}$ is generated by standard lifts of $\mathscr{G}_{\alpha}$ together with the (stage-$\alpha$) package automorphisms and the successor diagonal kernel elements $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ (Definition 5.23) acting on the new package coordinates, and $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter generated by pushforwards from $\tilde{\mathscr{F}}^{*}_{\alpha}$ together with the admissible subgroups $\Delta^{\uparrow}_{\alpha+1}(E,D)$.
-In particular, for each active package coordinate $d$ at stage $\alpha+1$ with countable support witnessed by $E_{d}$, the generator $\Delta^{\uparrow}_{\alpha+1}(E_{d},\{d\})$ belongs to $\tilde{\mathscr{F}}^{*}_{\alpha+1}$, so the canonical package generics are hereditarily symmetric by Lemmas 5.45 and 5.46.
-Thus each successor filter explicitly includes the diagonal subgroups protecting the representative coordinates of the active packages.
-At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group/filter as provided by the symmetric-iteration framework (API items (b)–(e), especially filter generation/completeness at limits, [5, Thm.
-5.5]).
+At successor stages, the group $\mathscr{G}_{\alpha+1}$ and filter $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ are those determined by the symmetric-iteration framework together with the diagonal infrastructure from Subsection 5.2 and the modified stage filters from Definition 5.50.
+Concretely, $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter generated by pullbacks $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$, together with admissible subgroups $\Delta_{\alpha+1}^{\uparrow}(E,D)$.
+In particular, for each active package coordinate $d$ at stage $\alpha+1$ whose countable support is witnessed by $E_{d}$, the generator $\Delta_{\alpha+1}^{\uparrow}(E_{d},\{d\})$ belongs to $\tilde{\mathscr{F}}^{*}_{\alpha+1}$, so the canonical package generics are hereditarily symmetric by Lemmas 5.47 and 5.48.
+At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group as provided by the symmetric-iteration framework (Remark 2.4(a),(b)).
+The limit-stage filter is $\tilde{\mathscr{F}}^{*}_{\lambda}$ as defined in Definition 5.50, and its normality and $\omega_{1}$-completeness are recorded in Lemma 5.54.
+For set-localization in the $\mathrm{Ord}$-length model we use only the stage-support lemma (Lemma 5.64) below.
 We write $\mathcal{M}$ for the final symmetric model of the $\mathrm{Ord}$-length iteration.
@@
 Indeed, the usual “two-step collapse” phenomena from forcing do not directly apply in the symmetric context, and the literature discusses serious obstacles to such a reduction.
-It remains open in general to what extent arbitrary symmetric iterations can be compressed to a single symmetric extension; see the discussion in [9, 18].
+It remains open in general to what extent arbitrary symmetric iterations can be compressed to a single symmetric extension; see the discussion in [9].
 5.6 What the iteration yields
 Let $\mathcal{M}$ denote the final symmetric model obtained from the iteration.
+Lemma 5.64 (Stage-localization of set-names at class length).
+Let $\dot{x}$ be a set-sized $\mathbb{P}_{\mathrm{Ord}}$-name, and define its stage-support
+$I(\dot{x})\ :=\ \bigcup\{\operatorname{supp}(p)\cup I(\dot{y}):\langle\dot{y},p\rangle\in\dot{x}\},$
+where $\operatorname{supp}(p)\subseteq\mathrm{Ord}$ is the (countable) support of the condition $p$. Then $I(\dot{x})$ is a set of ordinals, hence $\alpha:=\sup(I(\dot{x}))+1<\mathrm{Ord}$. Moreover:
+  1. 1.
+    $\dot{x}$ is in fact a $\mathbb{P}_{\alpha}$-name (all conditions appearing in $\dot{x}$ lie in $\mathbb{P}_{\alpha}$),
+  2. 2.
+    for any $V$-generic $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ and $G_{\alpha}:=G_{\infty}\cap\mathbb{P}_{\alpha}$, we have $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$.
+Proof.
+Define $I(\dot{x})$ by recursion on the rank of the name $\dot{x}$.
+Since $\dot{x}$ is set-sized, Replacement and Union yield that $I(\dot{x})$ is a set of ordinals, hence bounded in $\mathrm{Ord}$.
+Let $\alpha=\sup(I(\dot{x}))+1$.
+Every condition $p$ appearing in $\dot{x}$ has $\operatorname{supp}(p)\subseteq I(\dot{x})\subseteq\alpha$, so $p\in\mathbb{P}_{\alpha}$, hence $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name.
+The valuation recursion for $\dot{x}^{G_{\infty}}$ consults only conditions appearing in $\dot{x}$, all of which lie in $\mathbb{P}_{\alpha}$, hence depends only on $G_{\alpha}$.
+Therefore $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$.
+∎
 Proposition 5.65 (Final localized splitting $\mathsf{PP}^{\mathrm{split}}\!\restriction T$).
@@
 Proof.
-Work in the final symmetric model $\mathcal{M}$. Let $f:Y\twoheadrightarrow X$ be a surjection in $\mathcal{M}$ with $X,Y\subseteq T$.
-Choose a hereditarily symmetric name $\dot{f}$ such that $f=\dot{f}^{\,G_{\infty}}$.
-Let $I\subseteq\mathrm{Ord}$ be the countable stage support of $\dot{f}$ (API item (e)), and let $\alpha:=\sup(I)+1$.
-Then $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name and $f_{\alpha}:=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}$.
-Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name, its interpretation is unchanged by further forcing: $\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}$, so $f=f_{\alpha}$ in $\mathcal{M}$.
-Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$–name of bounded rank, and no forcing past stage $\alpha$ adds sets of that rank, Theorem 9.2 of Karagila [9] implies that $\dot{f}^{G_{\lambda}}=\dot{f}^{G_{\alpha}}$.
-Hence $f=f_{\alpha}$ in $\mathcal{M}$.
-If $f_{\alpha}$ already has a section in $M_{\alpha}$, then $f$ splits in $\mathcal{M}$ and we are done. Otherwise, $f_{\alpha}$ is eligible for scheduling by the bookkeeping.
-By the bookkeeping (Definition 5.59), at some successor stage we schedule the orbit package $\mathbb{Q}_{[f_{\alpha}]}$.
+Work in the final symmetric model $\mathcal{M}$. Recall Definition 4.8. Let $f:Y\twoheadrightarrow X$ be a surjection in $\mathcal{M}$ with $X,Y\subseteq T$.
+Choose a hereditarily symmetric set-name $\dot{f}$ such that $f=\dot{f}^{\,G_{\infty}}$.
+By Lemma 5.64, letting $\alpha:=\sup(I(\dot{f}))+1$, we may view $\dot{f}$ as a $\mathbb{P}_{\alpha}$-name and we have
+$f_{\alpha}:=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}\qquad\text{and}\qquad\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}.$
+Thus $f=f_{\alpha}\in M_{\alpha}\subseteq\mathcal{M}$.
+If $f_{\alpha}$ already has a section in $M_{\alpha}$, then $f$ splits in $\mathcal{M}$ and we are done.
+Otherwise, $f_{\alpha}$ is eligible for scheduling by the bookkeeping.
+By the final paragraph of Definition 5.61, there is some stage $\beta\geq\alpha$ such that either $f_{\alpha}$ already splits in $M_{\beta}$ (in which case we are done), or else the successor iterand at stage $\beta$ is the orbit $\mathsf{PP}$-package $\mathbb{Q}_{[f_{\alpha}]}$.
+Assume we are in the latter case, and fix such a stage $\beta$.
 Let $\dot{s}_{f_{\alpha}}$ be the canonical $\mathbb{Q}_{[f_{\alpha}]}$-name for the generic section in the $f_{\alpha}$-coordinate.
-By Lemma 5.45, $\dot{s}_{f_{\alpha}}$ is hereditarily symmetric for the package symmetric system $(\mathbb{Q}_{[f_{\alpha}]},\mathscr{G}_{[f_{\alpha}]},\mathscr{F}_{[f_{\alpha}]})$ in $M_{\beta}$.
-By Lemma 5.47, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{s}_{f_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence $s_{f_{\alpha}}:=\dot{s}_{f_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
+In the full package extension, this coordinate generic yields a section of $f_{\alpha}$ (cf.
+Proposition 4.15).
+By Lemma 5.47, $\dot{s}_{f_{\alpha}}$ is hereditarily symmetric for the package symmetric system $(\mathbb{Q}_{[f_{\alpha}]},\mathscr{G}_{[f_{\alpha}]},\mathscr{F}_{[f_{\alpha}]})$ over $M_{\beta}$.
+By Lemma 5.49, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{s}_{f_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence
+$s_{f_{\alpha}}:=\dot{s}_{f_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}.$
 Since $s_{f_{\alpha}}$ is a section of $f_{\alpha}$, it remains a section of $f$ in the final model $\mathcal{M}$.
@@
 Remark 5.66 (Countable-support names at limits of cofinality $\omega$).
-At limits $\lambda$ with $\operatorname{cf}(\lambda)=\omega$, a hereditarily symmetric name may have countable support that is cofinal in $\lambda$, so stage-bounding arguments do not apply.
-The iteration machinery of [5] handles such limits using limit-filter $\omega_{1}$-completeness together with the appropriate block/factorization analysis (cf.
-[5, Rem.
-3.9, Thm.
-5.5]), and we rely on those limit-stage conclusions.
+At limits $\lambda$ with $\operatorname{cf}(\lambda)=\omega$, a hereditarily symmetric name may have countable support cofinal in $\lambda$, so the usual stage-bounding arguments do not apply.
+In this manuscript we handle such limits by using the modified limit filter (Definition 5.50), whose normality and $\omega_{1}$-completeness are recorded in Lemma 5.54, together with the core admissibility lemma (Lemma 5.52).
+This matches the treatment of countable-cofinality limits in the general iteration framework ([5, §3.5–§3.6]).
 Proposition 5.67 ($A$ is not well-orderable).
@@
 Step 3: Choose a nontrivial transposition.
-Since $E$ is countable and $\omega_{1}$ is uncountable, there exist distinct $\beta,\gamma\in\omega_{1}\setminus E$.
+Since $E$ is countable, $\delta:=\sup(E)<\omega_{1}$.
+Choose distinct $\beta,\gamma\in\omega_{1}$ with $\delta<\beta<\gamma$.
+Then $\beta,\gamma\notin E$.
 Let $\pi=(\beta\ \gamma)\in\operatorname{Sym}(\omega_{1})$ be the transposition swapping $\beta$ and $\gamma$.
@@
 Proof.
-We argue by induction along the symmetric iteration. The base model $M_{0}=\mathcal{N}$ satisfies $\mathsf{DC}$ by Theorem 3.11.
-For the successor step, assume $M_{\alpha}\models\mathsf{ZF}+\mathsf{DC}$.
-The successor stage is obtained by forcing with the stage symmetric system at $\alpha$ and then taking the hereditarily symmetric submodel $M_{\alpha+1}$.
-By Lemma 5.6, each orbit package $\mathbb{S}_{[f]}$ used at stage $\alpha$ is $\omega_{1}$-closed (and hence adds no new $\omega$-sequences of ordinals), and the finite-support product $\dot{\mathbb{S}}_{\alpha}$ is again $\omega_{1}$-closed for the same reason.
-By Lemma 5.16, the relevant stage filter is $\omega_{1}$-complete.
-Therefore the $\mathsf{DC}$-preservation theorem for symmetric extensions applies (API item (c), cf.
-[5, Thm.
-3.20]), yielding $M_{\alpha+1}\models\mathsf{DC}$.
-At limit stages, $\mathsf{DC}$ is preserved by the countable-support symmetric iteration machinery (API item (c), and the limit-filter generation/completeness from API item (d); cf.
-[5, Thm.
-3.20] and [5, Thm.
-5.5] for the limit construction).
-Thus every stage model satisfies $\mathsf{DC}$, in particular the final model $\mathcal{M}\models\mathsf{DC}$. ∎
-Definition 5.66 (Supported subforcing and supported names).
-Fix a stage $\lambda$ and a globally admissible pair $(E,D)$ at stage $\lambda$. Define
-$\mathbb{P}_{\lambda}\restriction(E,D)=\{p\in\mathbb{P}_{\lambda}:p\restriction 0\in\operatorname{Add}(\omega,E)\ \text{and}\ \operatorname{supp}_{>0}(p)\subseteq D\},$
-where $\operatorname{supp}_{>0}(p):=\{c\in\mathrm{Coords}_{<\lambda}:p(c)\neq 1_{c}\}$ is the set of nontrivial higher-stage coordinates of $p$.
-A $\mathbb{P}_{\lambda}$-name $\dot{x}$ is $(E,D)$-supported if every condition appearing in the transitive closure $\operatorname{tc}(\dot{x})$ belongs to $\mathbb{P}_{\lambda}\restriction(E,D)$.
-Lemma 5.67 (Generic restriction to a supported subforcing).
-Let $(E,D)$ be globally admissible at stage $\lambda$, and let $\mathbb{Q}=\mathbb{P}_{\lambda}\restriction(E,D)$.
-Then every maximal antichain $A\subseteq\mathbb{Q}$ is also a maximal antichain in $\mathbb{P}_{\lambda}$.
-Consequently, if $G_{\lambda}\subseteq\mathbb{P}_{\lambda}$ is $V$-generic, then $G_{E,D}:=G_{\lambda}\cap\mathbb{Q}$ is $V$-generic for $\mathbb{Q}$, and $V[G_{E,D}]\subseteq V[G_{\lambda}]$.
-Proof.
-Let $A\subseteq\mathbb{Q}$ be maximal, and let $p\in\mathbb{P}_{\lambda}$.
-Put $p^{\prime}:=p\restriction(E,D)\in\mathbb{Q}$.
-Choose $q\in A$ compatible with $p^{\prime}$ in $\mathbb{Q}$, and let $r\in\mathbb{Q}$ witness this, i.e.
-$r\leq_{\mathbb{Q}}q,p^{\prime}$.
-Since $p$ extends $p^{\prime}$ on the coordinates that $r$ constrains, and $r$ is trivial off $E\cup D$, we can amalgamate $r$ with $p$ to obtain a common extension in $\mathbb{P}_{\lambda}$.
-Thus $q$ is compatible with $p$ in $\mathbb{P}_{\lambda}$, so $A$ is predense in $\mathbb{P}_{\lambda}$ and hence maximal there.
-Now $G_{\lambda}$ meets every maximal antichain in $\mathbb{P}_{\lambda}$, so $G_{\lambda}\cap A\neq\emptyset$.
-Since $A\subseteq\mathbb{Q}$, this gives $(G_{\lambda}\cap\mathbb{Q})\cap A\neq\emptyset$, proving $\mathbb{Q}$-genericity.
-∎
-Lemma 5.68 (Supported names are $\mathbb{Q}$-names and evaluate correctly).
-Let $(E,D)$ be globally admissible at stage $\lambda$, let $\mathbb{Q}=\mathbb{P}_{\lambda}\restriction(E,D)$, and let $G_{\lambda}\subseteq\mathbb{P}_{\lambda}$ be $V$-generic with $G_{E,D}=G_{\lambda}\cap\mathbb{Q}$.
-If $\dot{x}$ is $(E,D)$-supported, then $\dot{x}$ is (canonically) a $\mathbb{Q}$-name and
-$\dot{x}^{G_{\lambda}}=\dot{x}^{G_{E,D}}.$
-Proof.
-Since every condition appearing in $\operatorname{tc}(\dot{x})$ lies in $\mathbb{Q}$, we may view $\dot{x}$ as a $\mathbb{Q}$-name (by the usual identification of names with the same transitive closure).
-The valuation equality is proved by induction on name-rank: for each pair $(\dot{y},p)\in\dot{x}$ we have $p\in\mathbb{Q}$, hence $p\in G_{\lambda}$ iff $p\in G_{E,D}$, and the induction hypothesis applies to $\dot{y}$.
-∎
-Lemma 5.69 (Triviality on supported names).
-If $\dot{x}$ is $(E,D)$-supported and $g\in\Delta_{\lambda}^{\uparrow}(E,D)$, then $g\cdot\dot{x}=\dot{x}$.
-Proof.
-By induction on $\operatorname{rank}(\dot{x})$.
-If $p\in\mathbb{P}_{\lambda}\restriction(E,D)$, then $g(p)=p$ since the stage–0 projection of $g$ fixes $E$ pointwise and $g$ fixes every coordinate in $D$ pointwise.
-Now apply the recursive definition of the action on names.
+Let $A$ be a set in $\mathcal{M}$ and let $R\subseteq A\times A$ be a relation in $\mathcal{M}$ such that $\forall x\in A\,\exists y\in A\ (xRy)$.
+We show there is an $\omega$-sequence $\langle a_{n}:n<\omega\rangle$ in $\mathcal{M}$ with $a_{n}Ra_{n+1}$ for all $n$.
+Choose set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-names $\dot{A},\dot{R}\in\operatorname{HS}_{\mathrm{Ord}}$ such that $A=\dot{A}^{G_{\infty}}$ and $R=\dot{R}^{G_{\infty}}$.
+By Lemma 5.64, letting $\alpha:=\sup(I(\dot{A})\cup I(\dot{R}))+1$, we have $A,R\in M_{\alpha}$.
+By Lemma 5.54(iv), the stage filter $\tilde{\mathscr{F}}^{*}_{\alpha}$ is $\omega_{1}$-complete.
+Therefore, by Remark 2.4(d), $M_{\alpha}\models\mathsf{DC}$.
+Applying $\mathsf{DC}$ inside $M_{\alpha}$ to $(A,R)$, we obtain a sequence $\langle a_{n}:n<\omega\rangle\in M_{\alpha}\subseteq\mathcal{M}$ with $a_{n}Ra_{n+1}$ for all $n$.
+Hence $\mathcal{M}\models\mathsf{DC}$.
 ∎
@@
 Successor.
-Assume $M_{\alpha}\models\mathsf{SVC}(S)$ and let $X\in M_{\alpha+1}$ be nonempty.
-Choose a hereditarily symmetric $\mathbb{P}_{\alpha+1}$-name $\dot{X}\in\operatorname{HS}_{\alpha+1}$ (so $\dot{X}\in M_{\alpha}$) with $X=\dot{X}^{G_{\alpha+1}}$.
-By $\mathsf{SVC}(S)$ in $M_{\alpha}$, there exist an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow\dot{X}$ in $M_{\alpha}$.
-Fix $p_{0}\in\mathbb{P}_{\alpha+1}$ and a name $\dot{x}_{0}$ such that $p_{0}\Vdash\dot{x}_{0}\in\dot{X}$.
-Define a $\mathbb{P}_{\alpha+1}$-name $\dot{h}$ for a function $S\times\eta\to\dot{X}$ by: for $s\in S$ and $\xi<\eta$, write $e(s,\xi)=\langle\dot{y},p\rangle$ and set
-$\dot{h}(s,\xi)=\begin{cases}\dot{y}&\text{if }p\in\dot{G}_{\alpha+1},\\ \dot{x}_{0}&\text{otherwise.}\end{cases}$
-Let $h:=\dot{h}^{G_{\alpha+1}}$. Then $h:S\times\eta\to X$, and $h$ is surjective by construction.
-Moreover, $\dot{h}\in\operatorname{HS}_{\alpha+1}$: it is definable from parameters in $M_{\alpha}$, and any automorphism fixing $\dot{X}$ (hence lying in $\tilde{\mathscr{F}}^{*}_{\alpha+1}$) fixes $\dot{h}$ as well.
-Thus $h\in M_{\alpha+1}$, so $M_{\alpha+1}\models\mathsf{SVC}(S)$.
+Assume $M_{\alpha}\models\mathsf{SVC}(S)$, and let $X\in M_{\alpha+1}$ be nonempty.
+Choose $\dot{X}\in\operatorname{HS}_{\alpha+1}$ with $X=\dot{X}^{G_{\alpha+1}}$.
+Since $\dot{X}\in M_{\alpha}$ (HS-names are formed in the stage ground), we may apply $\mathsf{SVC}(S)$ in $M_{\alpha}$ to the set $\dot{X}$ and obtain an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow\dot{X}$ in $M_{\alpha}$.
+Fix $p_{0}\in\mathbb{P}_{\alpha+1}$ and a name $\dot{x}_{0}$ with $p_{0}\Vdash\dot{x}_{0}\in\dot{X}$.
+Work in the Boolean completion $\mathbb{B}_{\alpha+1}$ of $\mathbb{P}_{\alpha+1}$ in the background universe.
+For each $u=(s,\xi)\in S\times\eta$, write $e(u)=\langle\dot{y}_{u},p_{u}\rangle$.
+By the mixing lemma in $\mathbb{B}_{\alpha+1}$, there is a $\mathbb{P}_{\alpha+1}$-name $\dot{z}_{u}$ such that
+$1\Vdash\dot{z}_{u}=\begin{cases}\dot{y}_{u}&\text{if }p_{u}\in\dot{G}_{\alpha+1},\\ \dot{x}_{0}&\text{otherwise.}\end{cases}$
+Let $\dot{h}$ be the canonical name for the function $u\mapsto\dot{z}_{u}$ (so $\dot{h}$ is a $\mathbb{P}_{\alpha+1}$-name for a function $S\times\eta\to\dot{X}^{G_{\alpha+1}}$).
+Then $h:=\dot{h}^{G_{\alpha+1}}$ is a function $S\times\eta\to X$, and it is surjective: if $x\in X$, pick $\langle\dot{y},p\rangle\in\dot{X}$ with $p\in G_{\alpha+1}$ and $\dot{y}^{G_{\alpha+1}}=x$, and then pick $u$ with $e(u)=\langle\dot{y},p\rangle$; for that $u$, we have $h(u)=x$.
+Moreover, $\dot{h}\in\operatorname{HS}_{\alpha+1}$: it is built in the stage ground $M_{\alpha}$ from $e$ and $\dot{X}$, and any automorphism in $\mathscr{G}_{\alpha+1}$ fixing $\dot{X}$ fixes the resulting mixed values $\dot{z}_{u}$ and hence $\dot{h}$.
+Therefore $h\in M_{\alpha+1}$, so $M_{\alpha+1}\models\mathsf{SVC}(S)$.
 Limit.
 Let $\lambda$ be limit and assume $M_{\beta}\models\mathsf{SVC}(S)$ for all $\beta<\lambda$.
-Let $X\in M_{\lambda}$ be nonempty.
-By API item (e), $X$ has a hereditarily symmetric name with countable stage support bounded in $\lambda$, so $X\in M_{\beta}$ for some $\beta<\lambda$.
-By the inductive hypothesis, $M_{\beta}\models\mathsf{SVC}(S)$, hence $M_{\lambda}\models\mathsf{SVC}(S)$ as well.
+Let $X\in M_{\lambda}$ be nonempty, and choose $\dot{X}\in\operatorname{HS}_{\lambda}$ with $X=\dot{X}^{G_{\lambda}}$.
+By the countable-support nature of the iteration and the HS-support analysis ((cf.
+the countable-support analysis of HS-names in Section 3 and the preceding discussion of supports at limits), $\dot{X}$ depends on only countably many stages.
+Let $\langle\lambda_{n}:n<\omega\rangle$ be an increasing sequence of ordinals below $\lambda$ capturing those stages.
+For each $n$, let $X_{n}$ be the interpretation of the corresponding stage-truncation of $\dot{X}$; then $X_{n}\in M_{\lambda_{n}}$ and $X=\bigcup_{n<\omega}X_{n}$.
+By the inductive hypothesis, each $M_{\lambda_{n}}\models\mathsf{SVC}(S)$, so for each $n$ there is an ordinal $\eta_{n}$ and a surjection $e_{n}:S\times\eta_{n}\twoheadrightarrow X_{n}$ in $M_{\lambda_{n}}\subseteq M_{\lambda}$.
+Since $M_{\lambda}\models\mathsf{DC}$ (by Remark 2.4(d) and Lemma 5.54(iv)), $M_{\lambda}$ satisfies $\mathsf{AC}_{\omega}$, so we may choose such witnesses $\langle\eta_{n},e_{n}:n<\omega\rangle$ in $M_{\lambda}$.
+Let $\theta:=\sup_{n}\eta_{n}$ and define $e:S\times(\omega\cdot\theta)\twoheadrightarrow X$ by sending $(s,\omega\cdot\xi+n)$ to $e_{n}(s,\xi)$ when $\xi<\eta_{n}$, and to a fixed element of $X$ otherwise.
+Then $e\in M_{\lambda}$ and $e$ is surjective onto $X$.
+Hence $M_{\lambda}\models\mathsf{SVC}(S)$.
 Finally, $M_{\lambda}\models\mathsf{SVC}^{+}(T)$ follows from Fact 4.3. ∎
-Proposition 5.71 (Final $\mathsf{AC}_{\mathsf{WO}}$).
-$\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$.
+Corollary 5.70.
+Assume that $M_{\lambda}\models\mathsf{SVC}(S)$ for all $\lambda\in\mathrm{Ord}$.
+Then the final class-length model $\mathcal{M}$ satisfies $\mathsf{SVC}(S)$.
+Consequently, $\mathcal{M}\models\mathsf{SVC}^{+}(T)$ (for the relevant parameter $T$).
 Proof.
-Work in the final symmetric model $\mathcal{M}$.
-By Lemma 5.70, we have $\mathcal{M}\models\mathsf{SVC}(S)$.
-We show that $\mathcal{M}\models\mathsf{AC}_{<\aleph^{*}(S)^{M}}$; by the Ryan–Smith equivalences this implies $\mathsf{AC}_{\mathsf{WO}}$.
-Let $\lambda<\aleph^{*}(S)^{M}$ be an ordinal and let $f:Y\twoheadrightarrow\lambda$ be a surjection in $\mathcal{M}$.
-By $\mathsf{SVC}(S)$, fix an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow Y$ in $\mathcal{M}$.
-Let $g:=f\circ e:S\times\eta\twoheadrightarrow\lambda$.
-Let $\dot{g}$ be a hereditarily symmetric name such that $g=\dot{g}^{\,G_{\infty}}$.
-Let $I\subseteq\mathrm{Ord}$ be the countable stage support of $\dot{g}$ (API item (e)), and let $\alpha:=\sup(I)+1$.
-Then $\dot{g}$ is a $\mathbb{P}_{\alpha}$-name, and $g_{\alpha}:=\dot{g}^{\,G_{\alpha}}\in M_{\alpha}$.
-By the bookkeeping (Definition 5.59), there is some successor stage $\beta>\alpha$ at which the orbit package $\mathbb{R}_{[g_{\alpha}]}$ is scheduled.
-Let $\dot{r}_{g_{\alpha}}$ be the canonical $\mathbb{R}_{[g_{\alpha}]}$-name for the generic right inverse in the $g_{\alpha}$-coordinate.
-By Lemma 5.46, the canonical $\mathbb{R}_{[g_{\alpha}]}$-name $\dot{r}_{g_{\alpha}}$ is hereditarily symmetric for the package symmetric system in $M_{\beta}$.
-By Lemma 5.47, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{r}_{g_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence $r_{g_{\alpha}}:=\dot{r}_{g_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$, and $r_{g_{\alpha}}$ is a right inverse of $g$ in $\mathcal{M}$.
-Now define $t:=e\circ r_{g_{\alpha}}:\lambda\to Y$. Then for all $\xi<\lambda$,
-$f\bigl(t(\xi)\bigr)=f\bigl(e(r_{g_{\alpha}}(\xi))\bigr)=g\bigl(r_{g_{\alpha}}(\xi)\bigr)=\xi,$
-so $t$ is a right inverse to $f$.
-Thus every surjection $Y\twoheadrightarrow\lambda$ with $\lambda<\aleph^{*}(S)^{M}$ splits in $\mathcal{M}$, and therefore $\mathcal{M}\models\mathsf{AC}_{<\aleph^{*}(S)^{M}}$.
-Finally, since $\mathcal{M}\models\mathsf{SVC}(S)$, the Ryan–Smith theorem (see Remark 4.11) yields $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$. ∎
-Theorem 5.72 (Final model).
-There is a transitive model $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$.
+Let $x\in\mathcal{M}$.
+Choose a set-sized HS name $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$ with $x=\dot{x}^{G_{\infty}}$.
+By Lemma 5.64, letting $\alpha:=\sup(I(\dot{x}))+1$, we have $x=\dot{x}^{G_{\alpha}}\in M_{\alpha}$.
+Since $M_{\alpha}\models\mathsf{SVC}(S)$ (by Lemma 5.69), there is an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow x$ in $M_{\alpha}$, hence also in $\mathcal{M}$.
+Thus $\mathcal{M}\models\mathsf{SVC}(S)$.
+The final statement follows from Fact 4.3.
+∎
+Remark 5.71 (Structural characterization of $\mathsf{SVC}$ (context only)).
+Blass [2] proved that $\mathsf{SVC}$ is equivalent to: $\mathsf{AC}$ is forceable over the model by a set forcing.
+We do not use deeper structural characterizations here; see, e.g., work in set-theoretic geology such as Usuba [19] for related perspectives.
+In particular, Corollary 5.70 implies that $\mathsf{AC}$ is set-forceable over $\mathcal{M}$.
+Proposition 5.72 ($\mathsf{AC}_{\mathsf{WO}}$ in the final model).
+The final symmetric model $\mathcal{M}$ satisfies $\mathsf{AC}_{\mathsf{WO}}$.
 Proof.
-By Theorem 5.56, $\mathcal{M}\models\mathsf{ZF}$, and by Proposition 5.65, $\mathcal{M}\models\mathsf{DC}$.
-By Proposition 5.62, $\mathcal{M}\models\mathsf{PP}^{\mathrm{split}}\!\restriction T$, hence $\mathcal{M}\models\mathsf{PP}\!\restriction T$.
-By Proposition 5.71, $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$.
-By Lemma 5.70, $\mathcal{M}\models\mathsf{SVC}^{+}(T)$ for $T=\mathcal{P}(S)$, so the hypotheses of the Ryan–Smith localization theorem (Theorem 4.5, i.e., [17, Prop.
-3.17] applied with Ryan–Smith’s parameter $S$ taken to be our $T$) hold in $\mathcal{M}$.
-Therefore $\mathcal{M}\models\mathsf{PP}$.
-Finally, Proposition 5.64 gives $\mathcal{M}\models\neg\mathsf{AC}$. ∎
+Let $f:S\times\eta\twoheadrightarrow\lambda$ be a surjection in $\mathcal{M}$, where $\eta,\lambda\in\mathrm{Ord}$. We show $f$ has a right inverse in $\mathcal{M}$.
+Choose a set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{f}\in\operatorname{HS}_{\mathrm{Ord}}$ with $f=\dot{f}^{G_{\infty}}$.
+By Lemma 5.64, there is a set ordinal $\alpha$ such that $f\in M_{\alpha}$.
+If $f$ already has a right inverse in $M_{\alpha}$, we are done.
+Otherwise, $f$ is eligible for scheduling by the $\mathsf{AC}_{\mathsf{WO}}$-bookkeeping clause in Definition 5.61.
+Hence by the guarantee clause in that definition, there exists a stage $\beta\geq\alpha$ such that either $f$ already has a right inverse in $M_{\beta}$, or else the successor iterand at stage $\beta$ includes the orbit $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{[f]}$.
+Assume the latter case and fix such $\beta$.
+Let $\dot{r}_{f}$ be the canonical $\mathbb{R}_{[f]}$-name for the $f$-coordinate generic right inverse.
+In the full package extension, this coordinate generic yields a right inverse for $f$ (cf.
+Proposition 4.27).
+By Lemma 5.48, $\dot{r}_{f}$ is hereditarily symmetric for $(\mathbb{R}_{[f]},\mathscr{G}_{[f]},\mathscr{F}_{[f]})$ over $M_{\beta}$.
+By Lemma 5.49, the canonical $\mathbb{P}_{\beta+1}$-lift $\dot{r}_{f}^{*}$ is hereditarily symmetric at stage $\beta+1$.
+Therefore
+$r_{f}:=\dot{r}_{f}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$
+is a right inverse for $f$ in $\mathcal{M}$. Hence $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$. ∎
+Theorem 5.73 (Main theorem).
+Let $V\models\mathsf{ZFC}$.
+In the class forcing extension $V[G]$ obtained by the $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen-seed symmetric model $\mathcal{N}$ described above, the final symmetric model $\mathcal{M}$ satisfies
+$\mathsf{ZF}\ +\ \mathsf{DC}\ +\ \mathsf{PP}\ +\ \mathsf{AC}_{\mathsf{WO}}\ +\ \neg\mathsf{AC}.$
+Proof.
+$\mathsf{ZF}$ holds in $\mathcal{M}$ by Theorem 5.58.
+$\mathsf{DC}$ holds by Proposition 5.68.
+$\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.65 and 5.72.
+$\mathsf{AC}_{\mathsf{WO}}$ holds by Proposition 5.72.
+Finally, $\neg\mathsf{AC}$ holds by Proposition 5.67.
+∎
+Corollary 5.74 (Ordering Principle).
+The final model $\mathcal{M}$ satisfies the Ordering Principle: every set admits a linear order.
+Proof.
+Work in $\mathcal{M}$. Fix any set $X$. By Corollary 5.70, $\mathcal{M}\models\mathsf{SVC}(S)$, so choose an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow X$.
+By Theorem 4.5 together with Propositions 5.65 and 5.72, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
+Pulling back the lexicographic order on $S\times\eta$ along $i$ yields a linear order on $X$. ∎
+Remark 5.75 (Kinna–Wagner principles).
+For $n\geq 1$, the Kinna–Wagner principle $KWP_{n}$ states: for every set $X$ there exists an ordinal $\theta$ such that $X\hookrightarrow\mathcal{P}^{n}(\theta)$ (where $\mathcal{P}^{n}$ is the $n$-fold iterated powerset).
+(1) $KWP_{2}$ from $\mathsf{SVC}(S)$. Work in $\mathcal{M}$ and fix a set $X$. By Corollary 5.70, choose an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow X$. The map
+$x\ \longmapsto\ e^{-1}(\{x\})$
+is an injection of $X$ into $\mathcal{P}(S\times\eta)$.
+Since $S=A^{\omega}$ injects into $2^{\omega}$, we have $S\times\eta\hookrightarrow 2^{\omega}\times\eta$.
+Identifying $2^{\omega}$ with $\mathcal{P}(\omega)$, define an injection
+$\mathcal{P}(\omega)\times\eta\ \hookrightarrow\ \mathcal{P}(\omega\times\eta),\qquad(r,\xi)\longmapsto\{(n,\xi):n\in r\}.$
+Thus $S\times\eta\hookrightarrow\mathcal{P}(\omega\times\eta)$, so $\mathcal{P}(S\times\eta)\hookrightarrow\mathcal{P}(\mathcal{P}(\omega\times\eta))=\mathcal{P}^{2}(\omega\times\eta)$.
+Because $\omega\times\eta$ is well-orderable of type $\omega\cdot\eta$, we have $\mathcal{P}^{2}(\omega\times\eta)\cong\mathcal{P}^{2}(\omega\cdot\eta)$, and hence $X\hookrightarrow\mathcal{P}^{2}(\theta)$ for $\theta=\omega\cdot\eta$.
+This proves $KWP_{2}$ in $\mathcal{M}$.
+(2) $KWP_{1}$ using $\mathsf{PP}$.
+By Theorem 4.5 together with Propositions 5.65 and 5.72, $\mathcal{M}\models\mathsf{PP}$.
+Apply $\mathsf{PP}$ to the surjection $e:S\times\eta\twoheadrightarrow X$ to obtain an injection $i:X\hookrightarrow S\times\eta$.
+Composing with the injection $S\times\eta\hookrightarrow\mathcal{P}(\omega\times\eta)\cong\mathcal{P}(\omega\cdot\eta)$ from (1), we obtain $X\hookrightarrow\mathcal{P}(\theta)$ for $\theta=\omega\cdot\eta$.
+Thus $KWP_{1}$ holds in $\mathcal{M}$.
+6 Conclusion
+Starting from a ground model $V\models\mathsf{ZFC}$, we define an $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen symmetric seed model $\mathcal{N}$ and let $\mathcal{M}$ be the resulting final symmetric model.
+The iteration is arranged so that $\mathcal{M}$ satisfies $\mathsf{DC}$ (Proposition 5.68), satisfies $\mathsf{AC}_{\mathsf{WO}}$ (Proposition 5.72), and still contains the non-well-orderable Cohen set $A$ (Proposition 5.67).
+We force the localized splitting principle $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ for the fixed parameter $T$ (Proposition 5.65) and verify $\mathsf{SVC}^{+}(T)$ in $\mathcal{M}$ via persistence of $\mathsf{SVC}(S)$ (Lemma 5.69 and Corollary 5.70).
+Ryan–Smith localization (Theorem 4.5) then yields $\mathcal{M}\models\mathsf{PP}$, completing the proof that $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\mathsf{AC}_{\mathsf{WO}}+\neg\mathsf{AC}$ (Theorem 5.73).
+As recorded in the final section, this also implies the Ordering Principle (Corollary 5.74) and the stated Kinna–Wagner consequences (Remark 5.75).

### v5

@@
 Set-sized stages are treated in the usual set-theoretic background.
-To define and analyze the $\mathrm{Ord}$-length iteration as a definable class forcing (and to use the forcing theorem for $\mathbb{P}_{\mathrm{Ord}}$), we work in a background class theory supporting the relevant recursions, namely $\mathsf{GBC}+\mathsf{ETR}$, using a class well-order of $V$ only where it is explicitly invoked for bookkeeping; see Section 2 and Remark 5.59.
+To define and analyze the $\mathrm{Ord}$-length iteration as a definable class forcing (in particular, to carry out the relevant class recursions), we work in a background class theory supporting the relevant recursions, namely $\mathsf{GBC}+\mathsf{ETR}$, using a class well-order of $V$ only where it is explicitly invoked for bookkeeping; see Section 2 and Remark 5.60.
 Remark 1.1 (The axiom $\mathsf{AC}_{\mathsf{WO}}$).
-Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes choice for well-orderable sets: every well-orderable set of nonempty sets admits a choice function.
-Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds.
-This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used when applying Ryan–Smith localization (Theorem 4.5).
+Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes choice for well-orderable sets: every well-orderable family of nonempty sets admits a choice function (equivalently: for every family $\langle X_{i}:i\in I\rangle$ of nonempty sets indexed by a well-orderable set $I$, there is a choice function $c$ with $c(i)\in X_{i}$ for all $i\in I$).
+Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds. This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used when applying Ryan–Smith localization (Theorem 4.5).
 How $\mathsf{PP}$ is obtained.
@@
   3. (c)
-    Limit filters. In the relevant cofinality cases, the canonical limit filter generated from earlier-stage filters is normal and $\omega_{1}$-complete; see [5, §Limit filters].
+    Limit filters.
+The limit-stage filter used in this manuscript is the modified filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.50).
+Its normality and $\omega_{1}$-completeness are proved here (Lemma 5.55).
+We only use [5] for the general iteration template and projection/coherence infrastructure leading to the limit stage.
   4. (d)
     Bounded-stage $\mathsf{DC}$.
-For set ordinals $\lambda$, assuming $V\models\mathsf{ZFC}$, if the stage filter is $\omega_{1}$-complete then the stage symmetric model $M_{\lambda}=V(\mathbb{P}_{\lambda})_{\operatorname{HS}}$ satisfies $\mathsf{DC}$; see [5, §Dependent Choice and Choice Fragments].
+For any set ordinal $\lambda$, assuming $V\models\mathsf{ZFC}$, if the stage filter is $\omega_{1}$-complete then the stage symmetric model $M_{\lambda}=V(\mathbb{P}_{\lambda})_{\operatorname{HS}}$ satisfies $\mathsf{DC}$.
+This is proved in [5, Theorem (DC Preservation (set-length stages))].
 We stress that [5] does not assert $\mathsf{DC}$ for the final class-length symmetric model in general.
@@
 Subsection 5.5 invokes the symmetric-iteration framework to assemble the class-length iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$, with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and successor stages $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$ where $\dot{\mathbb{S}}_{\alpha}$ is the finite product of the scheduled orbit package factors.
+By Lemma 5.65, the canonical $\mathbb{P}_{\alpha}$-names for these orbit-package iterands are hereditarily symmetric, so the two-step presentation is well-defined in the stage ground $M_{\alpha}$.
   6. (6)
@@
 $\pi^{0}_{\lambda}(g):=\pi^{0}_{\alpha}\bigl(\pi^{\alpha}_{\lambda}(g)\bigr)\quad\text{for any }\alpha<\lambda.$
-This is well-defined by Remark 5.22. In particular, for $\beta<\lambda$, $\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}$.
+This is well-defined: if $\alpha<\beta<\lambda$, then by projection coherence (Remark 5.22),
+$\pi^{\alpha}_{\beta}\bigl(\pi^{\beta}_{\lambda}(g)\bigr)=\pi^{\alpha}_{\lambda}(g),$
+and applying $\pi^{0}_{\alpha}$ gives
+$\pi^{0}_{\alpha}\bigl(\pi^{\beta}_{\lambda}(g)\bigr)=\pi^{0}_{\alpha}\bigl(\pi^{\alpha}_{\beta}(\pi^{\beta}_{\lambda}(g))\bigr)=\pi^{0}_{\alpha}\bigl(\pi^{\alpha}_{\lambda}(g)\bigr),$
+so the value $\pi^{0}_{\alpha}(\pi^{\alpha}_{\lambda}(g))$ is independent of the chosen $\alpha$. In particular, for $\beta<\lambda$,
+$\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}.$
 Definition 5.24 (Package coordinate sets).
@@
   3. (iii)
-    $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for every $d\in D$.
+    $\operatorname{Fix}(E)\leq\operatorname{Stab}_{\mathscr{G}}(d)$ for every $d\in D$, where $\operatorname{Stab}_{\mathscr{G}}(d)$ denotes the stabilizer under the stage-0 action of $\mathscr{G}=\operatorname{Sym}(\omega_{1})$ on coordinates (Definition 5.24).
 When $D=\emptyset$, condition (iii) is vacuously satisfied, and $(E,\emptyset)$ is globally admissible for any countable $E$.
@@
 Therefore $k$ induces a bijection on $\mathrm{Coords}_{\beta+1}=\mathrm{Coords}_{\beta}\dot{\cup}\mathrm{NewCoords}_{\beta+1}$.
-At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf.
+At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are coherent inverse-limit elements whose restrictions to earlier stages determine them uniquely (cf.
 [5, Def.
@@
 Proof.
-For $d\in D_{1}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{1})\leq\operatorname{Stab}(d)$.
-For $d\in D_{2}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{2})\leq\operatorname{Stab}(d)$.
+For $d\in D_{1}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{1})\leq\operatorname{Stab}_{\mathscr{G}}(d)$.
+For $d\in D_{2}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{2})\leq\operatorname{Stab}_{\mathscr{G}}(d)$.
 ∎
@@
 Proof.
-For each $d\in D$, Lemma 5.9 provides a countable $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}(d)$.
+For each $d\in D$, Lemma 5.9 provides a countable $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}_{\mathscr{G}}(d)$.
 Let $E^{\ast}:=E\cup\bigcup_{d\in D}E_{d}$, which is a countable union of countable sets, hence countable.
-For each $d\in D$, we have $\operatorname{Fix}(E^{\ast})\leq\operatorname{Fix}(E_{d})\leq\operatorname{Stab}(d)$, so $(E^{\ast},D)$ is globally admissible.
+For each $d\in D$, we have $\operatorname{Fix}(E^{\ast})\leq\operatorname{Fix}(E_{d})\leq\operatorname{Stab}_{\mathscr{G}}(d)$, so $(E^{\ast},D)$ is globally admissible.
 The final statement follows from monotonicity: $E\subseteq E^{\ast}$ implies $\Delta_{\lambda}^{\uparrow}(E^{\ast},D)\leq\Delta_{\lambda}^{\uparrow}(E,D)$ by Lemma 5.36. ∎
@@
 For each $\beta\leq\lambda$ let $D_{\beta}:=D\cap\mathrm{Coords}_{\beta}$.
-By Definition 5.28, $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for every $d\in D$, hence $\pi\cdot d=d$ for all $d\in D$.
+By Definition 5.28, $\operatorname{Fix}(E)\leq\operatorname{Stab}_{\mathscr{G}}(d)$ for every $d\in D$, hence $\pi\cdot d=d$ for all $d\in D$.
 In particular, for each successor $\beta+1\leq\lambda$ the set
@@
 By the limit-stage group definition ([5, Def.
-3.1]), there is a unique direct-limit element $k_{\eta}\in\mathscr{G}_{\eta}$ with $\pi^{\beta}_{\eta}(k_{\eta})=k_{\beta}$ for all $\beta<\eta$.
+3.1]), elements of $\mathscr{G}_{\eta}$ are exactly the coherent threads $\langle h_{\beta}\rangle_{\beta<\eta}$ with $\pi^{\gamma}_{\beta}(h_{\beta})=h_{\gamma}$ for $\gamma<\beta<\eta$.
+Since $\langle k_{\beta}\rangle_{\beta<\eta}$ is such a coherent thread by construction, it defines a unique element $k_{\eta}\in\mathscr{G}_{\eta}$ with $\pi^{\beta}_{\eta}(k_{\eta})=k_{\beta}$ for all $\beta<\eta$.
 Since $\widehat{\pi}^{\,D_{\eta}}$ has the same restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$, we have $k_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
@@
 Let $\dot{\tau}^{*}$ denote the corresponding canonical $\mathbb{P}_{\beta+1}$-name under this two-step presentation.
-Assume $d$ has countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$ in $M_{\beta}$, so $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ in $M_{\beta}$.
+Assume $d$ has countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$ in $M_{\beta}$, so $\operatorname{Fix}(E)\leq\operatorname{Stab}_{\mathscr{G}}(d)$ in $M_{\beta}$.
 Then $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$; in particular, $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
 Proof.
-From $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ we get that $(E,\{d\})$ is globally admissible at stage $\beta+1$ (Definition 5.28).
+From $\operatorname{Fix}(E)\leq\operatorname{Stab}_{\mathscr{G}}(d)$ we get that $(E,\{d\})$ is globally admissible at stage $\beta+1$ (Definition 5.28).
 Hence $\Delta_{\beta+1}^{\uparrow}(E,\{d\})$ is one of the designated $\Delta^{\uparrow}$-generators of $\tilde{\mathscr{F}}^{*}_{\beta+1}$ (Definition 5.50).
@@
     $\Delta^{\uparrow}$-subgroups: $\Delta_{\lambda}^{\uparrow}(E,D)$ for countable $(E,D)$ globally admissible at stage $\lambda$.
+Remark 5.51 (Relation to the CSI limit filter).
+The companion paper [5] defines the limit-stage filter as the normal closure of head pullbacks.
+In this manuscript we work with the strengthened filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.50) obtained by additionally adjoining the subgroups $\Delta_{\lambda}^{\uparrow}(E,D)$ for globally admissible pairs $(E,D)$.
+All black-box iteration facts we import from [5] are used only in forms that require: (coherence of the projection system, normality, and $\omega_{1}$-completeness of the stage filters).
+These are verified for $\tilde{\mathscr{F}}^{*}_{\lambda}$ in Lemma 5.55 and its corollaries.
+In particular, the extra $\Delta^{\uparrow}$-generators are not present in the CSI framework and are introduced here solely to obtain the core admissibility lemma (Lemma 5.53) used in the $\neg\mathsf{AC}$ argument.
 Remark 5.52 (Admissibility in the filter definition).
@@
 To verify clause (iii) of Definition 5.28, fix $d\in D$ and choose $k$ with $d\in D_{k}$.
-Then $\operatorname{Fix}(E)\leq\operatorname{Fix}(E_{k})\leq\operatorname{Stab}(d)$, where the last inequality holds because $(E_{k},D_{k})$ is globally admissible.
+Then $\operatorname{Fix}(E)\leq\operatorname{Fix}(E_{k})\leq\operatorname{Stab}_{\mathscr{G}}(d)$, where the last inequality holds because $(E_{k},D_{k})$ is globally admissible.
 Therefore $(E,D)$ is globally admissible.
@@
 [5, §(Class-length iterations)] and Remark 5.60).
-In this setting, the forcing theorem holds for the definable class forcing $\mathbb{P}_{\mathrm{Ord}}$, and the class system $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$ supports the usual recursion defining hereditarily symmetric class-names.
+In this setting, definable class recursion suffices to define the valuation map for set-sized $\mathbb{P}_{\mathrm{Ord}}$-names and to carry out the usual recursion defining hereditarily symmetric class-names.
 For each set ordinal $\lambda$, the restriction to stage $\lambda$ yields the set-length symmetric system $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ and hence $M_{\lambda}\models\mathsf{ZF}$ by Theorem 5.58.
@@
 This arrangement ensures that every code is tested at unboundedly many stages, so a name that becomes a valid HS surjection instance at some stage $\beta$ will eventually be scheduled at some stage $\alpha>\beta$.
+Lemma 5.63 (Bookkeeping guarantee).
+Let $f\in\mathcal{M}$ be a surjection instance of one of our templates (PP or $\mathsf{AC}_{\mathsf{WO}}$).
+Choose $\dot{f}\in\operatorname{HS}$ with $f=\dot{f}^{G}$, and let $\beta:=\sup(I(\dot{f}))+1$ as in Lemma 5.67, so $f\in M_{\beta}$.
+Then there exists some $\alpha\geq\beta$ such that the bookkeeping schedule hits $(\beta,c_{\dot{f}})$ at stage $\alpha$, and hence:
+  • either $f$ already splits in $M_{\alpha}$, in which case $f$ splits in $\mathcal{M}$;
+  • or $f$ does not split in $M_{\alpha}$, and the successor-stage forcing at $\alpha$ uses the corresponding orbit package, so $f$ splits in $M_{\alpha+1}$ and hence in $\mathcal{M}$.
+Proof.
+By Definition 5.62 (unboundedness of the enumeration), there is some $\alpha\geq\beta$ with $\mathcal{B}(\alpha)=(\beta,c_{\dot{f}})$.
+At that stage, we run the bookkeeping test on $\dot{f}$ in $M_{\alpha}$.
+If it succeeds, the orbit package forcing at stage $\alpha$ adds a splitting object, which persists to all later stages.
+If it fails, then $f$ already has a splitting object in $M_{\alpha}$, and this also persists.
+In either case, $f$ splits in $\mathcal{M}$.
+∎
 Remark 5.64 (The $\aleph^{*}(S)$ horizon).
@@
 Then $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$.
+Lemma 5.65 (Stage iterands are hereditarily symmetric).
+For every stage $\alpha$, the $\mathbb{P}_{\alpha}$-name $\dot{\mathbb{S}}_{\alpha}$ is an element of $\operatorname{HS}_{\alpha}$.
+Proof.
+If the bookkeeping test fails at stage $\alpha$, then $\dot{\mathbb{S}}_{\alpha}$ is the canonical name for the trivial forcing, which is fixed by every automorphism and hence lies in $\operatorname{HS}_{\alpha}$.
+If the bookkeeping test succeeds, then $\dot{\mathbb{S}}_{\alpha}$ is (by construction) an orbit package forcing built from a scheduled surjection instance via the stage-0 action.
+For any $g\in\mathscr{G}_{\alpha}$, its stage-0 projection $\pi^{0}_{\alpha}(g)\in\mathscr{G}$ sends a surjection instance $f$ to $\pi^{0}_{\alpha}(g)\cdot f$, but preserves the orbit $[f]$.
+Consequently,
+$g\cdot\dot{\mathbb{S}}_{\alpha}=\dot{\mathbb{S}}_{\alpha}.$
+Thus $\mathscr{G}_{\alpha}\leq\operatorname{Stab}(\dot{\mathbb{S}}_{\alpha})$, so $\operatorname{Stab}(\dot{\mathbb{S}}_{\alpha})\in\tilde{\mathcal{F}}^{\ast}_{\alpha}$, and therefore $\dot{\mathbb{S}}_{\alpha}\in\operatorname{HS}_{\alpha}$.
+∎
 At successor stages, the group $\mathscr{G}_{\alpha+1}$ and filter $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ are those determined by the symmetric-iteration framework together with the diagonal infrastructure from Subsection 5.2 and the modified stage filters from Definition 5.50.
@@
 By Lemma 5.55(iv), the stage filter $\tilde{\mathscr{F}}^{*}_{\alpha}$ is $\omega_{1}$-complete.
-Therefore, by Remark 2.4(d), $M_{\alpha}\models\mathsf{DC}$.
+Since $\alpha$ is a set ordinal and $\tilde{\mathscr{F}}^{*}_{\alpha}$ is $\omega_{1}$-complete, the hypotheses of Remark 2.4(d) are satisfied, so $M_{\alpha}\models\mathsf{DC}$.
 Applying $\mathsf{DC}$ inside $M_{\alpha}$ to $(A,R)$, we obtain a sequence $\langle a_{n}:n<\omega\rangle\in M_{\alpha}\subseteq\mathcal{M}$ with $a_{n}Ra_{n+1}$ for all $n$.
@@
 ∎
+Definition 5.72 (Stage truncation of names).
+Let $\lambda$ be a limit stage and let $\alpha<\lambda$. For a $\mathbb{P}_{\lambda}$-name $\dot{x}$, define the $\mathbb{P}_{\alpha}$-name $\dot{x}\!\upharpoonright\!\alpha$ recursively by
+$\dot{x}\!\upharpoonright\!\alpha\;:=\;\Bigl\{\,\bigl\langle\dot{y}\!\upharpoonright\!\alpha,\ p\!\upharpoonright\!\alpha\bigr\rangle:\ \langle\dot{y},p\rangle\in\dot{x}\ \wedge\ \operatorname{supp}(p)\subseteq\alpha\,\Bigr\}.$
+(Thus we keep only those pairs whose witnessing condition already lies in $\mathbb{P}_{\alpha}$; in that case $p\upharpoonright\alpha=p$.)
+Lemma 5.73 (Union of truncations along an increasing stage sequence).
+Let $\lambda$ be a limit stage and let $\dot{X}\in\operatorname{HS}_{\lambda}$ with $1\Vdash_{\mathbb{P}_{\lambda}}$$\dot{X}$ is a set”.
+Suppose $\langle\lambda_{n}:n<\omega\rangle$ is an increasing sequence of ordinals below $\lambda$ such that for every $\langle\dot{y},p\rangle\in\operatorname{tc}(\dot{X})$ there is $n<\omega$ with $\operatorname{supp}(p)\subseteq\lambda_{n}$.
+Then for every $V$-generic $G_{\lambda}$,
+$\dot{X}^{G_{\lambda}}\;=\;\bigcup_{n<\omega}\bigl(\dot{X}\!\upharpoonright\!\lambda_{n}\bigr)^{G_{\lambda_{n}}}.$
+Proof.
+Fix $G_{\lambda}\subseteq\mathbb{P}_{\lambda}$ generic and write $G_{\lambda_{n}}:=G_{\lambda}\cap\mathbb{P}_{\lambda_{n}}$.
+For each $n<\omega$, set $X_{n}:=(\dot{X}\!\upharpoonright\!\lambda_{n})^{G_{\lambda_{n}}}$.
+We first note (by a straightforward induction on rank) that if $\operatorname{supp}(p)\subseteq\lambda_{n}$ and $\dot{y}$ is any $\mathbb{P}_{\lambda}$-name, then
+$(\dot{y}\!\upharpoonright\!\lambda_{n})^{G_{\lambda_{n}}}\;=\;\dot{y}^{G_{\lambda}}\quad\text{whenever }p\in G_{\lambda_{n}}.$
+($\subseteq$).
+Let $x\in\dot{X}^{G_{\lambda}}$.
+Choose $\langle\dot{y},p\rangle\in\dot{X}$ with $p\in G_{\lambda}$ and $x=\dot{y}^{G_{\lambda}}$.
+By hypothesis, pick $n<\omega$ with $\operatorname{supp}(p)\subseteq\lambda_{n}$.
+Then $p\in\mathbb{P}_{\lambda_{n}}$ and hence $p\in G_{\lambda_{n}}$.
+By definition of stage truncation, $\langle\dot{y}\!\upharpoonright\!\lambda_{n},\ p\rangle\in\dot{X}\!\upharpoonright\!\lambda_{n}$, so
+$x=\dot{y}^{G_{\lambda}}=(\dot{y}\!\upharpoonright\!\lambda_{n})^{G_{\lambda_{n}}}\in X_{n}.$
+($\supseteq$).
+Let $x\in X_{n}$.
+Then for some $\langle\dot{y}^{\prime},p\rangle\in\dot{X}\!\upharpoonright\!\lambda_{n}$ with $p\in G_{\lambda_{n}}$ we have $x=(\dot{y}^{\prime})^{G_{\lambda_{n}}}$.
+By definition of $\dot{X}\!\upharpoonright\!\lambda_{n}$, there is $\langle\dot{y},p\rangle\in\dot{X}$ with $\operatorname{supp}(p)\subseteq\lambda_{n}$ and $\dot{y}^{\prime}=\dot{y}\!\upharpoonright\!\lambda_{n}$.
+Since $p\in G_{\lambda_{n}}\subseteq G_{\lambda}$, we conclude
+$x=(\dot{y}\!\upharpoonright\!\lambda_{n})^{G_{\lambda_{n}}}=\dot{y}^{G_{\lambda}}\in\dot{X}^{G_{\lambda}}.$
+This proves $\dot{X}^{G_{\lambda}}=\bigcup_{n<\omega}X_{n}$. ∎
 Lemma 5.74 (Persistence of $\mathsf{SVC}(S)$).
@@
 Let $X\in M_{\lambda}$ be nonempty, and choose $\dot{X}\in\operatorname{HS}_{\lambda}$ with $X=\dot{X}^{G_{\lambda}}$.
-By the countable-support nature of the iteration and the HS-support analysis ((cf.
-the countable-support analysis of HS-names in Section 3 and the preceding discussion of supports at limits), $\dot{X}$ depends on only countably many stages.
-Let $\langle\lambda_{n}:n<\omega\rangle$ be an increasing sequence of ordinals below $\lambda$ capturing those stages.
-For each $n$, let $X_{n}$ be the interpretation of the corresponding stage-truncation of $\dot{X}$; then $X_{n}\in M_{\lambda_{n}}$ and $X=\bigcup_{n<\omega}X_{n}$.
+By the countable-support analysis of HS-names, fix an increasing sequence $\langle\lambda_{n}:n<\omega\rangle$ of ordinals below $\lambda$ satisfying the hypothesis of Lemma 5.73. For each $n$, set
+$X_{n}\;:=\;\bigl(\dot{X}\!\upharpoonright\!\lambda_{n}\bigr)^{G_{\lambda_{n}}}.$
+Then $X_{n}\in M_{\lambda_{n}}$ and $X=\bigcup_{n<\omega}X_{n}$ by Lemma 5.73.
 By the inductive hypothesis, each $M_{\lambda_{n}}\models\mathsf{SVC}(S)$, so for each $n$ there is an ordinal $\eta_{n}$ and a surjection $e_{n}:S\times\eta_{n}\twoheadrightarrow X_{n}$ in $M_{\lambda_{n}}\subseteq M_{\lambda}$.
@@
 $\mathsf{DC}$ holds by Proposition 5.71.
-$\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.65 and 5.72.
+$\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.68 and 5.77.
 $\mathsf{AC}_{\mathsf{WO}}$ holds by Proposition 5.77.
@@
 Work in $\mathcal{M}$. Fix any set $X$. By Corollary 5.75, $\mathcal{M}\models\mathsf{SVC}(S)$, so choose an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow X$.
-By Theorem 4.5 together with Propositions 5.65 and 5.72, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
+By Theorem 4.5 together with Propositions 5.68 and 5.77, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
 Pulling back the lexicographic order on $S\times\eta$ along $i$ yields a linear order on $X$. ∎
@@
 (2) $KWP_{1}$ using $\mathsf{PP}$.
-By Theorem 4.5 together with Propositions 5.65 and 5.72, $\mathcal{M}\models\mathsf{PP}$.
+By Theorem 4.5 together with Propositions 5.68 and 5.77, $\mathcal{M}\models\mathsf{PP}$.
 Apply $\mathsf{PP}$ to the surjection $e:S\times\eta\twoheadrightarrow X$ to obtain an injection $i:X\hookrightarrow S\times\eta$.

### v6

@@
 Assume $V\models\mathsf{ZFC}$.
-In a class forcing extension $V[G]$ obtained by an $\mathrm{Ord}$-length countable-support symmetric iteration over a Cohen symmetric seed model $\mathcal{N}$, the final symmetric model $\mathcal{M}$ satisfies
+Assume $V\models\mathsf{ZFC}$.
+Assume $V\models\mathsf{ZFC}$, and let $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic in the metatheoretic sense of Remark 2.1 for the $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen symmetric seed model $\mathcal{N}$.
+Then the final symmetric model $\mathcal{M}$ satisfies
 $\mathsf{ZF}\ +\ \mathsf{DC}\ +\ \mathsf{PP}\ +\ \mathsf{AC}_{\mathsf{WO}}\ +\ \neg\mathsf{AC}.$
@@
 Relation to countable-support symmetric iterations.
-The iteration is an instance of the general countable-support symmetric-iteration framework developed in [5]; we use its successor and limit constructions, the $\omega_{1}$-completeness and normality facts for the generated filters, and the $\mathsf{ZF}/\mathsf{DC}$ preservation theorems.
-What is specific to the present manuscript is the package bookkeeping for $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ and the orbit-symmetrization required to obtain genuine automorphisms of the successor iterands (Remark 2.4).
+The bounded-stage successor and limit constructions used here follow the countable-support symmetric-iteration scheme developed in [5].
+From that companion paper we use the stagewise group/projection infrastructure and the bounded-stage preservation results, especially $\mathsf{ZF}$ and $\mathsf{DC}$ at set stages.
+What is specific to the present manuscript is the explicit $\mathrm{Ord}$-length recursion for the present package forcings, together with the diagonal-lift/diagonal-cancellation machinery used to define the limit filters and to make the bookkeeping work for $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ (Remark 2.4).
 What is not claimed.
@@
 Organization.
-Section 2 fixes the metatheoretic conventions and records the countable-support symmetric-iteration API items imported from [5].
+Section 2 fixes the metatheoretic conventions, records the bounded-stage iteration facts imported from [5], and explains how the present manuscript carries out the specific $\mathrm{Ord}$-length recursion.
 Section 3 constructs the seed model $\mathcal{N}$ and establishes $\mathsf{DC}$, $\neg\mathsf{AC}$, and $\mathsf{SVC}(S)$ at stage 0.
@@
 2 Preliminaries and metatheoretic conventions
-We follow the forcing, automorphism, symmetric-extension, and countable-support symmetric-iteration framework developed in [5], based on Karagila’s iteration scheme [9].
-This section fixes the metatheory and notation used throughout, and records precisely which results are imported from [5].
+We follow the forcing, automorphism, symmetric-extension, and bounded-stage countable-support symmetric-iteration framework developed in [5], based on Karagila’s iteration scheme [9].
+This section fixes the metatheory and notation used throughout, records precisely which bounded-stage results are imported from [5], and explains how the present paper carries out the specific $\mathrm{Ord}$-length recursion needed for the package iteration.
 2.1 Background theory and conventions
 Remark 2.1 (Metatheory vs. object theory).
-All forcing notions, names, automorphism groups, filters, and supports are constructed in a fixed background universe $V$.
-  • For the set-forcing and symmetric-extension statements used at bounded stages, we work over $V\models\mathsf{ZF}$.
-  • Whenever we invoke preservation of dependent choice for a bounded stage symmetric model, we assume the background satisfies $V\models\mathsf{ZFC}$.
-Throughout this paper, $\mathsf{DC}$ means $\mathsf{DC}_{\omega}$ (dependent choice for $\omega$-sequences, in the standard set formulation used in [5, §3.5]).
-  • For the class-length recursion through $\mathrm{Ord}$, we assume a background theory adequate for class recursion (e.g.
-$V\models\mathsf{GBC}+\mathsf{ETR}$), and we assume Global Choice only when explicitly invoked for bookkeeping.
-When we say “countable”, $\omega_{1}$-complete”, etc., this is always with respect to the background universe $V$.
+All forcing notions, names, automorphism groups, filters, and supports are constructed over a fixed background set universe $V$.
+  • For the set-forcing and symmetric-extension statements used at bounded stages, it suffices to work over $V\models\mathsf{ZF}$.
+  • Whenever we invoke preservation of dependent choice for a bounded stage symmetric model, we assume $V\models\mathsf{ZFC}$.
+Throughout this paper, $\mathsf{DC}$ means $\mathsf{DC}_{\omega}$ (dependent choice for $\omega$-sequences, in the standard set formulation used in [5, Definition (Dependent Choice)]).
+  • For the class-length recursion through $\mathrm{Ord}$ (bookkeeping and the definition of the iteration itself), we work in a background two-sorted structure $(V,\mathcal{C})$ satisfying a class theory adequate for transfinite recursion along $\mathrm{Ord}$ (e.g.
+$(V,\mathcal{C})\models\mathsf{GBC}+\mathsf{ETR}_{\mathrm{Ord}}$, or simply $\mathsf{GBC}+\mathsf{ETR}$ in the standard sense).
+When convenient we fix a class well-order of $V$ (equivalently, assume Global Choice in the background), but we use such Global Choice only for bookkeeping/canonical enumerations and explicitly indicate when it is invoked.
+When we speak of a “$V$-generic” $G\subseteq\mathbb{P}_{\mathrm{Ord}}$, this is understood in the metatheory (equivalently, in an outer universe) as a subclass meeting every dense class $D\in\mathcal{C}$.
+We emphasize that we do not rely on a global forcing theorem for $\mathbb{P}_{\mathrm{Ord}}$: valuations of set-sized $\mathbb{P}_{\mathrm{Ord}}$-names localize to bounded forcing stages (Lemma 5.72), and set-sized final hereditarily symmetric names descend to bounded-stage hereditarily symmetric names (Lemma 5.73).
+Accordingly, set-theoretic arguments are carried out inside initial-segment extensions $V[G_{\alpha}]$ and their bounded-stage symmetric submodels.
+Finally, when we say “countable”, $\omega_{1}$-complete”, etc., this is always with respect to the background universe $V$.
 2.2 Restrictions, supports, and stage generics
 Remark 2.2 (Forcing vs. group restrictions).
-We follow the convention of [5, Remark 2.2] and do not overload restriction/projection notation.
+We follow the notation discipline for forcing versus group restrictions from [5, Remark (Notation discipline for later sections)] and do not overload restriction/projection notation.
   • For forcing, $p\upharpoonright\beta$ denotes the usual restriction of a condition $p\in\mathbb{P}_{\lambda}$ to an initial segment $\beta\leq\lambda$ (and when needed, $\pi_{\beta,\lambda}:\mathbb{P}_{\lambda}\to\mathbb{P}_{\beta}$ denotes the canonical projection map).
@@
 Notation 2.3 (Stage generics).
-Let $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic for the class-length iteration forcing.
-For each $\alpha\leq\mathrm{Ord}$, write $G_{\alpha}:=G\cap\mathbb{P}_{\alpha}$ for the induced $V$-generic filter on the initial segment forcing.
-2.3 Imported iteration results from [5]
+Fix (in the metatheory) a $V$-generic subclass $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ for the class-length iteration forcing.
+For each set ordinal $\alpha$, write $G_{\alpha}:=G\cap\mathbb{P}_{\alpha}$ for the induced $V$-generic filter on the initial-segment forcing.
+2.3 Imported bounded-stage iteration results from [5]
 Remark 2.4 (Imported iteration facts).
@@
   1. (a)
-    Existence of set- and class-length iterations.
-The countable-support symmetric iteration recursion can be carried out through any set ordinal, and (under the background class-recursion assumptions used there) through $\mathrm{Ord}$; see [5, §Countable-support symmetric iterations and §Class-length iterations].
+    Bounded-stage recursion and coherence.
+For each set ordinal $\lambda$, the countable-support symmetric iteration framework provides the bounded-stage forcing $\mathbb{P}_{\lambda}$, the group $\mathscr{G}_{\lambda}$, the stage filter, and the projection/coherence maps used throughout the set-stage analysis; see [5, §Limit Filters for Countable-Support Symmetric Iterations].
+The present paper does not import a general class-length existence theorem from [5]; instead, the $\mathrm{Ord}$-length iteration used here is defined explicitly by recursion in the background metatheory described in Remark 2.1.
   2. (b)
@@
 Its normality and $\omega_{1}$-completeness are proved here (Lemma 5.58).
-We only use [5] for the general iteration template and projection/coherence infrastructure leading to the limit stage.
+We only use [5] here for the bounded-stage projection/coherence infrastructure and the underlying countable-support limit construction.
   4. (d)
@@
 This is proved in [5, Theorem (DC Preservation (set-length stages))].
-We stress that [5] does not assert $\mathsf{DC}$ for the final class-length symmetric model in general.
-In this paper, $\mathsf{DC}$ for the final model is obtained by combining item (d) at bounded stages with the stage-localization lemma for set-sized $\mathbb{P}_{\mathrm{Ord}}$-names (Lemma 5.67).
+We stress that [5] is used here only at bounded stages.
+In particular, this paper does not import any general theorem asserting $\mathsf{DC}$ or $\mathsf{ZF}$ for the final class-length symmetric model.
+The final-model arguments are proved directly in the present manuscript: $\mathsf{DC}$ is obtained by combining item (d) with the stage-descent lemma for set-sized final hereditarily symmetric names (Lemma 5.73), and $\mathsf{ZF}$ for the final model is verified by the standard symmetric-extension argument applied to the final class system (Theorem 5.64 and Remark 5.63).
 Choice principles and the Ryan–Smith localization theorem are stated (and used) in Section 4.
@@
   5. (5)
-    Black-box symmetric iteration.
-Subsection 5.5 invokes the symmetric-iteration framework to assemble the class-length iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$, with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and successor stages $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$ where $\dot{\mathbb{S}}_{\alpha}$ is the finite product of the scheduled orbit package factors.
+    The $\mathrm{Ord}$-length recursion.
+Subsection 5.5 defines the specific $\mathrm{Ord}$-length iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ used in this paper, starting from $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and setting $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$ at successor stages, where $\dot{\mathbb{S}}_{\alpha}$ is the finite product of the scheduled orbit-package factors.
+The bounded-stage successor/limit infrastructure comes from Remark 2.4, while the transfinite recursion through $\mathrm{Ord}$ is carried out explicitly in the background metatheory from Remark 2.1.
 By Lemma 5.70, the canonical $\mathbb{P}_{\alpha}$-names for these orbit-package iterands are hereditarily symmetric, so the two-step presentation is well-defined in the stage ground $M_{\alpha}$.
@@
 We now describe the symmetric iteration that forces the two conjuncts $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ over the Cohen symmetric seed model $\mathcal{N}$, while maintaining $\neg\mathsf{AC}$ (witnessed by the non-well-orderability of $A$).
-All technical iteration infrastructure is imported via the API of Remark 2.4 (in particular, existence of the class-length iteration is used exactly under the hypotheses of [5, Theorem 3.25]).
+All technical bounded-stage iteration infrastructure is imported via the API of Remark 2.4.
+The $\mathrm{Ord}$-length iteration used in this section is then defined explicitly in the present manuscript by recursion in the background metatheory from Remark 2.1; it is not an instance of a general class-length existence theorem imported from [5].
 The purpose of this section is to isolate the package systems, the bookkeeping targets, and the verification that the final symmetric model satisfies the desired principles.
@@
 Remark 5.7 (Limit-stage $\omega_{1}$-completeness).
-The countable completeness of the diagonal stage filters is what allows the CSI iteration theorem to handle limit stages.
-Concretely, the limit-step existence theorem for iterations of symmetric systems requires $\omega_{1}$-complete filters (CSI, [5, Theorem 3.13]), and we have imported this requirement as item (c) in Remark 2.4.
+The countable completeness of the diagonal stage filters is what makes the bounded-stage countable-support limit construction compatible with the hereditary-symmetry and $\mathsf{DC}$-preservation arguments.
+In [5], $\omega_{1}$-completeness is a key property of the stage filters at limits (see the limit-filter construction and the theorem on limit filter properties), and we import that requirement here as item (c) in Remark 2.4.
 Lemma 5.8 (Well-definedness of the orbit package actions).
@@
 Proof.
-This is exactly the successor-step group structure supplied by the countable-support symmetric iteration framework (Remark 2.4, cf. [5, Def. 3.1]). Concretely:
+This is exactly the successor-step group structure supplied by the bounded-stage countable-support symmetric iteration framework (Remark 2.4; cf. [5, §Framework and Setup]). Concretely:
   • $\pi^{\alpha}_{\alpha+1}$ is restriction to the $\mathbb{P}_{\alpha}$-part;
@@
 For $\gamma<\beta$ we write $\pi^{\gamma}_{\beta}:\mathscr{G}_{\beta}\to\mathscr{G}_{\gamma}$ for the canonical projection homomorphisms of the iteration groups (cf.
-[5, Def.
-3.1]).
+[5, Definition (Countable-support iteration setup)]).
 These satisfy coherence: for all $\gamma<\beta<\lambda$,
@@
 Proof.
-(i) This is immediate from Definition 5.23 together with the coherence of the projections (Remark 5.22); cf. also [5, Def. 3.1].
+(i) This is immediate from Definition 5.23 together with the coherence of the projections (Remark 5.22); cf. also [5, Definition (Countable-support iteration setup)].
 (ii) We prove by induction on stages that each $k\in\mathscr{G}_{\beta}$ determines a bijection of $\mathrm{Coords}_{\beta}$, and hence sends countable sets of coordinates to countable sets.
 At stage 0 there are no package coordinates.
-At a successor stage $\beta+1$, write $k=\mathrm{st}(h)\cdot u$ as in Lemma 5.21, where $h=\pi^{\beta}_{\beta+1}(k)\in\mathscr{G}_{\beta}$ and $u\in K_{\beta+1}$ acts trivially on the $\mathbb{P}_{\beta}$-part.
+At a successor stage $\beta+1$, write $k=\mathrm{spl}(h)\cdot u$ as in Lemma 5.21, where $h=\pi^{\beta}_{\beta+1}(k)\in\mathscr{G}_{\beta}$ and $u\in K_{\beta+1}$ acts trivially on the $\mathbb{P}_{\beta}$-part.
 By the inductive hypothesis, $h$ induces a bijection of $\mathrm{Coords}_{\beta}$.
-The standard lift $\mathrm{st}(h)$ extends this action and fixes every element of $\mathrm{NewCoords}_{\beta+1}$.
+The standard lift $\mathrm{spl}(h)$ extends this action and fixes every element of $\mathrm{NewCoords}_{\beta+1}$.
 On the other hand, $u$ fixes $\mathrm{Coords}_{\beta}$ pointwise and acts on the successor iterand $\dot{\mathbb{Q}}_{\beta}$ by compositions of package automorphisms (Lemma 5.8) and diagonal-cancellation automorphisms (Definition 5.11), each of which induces a bijection on the new coordinate index set.
@@
 At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are coherent inverse-limit elements whose restrictions to earlier stages determine them uniquely (cf.
-[5, Def.
-3.1]).
+[5, Definition (Countable-support iteration setup)]).
 The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection.
@@
 Let $k\in\Delta_{\lambda}^{\uparrow}(E,D)$. By Lemma 5.43, $\pi^{\beta}_{\lambda}(k)\in\Delta_{\beta}^{\uparrow}(E,D)\subseteq H$. Hence $k\in(\pi^{\beta}_{\lambda})^{-1}[H]$. ∎
+Lemma 5.45 (Tail extension by identity).
+Let $\beta\leq\lambda$ be stages of the iteration, where $\lambda$ may also be $\mathrm{Ord}$ in the background class metatheory.
+For every $k\in\mathscr{G}_{\beta}$, there exists $\hat{k}\in\mathscr{G}_{\lambda}$ such that:
+  1. (i)
+    $\pi^{\beta}_{\lambda}(\hat{k})=k$;
+  2. (ii)
+[…]
+    $\hat{k}_{\gamma+1}=\mathrm{spl}(\hat{k}_{\gamma});$
+  3. (iii)
+    equivalently, $\hat{k}$ acts as $k$ on the $\mathbb{P}_{\beta}$-part and trivially on every iterand added at stages in $\beta,\lambda)$.
+Proof.
+If $\beta=\lambda$, take $\hat{k}:=k$. Assume $\beta<\lambda$.
+We build by transfinite recursion a coherent thread $\langle k_{\gamma}:\beta\leq\gamma\leq\lambda\rangle$ such that $k_{\beta}=k$, each $k_{\gamma}\in\mathscr{G}_{\gamma}$, and $\pi^{\delta}_{\gamma}(k_{\gamma})=k_{\delta}$ whenever $\beta\leq\delta<\gamma\leq\lambda$.
+Base stage. Set $k_{\beta}:=k$.
+Successor stage $\gamma+1$. Given $k_{\gamma}\in\mathscr{G}_{\gamma}$, define
+$k_{\gamma+1}:=\mathrm{spl}(k_{\gamma})\in\mathscr{G}_{\gamma+1},$
+where $\mathrm{spl}$ is the splitting section from Lemma 5.21. Then
+$\pi^{\gamma}_{\gamma+1}(k_{\gamma+1})=\pi^{\gamma}_{\gamma+1}(\mathrm{spl}(k_{\gamma}))=k_{\gamma},$
+so coherence is preserved. By construction, $k_{\gamma+1}$ acts as $k_{\gamma}$ on the $\mathbb{P}_{\gamma}$-part and trivially on the new iterand at stage $\gamma$.
+Limit stage $\eta\leq\lambda$.
+Assume $\langle k_{\gamma}:\beta\leq\gamma<\eta\rangle$ has been constructed coherently.
+By the limit-stage group definition used in the iteration, coherent threads determine unique elements of $\mathscr{G}_{\eta}$.
+Thus there is a unique $k_{\eta}\in\mathscr{G}_{\eta}$ such that
+$\pi^{\gamma}_{\eta}(k_{\eta})=k_{\gamma}\qquad(\beta\leq\gamma<\eta).$
+If $\eta=\mathrm{Ord}$, this is carried out in the same background metatheory used for the class-length bookkeeping (Remark 5.65).
+Now set $\hat{k}:=k_{\lambda}$.
+Then (i) holds by construction.
+Clause (ii) is exactly the successor clause of the recursion.
+Clause (iii) is the equivalent operational description: at each successor step we extend by the splitting section, which acts trivially on the new iterand and preserves the previously constructed action.
+∎
+Corollary 5.46 (Identity extension for $\Delta^{\uparrow}$-groups).
+Let $\beta\leq\lambda$, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{\beta}$.
+If $k\in\Delta_{\beta}^{\uparrow}(E,D)$, then there exists $\hat{k}\in\Delta_{\lambda}^{\uparrow}(E,D)$ such that
+$\pi^{\beta}_{\lambda}(\hat{k})=k.$
+Proof.
+Let $\hat{k}\in\mathscr{G}_{\lambda}$ be the identity-tail extension of $k$ given by Lemma 5.45.
+For the stage-0 part, by Definition 5.23 and projection coherence (Remark 5.22),
+$\pi^{0}_{\lambda}(\hat{k})=\pi^{0}_{\beta}(\pi^{\beta}_{\lambda}(\hat{k}))=\pi^{0}_{\beta}(k)\in\operatorname{Fix}(E).$
+For the coordinate part, let $d\in D\subseteq\mathrm{Coords}_{\beta}$.
+At stage $\beta$, the element $k$ fixes $d$ pointwise.
+At every successor step above $\beta$, the extension is obtained by the splitting section, which preserves the action on all previously existing coordinates and acts trivially on the new iterand.
+By induction on stages from $\beta$ up to $\lambda$, it follows that $\hat{k}$ fixes $d$ pointwise as well.
+Hence $\hat{k}\in\Delta_{\lambda}^{\uparrow}(E,D)$ by Definition 5.29(A). ∎
+Lemma 5.47 (Action on bounded names factors through projection).
+Let $\beta\leq\lambda$ be stages, let $\tau$ be a $\mathbb{P}_{\beta}$-name, and let $k\in\mathscr{G}_{\lambda}$. Viewing $\tau$ as a $\mathbb{P}_{\lambda}$-name by canonical inclusion, one has
+$k\cdot\tau\;=\;\pi^{\beta}_{\lambda}(k)\cdot\tau.$
+Proof.
+We argue by induction on the rank of $\tau$.
+Let $\langle\sigma,p\rangle\in\tau$.
+Since $\tau$ is a $\mathbb{P}_{\beta}$-name, we have $p\in\mathbb{P}_{\beta}$.
+The projection $\pi^{\beta}_{\lambda}(k)$ is, by definition, the restriction of $k$ to the $\mathbb{P}_{\beta}$-part, so
+$k\cdot p\;=\;\pi^{\beta}_{\lambda}(k)\cdot p.$
+By the induction hypothesis,
+$k\cdot\sigma\;=\;\pi^{\beta}_{\lambda}(k)\cdot\sigma.$
+Therefore the image of each pair $\langle\sigma,p\rangle$ under $k$ is exactly the same as its image under $\pi^{\beta}_{\lambda}(k)$, and hence $k\cdot\tau=\pi^{\beta}_{\lambda}(k)\cdot\tau$. ∎
 Lemma 5.48 (Iteration-level diagonal lifts lie in $\mathscr{G}_{\lambda}$).
@@
 Let $\eta$ be limit and assume $\langle k_{\beta}:\beta<\eta\rangle$ has been constructed coherently.
-By the limit-stage group definition ([5, Def.
-3.1]), elements of $\mathscr{G}_{\eta}$ are exactly the coherent threads $\langle h_{\beta}\rangle_{\beta<\eta}$ with $\pi^{\gamma}_{\beta}(h_{\beta})=h_{\gamma}$ for $\gamma<\beta<\eta$.
+By the limit-stage group clause in the countable-support iteration setup ([5, Definition (Countable-support iteration setup)]), elements of $\mathscr{G}_{\eta}$ are exactly the coherent threads $\langle h_{\beta}\rangle_{\beta<\eta}$ with $\pi^{\gamma}_{\beta}(h_{\beta})=h_{\gamma}$ for $\gamma<\beta<\eta$.
 Since $\langle k_{\beta}\rangle_{\beta<\eta}$ is such a coherent thread by construction, it defines a unique element $k_{\eta}\in\mathscr{G}_{\eta}$ with $\pi^{\beta}_{\eta}(k_{\eta})=k_{\beta}$ for all $\beta<\eta$.
@@
 Remark 5.54 (Relation to the CSI limit filter).
-The companion paper [5] defines the limit-stage filter as the normal closure of head pullbacks.
+The companion paper [5] defines the limit-stage filter from head pullbacks alone: at uncountable cofinality as the normal filter generated by those pullbacks, and at countable cofinality as the smallest normal $\omega_{1}$-complete filter containing them.
 In this manuscript we work with the strengthened filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.53) obtained by additionally adjoining the subgroups $\Delta_{\lambda}^{\uparrow}(E,D)$ for globally admissible pairs $(E,D)$.
@@
 ∎
-Theorem 5.59 ($\mathcal{M}\models\mathsf{ZF}$).
-Let $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic and $G_{\lambda}:=G_{\infty}\cap\mathbb{P}_{\lambda}$.
-Let $\operatorname{HS}_{\mathrm{Ord}}$ be the class of hereditarily symmetric (cf.
-Definition 5.56) $\mathbb{P}_{\mathrm{Ord}}$-names (class-names) with respect to $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$, and define
+Lemma 5.62 (Upward persistence of hereditarily symmetric names).
+Let $\beta\leq\lambda$ be stages, where $\lambda$ may also be $\mathrm{Ord}$.
+If $\tau\in\operatorname{HS}_{\beta}$, then, viewed as a $\mathbb{P}_{\lambda}$-name via canonical inclusion, $\tau\in\operatorname{HS}_{\lambda}$.
+Proof.
+We argue by induction on the rank of $\tau$.
+Since $\tau\in\operatorname{HS}_{\beta}$, every subname $\sigma$ of $\tau$ belongs to $\operatorname{HS}_{\beta}$.
+By the induction hypothesis, each such $\sigma$ also belongs to $\operatorname{HS}_{\lambda}$.
+It remains to prove that $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}$. Because $\tau\in\operatorname{HS}_{\beta}$, we have
+$\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\beta}.$
+By Definition 5.53, the pullback
+$(\pi^{\beta}_{\lambda})^{-1}\!\big[\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)\big]$
+belongs to $\tilde{\mathscr{F}}^{*}_{\lambda}$.
+We claim that
+$(\pi^{\beta}_{\lambda})^{-1}\!\big[\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)\big]\subseteq\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau).$
+Indeed, if $k\in\mathscr{G}_{\lambda}$ satisfies $\pi^{\beta}_{\lambda}(k)\in\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)$, then by Lemma 5.47,
+$k\cdot\tau=\pi^{\beta}_{\lambda}(k)\cdot\tau=\tau.$
+So $k\in\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)$, proving the claim.
+By upward closure of $\tilde{\mathscr{F}}^{*}_{\lambda}$, it follows that $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}$.
+Therefore $\tau\in\operatorname{HS}_{\lambda}$.
+∎
+Remark 5.63 (Why the standard symmetric-extension proof applies at length $\mathrm{Ord}$).
+Although $\mathbb{P}_{\mathrm{Ord}}$ is a class forcing, every $\mathbb{P}_{\mathrm{Ord}}$-name used in the final model construction is a set.
+The usual proof of the symmetric extension theorem (for example, [8, Lemma 15.51]) is a recursion on the ranks of set-sized names and uses only:
+  1. (i)
+    the valuation map for set-sized names,
+  2. (ii)
+    the action of the automorphism group on names, and
+  3. (iii)
+    that the filter of subgroups is proper and normal.
+In the present setting these ingredients are available for $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$: valuation is defined in the background metatheory from Remark 5.65, the group action is defined stagewise by the iteration construction, and properness/normality are provided by Lemma 5.58.
+Thus the same proof applies to the final class of hereditarily symmetric names.
+Theorem 5.64 (The final class hereditarily symmetric model).
+Let $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic.
+Let $\operatorname{HS}_{\mathrm{Ord}}$ be the class of hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-names (cf.
+Definition 5.59) with respect to $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$, and define
 $\mathcal{M}\ :=\ \{\dot{x}^{G_{\infty}}:\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}\}.$
@@
 Proof.
-Work in the background metatheory used to formalize class-length iterations (cf.
-[5, §(Class-length iterations)] and Remark 5.60).
-In this setting, definable class recursion suffices to define the valuation map for set-sized $\mathbb{P}_{\mathrm{Ord}}$-names and to carry out the usual recursion defining hereditarily symmetric class-names.
-For each set ordinal $\lambda$, the restriction to stage $\lambda$ yields the set-length symmetric system $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ and hence $M_{\lambda}\models\mathsf{ZF}$ by Theorem 5.58.
-Moreover, every set-sized $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{x}$ is supported in some stage $\lambda$, so $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\lambda}}\in M_{\lambda}$.
-Consequently,
-$\mathcal{M}=\{\dot{x}^{G_{\infty}}:\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}\}=\bigcup_{\lambda\in\mathrm{Ord}}M_{\lambda},$
-so $\mathcal{M}$ is a transitive class.
-The verification of the $\mathsf{ZF}$ axioms in $\mathcal{M}$ follows by the standard class-length symmetric iteration argument (ibid.): each axiom instance with set parameters from $\mathcal{M}$ is witnessed at some stage $\lambda$ and hence holds in $M_{\lambda}$, and therefore holds in $\mathcal{M}$.
+Work in the background metatheory from Remark 5.65.
+By Remark 5.63, the standard proof of the symmetric extension theorem applies to the class system $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$.
+Hence the class of interpretations of hereditarily symmetric names,
+$\mathcal{M}=\{\dot{x}^{G_{\infty}}:\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}\},$
+satisfies $\mathsf{ZF}$.
+For transitivity, let $x\in\mathcal{M}$ and $y\in x$.
+Choose $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$ with $x=\dot{x}^{G_{\infty}}$.
+Then there is some pair $\langle\dot{y},p\rangle\in\dot{x}$ with $p\in G_{\infty}$ and $y=\dot{y}^{G_{\infty}}$.
+Since $\dot{x}$ is hereditarily symmetric, $\dot{y}\in\operatorname{HS}_{\mathrm{Ord}}$, so $y\in\mathcal{M}$.
 ∎
 5.4 Bookkeeping: names versus interpretations
-Remark 5.60 (Background metatheory for class-length constructions).
-To speak formally about an $\mathrm{Ord}$-length bookkeeping function and an $\mathrm{Ord}$-length symmetric iteration as class objects, we work in a background theory of classes supporting definable class recursion for the iteration, e.g.
-$\mathsf{GBC}+\mathsf{ETR}$, and we fix a class well-order of $V$ (equivalently, assume Global Choice in the background).
-Its set universe is denoted $V$ and satisfies $\mathsf{ZFC}$.
-All symmetric models $\mathcal{N}$, intermediate stages $M_{\alpha}$, and the final model $\mathcal{M}$ are constructed inside this set universe; only the bookkeeping map $\mathcal{B}$ and the iteration itself are treated as class objects in the ambient background.
+Remark 5.65 (Metatheory for bookkeeping).
+We continue with the metatheoretic conventions of Remark 2.1.
+For the $\mathrm{Ord}$-length bookkeeping in Section 5.4 (in particular, to define the class function $\mathcal{B}:\mathrm{Ord}\to V$ in Definition 5.67), we work in a background two-sorted structure $(V,\mathcal{C})\models\mathsf{GBC}+\mathsf{ETR}$ and, only for this bookkeeping/canonical-enumeration purpose, fix a class well-order of $V$ (equivalently, assume Global Choice in the background).
+All forcing semantics for $\mathbb{P}_{\mathrm{Ord}}$ are understood externally exactly as in Remark 2.1 (in particular, we do not assume a global forcing theorem for $\mathbb{P}_{\mathrm{Ord}}$).
 We now fix a bookkeeping device which ensures that every relevant surjection instance appearing in the final model is handled at some successor stage.
@@
 Lemma 5.68 (Bookkeeping guarantee).
-Let $f\in\mathcal{M}$ be a surjection instance of one of our templates (PP or $\mathsf{AC}_{\mathsf{WO}}$).
-Choose $\dot{f}\in\operatorname{HS}$ with $f=\dot{f}^{G}$, and let $\beta:=\sup(I(\dot{f}))+1$ as in Lemma 5.67, so $f\in M_{\beta}$.
-Then there exists some $\alpha\geq\beta$ such that the bookkeeping schedule hits $(\beta,c_{\dot{f}})$ at stage $\alpha$, and hence:
-  • either $f$ already splits in $M_{\alpha}$, in which case $f$ splits in $\mathcal{M}$;
-  • or $f$ does not split in $M_{\alpha}$, and the successor-stage forcing at $\alpha$ uses the corresponding orbit package, so $f$ splits in $M_{\alpha+1}$ and hence in $\mathcal{M}$.
+Let $f\in\mathcal{M}$ be a surjection instance of one of our two templates ($\mathsf{PP}$ or $\mathsf{AC}_{\mathsf{WO}}$).
+Choose a set-sized $\dot{f}\in\operatorname{HS}_{\mathrm{Ord}}$ with $f=\dot{f}^{G_{\infty}}$.
+By Lemma 5.73, there is $\beta<\mathrm{Ord}$ such that $\dot{f}$, viewed as a $\mathbb{P}_{\beta}$-name, belongs to $\operatorname{HS}_{\beta}$ and
+$f=\dot{f}^{G_{\beta}}\in M_{\beta}.$
+Then there exists some $\alpha\geq\beta$ such that the bookkeeping schedule hits $(\beta,c_{\dot{f}})$ at stage $\alpha$, and at that stage:
+  • either $f$ already has the required splitting object in $M_{\alpha}$ (a section in the $\mathsf{PP}$-case, a right inverse in the $\mathsf{AC}_{\mathsf{WO}}$-case), or
+  • the successor-stage forcing at $\alpha$ uses the corresponding orbit package.
+Consequently, $f$ has the required splitting object in $\mathcal{M}$.
 Proof.
-By Definition 5.62 (unboundedness of the enumeration), there is some $\alpha\geq\beta$ with $\mathcal{B}(\alpha)=(\beta,c_{\dot{f}})$.
-At that stage, we run the bookkeeping test on $\dot{f}$ in $M_{\alpha}$.
-If it succeeds, the orbit package forcing at stage $\alpha$ adds a splitting object, which persists to all later stages.
-If it fails, then $f$ already has a splitting object in $M_{\alpha}$, and this also persists.
-In either case, $f$ splits in $\mathcal{M}$.
+By the unboundedness clause in Definition 5.67, there exists $\alpha\geq\beta$ with $\mathcal{B}(\alpha)=(\beta,c_{\dot{f}})$.
+View $\dot{f}$ as a $\mathbb{P}_{\alpha}$-name via canonical inclusion.
+By Lemma 5.62, this included name belongs to $\operatorname{HS}_{\alpha}$.
+Since $\dot{f}$ is already a $\mathbb{P}_{\beta}$-name, its interpretation at stage $\alpha$ is still $f$:
+$\dot{f}^{G_{\alpha}}=\dot{f}^{G_{\beta}}=f.$
+Thus, when the bookkeeping test is run on $(\beta,c_{\dot{f}})$ at stage $\alpha$, clauses (i), (ii), and (iii) of Definition 5.67 all hold exactly when $f$ is a genuine stage-$\alpha$ instance of the required template.
+If $f$ already has the required splitting object in $M_{\alpha}$, we are done.
+Otherwise the bookkeeping test succeeds, so the stage-$\alpha$ iterand is the corresponding orbit package $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$.
+In the package extension this adds the required splitting object for $f$, and by the hereditary-symmetry lifting lemmas used at successor stages, that object belongs to $M_{\alpha+1}\subseteq\mathcal{M}$.
+Hence $f$ has the required splitting object in $\mathcal{M}$.
 ∎
@@
 Since every code is retested at unboundedly many stages (Definition 5.67), no a priori stability theorem for $\aleph^{*}(S)$ is needed.
-5.5 The symmetric iteration (black box)
-We now invoke the iteration framework via Remark 2.4.
-Working in the background metatheory, we form a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$.
-Concretely, we build an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [5, §3.6], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$.
+5.5 The $\mathrm{Ord}$-length symmetric iteration
+We now carry out the specific $\mathrm{Ord}$-length recursion used in this paper.
+Working in the background metatheory from Remark 2.1, we define by transfinite recursion on $\alpha\in\mathrm{Ord}$ an iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$, starting from $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and producing intermediate symmetric models $M_{\alpha}$.
+At successor stages we use the orbit-package iterands described below; at limit stages we use the countable-support forcing limit and the corresponding group/projection infrastructure from the bounded-stage framework, together with the modified limit filter from Definition 5.53.
+No general class-length theorem from [5] is invoked at this point; only the bounded-stage infrastructure summarized in Remark 2.4 is used.
 The next subsection states the properties of the final symmetric model that we prove.
@@
 Concretely, $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter generated by pullbacks $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$, together with admissible subgroups $\Delta_{\alpha+1}^{\uparrow}(E,D)$.
-In particular, for each active package coordinate $d$ at stage $\alpha+1$ whose countable support is witnessed by $E_{d}$, the generator $\Delta_{\alpha+1}^{\uparrow}(E_{d},\{d\})$ belongs to $\tilde{\mathscr{F}}^{*}_{\alpha+1}$, so the canonical package generics are hereditarily symmetric by Lemmas 5.47 and 5.48.
+In particular, for each active package coordinate $d$ at stage $\alpha+1$ whose countable support is witnessed by $E_{d}$, the generator $\Delta_{\alpha+1}^{\uparrow}(E_{d},\{d\})$ belongs to $\tilde{\mathscr{F}}^{*}_{\alpha+1}$, so the canonical package generics are hereditarily symmetric by Lemmas 5.50 and 5.51.
 At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group as provided by the symmetric-iteration framework (Remark 2.4(a),(b)).
 The limit-stage filter is $\tilde{\mathscr{F}}^{*}_{\lambda}$ as defined in Definition 5.53, and its normality and $\omega_{1}$-completeness are recorded in Lemma 5.58.
-For set-localization in the $\mathrm{Ord}$-length model we use only the stage-support lemma (Lemma 5.67) below.
+For set-localization in the $\mathrm{Ord}$-length model we use the forcing-localization lemma (Lemma 5.72) together with the descent lemma for set-sized final hereditarily symmetric names (Lemma 5.73) below.
 We write $\mathcal{M}$ for the final symmetric model of the $\mathrm{Ord}$-length iteration.
@@
 Let $\mathcal{M}$ denote the final symmetric model obtained from the iteration.
-Lemma 5.67 (Stage-localization of set-names at class length).
+Lemma 5.72 (Forcing-localization of set-sized class-length names).
 Let $\dot{x}$ be a set-sized $\mathbb{P}_{\mathrm{Ord}}$-name, and define its stage-support
 $I(\dot{x})\ :=\ \bigcup\{\operatorname{supp}(p)\cup I(\dot{y}):\langle\dot{y},p\rangle\in\dot{x}\},$
-where $\operatorname{supp}(p)\subseteq\mathrm{Ord}$ is the (countable) support of the condition $p$. Then $I(\dot{x})$ is a set of ordinals, hence $\alpha:=\sup(I(\dot{x}))+1<\mathrm{Ord}$. Moreover:
+where $\operatorname{supp}(p)\subseteq\mathrm{Ord}$ is the (countable) support of the condition $p$. Then $I(\dot{x})$ is a set of ordinals, hence
+$\alpha:=\sup(I(\dot{x}))+1<\mathrm{Ord}.$
+Moreover:
   1. 1.
-    $\dot{x}$ is in fact a $\mathbb{P}_{\alpha}$-name (all conditions appearing in $\dot{x}$ lie in $\mathbb{P}_{\alpha}$),
+    every subname of $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name, and in particular $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name;
   2. 2.
-    for any $V$-generic $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ and $G_{\alpha}:=G_{\infty}\cap\mathbb{P}_{\alpha}$, we have $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$.
+    for any $V$-generic $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ and $G_{\alpha}:=G_{\infty}\cap\mathbb{P}_{\alpha}$, one has $\dot{y}^{G_{\infty}}=\dot{y}^{G_{\alpha}}$ for every subname $\dot{y}$ of $\dot{x}$, and in particular $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$.
 Proof.
 Define $I(\dot{x})$ by recursion on the rank of the name $\dot{x}$.
-Since $\dot{x}$ is set-sized, Replacement and Union yield that $I(\dot{x})$ is a set of ordinals, hence bounded in $\mathrm{Ord}$.
+Since $\dot{x}$ is set-sized, Replacement and Union imply that $I(\dot{x})$ is a set of ordinals, hence bounded in $\mathrm{Ord}$.
 Let $\alpha=\sup(I(\dot{x}))+1$.
-Every condition $p$ appearing in $\dot{x}$ has $\operatorname{supp}(p)\subseteq I(\dot{x})\subseteq\alpha$, so $p\in\mathbb{P}_{\alpha}$, hence $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name.
-The valuation recursion for $\dot{x}^{G_{\infty}}$ consults only conditions appearing in $\dot{x}$, all of which lie in $\mathbb{P}_{\alpha}$, hence depends only on $G_{\alpha}$.
-Therefore $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$.
+If $\dot{y}$ is a subname of $\dot{x}$, then by construction $I(\dot{y})\subseteq I(\dot{x})\subseteq\alpha$.
+Hence every condition appearing in $\dot{y}$ has support contained in $\alpha$, so $\dot{y}$ is a $\mathbb{P}_{\alpha}$-name.
+This proves (1).
+For (2), fix a subname $\dot{y}$ of $\dot{x}$.
+Since all conditions appearing in $\dot{y}$ lie in $\mathbb{P}_{\alpha}$, the valuation recursion for $\dot{y}^{G_{\infty}}$ consults only $G_{\alpha}$, so $\dot{y}^{G_{\infty}}=\dot{y}^{G_{\alpha}}$.
+The special case $\dot{y}=\dot{x}$ gives the final statement.
+∎
+Lemma 5.73 (Stage descent of set-sized hereditarily symmetric names).
+Let $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$ be set-sized. Then there exists an ordinal $\alpha<\mathrm{Ord}$ such that:
+  1. 1.
+    every subname of $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name;
+  2. 2.
+    every subname of $\dot{x}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$;
+  3. 3.
+    for every $V$-generic $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$, with $G_{\alpha}:=G_{\infty}\cap\mathbb{P}_{\alpha}$, one has
+    $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}.$
+In particular, $\dot{x}^{G_{\infty}}\in M_{\alpha}$.
+Proof.
+For each subname $\dot{y}\in\operatorname{tc}(\dot{x})\cup\{\dot{x}\}$, since $\dot{y}\in\operatorname{HS}_{\mathrm{Ord}}$, we have
+$\operatorname{Stab}_{\mathscr{G}_{\mathrm{Ord}}}(\dot{y})\in\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}}.$
+By Lemma 5.60, choose a countable globally admissible pair $(E_{\dot{y}},D_{\dot{y}})$ at stage $\mathrm{Ord}$ such that
+$\Delta_{\mathrm{Ord}}^{\uparrow}(E_{\dot{y}},D_{\dot{y}})\subseteq\operatorname{Stab}_{\mathscr{G}_{\mathrm{Ord}}}(\dot{y}).$
+For each such $\dot{y}$, let
+$S(\dot{y}):=\{\xi<\mathrm{Ord}:\exists z\ \langle\xi,z\rangle\in D_{\dot{y}}\}.$
+Since $D_{\dot{y}}$ is countable and consists of package coordinates, $S(\dot{y})$ is a countable set of ordinals.
+Define
+$J(\dot{x}):=I(\dot{x})\cup\bigcup_{\dot{y}\in\operatorname{tc}(\dot{x})\cup\{\dot{x}\}}\bigl(E_{\dot{y}}\cup S(\dot{y})\bigr).$
+Because $\dot{x}$ is set-sized, the collection $\operatorname{tc}(\dot{x})\cup\{\dot{x}\}$ is a set, and each $E_{\dot{y}}\cup S(\dot{y})$ is countable. Hence $J(\dot{x})$ is a set of ordinals. Let
+$\alpha:=\sup(J(\dot{x}))+1<\mathrm{Ord}.$
+By Lemma 5.72, every subname of $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name, and the valuation of every subname depends only on $G_{\alpha}$.
+This proves (1) and (3), once we show that every subname belongs to $\operatorname{HS}_{\alpha}$.
+We prove by induction on rank that every subname $\dot{y}$ of $\dot{x}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$.
+So fix such a $\dot{y}$, and assume inductively that every proper subname of $\dot{y}$ belongs to $\operatorname{HS}_{\alpha}$.
+Because $E_{\dot{y}}\subseteq J(\dot{x})\subseteq\alpha$, and every coordinate in $D_{\dot{y}}$ has stage index in $S(\dot{y})\subseteq J(\dot{x})\subseteq\alpha$, we have $D_{\dot{y}}\subseteq\mathrm{Coords}_{\alpha}$.
+Moreover, the defining admissibility condition $\operatorname{Fix}(E_{\dot{y}})\leq\operatorname{Stab}_{\mathscr{G}}(d)$ for $d\in D_{\dot{y}}$ is independent of the ambient stage once the coordinates lie below that stage.
+Hence $(E_{\dot{y}},D_{\dot{y}})$ is globally admissible at stage $\alpha$.
+Let
+$k\in\Delta_{\alpha}^{\uparrow}(E_{\dot{y}},D_{\dot{y}}).$
+By Corollary 5.46, there exists
+$\hat{k}\in\Delta_{\mathrm{Ord}}^{\uparrow}(E_{\dot{y}},D_{\dot{y}})$
+with $\pi^{\alpha}_{\mathrm{Ord}}(\hat{k})=k$. Since
+$\Delta_{\mathrm{Ord}}^{\uparrow}(E_{\dot{y}},D_{\dot{y}})\subseteq\operatorname{Stab}_{\mathscr{G}_{\mathrm{Ord}}}(\dot{y}),$
+we have $\hat{k}\cdot\dot{y}=\dot{y}$. Now $\dot{y}$ is a $\mathbb{P}_{\alpha}$-name, so by Lemma 5.47,
+$k\cdot\dot{y}=\pi^{\alpha}_{\mathrm{Ord}}(\hat{k})\cdot\dot{y}=\hat{k}\cdot\dot{y}=\dot{y}.$
+Hence
+$\Delta_{\alpha}^{\uparrow}(E_{\dot{y}},D_{\dot{y}})\subseteq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{y}).$
+By Lemma 5.60, this implies $\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{y})\in\tilde{\mathscr{F}}^{*}_{\alpha}$.
+Together with the induction hypothesis for the subnames of $\dot{y}$, this shows $\dot{y}\in\operatorname{HS}_{\alpha}$. The induction is complete, so (2) holds.
+Finally, (3) is exactly the valuation clause from Lemma 5.72, and therefore
+$\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}\in M_{\alpha}.$
+Proposition 5.74 (Stage capture).
+For every $x\in\mathcal{M}$, there exists $\alpha<\mathrm{Ord}$ such that $x\in M_{\alpha}$. Equivalently,
+$\mathcal{M}=\bigcup_{\alpha\in\mathrm{Ord}}M_{\alpha}.$
+Proof.
+Write $x=\dot{x}^{G_{\infty}}$ for some set-sized $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$.
+By Lemma 5.73, there is $\alpha<\mathrm{Ord}$ such that $\dot{x}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$ and $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$.
+Hence $x\in M_{\alpha}$.
 ∎
@@
 Work in the final symmetric model $\mathcal{M}$. Recall Definition 4.8. Let $f:Y\twoheadrightarrow X$ be a surjection in $\mathcal{M}$ with $X,Y\subseteq T$.
-Choose a hereditarily symmetric set-name $\dot{f}$ such that $f=\dot{f}^{\,G_{\infty}}$.
-By Lemma 5.67, letting $\alpha:=\sup(I(\dot{f}))+1$, we may view $\dot{f}$ as a $\mathbb{P}_{\alpha}$-name and we have
-$f_{\alpha}:=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}\qquad\text{and}\qquad\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}.$
-Thus $f=f_{\alpha}\in M_{\alpha}\subseteq\mathcal{M}$.
-If $f_{\alpha}$ already has a section in $M_{\alpha}$, then $f$ splits in $\mathcal{M}$ and we are done.
-Otherwise, $f_{\alpha}$ is eligible for scheduling by the bookkeeping.
-By the final paragraph of Definition 5.62, there is some stage $\beta\geq\alpha$ such that either $f_{\alpha}$ already splits in $M_{\beta}$ (in which case we are done), or else the successor iterand at stage $\beta$ is the orbit $\mathsf{PP}$-package $\mathbb{Q}_{[f_{\alpha}]}$.
-Assume we are in the latter case, and fix such a stage $\beta$.
-Let $\dot{s}_{f_{\alpha}}$ be the canonical $\mathbb{Q}_{[f_{\alpha}]}$-name for the generic section in the $f_{\alpha}$-coordinate.
-In the full package extension, this coordinate generic yields a section of $f_{\alpha}$ (cf.
+Choose a set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{f}\in\operatorname{HS}_{\mathrm{Ord}}$ such that $f=\dot{f}^{\,G_{\infty}}$.
+By Lemma 5.73, there is $\alpha<\mathrm{Ord}$ such that $\dot{f}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$ and
+$f=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}.$
+If $f$ already has a section in $M_{\alpha}$, we are done.
+By Lemma 5.68, there is a stage $\beta\geq\alpha$ such that the successor iterand at stage $\beta$ is the orbit $\mathsf{PP}$-package $\mathbb{Q}_{[f]}$. Fix such a stage $\beta$.
+Let $\dot{s}_{f}$ be the canonical $\mathbb{Q}_{[f]}$-name for the generic section in the $f$-coordinate.
+In the full package extension, this coordinate generic yields a section of $f$ (cf.
 Proposition 4.15).
-By Lemma 5.47, $\dot{s}_{f_{\alpha}}$ is hereditarily symmetric for the package symmetric system $(\mathbb{Q}_{[f_{\alpha}]},\mathscr{G}_{[f_{\alpha}]},\mathscr{F}_{[f_{\alpha}]})$ over $M_{\beta}$.
-By Lemma 5.49, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{s}_{f_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence
-$s_{f_{\alpha}}:=\dot{s}_{f_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}.$
-Since $s_{f_{\alpha}}$ is a section of $f_{\alpha}$, it remains a section of $f$ in the final model $\mathcal{M}$.
+By Lemma 5.50, $\dot{s}_{f}$ is hereditarily symmetric for the package symmetric system $(\mathbb{Q}_{[f]},\mathscr{G}_{[f]},\mathscr{F}_{[f]})$ over $M_{\beta}$.
+By Lemma 5.52, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{s}_{f}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence
+$s_{f}:=\dot{s}_{f}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}.$
+Since $s_{f}$ is a section of $f$, it remains a section of $f$ in the final model $\mathcal{M}$.
 Therefore every $f:Y\twoheadrightarrow X\subseteq T$ in $\mathcal{M}$ splits, i.e.
@@
 In this manuscript we handle such limits by using the modified limit filter (Definition 5.53), whose normality and $\omega_{1}$-completeness are recorded in Lemma 5.58, together with the core admissibility lemma (Lemma 5.56).
-This matches the treatment of countable-cofinality limits in the general iteration framework ([5, §3.5–§3.6]).
+This matches the bounded-stage treatment of countable-cofinality limits in [5, Limit filter and Remark (Failure of stage-bounding at countable cofinality)].
 Proposition 5.77 ($A$ is not well-orderable).
@@
 Step 1: Locate $w$ at a bounded stage.
-Since $\mathcal{M}=\bigcup_{\lambda\in\mathrm{Ord}}M_{\lambda}$ (Theorem 5.59), there exists an ordinal $\lambda$ such that $w\in M_{\lambda}$.
+Since $\mathcal{M}=\bigcup_{\lambda\in\mathrm{Ord}}M_{\lambda}$ (Proposition 5.74), there exists an ordinal $\lambda$ such that $w\in M_{\lambda}$.
 Fix a hereditarily symmetric $\mathbb{P}_{\lambda}$-name $\dot{w}\in\operatorname{HS}_{\lambda}$ with $\dot{w}^{G_{\lambda}}=w$.
@@
 Choose set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-names $\dot{A},\dot{R}\in\operatorname{HS}_{\mathrm{Ord}}$ such that $A=\dot{A}^{G_{\infty}}$ and $R=\dot{R}^{G_{\infty}}$.
-By Lemma 5.67, letting $\alpha:=\sup(I(\dot{A})\cup I(\dot{R}))+1$, we have $A,R\in M_{\alpha}$.
+By Lemma 5.73, there are ordinals $\alpha_{A},\alpha_{R}<\mathrm{Ord}$ such that $\dot{A}$, viewed as a $\mathbb{P}_{\alpha_{A}}$-name, belongs to $\operatorname{HS}_{\alpha_{A}}$ and $\dot{R}$, viewed as a $\mathbb{P}_{\alpha_{R}}$-name, belongs to $\operatorname{HS}_{\alpha_{R}}$, with
+$A=\dot{A}^{G_{\alpha_{A}}}\qquad\text{and}\qquad R=\dot{R}^{G_{\alpha_{R}}}.$
+Let $\alpha:=\max\{\alpha_{A},\alpha_{R}\}$.
+By Lemma 5.62, the canonical inclusions of $\dot{A}$ and $\dot{R}$ belong to $\operatorname{HS}_{\alpha}$, and their interpretations at stage $\alpha$ are still $A$ and $R$.
+Hence $A,R\in M_{\alpha}$.
 By Lemma 5.58(iv), the stage filter $\tilde{\mathscr{F}}^{*}_{\alpha}$ is $\omega_{1}$-complete.
@@
 Choose a set-sized HS name $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$ with $x=\dot{x}^{G_{\infty}}$.
-By Lemma 5.67, letting $\alpha:=\sup(I(\dot{x}))+1$, we have $x=\dot{x}^{G_{\alpha}}\in M_{\alpha}$.
+By Lemma 5.73, there is $\alpha<\mathrm{Ord}$ such that $\dot{x}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$ and
+$x=\dot{x}^{G_{\alpha}}\in M_{\alpha}.$
 Since $M_{\alpha}\models\mathsf{SVC}(S)$ (by Lemma 5.81), there is an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow x$ in $M_{\alpha}$, hence also in $\mathcal{M}$.
@@
 Choose a set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{f}\in\operatorname{HS}_{\mathrm{Ord}}$ with $f=\dot{f}^{G_{\infty}}$.
-By Lemma 5.67, there is a set ordinal $\alpha$ such that $f\in M_{\alpha}$.
+By Lemma 5.73, there is $\alpha<\mathrm{Ord}$ such that $\dot{f}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$ and
+$f=\dot{f}^{G_{\alpha}}\in M_{\alpha}.$
 If $f$ already has a right inverse in $M_{\alpha}$, we are done.
-Otherwise, $f$ is eligible for scheduling by the $\mathsf{AC}_{\mathsf{WO}}$-bookkeeping clause in Definition 5.62.
-Hence by the guarantee clause in that definition, there exists a stage $\beta\geq\alpha$ such that either $f$ already has a right inverse in $M_{\beta}$, or else the successor iterand at stage $\beta$ includes the orbit $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{[f]}$.
-Assume the latter case and fix such $\beta$.
+By Lemma 5.68, there is a stage $\beta\geq\alpha$ such that the successor iterand at stage $\beta$ is the orbit $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{[f]}$. Fix such $\beta$.
 Let $\dot{r}_{f}$ be the canonical $\mathbb{R}_{[f]}$-name for the $f$-coordinate generic right inverse.
@@
 Theorem 5.85 (Main theorem).
-Let $V\models\mathsf{ZFC}$.
-In the class forcing extension $V[G]$ obtained by the $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen-seed symmetric model $\mathcal{N}$ described above, the final symmetric model $\mathcal{M}$ satisfies
+Let $V\models\mathsf{ZFC}$, and let $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic in the metatheoretic sense of Remark 2.1 for the $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen-seed symmetric model $\mathcal{N}$ described above.
+Then the final symmetric model $\mathcal{M}$ satisfies
 $\mathsf{ZF}\ +\ \mathsf{DC}\ +\ \mathsf{PP}\ +\ \mathsf{AC}_{\mathsf{WO}}\ +\ \neg\mathsf{AC}.$
@@
 $\mathsf{DC}$ holds by Proposition 5.78.
-$\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.68 and 5.77.
+$\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.75 and 5.84.
 $\mathsf{AC}_{\mathsf{WO}}$ holds by Proposition 5.84.
@@
 Work in $\mathcal{M}$. Fix any set $X$. By Corollary 5.82, $\mathcal{M}\models\mathsf{SVC}(S)$, so choose an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow X$.
-By Theorem 4.5 together with Propositions 5.68 and 5.77, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
+By Theorem 4.5 together with Propositions 5.75 and 5.84, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
 Pulling back the lexicographic order on $S\times\eta$ along $i$ yields a linear order on $X$. ∎
@@
 (2) $KWP_{1}$ using $\mathsf{PP}$.
-By Theorem 4.5 together with Propositions 5.68 and 5.77, $\mathcal{M}\models\mathsf{PP}$.
+By Theorem 4.5 together with Propositions 5.75 and 5.84, $\mathcal{M}\models\mathsf{PP}$.
 Apply $\mathsf{PP}$ to the surjection $e:S\times\eta\twoheadrightarrow X$ to obtain an injection $i:X\hookrightarrow S\times\eta$.